A Complete Bibliography of Publications in the Journal of the Australian Mathematical

Society: Series B, Applied Mathematics and the The ANZIAM Journal. The Australian &

New Zealand Industrial and Applied Mathematics Journal

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

25 April 2024 Version 1.05

Title word cross-reference

(2 + 1) [1601, 1511]. (N) [641]. (n · p) [902]. 0 < α < 1 [1781]. 17 [1333]. 2 [1303, 87, 1150, 759, 1313, 715]. 3 [1036, 1764, 1803]. 4 [1352]. [0, 1] [548]. _{0} [625]. A [101]. C[a, b] [665]. C^{0} [830, 857]. C^{1} [946]. d [1665]. Δ^{2}V > 0 [76]. η [1287, 1395]. exp(e^{t}) [803]. F_{1} [883]. h [1716]. H = T + V [698]. H_{∞} [1169, 1263, 1430]. i[Y, Z] = 2Y [232]. I_{∞}(M) [1339]. k [1531, 1342]. K_{2} [1131]. L [1391, 1141]. L^{2} [1609]. L^{r} [1232]. L_{2} [1424, 1402]. L_{p} [29]. LU [1404]. R^{n} [1746]. n [1168, 513, 1149]. N = 3 [1297]. N: G [1342]. o(1/k)

[1774]. o(p, q) [655]. p [761, 1412]. p(x) [1340, 1375]. P_{2} [1131]. q [1020]. QR [1404]. R^{2} [577]. S^{1} [1133]. σ [1715]. tanh^{3} y [1098]. V [1130]. W_{2}^{(1)} [767]. Y [232]. y′′ = f(x, y, y′) [355].

* [30].

-approximation [1287, 1395]. -collocation-like [857]. -complexes [1531]. -D [1303]. -data [946]. -Dimensional [1149, 1601, 1511, 759, 1313, 715]. -exponential [1232]. -fields [1715]. -implications [1716]. -integers [1020]. -Laplace [761]. -Laplacian [1340, 1375, 1412]. -matrix [1391]. -out-of- [1342]. -parametric [1036]. -point [1352, 1333]. -topology [101].

1 [582, 273, 923, 1212, 1264, 200, 1275, 931]. 1-2 [1923, 1924]. 19 [1905].

2 [583, 1276]. 2d [1849]. 2nd [823]. 2nd-order [823].

3 [584]. 3d [1894].

abandonment [1821]. Abel [1089]. Abel-type [1089]. Ablowitz [1127]. ABO [1697]. ABO-blood [1697]. above [913, 827]. ABS [1242]. absolute [1428]. absolutely [232]. abstract [1314, 1165]. Accelerated [1004, 1034]. accelerating [498]. account [841]. accumulation [536]. accuracy [458, 1663]. accurate [1864, 710]. achievements [1838]. acoustic [906, 54]. activated [1780]. Active [1925, 1372]. active-set [1372]. activity [1446]. adapted [1664]. Adaptive [856, 1789, 1624, 112, 424]. addition [797]. additional [1051]. Additive [1146]. adiabatic [1225, 1073]. adjacent [519, 1084]. adjoint [1132]. admissible [1174]. Adomian [1482, 1753, 1595, 1635]. adopting [1907]. advection [1885, 1251]. Adventures [1629]. Aedes [1551]. aegypti [1551]. aerial [1580]. Aeronautical [532]. Aesthetic [1028]. affine [1228]. after [1336]. age [1382, 1185, 1738]. age-structured [1185]. Aggregation [22]. Aggregation-decomposition [22]. aimed [692]. air [1885, 267, 394, 5]. aircraft [445]. aircraft-like [445]. Airfoils [394, 467]. AKNS [524]. algebra [698, 12]. algebraic [90, 1697, 952]. algebras [78, 348, 1744]. Algorithm [1872, 216, 1696, 1500, 126, 146, 550, 1576, 311, 1212, 1539, 1392, 79, 30, 716, 1481, 1253, 1240, 1311, 676, 132, 1907, 1467, 779, 1228]. algorithms [1666, 1351, 1264, 1786, 1179, 1490, 815, 338, 482, 1711, 961]. Allee [1515]. allocation [1592]. allometric [865]. alluvial [638]. almost [1, 395, 1140, 1724]. almost-periodic [395]. along [392, 899, 1359]. alpha [1863]. alpha-beta-rho [1863]. alternate [561]. alternating [1774]. alternative [1841, 1709]. Alzer [1195]. Ambiguity [1427]. American [1472, 786, 1497, 1861]. American-style [1861]. amount [1407]. amplitude [906, 1685, 724, 458, 308, 373]. analogue [1118]. analyses [1796]. analysing [1730]. Analysis [616, 741, 1490, 1094, 1854, 1288, 1005, 1455, 1269, 924, 194, 299, 1570, 332, 1688, 1884, 724, 1278, 536, 1321, 531, 1519, 1725, 996, 555, 1849, 81, 1607, 1595, 991, 695, 1294, 759, 815, 1193, 1282, 1515, 246, 112, 1122, 775, 1304, 288, 406, 460, 1342, 1922, 624]. Analytic [921, 6, 780, 561, 1367, 1445]. Analytical [862, 1657, 919, 1739, 1741, 1732, 1885, 1893, 1813, 1860, 1497, 1675, 1560, 1687, 1862, 1674, 892, 1659, 1309]. angle [77, 316, 1281, 585, 1313]. angles [467]. angular [1797]. anharmonic [795, 853]. animals [1175]. anisotropic [602, 566, 115, 184, 1154, 1649]. annular [277]. antecedents [681]. anti [292, 37, 49]. anti-cyclonic [292]. anti-plane [37, 49]. antibodies [1265]. antiplane [1508]. Anz [1855, 1856, 1866, 1867, 1873, 1874, 1881, 1882, 10, 11, 23, 24, 34, 35, 47, 48, 59, 60, 72, 73, 85, 86, 95, 96, 108, 109, 116, 117, 130, 131, 156, 157, 169, 170, 144, 145, 181, 182, 197, 198, 206, 207, 214, 215, 224, 225, 237, 238, 247, 248, 258, 259, 269, 270, 284, 285, 294, 295, 305, 306, 314, 315, 325, 326, 335, 336, 345, 346, 356, 357, 369, 370, 379, 380, 388, 389, 399, 400, 410, 411, 422, 423, 435, 436]. ANZ [446, 447, 455, 456, 464, 465, 474, 475, 485, 486, 496, 497, 507, 508, 517, 518, 528, 529, 541, 542, 551, 552, 559, 560, 568, 569, 580, 581, 589, 590, 598, 599, 608, 609, 617, 618, 629, 630, 639, 640, 648, 649, 660, 661, 670, 671, 679, 680, 690, 691, 700, 701, 708, 709, 717, 718, 728, 729, 739, 740, 752, 753, 762, 763, 771, 772, 781, 782, 790, 791, 801, 802, 813, 814, 825, 826, 836, 837, 850, 851, 860, 861, 873, 874]. ANZ [884, 885, 894, 895, 904, 905, 915, 916, 926, 927, 936, 937, 947, 948, 962, 963, 972, 973, 985, 986, 997, 998, 1010, 1011, 1030, 1031, 1040, 1041, 1052, 1053, 1065, 1066, 1081, 1082, 1090, 1091, 1101, 1102, 1113, 1114, 1134, 1144, 1145, 1157, 1158, 1171, 1172, 1183, 1184, 1198, 1199, 1210, 1211, 1223, 1224, 1237, 1238, 1248, 1249, 1260, 1261, 1271, 1272, 1284, 1285, 1295, 1296, 1305, 1306, 1318, 1319, 1331, 1332, 1345, 1346, 1357, 1358, 1370, 1371, 1380, 1381, 1393, 1394, 1449]. ANZ [1450, 1460, 1461, 1470, 1471, 1483, 1484, 1491, 1492, 1501, 1502, 1512, 1513, 1521, 1522, 1529, 1530, 1535, 1536, 1541, 1542, 1547, 1548, 1554, 1555, 1561, 1562, 1572, 1573, 1582, 1583, 1589, 1590, 1596, 1597, 1604, 1605, 1612, 1613, 1620, 1621, 1627, 1628, 1636, 1637, 1644, 1645, 1651, 1652, 1660, 1661, 1669, 1670, 1682, 1683, 1691, 1692, 1699, 1700, 1706, 1707, 1735, 1736, 1747, 1748, 1756, 1757, 1767, 1768, 1776, 1777, 1783, 1784, 1791, 1792, 1799, 1800, 1806, 1807, 1814, 1815, 1822, 1823, 1828]. ANZ [1829, 1836, 1837, 1847, 1848, 1888, 1889, 1896, 1897, 1903, 1904, 1911, 1912, 1923, 1924, 1930, 1931]. AOR [1391]. AOR-type [1391]. aorta [171]. aperture [1078]. Apex [421]. apparatus [1922]. apparatus-induced [1922]. Apparent [980, 1844]. Appell [883]. Application [1786, 1704, 428, 509, 685, 319, 990, 922, 1046, 1363, 311, 1789, 516, 1854, 774, 746, 819, 183, 494, 445, 561, 855, 993]. Applications [211, 558, 1107, 1708, 1891, 1382, 810, 1423, 354, 612, 822, 1841, 1340, 300, 1459, 823, 1473, 1663, 1657, 1456, 1805, 968, 1366, 809, 1452, 944, 1267, 1465]. Applied [582, 194, 923, 1549, 1609, 601, 868, 1265]. approach [1704, 1732, 880, 1024, 426, 1788, 1908, 437, 1653, 1230, 1630, 1753, 161, 261, 1743, 730, 1473, 1745, 1851, 1282, 1534, 1487, 1674, 74, 1127, 1359, 917]. approaches [1740]. Appropriate [545, 1753]. Approximate [1673, 1927, 1023, 946, 245, 1833, 1203, 118, 1462, 667, 1495, 1845]. approximately [113, 1526]. Approximating [1588, 378]. Approximation [1712, 1812, 1287, 1395, 491, 758, 344, 114, 1841, 1375, 1893, 1860, 1036, 1731, 162, 202, 226, 1311, 1021, 875, 187, 448]. approximations [449, 607, 751, 745, 859, 883, 172, 1497, 849, 882, 561, 766, 842]. arbitrage [1863]. arbitrarily [602, 483]. arbitrarily-oriented [602]. arbitrary [332, 800, 1463, 585, 637, 736, 467, 577]. arc [19, 693]. arcs [1021]. area [1546]. argument [820]. arguments [1425, 79]. arise [256]. arising [1003, 1666, 533, 900, 1050, 534, 249, 1545, 1075, 1201, 1506, 1493, 1574]. Arithmetic [1425]. arms [888, 543]. arrest [631]. Arrhenius [219]. arrival [1886]. arterial [100]. artery [1698]. ascent [1032]. ashfall [1396]. Asian [1677]. Aspects [755, 111, 1566, 1080]. assembly [1070]. assessed [1565]. Assessing [499]. asset [1827, 1676, 1862, 1453, 1727]. asset-liability [1827]. assets [1827, 1740]. assignment [110, 1895]. associated [299, 313, 754, 1427]. assumptions [1151]. Asymmetric [122, 150]. Asymmetrical [1853]. Asymptotic [547, 82, 788, 1062, 1387, 1849, 669, 365, 1826, 390, 1236, 45, 307, 789, 234, 240, 533, 374, 395, 839, 531, 1459, 178, 545, 256, 1329, 515]. asymptotically [371]. Asymptotics [1843, 1303, 651, 1865]. Asymptotology [1119]. atherosclerotic [1557]. atmospheric [1685, 265, 1281]. attaching [1045, 1265]. attachment [404]. attack [1281]. attenuance [1263]. attractivity [838, 1008, 1194, 1164]. attractors [1189, 756]. AUD [1678]. AUD/usd [1678]. Australia [1905]. Australian [747]. automatic [550]. autonomous [668, 734, 698]. availability [1834]. average [1108]. Averaging [993]. Avian [1494]. axi [1543]. axi-symmetric [1543]. axial [480, 737, 1463, 32]. axially [13, 80]. Axisymmertic [300]. Axisymmetric [1793, 1527, 1155, 1615, 1552, 1870, 984]. azimuthal [1638].

B [471, 1069, 1067, 1400, 1569, 1698]. B-spline [1400]. Back [11, 24, 35, 48, 60, 73, 86, 96, 109, 117, 131, 157, 170, 145, 182, 198, 207, 215, 225, 238, 248, 259, 270, 285, 295, 306, 315, 326, 336, 346, 357, 370, 380, 389, 400, 411, 423, 436, 447, 456, 465, 475, 486, 497, 508, 518, 529, 542, 552, 560, 569, 581, 590, 599, 609, 618, 630, 640, 649, 661, 671, 680, 691, 701, 709, 718, 729, 740, 753, 763, 772, 782, 791, 802, 814, 826, 837, 851, 861, 874, 885, 895, 905, 916]. Back [927, 937, 948, 963, 973, 986, 998, 1011, 1031, 1041, 1053, 1066, 1082, 1091, 1102, 1114, 1145, 1158, 1172, 1184, 1199, 1211, 1224, 1238, 1249, 1261, 1272, 1285, 1296, 1306, 1319, 1332, 1346, 1358, 1371, 1381, 1394, 1450, 1461, 1471, 1484, 1492, 1502, 1513, 1522, 1530, 1536, 1542, 1548, 1555, 1561, 1573, 1582, 1590, 1596, 1605, 1613, 1621, 1628, 1637, 1645, 1652, 1661, 1670, 1683, 1692, 1700, 1707, 1736, 1748, 1757, 1768, 1777, 1784, 1792, 1800, 1807, 1815, 1823, 1829, 1837, 1848, 1856, 1867]. back [1874, 1882, 1889, 1897, 1904, 1912, 1924, 1931, 1253]. backgammon [137]. backgammon-like [137]. Bäcklund [685, 1128, 1130]. backtesting [1678]. backtracking [1228]. backward [1454, 794]. bacteriophage [1844]. bacterium [199]. bagasse [1071, 1917]. ball [1244]. Banach [1314]. band [51]. bandlimited [748]. bandwidth [1702, 1902]. bang [1624]. bang-bang [1624]. bargaining [213]. barrier [81, 123, 929, 1610, 459, 1787, 155, 1002]. barriers [880, 1729, 1478, 1327, 896]. Barzilai [1789]. based [1466, 1879, 1352, 1908, 1500, 1169, 1540, 161, 261, 1539, 565, 1468, 1479, 1830, 1254, 1545, 1901, 1268, 1020, 1481, 1516, 1414]. basic [194, 724]. basin [1701, 155]. basis [1879, 1863]. basket [1080]. basket-heating [1080]. Batch [1667, 1073]. Batchelor [1334]. batteries [1361]. bauxite [1919]. Bayesian [1294]. be [1320, 967, 1265]. beam [1463, 1137]. beams [877, 1907]. bearing [1208, 153]. bearings [956, 1235]. bearings-only [956]. became [1068]. behavior [547, 839]. behaviour [1602, 234, 1080, 82, 387, 788, 1062, 1826, 1236]. behind [1281]. bell [1407]. Bellman [1592]. below [492]. Bénard [1096, 1431]. benchmark [1495]. beneath [600, 1289]. benefit [1625]. Bernstein [1020]. Bessel [1017, 733, 397, 244, 180, 333, 405]. Best [226, 1389, 1788, 1386, 1036, 1428]. beta [1168, 1863]. between [173, 1754, 1584, 871, 1765, 520, 179, 1646, 586, 1096, 848, 971, 644, 523, 1154, 32, 591, 834]. beyond [1537, 1553, 1581]. Bézier [1188, 775]. BFGS [1215, 550]. BH3 [199]. Bi [1745, 1788, 1895]. Bi-level [1745]. bi-objective [1788, 1895]. Bianchi [983, 625]. Biased [1598]. bicriterion [945]. bifuraction [311]. Bifurcation [1688, 97, 1043, 1200, 1257, 706, 1884, 564, 632, 1607, 122]. bifurcations [453, 373]. big [1711]. biharmonic [1088]. Bilateral [1057, 1162, 891, 431, 1110, 1156]. bilayer [1655]. Bilinear [891, 1587, 1110]. binaries [1453]. Binary [1129]. bio [1523]. bio-mathematical [1523]. biodegradation [1780]. bioeconomic [1209, 1247]. Biological [1759, 1650, 521, 1586, 918, 1033]. biology [1758]. Biorthogonal [1376, 1233, 1898, 1311]. Biot [1075]. biphasic [1925]. birth [933, 792]. Bishop [135]. bivariate [1006]. Black [1123, 979, 1453, 1864]. Blake [1759]. block [1212, 1840, 1455]. block-GI [1212]. block-GI/M/1 [1212]. block-sor [1455]. blood [1697, 100, 4, 1698]. Blow [1601, 1416]. Blow-up [1601, 1416]. blowflies [1909]. Blowup [1323]. BMAP [931]. BMAP/G/1 [931]. BMP [1322]. bodies [1281, 1313, 1550, 1359]. body [392, 1525, 949]. Boltzmann [41]. Boltzmann-type [41]. Bond [1771]. Bondi [1374]. bonds [1893, 1309]. Bonhoeffer [862]. boosting [1567]. Booth [1273]. BOPS [1705]. Borel [1715]. borrowing [907]. Borwein [1178, 1789]. Bose [164, 158]. both [1704, 1389]. bottom [364, 1640]. Bound [1741, 1432, 1599]. bound-constrained [1599]. boundaries [340, 881]. Boundary [526, 667, 405, 493, 1161, 1026, 106, 272, 641, 738, 228, 685, 988, 1227, 856, 1049, 329, 1038, 89, 327, 165, 184, 360, 437, 1291, 1804, 547, 179, 439, 1673, 350, 623, 727, 1365, 1100, 877, 1132, 1559, 1229, 253, 1786, 519, 1497, 1397, 1663, 1143, 383, 601, 1043, 1487, 705, 321, 183, 694, 787, 1262, 1084, 1578, 1526, 1093, 1218, 1412, 262, 98, 1201, 1458, 1266, 902, 461, 1379, 479, 210]. boundary-element [856]. boundary-integral [601]. boundary-value [641, 685, 179, 877, 902, 210]. bounded [414]. boundedness [22]. Bounds [1677, 304, 1520, 1878, 1389, 272, 1300, 553, 922, 340, 1307, 172, 1047, 561, 855, 1037]. Boussinesq [1803]. Bow [707, 1045, 1173]. bows [404]. Braess [702]. Bragg [807]. branches [273]. Branching [1895, 1764]. breadth [1538]. breakdown [673]. breaking [706]. breakthrough [1207]. Brenner [1195]. Brian [1068]. bridging [965]. brief [1029]. British [1673]. broadband [554]. broadcasts [1255]. broader [1702, 1902]. broken [1163]. Brown [990, 746, 819]. Brownian [1908, 1451, 780]. Broyden [418]. Broyden-like [418]. Bruno [1349]. Brusselator [1739, 453]. bubble [228, 509, 329, 1204, 1299]. Bubbles [899, 300]. Buckling [332, 566]. budgets [1055]. buoyant [653]. Burgers [1579]. burning [711, 1850]. Bursting [1918]. Burundi [1877]. bushfire [887]. business [1197]. buy [1820]. buy-online [1820].

C [1838, 1069]. cable [920]. Calculating [1377]. Calculation [1755, 1099, 363, 1095]. calculations [500]. Calculus [1574, 1584, 1676]. calibration [1696]. call [1810]. calls [1675]. calorimeter [178]. can [967, 1507, 1265]. cancer [877, 1574]. canonical [353]. capacity [1834]. capillaries [513]. capillarity [868]. capillary [77, 1733, 1106, 1702, 1902, 1928]. capillary-gravity [1702]. capital [536, 1600, 343, 406]. carbon [1584]. carcass [865]. cardiac [631]. Carlo [1553, 1581, 1537, 1740]. carrying [1057, 1834]. Carslaw [1559]. CARTOPT [1599]. case [1523, 1843, 1505]. cases [274]. cash [1672]. cash-in [1672]. cash-out [1672]. caster [719]. catalytic [167]. catchment [1701]. Cauchy [1146, 846, 768, 1039, 1021]. Cauchy-type [1021]. causal [127]. caused [597]. cautionary [1119]. cavitation [509, 329, 1429]. Cavities [1281, 92, 129, 334, 70, 93]. cavity [789, 267, 1078, 1401]. celebrating [1838]. cell [743, 468, 1382, 1762, 328, 363, 534, 1094, 1854, 1845, 1506, 1493, 1664, 1750, 1574, 1853]. cell-types [468]. cells [575, 1876, 487]. central [1410, 1714]. Centre [613, 494, 674, 833]. centre-manifold [494]. Certain [1110, 969, 615, 6, 344, 678, 183, 1156]. Cesari [1180]. Cesari-type [1180]. chain [1754, 254, 1212, 1264, 1122]. chain-thermal [254]. chains [319, 999, 1086, 1405, 1177]. Chakalov [1099]. channel [1143, 970, 1560, 1640, 1805, 653, 1002, 282]. channels [833]. Chaos [457, 795, 1048]. Chaotic [673, 1602, 1842]. characterisation [1337]. Characteristic [348, 88, 78, 606]. characteristics [1121, 100, 1373]. characterization [902]. Characterizations [1715, 99]. charged [439, 279]. charging [1361]. Charwat [298]. Chebyshev [1188, 1022, 675, 1499]. chemical [167, 816, 878]. chemistry [219]. chemostat [620]. chemotherapy [968]. Cherry [531]. chiral [1773, 1297, 1486]. chlorophyll [1648]. choice [907]. cholera [1477]. cholesteric [32]. CIR [1813]. circle [1427, 260]. circuit [441]. circuits [257]. circular [19, 725, 1334, 770, 290, 32, 1231, 243, 1480, 271]. claim [1407]. clamped [65]. class [449, 106, 103, 1049, 524, 656, 184, 82, 212, 760, 761, 1097, 1229, 1473, 1194, 366, 79, 193, 879, 80, 321, 1042, 1156, 968, 1626, 419, 1452, 322, 776, 840, 430]. classes [1057]. classical [1804, 337, 1229, 180]. classification [1711]. clearance [1749]. clipped [27, 161, 261]. close [394]. Closed [439, 457, 1622, 1744, 1618, 1309]. closed-form [1622, 1618, 1309]. closely [359]. closure [1518]. Cloud [1429]. clouds [1871]. clue [1568]. Cluster [1034, 1849, 1658]. coarse [299]. coaxial [1920]. cobweb [1217, 1230]. cocoercive [1385]. coded [1696]. coefficient [1039, 1390]. coefficients [165, 255, 1356, 1329]. cognitive [1717]. Coherent [164]. cohesionless [1205]. Coifman [1233]. coil [1095]. cointegrated [1827]. collapse [1374]. collapsible [1805]. collision [1525]. collisions [538]. Collocation [190, 675, 1027, 1049, 623, 727, 823, 903, 565, 857, 383, 1689, 1634, 1400]. combination [616]. combinatorial [604]. Combined [1507, 1630, 1213, 1624, 527]. Combustion [1077, 864, 1067, 1176, 845, 1079, 1072, 1755, 1225, 1639, 1068, 163, 204, 241, 281, 403, 546, 621, 867, 1312]. common [38, 1703]. communication [1808]. community [1160, 1844]. commutation [232]. commuting [730]. comparative [764, 1523, 842, 1740, 1379]. Comparison [612, 1497, 1061, 192, 866, 1107, 1532, 678, 1290, 1312, 1899, 943, 213]. competing [428, 1247]. competition [1602, 159, 252, 1096, 620, 1246, 1033]. Competitive [1725, 468, 1591, 605, 1842]. complementarity [1643, 313, 1409, 731]. Complementary [106]. complete [147]. completely [1812, 210, 1364, 200]. Completion [1614, 834]. Completion-Time [834]. completions [1136]. Complex [1730, 768, 1720, 1774, 1755, 1713, 1831]. complexes [1531]. component [337, 815]. Components [1442]. composed [556]. composing [1310]. composition [865]. compressible [1072, 1915, 49, 1639]. Computable [1405]. Computation [241, 1852, 845, 706, 1439, 841, 1018]. Computational [1313, 1721, 730, 1473, 79, 1411, 132, 445, 527, 776, 1671]. computations [1025]. Computer [1333, 662]. Computer-controlled [662]. Computing [935, 1182, 1746]. concentration [26, 183]. Concentric [19]. concept [1709]. conceptual [1446]. concerning [665]. Concise [417]. condensation [1408]. condensed [1654]. condensed-matter [1654]. condenser [1231]. condition [1153, 547, 1559, 443, 1487, 1526]. conditional [1044, 964, 322]. conditioning [500]. Conditions [1033, 1206, 1227, 1038, 89, 349, 42, 526, 983, 1151, 1007, 350, 1070, 1100, 1132, 1307, 946, 820, 330, 645, 1180, 1436, 1225, 1901, 481, 1877, 705, 545, 667, 1909, 882, 98, 1329, 945, 1672]. conducting [1243, 371, 1084, 1056]. conduction [1559]. conductive [1773]. conductivity [800, 650, 546]. conductor [205, 827]. conductors [1214, 736]. cone [481]. configurations [1868]. Conformal [1549, 1830]. congruency [924]. coning [1801, 1615, 601]. conjecture [1413]. conjugate [1796, 550, 1576, 989]. conjugated [1441]. connected [155, 577]. connection [871, 1127]. connections [1055]. connector [1835]. conquer [793]. Conservation [1202, 1696, 1546, 1373, 1684]. conservative [1797, 57]. consistent [1625]. consisting [567]. constant [1298, 652, 149, 1386, 623, 727, 1039, 255, 1045, 817, 1326, 1309]. constant-coefficient [1039]. constant-rate [652]. constant-sign [1298]. constants [276, 1165]. constrained [1188, 745, 421, 1372, 820, 1592, 1179, 1129, 1830, 705, 1599, 592, 657, 941, 960, 502, 1228]. Constraint [525, 658, 1239, 160, 943]. constraints [480, 208, 1643, 1617, 114, 1340, 1307, 1633, 1631, 1313, 481, 1015, 99, 527, 809, 1680]. construct [1310, 1872]. constructing [1240, 1311]. construction [833]. constructive [875]. consumer [1304]. consumption [1790, 907, 1869]. consumption/portfolio [907]. Contact [1517, 20, 77, 1359]. contacts [804]. containers [1312]. containing [1801, 1152]. containment [1085]. contaminant [519]. contaminants [307]. Content [1009, 1080]. Contents [46, 94, 143, 196, 236, 283, 324, 368, 409, 454, 495, 540, 579, 659, 699, 872, 914, 1064, 1112, 1170, 1222, 1270, 1317, 1369]. continental [51]. Continuation [626, 63, 614, 1624]. Continuity [756, 273, 450]. Continuous [822, 642, 988, 6, 999, 232, 536, 174, 212, 213, 1108, 291, 1405, 952, 404, 1140, 52, 99, 966, 677, 1348, 719, 479]. Continuous-time [822, 999, 1405, 952, 966, 677]. continuously [858, 1203, 1200, 1257]. continuum [1122]. Contois [1650, 1602]. contour [616]. contours [329]. contribution [1625]. Control [642, 44, 767, 966, 1177, 473, 1255, 28, 62, 208, 633, 89, 233, 1162, 107, 526, 1286, 396, 587, 1315, 1007, 1488, 1169, 1540, 788, 1372, 576, 1778, 750, 841, 765, 330, 1229, 888, 1239, 1384, 1473, 1479, 1851, 1631, 1624, 79, 381, 538, 105, 1634, 338, 1060, 879, 1587, 1277, 935, 74, 1414, 424, 676, 1368, 132, 419, 445, 527, 592, 657, 941, 1103, 809, 908, 1001, 405, 322, 878, 1558, 794, 840, 945, 1690, 502, 1430, 1680]. control-state [1372]. Controllability [663, 749, 502, 808, 1343, 493]. controllable [634]. controlled [634, 840, 662]. controller [1720, 1580, 1781]. controls [1180, 52, 98, 592, 890]. controversy [260]. convection [304, 1096, 1507, 1898, 646, 160, 312, 674, 647, 876, 1431, 16]. convective [265, 1458]. Convergence [745, 696, 759, 39, 17, 71, 1027, 347, 195, 418, 1372, 1726, 123, 1774, 1405, 112, 1392, 459, 893, 784]. convergent [38, 242]. Converging [256]. convertible [1893, 1309]. convex [414, 42, 666, 250, 1633, 835, 1425, 1274, 99, 367, 1359, 943]. convexity [758, 413, 586, 386, 1216]. convolution [934, 1253]. Cooke [25, 84]. Cooling [654, 832, 1507]. cooperation [1808]. cooperative [1062]. coordinate [1852, 1269]. coordinate-free [1852]. coordinates [1752, 132]. coordination [1718]. coplanar [602]. coral [1568]. core [1773, 832]. corner [1731]. corners [1399]. corrected [1301]. Correction [537, 1469, 1553, 1581, 191, 71]. corrections [189, 1647]. corresponding [1131, 1376]. Corrigendum [1344]. corrugated [1154]. cosine [1322, 1308]. cosmic [978, 40, 64]. cosmic-ray [40, 64]. cosmological [1891, 625, 979]. cosmologies [58]. Cost [737, 1694, 28, 822, 1315, 1488, 1108, 750, 1820, 961]. Cost-time [737]. costs [330, 676]. Couette [268, 925, 1662, 1914].

Coulomb [337]. Couple [1751]. coupled [1511, 1641, 1316, 1478, 1610, 366, 920, 1575, 245]. coupler [457]. coupling [533, 1468]. covariant [125]. Cover [10, 11, 23, 24, 34, 35, 47, 48, 59, 60, 72, 73, 85, 86, 95, 96, 108, 109, 116, 117, 130, 131, 156, 157, 169, 170, 144, 145, 181, 182, 197, 198, 206, 207, 214, 215, 224, 225, 237, 238, 247, 248, 258, 259, 269, 270, 284, 285, 294, 295, 305, 306, 314, 315, 325, 326, 335, 336, 345, 346, 356, 357, 369, 370, 379, 380, 388, 389, 399, 400, 410, 411, 422, 423, 435, 436, 446, 447, 455, 456, 464, 465, 474, 475]. Cover [485, 486, 496, 497, 507, 508, 517, 518, 528, 529, 541, 542, 551, 552, 559, 560, 568, 569, 580, 581, 589, 590, 598, 599, 608, 609, 617, 618, 629, 630, 639, 640, 648, 649, 660, 661, 670, 671, 679, 680, 690, 691, 700, 701, 708, 709, 717, 718, 728, 729, 739, 740, 752, 753, 762, 763, 771, 772, 781, 782, 790, 791, 801, 802, 813, 814, 825, 826, 836, 837, 850, 851, 860, 861, 873, 874, 884, 885, 894, 895, 904, 905, 915, 916]. Cover [926, 927, 936, 937, 947, 948, 962, 963, 972, 973, 985, 986, 997, 998, 1010, 1011, 1030, 1040, 1052, 1065, 1081, 1082, 1090, 1091, 1101, 1102, 1113, 1114, 1134, 1144, 1145, 1157, 1158, 1171, 1172, 1183, 1184, 1198, 1199, 1210, 1211, 1223, 1224, 1237, 1238, 1248, 1249, 1260, 1261, 1271, 1272, 1284, 1285, 1295, 1296, 1305, 1306, 1318, 1319, 1331, 1332, 1345, 1346, 1357, 1358, 1370, 1371, 1380, 1381, 1393, 1394, 1449, 1450, 1460, 1461, 1470, 1471, 1483, 1484, 1491, 1492, 1501, 1502, 1512]. Cover [1513, 1521, 1522, 1529, 1530, 1535, 1536, 1541, 1542, 1547, 1548, 1554, 1555, 1561, 1562, 1572, 1573, 1582, 1583, 1589, 1590, 1596, 1597, 1604, 1605, 1612, 1613, 1620, 1621, 1627, 1628, 1636, 1637, 1644, 1645, 1651, 1652, 1660, 1661, 1669, 1670, 1682, 1683, 1691, 1692, 1699, 1700, 1706, 1707, 1735, 1736, 1747, 1748, 1756, 1757, 1767, 1768, 1776, 1777, 1783, 1784, 1791, 1792, 1799, 1800, 1806, 1807, 1814, 1815, 1822, 1823, 1828, 1829, 1836, 1837, 1847, 1848, 1855, 1856, 1866, 1867, 1873, 1874, 1881]. cover [1882, 1888, 1889, 1896, 1897, 1903, 1904, 1911, 1912, 1923, 1924, 1930, 1931]. Coverage [1717]. covers [1136]. COVID [1905]. COVID-19 [1905]. coxial [32]. CPDO [1672]. crack [19, 115, 1508]. cracks [602]. credit [1860]. Creeping [1526, 1646]. crested [463]. cresting [712, 714]. criteria [1287, 1498, 549, 961, 430]. criterion [1108, 1625, 1250]. Critical [1801, 442, 387, 645, 1575, 1817, 1057, 1160, 854, 1386, 1070, 1532, 1427, 1436, 1755, 1878, 484, 241, 892]. Criticality [403, 546, 438]. cross [725, 1080, 1463, 637, 193]. cross-section [725]. cross-validation [193]. crossing [1775, 977, 1074]. cruise [1479]. crystal [32]. CSTR [1073]. cubic [988, 1283, 1181]. Current [1702, 724, 1902]. Current-modified [1702]. currents [489, 919, 1928]. curve [1694, 967, 996, 1193, 716]. curved [1785, 1093, 949, 1266]. curves [1188, 992, 1220, 377, 775]. curvilinear [1695]. Cusp [364, 317]. Cusp-like [364, 317]. cusped [433]. customers [1821]. cutoff [492]. cycle [1197]. cycles [1382, 159]. Cyclic [286, 1361]. cyclone [289]. cyclones [747]. cyclonic [292]. cylinder [1334, 827, 770]. cylinders [55, 848, 971, 32]. cylindrical [566, 832, 92, 129, 334, 628, 684]. cytoskeleton [1762].

D [1303, 87, 1150, 1764, 1803]. Dai [1796]. dam [43]. damped [582, 583, 584, 673]. damping [155]. Darboux [1879, 809]. Darling [1701]. data [764, 988, 1710, 522, 1321, 1609, 967, 1833, 946, 1474, 1713, 163, 403, 438, 546, 1711]. Dean [511]. death [933, 792]. Deblurring [1383]. decay [1137]. decentralized [107]. decision [1705, 1108, 166, 869]. Decomposition [1197, 1360, 1482, 1753, 22, 1635, 1275, 1276]. dedicated [1719]. deduction [1467]. deep [930, 286, 490, 1106, 1902, 603, 585]. deep-water [585]. deeply [290]. default [1860]. defective [863]. Deferred [191, 189]. defined [323, 505, 1625]. definite [1825, 1182, 1213, 1545, 1263]. definitions [1532]. deformable [686, 1244]. deformation [738, 55, 1463, 276, 1359]. deformations [61]. deformed [1741, 1907]. deformity [1163]. degeneracy [91]. Degenerate [563, 453, 123, 1152, 1510, 39, 1246]. degenerate-kernel [39]. Degree [1188, 775, 428]. degrees [1292]. Delay [1378, 1430, 264, 1688, 1884, 233, 611, 1169, 1616, 605, 696, 1468, 1790, 664, 1324, 1194, 1290, 98, 1574, 1558, 1267, 1328, 1191, 1221]. Delay-dependent [1378, 1430, 1468]. delay-differential [98]. Delayed [320, 838, 1540, 1600, 79, 1001, 1234]. delays [1754, 1315, 159, 309, 1008, 1314, 1476, 1607, 138, 1417, 1140, 1465, 1430]. delivery [1718, 1820]. demand [1745, 1681, 1820]. demands [1725]. Demianski [506]. Demianski-type [506]. demodulation [1730]. demon [1737]. dengue [1551]. denoising [1383]. density [316, 933, 1638, 900, 401, 1175, 870, 636, 1922]. density-dependent [933, 1922]. Denumerable [1108]. dependence [1703, 403, 438, 546, 621]. dependency [988]. dependent [1104, 1496, 234, 933, 319, 1638, 1607, 1410, 56, 1468, 1378, 1862, 931, 636, 546, 1368, 1191, 1922, 1430]. depletion [1077]. deposit [907]. deposition [1396]. depth [913, 6, 1798, 811, 628, 1902, 707, 1135, 1775, 463, 466, 688, 1075, 1338, 1928, 721, 896, 911]. depth-uniform [1902]. Derivation [1175, 62, 490, 33, 373, 1428]. Derivative [1032, 1901, 1887]. Derivative-type [1032]. Derivatives [797, 1927]. describe [1925]. describing [1570]. description [594, 1809]. descriptions [545]. descriptor [1315]. Design [1656, 704, 554, 331, 1580, 1219, 1414, 1095, 1921, 1452, 1722, 1781]. destructive [209]. Detachment [218, 217]. Detecting [1531]. Detection [257, 924]. determinant [780]. determinantal [1486]. Determination [1584, 633, 634, 441, 1427, 56, 1436]. determined [66]. Determining [477, 1163, 575, 1264]. Deterministic [1477, 1396, 1565]. developed [637]. developing [798]. development [1611, 1316]. deviating [820]. deviation [149]. deviations [933, 1428]. Devil [1570]. Diabatic [1200]. diagnostics [1679]. Diagonal [1026]. diagram [1139]. diameter [110]. dichotomy [1817]. diet [1538]. Diffeomorphisms [1133]. difference [537, 866, 1044, 1107, 1300, 231, 31, 203, 339, 318, 362, 382, 442, 1003, 656, 1797, 1291, 347, 1419, 1469, 82, 839, 280, 351, 678, 759, 121, 202, 226, 407, 1364, 176, 849, 882, 151, 168, 1109, 1236, 1259, 1740, 1061]. differences [1520]. differencing [1850]. different [1154, 1577]. differentiability [414]. differentiable [1595, 1203]. Differential [890, 969, 272, 866, 1058, 1087, 1044, 1107, 323, 468, 231, 31, 264, 1339, 633, 1027, 1189, 1003, 89, 656, 533, 611, 1609, 1008, 741, 1050, 1314, 1046, 534, 575, 757, 539, 1598, 1343, 669, 1190, 765, 903, 957, 50, 1131, 443, 1786, 730, 1239, 353, 695, 1397, 1194, 1290, 1456, 200, 210, 222, 1929, 366, 105, 964, 935, 1454, 667, 1499, 1845, 98, 1202, 1506, 1493, 1124, 333, 405, 1267, 430, 1221]. differential-difference [866, 1044, 1107, 231]. diffraction [92, 129, 334]. Diffusion [720, 1265, 1739, 641, 505, 685, 1906, 1051, 273, 274, 376, 1885, 1858, 387, 651, 954, 1363, 1532, 152, 375, 1251, 644, 1898, 1731, 1316, 275, 759, 1524, 249, 1071, 1876, 1510, 1514, 1824, 1487, 705, 221, 1870, 646, 778, 824, 987, 1464, 647, 876, 620, 1018, 715, 910, 1323, 1577, 1465, 1869, 1680]. diffusion-convection [647, 876]. diffusive [304, 696, 1086]. diffusivities [375, 817]. diffusivity [152]. digital [556, 1229]. dilation [818]. Dilaton [979]. dimension [1189]. Dimensional [1149, 75, 359, 239, 764, 328, 363, 1603, 111, 1885, 100, 711, 887, 1601, 1858, 190, 1100, 1281, 665, 83, 227, 1511, 991, 1537, 1553, 1581, 899, 759, 1078, 1313, 221, 901, 153, 715, 1411, 692, 1842, 1840, 1654, 1445, 943, 1232, 1455, 747]. dimensions [1665, 1906, 1817, 1054, 993]. dinner [1069]. dipole [9]. Dirac [1654]. Direct [1695, 1087, 1351, 788, 56, 1742, 1379]. directed [257]. Direction [1250, 550, 1774]. directional [1713]. directions [989]. Dirichlet [1032, 1375, 1100]. disappearance [1844]. discontinuity [642, 1038]. discontinuous [1161, 1339, 82, 632, 1474, 1824, 218, 1348]. discounted [1835]. discrepancy [760]. Discrete [76, 1131, 610, 1654, 1825, 299, 745, 1476, 152, 1778, 1726, 959, 1378, 1140, 79, 338, 1322, 693, 501, 1118, 74, 966, 543, 1329, 945, 1280, 792]. discrete-time [1476, 1378, 338, 501, 792]. discretely [1622, 1674]. discretely-sampled [1622]. discretization [957]. disease [1570, 1765, 1563, 1368, 1568]. diseases [1566, 1160, 1730]. disk [1243, 65]. disks [483]. Dispersal [1883]. dispersion [75, 307, 1418, 1885, 1251, 833, 1922]. dispersionless [1123]. displacement [328, 363]. dissimilar [684]. dissipation [1035, 595, 1356, 16]. dissipational [1445]. dissipative [564]. distance [1722]. distant [523]. distinguishability [900]. distortion [161]. Distributed [1633, 820, 1407, 645, 1137, 1711]. distribution [1016, 863, 1451, 822, 392, 1387, 669, 1407, 1417, 1219, 844, 1353, 855, 1037, 1005]. distributionally [1851]. distributions [1544, 66, 922, 575, 959]. disturbance [1341]. disturbances [1638, 1098]. divergence [1275]. divergent [18]. diverging [970]. Diversity [1294]. divide [793]. divide-and-conquer [793]. divided [1364]. dividend [1600, 1309]. divisible [319]. division [1845, 1664, 1750, 1853]. DNLS [1123]. Do [1877, 1892]. docking [1851]. doesn’t [1771]. domain [1286, 981, 1731, 1774, 1181]. domains [327, 1386, 1849, 647, 1232]. dominance [477]. dominant [490]. Dominated [514]. dominating [696]. Dominici [1459]. Double [1915, 827, 1104, 1148, 453, 304, 152, 249]. double-diffusive [304]. Doubling [137]. doubly [1510, 1275, 1276, 577]. doubly-connected [577]. Douglas [1614]. drain [1801]. Drainage [1336].

Draping [1839]. drift [999]. drink [1507, 1737]. drive [1877]. driven [1401, 1454, 1794]. driving [858, 744]. drug [1505]. drum [267, 277]. Dual [459, 1500, 1372, 1833, 1143, 1517, 1219, 381, 1578, 1379]. dual-reciprocity [1578]. Duality [413, 658, 523, 1256, 578, 1287, 361, 713, 415, 421, 1264, 1083, 250, 386, 549, 74, 133, 367, 677, 682]. duct [492, 14, 290]. ducts [1785, 637]. due [1801, 978, 1243, 228, 234, 433, 364, 800, 603, 688, 868, 317, 911]. Duffing [458, 202, 226]. Dufort [1850]. duopolists [503]. durability [426]. dye [1876]. dye-sensitized [1876]. Dynamic [1754, 1269, 1226, 394, 1592, 634, 1304, 406]. Dynamical [1570, 613, 1302, 1488, 595, 1734, 1076, 175, 545, 875, 502]. Dynamics [219, 1001, 1724, 1825, 1751, 1350, 1107, 1564, 509, 755, 1567, 1230, 1525, 1679, 1083, 1616, 1844, 1187, 1782, 301, 1863, 1201, 1429, 1267].

E. [1231]. Early [1056, 1497]. eccentric [848, 971, 841]. ecologically [1565]. ecology [1594]. economic [212]. economy [1453]. ecosystem [309, 1360]. edge [393, 467]. Editorial [1485, 1503, 1816, 1890, 1913, 1708, 1653, 1758, 1838, 1563, 1671, 1857]. Effect [1775, 32, 1533, 930, 912, 267, 1686, 888, 1125, 1257, 1516, 1569, 650, 155, 16]. Effective [1443, 1659]. Effects [205, 1155, 1616, 1538, 1235, 1668, 359, 1079, 1214, 689, 1760, 113, 1204, 1844, 1078, 1515, 1480, 1368, 624]. Efficiency [1216, 1457, 1178, 487]. Efficient [1852, 1666, 1351, 859, 1473, 1428]. eigenfunction [1176, 1084]. Eigenfunctions [1506]. eigenproblem [1104]. eigensolutions [596]. Eigenvalue [902, 1004, 1025, 191, 1421, 453, 614, 172, 97, 516, 1493]. Eigenvalues [350, 478, 695, 461, 1034, 537, 136, 128, 1325, 1399, 515]. Einstein [125, 33, 981, 13, 80, 506]. Elastic [1282, 602, 789, 19, 921, 92, 129, 334, 1508, 372, 684, 37, 544, 1154, 1517, 246, 3, 538, 1533, 1649, 20, 276, 1359]. elastic-plastic [1359]. elasticity [1153, 26, 184, 183]. elastodynamic [87]. elastodynamics [1841]. elastomers [1649]. elastoplastic [738]. elastostatics [49]. electricity [1681]. electro [1056]. electro-magnetic [1056]. electromagnetic [1303, 371, 736, 1126]. electron [1121]. Electrophysiology [920]. electrostatic [1593]. element [1026, 1025, 856, 352, 462, 957, 1743, 1731, 1663, 1143, 1181, 705, 716, 694, 1578, 622, 1819, 719, 1690, 1379]. elementary [62, 282]. elements [616, 88, 12, 1088, 655]. elliptic [45, 26, 334, 1162, 165, 360, 859, 242, 1007, 1372, 823, 1152, 857, 1180, 321, 1262, 1819, 1390]. elliptic-type [859]. embedded [674]. Emden [615, 626, 1635]. emergent [1751]. emerging [831]. enclosed [267]. End [359]. endemic [683]. Endex [1073]. endothermic [1842]. endpoint [809]. Energy [57, 296, 924, 1717, 573, 1427, 705, 1137, 1356, 1577, 1831]. energy-preserving [1831]. engine [1408]. engineering [1663]. enhanced [956]. enhancing [941]. enrichment [1844]. enrolment [2]. entering [1658]. entrance [14]. entrant [1731]. entropies [959]. entropy [1059, 1294, 1811]. Entry [243]. envelope [1619]. environment [1191]. environmental [1877]. environments [1648]. enzyme [1288]. Epidemic [1564, 664, 683, 1687, 1187, 1571, 1569, 1328]. epidemics [1298]. epsilon [216]. epsilon-algorithm [216]. equal [1546]. equal-area [1546]. equation [1825, 1017, 1879, 1146, 194, 615, 264, 1579, 203, 339, 318, 362, 382, 429, 442, 768, 988, 818, 491, 633, 1027, 188, 1418, 1733, 191, 165, 184, 360, 1797, 328, 1451, 1702, 724, 149, 1419, 1469, 1182, 1885, 533, 458, 1340, 1150, 741, 954, 799, 1050, 1100, 596, 1902, 534, 1363, 757, 351, 1251, 1283, 1592, 903, 1019, 588, 1743, 1509, 1474, 1239, 353, 1635, 1518, 1088, 1876, 1920, 1772, 703, 1514, 1545, 401, 1846, 705, 41, 1587, 221, 1870, 202, 226, 984, 982, 693, 935, 1231, 321, 176, 183, 1878, 1400, 876, 1130, 1042, 151, 168, 675, 620, 1018, 1127, 622]. equation [1845, 868, 710, 187, 1506, 1493, 1750, 910, 993, 40, 64, 493, 1165, 1831, 1061, 1489, 1379]. Equations [203, 17, 71, 63, 1482, 470, 1139, 969, 272, 192, 185, 1153, 103, 866, 1058, 1087, 1044, 1107, 323, 468, 1300, 889, 231, 31, 1117, 1906, 273, 274, 376, 90, 1189, 402, 563, 1049, 1003, 524, 327, 656, 6, 360, 437, 1291, 1106, 1351, 195, 303, 612, 111, 82, 100, 1918, 611, 1609, 839, 1008, 190, 623, 727, 33, 651, 1220, 1314, 1046, 575, 154, 539, 1752, 280, 761, 1598, 535, 669, 1190, 1039, 765, 1511, 823, 957, 1097, 644, 981, 50, 1641, 1131, 443, 565, 408, 678, 1786, 1898, 255, 695, 857, 759, 1397, 1194, 1456, 1180, 25, 84]. equations [121, 952, 1129, 200, 210, 222, 1510, 1929, 1276, 366, 1811, 105, 953, 1123, 13, 80, 287, 407, 964, 189, 1089, 1454, 373, 667, 852, 886, 627, 849, 882, 1109, 1236, 1259, 968, 1499, 1018, 715, 1411, 98, 186, 1202, 1373, 1124, 245, 593, 779, 1577, 1267, 1167, 1390, 430, 1232, 1221, 784, 1250, 1684, 448]. equi [22]. equi-ultimate [22]. equilibria [174, 212, 213, 252, 1398]. Equilibrium [158, 1055, 1827, 1230, 683]. equity [1908]. equivalence [384, 361, 385, 732]. equivalent [1364]. Erdos [1333]. ergodic [1405]. ergodicity [792]. Error [1459, 1262, 462, 1419, 1469, 172, 1585, 1367, 775, 561, 1711]. errors [1025]. eruptions [1442]. Estimates [1845, 124, 1057, 758, 462, 1609, 154, 365, 1459, 1262, 1259]. estimating [1054]. estimation [67, 1832, 1179, 128]. estimators [112, 193]. Euclidean [1343, 901]. Euler [1352, 1258, 1528, 1355, 1138, 1000, 1268]. European [1813, 1753, 1585, 1647, 1740]. European-style [1753]. evaluating [242]. Evaluation [54, 1023, 558]. evaporation [1408, 913, 687]. events [681]. Evolution [1798, 1373, 889, 524, 1733, 1106, 1702, 724, 171, 703, 494, 545, 674, 1098, 281, 621, 868, 1165]. Evolutionary [1825, 514, 548]. Exact [1496, 652, 1906, 524, 1608, 154, 393, 1687, 646, 202, 647, 615, 1117, 1511, 56, 80, 964, 1127, 910, 282]. exactly [337]. examination [1921]. example [1080]. Examples [1166, 373]. Exchange [252, 167, 513]. excited [805, 1805]. Exclusion [136, 468, 1544]. exercise [1472, 786, 1673, 1497]. Existence [1111, 473, 738, 278, 1227, 402, 1804, 1279, 761, 981, 1180, 3, 105, 1277, 968, 98, 593, 1602, 273, 656, 265, 536, 1150, 1846, 52, 627, 1232, 1280, 1416]. exit [1451]. exothermic [1842]. Expansion [1407, 398, 1176, 978, 789, 1006, 1403]. expansions [1302, 531, 1349, 397, 1084]. expansive [1474]. expectancy [1567]. expected [1498]. experiments [213, 774]. expiry [1472]. Explicit [1300, 1518, 1887, 1325, 1399, 1349, 759, 1901, 1899]. explosion [254, 438]. explosions [645]. Exponential [1137, 764, 735, 1314, 1036, 393, 1326, 1853, 1232]. exponentially [575]. exponentials [1588]. exponents [1802, 1386]. expression [652, 1695]. expressions [1687]. extendable [1810]. Extended [174, 212, 213, 211]. extendible [1647]. Extending [970]. Extension [1363, 1466, 810, 1875, 1233, 193]. extensions [871, 1330, 1647]. exterior [981]. extinction [219, 651, 1225, 567]. Extrapolation [216, 1049, 81, 1029]. extremal [1151, 665, 1307]. Extreme [1140]. extremum [444].

F [1069, 1067]. Faà [1349]. facial [1570]. Facility [1348]. factor [1790, 1635]. factorisation [636]. factorization [450]. factorizations [451, 1404]. factors [1028, 256]. fair [1310]. fall [951]. falling [899]. Families [355]. family [1350, 735, 1233]. Fast [17, 1051, 778, 824, 849, 1372, 1240, 71]. FBM [1254]. FBM-based [1254]. Feasibility [1614]. feature [924]. Feedback [1631, 1141, 1766, 788, 1137, 1001]. feedbacks [696]. Fejér [1022]. fermentation [1616]. ferrofluids [928]. ferromagnetic [1126]. fertilizer [427]. fibre [209, 457]. fictitious [1619]. Field [279, 978, 457, 769, 490, 1447, 33, 535, 201, 472, 1377, 7, 32, 80, 1208, 886, 1840]. fields [1891, 1715, 754, 1410, 1275, 1276, 506, 279, 1056, 1275]. Fifth [1733]. Fifth-order [1733]. filds [139]. film [148]. filter [1159, 1832]. filtering [1476]. filtration [1416]. Final [989, 1687]. finance [1671, 1857]. financial [1680, 1857]. finding [1, 1212, 91]. fingers [1685]. Finite [462, 1731, 1861, 537, 1025, 958, 652, 616, 913, 352, 1797, 347, 724, 1419, 1469, 811, 1850, 628, 665, 1363, 707, 1135, 227, 1849, 511, 957, 37, 49, 1775, 1743, 759, 1663, 463, 466, 1140, 1928, 705, 647, 849, 882, 622, 1819, 719, 1672, 1740, 1690, 448]. finite-difference [882, 1740]. finite-element [957]. finite-section [448]. fires [235]. first [1161, 192, 492, 1451, 1151, 698, 623, 727, 1849, 1630, 869, 765, 1585, 1122, 221, 98, 1506, 1149, 1359, 593, 1690, 1221]. first-kind [623, 727]. first-order [1161, 1151, 1630, 1585, 1506, 1690, 1221]. first-passage [869]. fish [7, 1247]. Fisher [1579, 1117, 1878, 622]. fishery [1209]. fit [967]. five [1245, 487]. five-parameter [1245]. five-spot [487]. Fixed [557, 1141, 501, 1055, 343, 406]. fixed-interval [501]. flame [1852, 741, 572]. flames [504]. flaps [467]. flat [1214, 364, 14, 1458, 1173]. flat-plate [1214]. Flexibacter [199]. flexible [1743, 503, 230, 1606]. Flight [1920]. Flow [173, 597, 689, 754, 1219, 1496, 1527, 641, 1243, 480, 686, 75, 307, 951, 302, 512, 15, 725, 462, 1761, 68, 268, 925, 822, 100, 1623, 1646, 1662, 1728, 1875, 1914, 1136, 1504, 1050, 1785, 300, 848, 971, 600, 1632, 1686, 1775, 253, 404, 970, 1560, 1854, 637, 1398, 688, 1181, 919, 1586, 723, 1640, 32, 148, 1401, 716, 487, 243, 484, 1480, 1526, 1289, 5, 317, 260, 911, 1842, 1149, 653, 1002, 1698]. flows [831, 909, 951, 433, 1638, 1279, 177, 697, 364, 707, 1334, 991, 298, 1143, 1129, 1045, 1868, 1764, 886, 1495, 1173, 282]. Flowshop [1293]. Flowshop/no [1293]. Flowshop/no-idle [1293]. flowtime [1293]. fluctuations [457]. fluid [1496, 1243, 654, 832, 505, 1557, 437, 68, 1556, 328, 363, 854, 1278, 597, 689, 811, 1875, 1894, 1096, 1135, 1422, 1552, 807, 688, 1192, 1764, 723, 1640, 1782, 625, 1619, 1668, 342, 102, 1084, 1759, 484, 1480, 1201, 1803, 1698, 847, 892]. fluid-structural [1557]. fluids [1186, 1793, 1646, 1760, 148, 627]. fluorescence [1648]. Flux [521, 57, 1373, 391]. focus [1654]. foliage [316, 401]. Foliations [976]. follower [548]. food [1754, 1086, 1834]. Force [392]. Forced [988, 582, 583, 584, 673, 613, 293, 458, 147, 1489]. forcing [583, 584, 458]. forearms [1163]. forecasting [1681]. forgetting [1826]. form [1482, 1773, 1486, 490, 1766, 439, 1072, 698, 353, 774, 1622, 1744, 1618, 1309, 1684]. formal [192]. formalism [1374]. Formation [854, 1685, 711, 919, 949]. formed [1359]. forming [1320, 1002]. forms [563, 1351, 655]. formula [18, 1813, 1860, 1128, 1459, 1595, 1355, 1585, 1349, 1462, 1862, 251, 220, 1301, 1622, 1453, 530, 1165, 1618]. formulae [1466, 1693, 1352, 1258, 1138, 398, 1268, 515]. formulas [1032, 1528]. formulation [1269, 490, 375, 1772, 1181, 1452]. forward [794]. four [339, 964, 1399]. Fourier [1677, 299, 785, 635, 844, 1831]. Fourth [889, 1106, 1641, 1804, 724, 351, 1131, 1474, 868]. Fourth-order [889, 1641, 1804, 1131, 1474, 868]. Fowler [615, 626, 1635]. fractal [793, 996, 1193]. fractals [990, 746, 819]. Fractional [934, 133, 1017, 1884, 426, 1418, 1908, 1720, 1859, 1927, 1251, 1595, 1147, 1731, 1876, 1514, 1824, 635, 844, 677, 1499, 1626, 1864, 578, 1781, 955]. fractional-order [1884, 1720, 1499, 1781]. framelets [1713]. framework [1727]. Frankel [1850]. Fredholm [188, 1411]. Free [688, 1782, 1852, 831, 1564, 433, 547, 1603, 597, 300, 364, 600, 1686, 1427, 510, 1487, 1289, 445, 5, 317, 911]. Free-surface [688, 1782, 831, 433, 364, 1686, 317, 911]. freedom [1292]. Freely [1320]. freqencies [155]. frequencies [1214]. frequency [1303, 789]. friction [804]. Fritz [42]. Front [10, 23, 34, 47, 59, 72, 85, 95, 108, 116, 130, 156, 169, 144, 181, 197, 206, 214, 224, 237, 247, 258, 269, 284, 294, 305, 314, 325, 335, 345, 356, 369, 379, 388, 399, 410, 422, 435, 446, 455, 464, 474, 485, 496, 507, 517, 528, 541, 551, 559, 568, 580, 589, 598, 608, 617, 629, 639, 648, 660, 670, 679, 690, 700, 708, 717, 728, 739, 752, 762, 771, 781, 790, 801, 813, 825, 836, 850, 860, 873, 884, 894, 904, 915]. Front [926, 936, 947, 962, 972, 985, 997, 1010, 1030, 1040, 1052, 1065, 1081, 1090, 1101, 1113, 1134, 1144, 1157, 1171, 1183, 1198, 1210, 1223, 1237, 1248, 1260, 1271, 1284, 1295, 1305, 1318, 1331, 1345, 1357, 1370, 1380, 1393, 1449, 1460, 1470, 1483, 1491, 1501, 1512, 1521, 1529, 1535, 1541, 1547, 1554, 1562, 1572, 1583, 1589, 1597, 1604, 1612, 1620, 1627, 1636, 1644, 1651, 1660, 1669, 1682, 1691, 1699, 1706, 1735, 1747, 1756, 1767, 1776, 1783, 1791, 1799, 1806, 1814, 1822, 1828, 1836, 1847, 1855]. front [1866, 1873, 1881, 1888, 1896, 1903, 1911, 1923, 1930]. fronts [1852, 741, 572]. Froude [150, 433, 488, 510]. fuel [1479]. fulid [293]. full [1539]. Fully [951, 1894, 1803, 637]. function [1168, 1389, 1153, 1022, 234, 1500, 1059, 883, 1843, 81, 123, 1459, 606, 1367, 29, 65, 162, 1175, 1364, 1089, 1103, 276, 1588, 1834]. Functional [575, 1146, 1058, 1087, 1227, 1339, 1488, 534, 1410, 1456, 1845, 1506, 1493, 1234, 430]. functional-differential [1058, 1087]. functionally [1508]. functionals [316, 323]. functions [1350, 1879, 1044, 921, 733, 810, 871, 1712, 349, 66, 354, 742, 76, 242, 900, 1142, 425, 1832, 1196, 250, 1147, 397, 1203, 1006, 1812, 244, 377, 932, 891, 1425, 459, 431, 1274, 469, 1089, 1376, 1110, 1156, 180, 1863, 1366, 1348, 797, 1448]. fundamental [1418]. furnace [1667]. Further [777, 1100, 1456]. future [1568]. fuzzy [1715, 756].

G [923, 1141, 1838, 931, 1069]. Gâteaux [1203]. gain [987]. Galerkin [1026, 1879, 327, 1351, 303, 1609, 1824, 622]. Game [1820, 361, 426, 732, 890]. games [1111, 174, 212, 213, 1310, 137]. gamma [871, 1245, 1674, 810]. gamma-Weibull [1245]. gap [898, 1610]. gaps [821, 1787]. Gas [712, 110, 158, 1072, 1871, 1657, 1639, 1208, 513, 153]. gas-lubricated [1208, 153]. gate [831, 1656]. gauge [53]. Gauss [1843, 1367, 1089, 1042]. Gaussian [1330, 1618]. gene [1117]. General [1528, 787, 1385, 1482, 139, 1466, 1352, 1206, 208, 1488, 698, 1132, 375, 280, 1585, 1094, 1225, 398, 1268, 1887, 279, 176, 1878, 151, 168, 527, 43, 1680, 174]. Generalisation [1301, 27]. generalised [1269, 1297, 614, 859, 626, 1150, 1142, 535, 666, 1196, 1147, 516, 393, 835, 549, 417, 482, 500, 1216, 502]. generalization [220, 1020, 1909]. Generalizations [1366, 995]. Generalized [1352, 426, 354, 1409, 61, 1362, 843, 939, 1496, 969, 378, 216, 1579, 1022, 810, 871, 413, 994, 201, 891, 1586, 193, 1716, 964, 1401, 731, 1698]. generate [574]. generated [177]. generating [1389, 891, 431, 1110, 1156]. generation [310, 289]. Generic [414]. genetic [1696]. genetics [1324]. genotypes [567]. geodesic [149, 390]. Geometric [166, 792, 91, 1028, 133, 558]. geometries [1755]. geometry [1079, 793, 596, 1130, 1426]. geophysical [292]. geothermal [1447, 1446]. Gerchberg [785]. gestation [1754]. Gevrey [1097]. GI [553, 1264]. GI/M/1 [1212, 1264]. GI/M/n/n [553]. Ginzburg [768, 1150]. glass [1667]. gliding [199]. Global [838, 1150, 374, 395, 1008, 1086, 683, 1194, 852, 1164, 1416, 1488, 1432, 418, 1307, 664, 965, 1599, 1234]. GMRES [1392]. gold [1655]. governed [765, 98]. grade [1782]. graded [1508]. Gradient [815, 1796, 858, 1576, 765, 1633, 322, 1445]. gradients [1710, 417, 989]. Graeme [1913]. granular [1205]. graph [257, 441, 1723]. graphical [141]. graphs [1132, 1723]. grass [235]. gravitation [266]. gravity [889, 930, 286, 1733, 1106, 1702, 489, 1902, 888, 1478, 1610, 919, 1928, 975]. gravity-capillary [1733, 1106, 1928]. Gray [1068, 1069, 1067]. Greeks [1676]. Green [427, 1153, 87, 65]. grids [347, 1321]. Gronwall [1423]. Gross [1797]. Ground [638]. group [615, 1697, 147, 1413, 574, 1124]. Groups [505, 1320, 88, 211]. growing [575]. Growth [351, 1109, 1259, 429, 1557, 612, 1766, 304, 534, 1763, 1809, 1444, 1437, 427, 1506, 1493, 1664, 1750, 1724, 1853]. guaranteed [1315]. guidance [956]. Guide [1302]. Guinea [1505].

H [530]. Haar [1634]. Hadamard [1352, 780]. Hadley [265]. haemoglobin [513]. half [1032, 87, 1155, 113, 643, 1154, 297, 1533, 687, 448]. half-line [448]. half-plane [113, 687]. half-space [1032, 87, 1155, 643]. half-spaces [1154, 1533]. hall [1480]. Halley [784]. Hamilton [1592, 982, 993]. Hamiltonian [378, 573, 1136, 698, 257, 441, 211, 1926, 1280]. Hamiltonians [655]. Hammerstein [565]. Hamming [1722]. Hankel [969]. Hanoi [843]. hard [896]. hardy [1520]. Harmless [309]. harmonic [722, 56, 147, 1095]. Harvesting [1209, 1247]. Hausdorff [1054]. having [990, 597, 32, 746, 819, 1373]. Hawkes [1795]. health [1765]. Hearing [277, 1523, 577]. Heat [14, 1243, 1408, 1914, 1204, 1559, 1545, 1458, 391, 1440, 493, 719]. heated [1527, 302, 512, 1093, 1266]. heating [1080, 643, 1078, 1075]. Heisenberg [1122]. Hele [328, 363, 1868, 487]. helical [32]. helicopters [1920]. helium [308]. Helmholtz [1646, 1509, 1379]. hemivariational [1277]. hereditary [757]. Hermite [1022, 1365, 1459]. Hermitian [1926]. Hessian [946]. Hessians [1203]. Heston [1813, 1900, 1622, 1674]. heterogeneity [1765, 1844, 1565]. Heterogeneous [1665, 1154, 1434, 1246]. heuristic [441]. hierarchies [1131]. hierarchy [1128]. High [1901, 1690, 1, 268, 458, 1537, 1553, 1581, 1468, 1824, 16, 1577]. high-dimensional [1537, 1553, 1581]. High-order [1690, 1, 1468, 1824]. higher [846, 758, 922, 408, 1311, 1817, 907]. higher-order [408]. Hilary [1273]. Hilbert [1553, 1581, 1159, 1832, 1355, 1537, 1165, 890]. Hill [967]. Historical [1678]. histories [1568]. HIV [1187, 1505]. Hocking [1913]. hold [1771]. Holder [1647, 1810, 995]. holder-extendable [1810]. Holder-extendible [1647]. hole [26]. holes [1386, 1123, 979]. Holling [1884, 1234]. Holling-type [1234]. Homogeneous [58, 121, 200, 627, 151, 168, 1109]. Homotopy [614, 516, 1624]. honour [1913, 1067]. honouring [764]. hopeless [1014]. Hopf [1017, 1884, 450, 632, 311, 1607]. horizon [536, 1634]. horizons [980]. Horizontal [583, 1244, 712]. hot [204]. hotspot [619]. hotspots [650]. Howarth [1518]. Hu [1413]. hull [150]. human [1697, 1834]. humanoid [1083]. hunting [311]. Huxley [1117]. Hwang [1413]. hybrid [1754, 1566, 437, 1488, 558]. hydraulic [951]. hydrodynamics [1907, 530]. hydrogen [1749, 1883]. hydromagnetic [160]. hydrothermal [1442]. Hyper [1420]. Hyper-Wiener [1420]. hyperbolic [1906, 1277, 935, 1690]. Hypercircle [124, 103, 154]. hyperelastic [61]. Hyperelliptic [1297]. hypergeometric [1843, 932, 398, 341, 1330, 1089]. hypersurfaces [1415]. hypothesis [865].

Iang [1423]. ice [1437]. Ideal [1875, 68, 1205]. identical [1245, 1359]. identification [1531, 917]. identities [983, 88, 78, 348, 440, 604]. identity [1466, 12]. idle [1293]. ignition [1436, 262, 621, 1074]. II [385, 1034, 382, 274, 212, 1809, 210, 1200, 819, 1084, 218, 1037]. III [376, 990, 213, 222, 1257]. ill [420, 760, 1833, 1475]. ill-posed [420, 760, 1833, 1475]. Image [1713, 1830]. images [505, 1383, 300, 1322]. imcompressible [37]. IMEX [1908]. IMEX-based [1908]. immersed [1819]. immigration [1354]. immiscible [328, 363]. immobilised [1288]. immune [1567]. impact [1795, 531]. Impaired [480]. impedance [54]. Implementation [1824, 146, 550]. implementations [1215]. implications [1716]. implicit [323, 1831]. imposed [1920]. improbable [681]. Improved [340, 714, 1579, 123, 1479]. improving [1457]. impulse [228, 509]. impulses [1194]. Impulsive [1255, 866, 1058, 1087, 1044, 1107, 612, 1151, 1397, 1290, 1341, 1456, 1558, 1724, 1263, 1267]. incident [638, 585, 1533]. inclined [899]. included [113]. Including [1299, 1505]. incompatibility [1695]. incomplete [810, 871, 896]. incompressible [372, 61, 32, 627]. incompressinle [276]. increasing [1567]. independence [1372]. Independent [1245, 1726]. Index [1880, 1910, 480, 1544, 1205, 1420, 1076]. indexation [331]. indicator [1811]. indirect [1755]. induced [1496, 167, 689, 600, 1686, 510, 484, 1289, 1922]. industrial [1650]. inelastic [1125]. Inequalities [1168, 939, 1466, 1352, 1423, 959, 1152, 1105, 1277, 1195, 1268, 1594, 855, 1037, 1385]. inequality [1148, 1432, 900, 922, 1307, 1355, 1335, 1633, 1424, 1402, 995, 774, 796, 1364, 1366]. inert [1257]. Inertial [620, 628, 113, 434, 310, 296]. inexact [1490, 358, 779]. inextensible [3]. Inf [882]. Inf-sup [882]. infeasible [1539]. infectious [1566, 1160, 1730, 1563, 1568]. infiltration [1578]. Infinite [488, 654, 19, 150, 433, 1286, 1798, 536, 425, 1343, 510, 1213, 688, 1181, 1634, 1338, 721, 896, 69, 911, 993, 943, 1075]. infinite-depth [1798]. infinite-dimensional [943]. infinite-Froude-number [433]. infinite-horizon [1634]. Infinitely [319]. infinitesimal [1638]. infinity [1121, 1051, 632, 778, 824]. infinity-additional [1051]. influence [1567, 1192, 163]. influenza [1494]. information [1215, 1047, 1322]. inheritance [1697]. inhibiting [1257]. inhibition [1586]. inhomogeneous [1798, 280]. Initial [1374, 1359, 1161, 75, 1206, 566, 1595, 1229, 645, 1474, 545, 1533, 163, 1474]. initial-value [1474]. Initiation [1079, 438]. injection [725, 783]. injections [1600]. innovations [1410]. inpainting [1713]. Input [227, 107, 525]. inputs [27]. insect [520, 591]. inserting [1322]. instabilities [1543, 1850]. instability [51, 1775, 1803]. Instantaneous [987, 329, 696]. insulated [832]. Insulating [271]. integer [1788, 297]. integers [1291, 1020]. Integrability [983, 1120, 910]. integrable [1133, 469, 868, 1124]. Integral [1019, 1772, 470, 1466, 192, 1017, 185, 1153, 1879, 194, 15, 1049, 327, 188, 733, 1423, 191, 165, 184, 360, 437, 328, 1451, 195, 303, 179, 190, 623, 727, 1046, 1039, 519, 1349, 397, 25, 84, 601, 1268, 693, 1089, 1231, 321, 183, 675, 968, 1727, 1411, 710, 187, 186, 855, 1037, 1149, 448]. integral/Taylor [437]. integrals [1148, 735, 859, 242, 698, 1023, 604, 1122, 221, 1021, 180, 655]. Integrated [1580, 1351, 554]. integrating [1635]. Integration [901, 223, 1283, 1537, 1553, 1581, 118, 984, 469, 1014]. integrators [1438]. Integro [468, 1027, 1003, 533, 1609, 903, 957, 1872, 1499]. Integro-differential [468, 1027, 533, 1609, 903, 957, 1499]. integro-differential-difference [1003]. integrodifferential [663, 749, 808]. intense [438, 289]. intensity [296]. interacting [1070]. Interaction [490, 291, 1871, 653, 1557, 573, 1125, 290, 1571]. interactions [158, 1765, 9]. Interconnected [830, 767]. interconnections [142]. interdependent [1298]. interface [1646, 392, 1154, 1819]. interfacial [724, 1608, 721]. interference [1754, 520, 591]. interior [1104, 1389, 45, 1500, 91, 1539, 570, 989, 1228]. interior-point [1389, 1500, 1539]. intermediate [1074]. Internal [870, 684, 291]. interpolation [1022, 1365]. Interpolatory [469, 1376]. interpretation [1593, 1448]. interpreting [240]. interrelated [1725]. intersection [1746]. interstellar [1871]. interval [1039, 1635, 501, 69, 675, 1308]. intervals [1575]. intractable [1014]. intriguing [953]. introduction [1551]. Invariance [119, 615]. invariant [615, 134, 1629, 56, 635]. inventory [1725, 1745]. Inverse [1132, 70, 93, 1390, 785, 1019, 773, 162, 301, 780, 955]. Inverses [1362, 1325, 1399]. Inversion [297, 1693, 1462]. Investigating [425]. investigation [150, 1204]. investment [1790, 1524, 1464, 1680]. invex [1147]. invexity [412, 586, 549]. Inviscid [1552, 230]. involving [1607, 604, 1147, 1239, 1517, 1515, 79, 1364, 1089, 1156, 419, 592, 1493, 405]. ion [1441, 1658]. ion-water [1658]. ionisation [1299]. iron [205]. irrational [921]. island [1914, 1770, 1446]. Isolated [1723]. Isospectral [697]. isotropic [19, 1155, 61, 3, 276]. isotropy [574]. Issue [1380, 1381, 1013, 1115, 10, 11, 23, 24, 34, 35, 47, 48, 59, 60, 72, 73, 85, 86, 95, 96, 108, 109, 116, 117, 130, 131, 156, 157, 169, 170, 144, 145, 181, 182, 197, 198, 206, 207, 214, 215, 224, 225, 237, 238, 247, 248, 258, 259, 269, 270, 284, 285, 294, 295, 305, 306, 314, 315, 325, 326, 335, 336, 345, 346, 356, 357, 369, 370, 379, 380, 388, 389, 399, 400, 410, 411, 422, 423, 435, 436, 446, 447, 455, 456]. issue [464, 465, 474, 475, 485, 486, 496, 497, 507, 508, 517, 518, 528, 529, 541, 542, 551, 552, 559, 560, 568, 569, 580, 581, 589, 590, 598, 599, 608, 609, 617, 618, 629, 630, 639, 640, 648, 649, 660, 661, 670, 671, 679, 680, 690, 691, 700, 701, 708, 709, 717, 718, 728, 729, 739, 740, 752, 753, 762, 763, 771, 772, 781, 782, 790, 791, 801, 802, 813, 814, 825, 826, 836, 837, 850, 851, 860, 861, 873, 874, 884, 885, 894, 895]. issue [904, 905, 915, 916, 926, 927, 936, 937, 947, 948, 962, 963, 972, 973, 985, 986, 997, 998, 1010, 1011, 1030, 1031, 1040, 1041, 1052, 1053, 1065, 1066, 1081, 1082, 1090, 1091, 1101, 1102, 1113, 1114, 1134, 1144, 1145, 1157, 1158, 1171, 1172, 1183, 1184, 1198, 1199, 1210, 1211, 1223, 1224, 1237, 1238, 1248, 1249, 1260, 1261, 1271, 1272, 1284, 1285, 1295, 1296, 1305, 1306, 1318, 1319, 1331, 1332, 1345, 1346, 1357, 1358, 1370, 1371, 1393, 1394, 1449, 1450, 1460, 1461, 1470, 1471, 1483]. issue [1484, 1491, 1492, 1501, 1502, 1512, 1513, 1521, 1522, 1529, 1530, 1535, 1536, 1541, 1542, 1547, 1548, 1554, 1555, 1561, 1562, 1572, 1573, 1582, 1583, 1589, 1590, 1596, 1597, 1604, 1605, 1612, 1613, 1620, 1621, 1627, 1628, 1636, 1637, 1644, 1645, 1651, 1652, 1660, 1661, 1669, 1670, 1682, 1683, 1691, 1692, 1699, 1700, 1706, 1707, 1735, 1736, 1747, 1748, 1756, 1757, 1767, 1768, 1776, 1777, 1783, 1784, 1791, 1792, 1799, 1800, 1806, 1807, 1814, 1815, 1822, 1823, 1828, 1829, 1836, 1837, 1847, 1848, 1855]. issue [1856, 1866, 1867, 1873, 1874, 1881, 1882, 1888, 1889, 1896, 1897, 1903, 1904, 1911, 1912, 1923, 1924, 1930, 1931, 1913, 1435, 1838, 1857]. issues [865]. Itô [1727]. Iterated [1475, 188, 756]. iteration [1389, 773, 1412, 1455]. iterations [562, 989]. Iterative [255, 1545, 1350, 299, 1182, 588, 1734, 498, 787, 893, 1391].

J. [471, 530]. Jacobi [1592, 397, 982, 993]. Jaeger [1559]. James [1771]. Jensen [922, 774]. Jessen [1432]. jet [722, 798, 692]. jets [899, 1552, 1002]. job [1718]. John [42, 1759]. join [1584]. Joined [684]. joint [1693, 1269, 460]. journal [1235]. journey [1694, 744]. jump [1476, 1252, 1524, 1631, 1378, 1534, 1464, 1526, 776, 1869, 1680, 432]. jump-diffusion [1869, 1680]. jump-linear [1534]. jumper [624]. jumping [1315]. jumps [1704, 806, 966]. just [1892].

Kármán [1518, 1587]. Kelvin [509, 1646]. Kepler [211]. kernel [1500, 903, 957, 118, 693, 39, 187]. kernels [1032, 1089]. Kerr [982]. kind [1188, 192, 188, 623, 727, 189, 1411]. kinetic [547, 111, 1616, 41]. kinetics [1650, 1602, 323, 396, 1193, 513]. Kirchhoff [1841]. knee [1336]. Knots [975]. known [1389, 1033]. KOBOL [1779]. Kohlrausch [1588]. Kolmogorov [1164]. Korteweg [1641, 1831]. KPP [1878]. Krein [278]. Krylov [773]. Kuramoto [1802]. Kutta [355, 194, 1439, 1899].

Laboratories [532]. lag [264, 132, 419, 322]. lagged [142]. Lagrange [1388, 1517]. Lagrangian [425, 1434, 301]. Lagrangians [125]. lags [44]. Laguerre [1593]. lakes [1794]. laminar [1623, 970, 1842]. Lanczos [773]. Landau [768, 1150]. Lanier [1921]. Laplace [1469, 742, 803, 1419, 698, 922, 761, 1545, 1769]. Laplacian [1340, 1375, 350, 478, 1412, 461, 515, 1455]. Large [933, 1635, 1795, 1389, 566, 1685, 887, 1321, 1843, 142, 1474, 595, 1789, 1463, 1290, 1734, 1534, 1440, 358]. large-amplitude [1685]. large-scale [887, 142, 1789, 1290, 1734, 1440]. large-time [1474]. large-update [1389]. largely [1907]. laser [755]. late [371]. Lattice [1035]. Lavrentiev [1475]. law [702, 27, 70, 93, 1684]. laws [1696, 1546, 1202, 1373]. layer [45, 1556, 1096, 1422, 1883, 807, 919, 1640, 1668, 1098, 1495, 342, 1084, 484, 1266]. layered [906, 958, 1915, 827, 371, 721, 1698, 847, 892]. layers [1093, 1458]. leaching [754]. leading [1743]. leaky [920]. Learning [1710, 1709]. least [126, 146, 114, 1036, 383, 30, 1428, 358]. least-squares [114]. Lebesgue [1022]. Legendre [604, 682, 797]. Leipnik [1353]. Lemaître [1374]. lemmas [990, 746, 819]. length [387, 1443, 931, 1817, 1659]. Lenz [698]. Leontief [406]. level [1694, 854, 1900, 555, 1778, 1745, 1868, 1577]. level-set [1868].

Levenberg [30, 358]. Lévy [1676, 1454]. Lewis [1225]. liability [1827]. Liao [1796]. Liapunov [76]. lid [1401]. lid-driven [1401]. Lie [685, 698, 78, 348, 440, 12, 211, 1744]. Liénard [1482, 1489]. Liénard-type [1489]. Life [1568, 1567]. lift [1920]. lifting [710]. ligand [1265]. light [438]. like [418, 364, 544, 857, 1415, 445, 137, 317]. likelihood [261, 1179]. likely [681]. Limit [159, 402, 1410, 9]. limitations [1920]. limited [1215, 1763, 1809]. limiting [1205, 1076]. limits [744]. line [1801, 433, 689, 418, 600, 1422, 827, 1434, 1377, 716, 1428, 847, 892, 448]. Linear [316, 1242, 1153, 231, 521, 339, 318, 382, 89, 233, 1788, 1797, 477, 1100, 57, 697, 136, 539, 1679, 280, 351, 1252, 227, 1159, 1832, 1307, 1833, 511, 1410, 595, 138, 1384, 121, 366, 1545, 1780, 1534, 459, 570, 571, 1887, 901, 498, 1481, 301, 1256, 1042, 1414, 1575, 151, 168, 1626, 1411, 98, 419, 655, 875, 276, 1329, 1750, 16, 1124, 1266, 776, 1165, 1391, 961, 479, 99]. linearisation [1262]. Linearised [913, 687]. linearity [231]. Linearized [217, 218, 6, 1061]. Linearly [1831]. liner [176]. liners [1916]. lines [390]. link [888]. linkages [1269]. Liouville [1104, 537, 128, 1218]. lipid [1655]. Lipschitz [1058]. liquid [628, 253, 32, 721]. liver [323, 468, 396]. LMI [1468]. load [1463]. loans [1861]. Lobatto [1099]. Local [644, 17, 71, 1879, 924, 1372, 1178, 1898, 1678, 162, 1253]. localisation [1174]. localised [1323]. Localized [1863, 1693]. locally [1643, 1137, 1415]. located [483]. Locating [805]. location [390, 1835, 1348]. locomotion [1760]. logarithmic [735, 995, 693]. logarithmic-kernel [693]. logic [1656]. logistic [203, 1688]. logistics [1721]. lognormal [606, 1681]. Lommel [733]. Long [909, 334, 490, 51, 489, 1900, 308, 545, 432, 624]. long-jumper [624]. long-run [1900]. longitudinal [684]. loop [457]. looped [1005]. Lorentz [1415]. Lorenz [755]. loss [1051, 557, 778, 824, 987]. losses [1347]. Lotka [1048, 252, 374, 395, 1128, 605, 696, 1515]. Love [179, 1231]. Low [789, 302, 1303, 512, 1080, 1654]. low-dimensional [1654]. lower [1665, 1166, 340, 944]. lubricated [1208, 153]. lumped [352, 219]. lung [513]. Lyapunov [866, 1087, 1044, 1802].

M [553, 531, 1212, 1264, 1141]. M/G/ [1141]. Mach [1057]. machine [1718]. machines [1719]. Maclaurin [1355]. macro [1765]. macro- [1765]. macroscale [1751]. magic [1892]. magnet [1840]. magnetic [533, 32, 1056, 1840]. magnetization [1486]. magnetized [928]. Magnetohydrodynamic [1480, 722, 1143, 852]. magnetohydrodynamics [1097]. Mahony [953]. maker [912, 628]. Maki [1826]. making [1705]. Malliavin [1676]. managed [1745]. management [1827, 554]. manifold [494, 674, 620]. manifolds [613, 833]. manure [427]. map [1601, 1427, 923]. MAP/G/1 [923]. mappings [1433, 574]. maps [1337, 1549]. Marangoni [1096]. marine [1844, 1648]. market [1217, 1681, 1680]. Markov [319, 999, 1315, 1108, 1212, 1264, 869, 1405, 1631, 1744, 1177]. Markovian [1886, 1476, 1252, 1387, 1417, 1378, 1821, 1869]. Marquardt [358]. masks [1915]. mass [352, 954, 13]. matched [789, 833]. matching [955]. material [845, 652, 544, 276]. materials [958, 1508, 1205, 61, 37, 49, 1517, 276]. maternity [234]. Mathematical [1447, 532, 1780, 1708, 1523, 1643, 349, 1765, 1770, 554, 171, 666, 1905, 481, 1877, 99, 513, 1368, 1440, 624]. mathematics [1549, 1857]. Matrices [1362, 1026, 451, 12, 556, 946, 636, 780, 1120, 1325, 1399]. Matrix [1796, 1614, 806, 451, 429, 361, 1182, 1488, 88, 1132, 1212, 1264, 1307, 669, 1190, 353, 1462, 1818, 732, 1391, 1326, 558]. matrix-geometric [558]. matrix-Riccati [669]. matter [10, 11, 23, 24, 34, 35, 47, 48, 59, 60, 72, 73, 85, 86, 95, 96, 108, 109, 116, 117, 130, 131, 156, 157, 169, 170, 144, 145, 181, 182, 197, 198, 206, 207, 214, 215, 224, 225, 237, 238, 247, 248, 258, 259, 269, 270, 284, 285, 294, 295, 305, 306, 314, 315, 325, 326, 335, 336, 345, 346, 356, 357, 369, 370, 379, 380, 388, 389, 399, 400, 410, 411, 422, 423, 435, 436, 446, 447, 455, 456, 464, 465, 474, 475]. matter [485, 486, 496, 497, 507, 508, 517, 518, 528, 529, 541, 542, 551, 552, 559, 560, 568, 569, 580, 581, 589, 590, 598, 599, 608, 609, 617, 618, 629, 630, 639, 640, 648, 649, 660, 661, 670, 671, 679, 680, 690, 691, 700, 701, 708, 709, 717, 718, 728, 729, 739, 740, 752, 753, 762, 763, 771, 772, 781, 782, 790, 791, 801, 802, 813, 814, 825, 826, 836, 837, 850, 851, 860, 861, 873, 874, 884, 885, 894, 895, 904, 905, 915, 916]. matter [926, 927, 936, 937, 947, 948, 962, 963, 972, 973, 985, 986, 997, 998, 1010, 1011, 1030, 1031, 1040, 1041, 1052, 1053, 1065, 1066, 1081, 1082, 1090, 1091, 1101, 1102, 1113, 1114, 1134, 1144, 1145, 1157, 1158, 1171, 1172, 1183, 1184, 1198, 1199, 1210, 1211, 1223, 1224, 1237, 1238, 1248, 1249, 1260, 1261, 1271, 1272, 1284, 1285, 1295, 1296, 1305, 1306, 1318, 1319, 1331, 1332, 1345, 1346, 1357, 1358, 1370, 1371, 1380, 1381, 1393, 1394, 1449, 1450, 1460, 1461, 1470, 1471, 1483, 1484, 1491]. matter [1492, 1501, 1502, 1512, 1513, 1521, 1522, 1529, 1530, 1535, 1536, 1541, 1542, 1547, 1548, 1554, 1555, 1561, 1562, 1572, 1573, 1582, 1583, 1589, 1590, 1596, 1597, 1604, 1605, 1612, 1613, 1620, 1621, 1627, 1628, 1636, 1637, 1644, 1645, 1651, 1652, 1660, 1661, 1669, 1670, 1682, 1683, 1691, 1692, 1699, 1700, 1706, 1707, 1735, 1736, 1747, 1748, 1756, 1757, 1767, 1768, 1776, 1777, 1783, 1784, 1791, 1792, 1799, 1800, 1806, 1807, 1814, 1815, 1822, 1823, 1828, 1829, 1836, 1837, 1847, 1848, 1855, 1856, 1866]. matter [1867, 1873, 1874, 1881, 1882, 1888, 1889, 1896, 1897, 1903, 1904, 1911, 1912, 1923, 1924, 1930, 1931, 75, 1654]. maturation [1094]. maturity [1861]. maxima [367]. maximal [1190, 127, 1672]. Maximising [1207]. maximizing [1498, 445]. Maximum [77, 62, 1419, 1469, 1136, 757, 275, 1239, 1179, 794, 840]. Maxwell [1496, 402, 563, 506]. Mayer [378]. Mean [1827, 177, 484, 1810, 1849, 1625, 995, 1703, 1425, 1293, 1862, 1618]. mean-reverting [1810, 1862, 1618]. Mean-variance [1827, 1625, 1703]. measurement [4]. measurements [956]. Mechanical [1766]. mechanics [1523, 437, 1758, 522, 211, 1759]. mechanism [1826, 1288]. media [152, 375, 253, 519, 1922]. median [523]. medium [1801, 654, 832, 328, 363, 1422, 1615, 371, 1356, 621, 867, 1126, 847, 892]. meeting [1067, 1002]. Mellin [1377, 932, 1714]. Membrane [1610, 743, 1478]. Membrane-coupled [1610]. membranes [521, 684]. Memoriam [1273]. Memory [1217, 703]. meningococcal [1569]. Menten [996, 1288]. Mesh [1372, 299]. meshless [1879]. meson [13, 279]. metallofullerene [1656]. method [1482, 1026, 1139, 1242, 1796, 845, 1287, 1395, 103, 1879, 866, 1087, 1107, 378, 789, 1579, 1457, 1027, 1049, 327, 352, 1643, 923, 165, 184, 360, 785, 1197, 614, 1182, 303, 1546, 100, 1609, 1540, 1046, 1785, 1372, 1730, 1833, 1039, 123, 441, 1753, 1592, 823, 957, 201, 1241, 565, 519, 1509, 56, 773, 857, 759, 1213, 1633, 1663, 1774, 516, 1635, 965, 1143, 952, 601, 1755, 637, 1545, 1734, 1868, 1634, 570, 705, 984, 1742, 693, 1401, 321, 183, 1599, 498, 694, 1400, 1414, 1095, 1578, 804, 893, 1499, 622, 1411, 445, 527, 1709, 1819, 322, 776, 358, 719]. method [1391, 1167, 960, 1690, 784, 1379]. Methods [1614, 17, 71, 63, 1677, 1389, 192, 194, 299, 1, 1708, 616, 1439, 856, 355, 242, 190, 418, 425, 81, 903, 1641, 1019, 1229, 1537, 1553, 1581, 1088, 1901, 1887, 498, 1428, 787, 39, 1899, 1452, 989, 1606, 1721, 1671]. metric [978, 982, 1926]. metrics [506]. MHD [1527, 1232]. Michaelis [996, 1288]. Michell [530]. micro [1765]. micro-vessels [1765]. microbial [1602, 1616]. microorganisms [1761]. micropolar [26, 1192, 1619, 1356, 1480]. micropollutants [1780]. micropore [1363]. microscale [1751]. microscope [1759]. microwave [619, 643, 1078, 1075]. microwaves [1919]. midpoint [1258]. milk [1080]. Miller [471]. mine [682]. mineral [754]. mines [1916]. Minimal [1190, 1203, 127]. Minimax [1162, 1147, 893, 578]. Minimisation [323, 1307]. minimise [1293]. Minimization [834, 1383, 745, 1743, 367]. minimizing [1498]. minimum [77, 587, 161, 104, 1664, 1722]. mining [1729]. Minkowski [101, 574]. minmax [1473]. misfit [19]. Missile [1414, 1479]. mix [1452]. Mixed [1468, 1788, 179, 439, 1409, 1372, 1443]. mixed-slip [1443]. mixing [1016, 1098]. Mnemonic [1362]. mode [1705, 1479, 1820, 610, 1414]. model [1754, 1739, 1704, 1297, 1486, 1570, 1085, 1566, 1441, 2, 1884, 1382, 1779, 1765, 1762, 1197, 554, 711, 887, 1858, 1525, 1859, 1927, 1813, 1900, 152, 1763, 1809, 291, 869, 1036, 1209, 1616, 1676, 1063, 1479, 1826, 1851, 1396, 664, 1324, 1524, 1678, 1518, 249, 1071, 1681, 1883, 1854, 714, 1446, 1687, 1187, 1734, 1299, 427, 1076, 1780, 1185, 1820, 591, 920, 1738, 1247, 1571, 1622, 1674, 816, 1909, 513, 424, 567, 620, 798, 1246, 1845, 1368, 1864, 1506, 1493, 1664, 911, 1440, 1001, 783, 1648, 343, 406, 1724, 1328, 1698, 747]. Modelling [235, 1701, 1749, 1444, 1437, 1551, 1919, 1401, 1569, 1658, 1505, 1606, 1834, 1186, 1298, 1523, 1788, 1080, 1653, 1623, 1728, 1447, 534, 1876, 1929, 1442, 1181, 1877, 1780, 1028, 1563, 875, 1574]. Models [961, 1852, 924, 769, 1908, 520, 612, 536, 171, 1477, 212, 252, 557, 1360, 1552, 683, 1094, 1254, 1140, 1219, 1313, 625, 1304, 1744, 1312, 1565, 833, 1654, 1033]. modes [573, 1096, 1668]. modification [1902]. modified [1153, 1544, 1702, 1900, 1576, 123, 1635, 989, 1819, 1391, 1831, 1360]. modular [472]. modulation [489, 1928, 1431]. modulational [1775]. moisture [1080]. molecular [1016]. moment [869, 256, 245]. momenta [655, 229]. moments [922, 855, 1037]. momentum [66, 1797]. monoenergetic [64]. monoenergetic-source [64]. monopolies [672, 828]. Monotone [562, 1812, 1274, 548]. monotonic [1364, 1366]. Monotonicity [244, 1367, 1059]. Monte [1537, 1553, 1581, 1740]. Montgomery [1466]. Morrison [1462]. mortgages [331]. mosquitoes [1551]. most [698]. Motion [1619, 1017, 1908, 638, 1451, 149, 265, 1204, 340, 298, 1192, 199, 721, 1084]. motions [925, 1220]. moulding [783]. mounted [391]. movable [1116]. movement [841]. Moving [253, 113, 394, 340, 888]. MPOD [1360]. multi [906, 480, 713, 1307, 1229, 1676, 888, 843, 835, 193, 672, 1453, 1727]. multi-asset [1676, 1453, 1727]. multi-extremal [1307]. multi-index [480]. multi-layered [906]. multi-link [888]. multi-national [672]. multi-objective [713, 835]. multi-parameter [1229, 193]. multi-peg [843]. multi-period [1453]. multiconnection [556]. multifractionality [1891]. multifunctions [940]. Multigrid [1088]. multilayer [1532]. multilevel [1818]. multimodal [1294]. multinational [828]. multiobjective [1287, 1395, 658]. multiphase [1763, 1809]. Multiple [883, 1014, 859, 1540, 1726, 806, 27, 1254, 1414, 804, 1018, 1626, 1574, 961, 1430]. multiple-order [806]. Multiple-term [883, 859]. multiple-time [1540]. multiplicative [1630, 1256]. Multiplicatively [1744]. multiplicity [376, 1075]. multipliers [950, 1774, 1517]. multiply [1892]. multipoint [1412]. multiscale [1763, 1809, 1865]. multiscaling [1376]. Multiserver [1347]. multistep [17, 71]. Multivariable [112, 1330]. multivariate [344, 1349, 636]. multiwavelets [1240, 1311, 1376]. Murray [1701]. muscle [1175]. must [1744]. mutual [1754]. mutualism [1063].

n [553, 553]. nanoparticle [1655]. nanoscale [1653, 1443]. nanostructures [1584, 1657, 1444]. nanotori [1656]. nanotori-metallofullerene [1656]. nanotubes [1420, 1658]. Nash [1892]. national [1681, 672]. Natural [1181, 1507, 155]. Navier [491, 1785, 1752, 849, 1250]. navigation [956]. Near [978, 228, 329, 1472, 453, 1377, 1073]. near-adiabatic [1073]. Near-field [978]. nearly [726, 812, 897, 770, 1327, 896]. Necessary [1007, 820, 330, 443, 481, 945]. need [331]. negative [1383, 573, 696]. nematic [1649]. neoclassical [1724]. neodymium [1840]. nested [367]. net [234, 555]. nets [555]. Network [1722, 554, 523, 1854, 1219, 961]. networks [1564, 110, 863, 607, 751, 702, 1712, 1717, 822, 1720, 556, 557, 1468, 1687, 561, 1005]. Neumann [350, 1043, 1201]. neural [1712, 1468]. neuronal [1140]. neurons [920]. neutral [663, 749, 808, 656, 767, 839, 1343, 605, 443, 408, 678, 593, 430]. neutral-delay [605]. neutral-type [839]. neutralized [1900]. Neutrally [925]. neutron [596, 877]. Newton [1139, 1337, 1046, 1742, 1481, 779, 917, 1416]. Newtonian [1243, 1401, 58]. Nicholson [1909]. nine [730, 567]. nitrogenous [427]. NLS [1123]. NMR [1095]. no [1863]. no-arbitrage [1863]. no-idle [1293]. nodal [390]. Nodera [1392]. nodes [1022]. noise [1630]. noisy [1713]. Non [209, 1243, 1225, 180, 276, 231, 339, 318, 382, 668, 1595, 1239, 1245, 138, 7, 627, 875, 16, 817, 1416]. Non-adiabatic [1225]. non-autonomous [668]. Non-classical [180]. non-constant [817]. Non-destructive [209]. non-differentiable [1595]. non-homogeneous [627]. non-identical [1245]. Non-linear [276, 231, 339, 318, 382, 138, 875, 16]. non-Newton [1416]. Non-Newtonian [1243]. non-uniform [7]. non-well-posed [1239]. nonautonomous [1189, 1062]. nonblocking [556]. noncoercive [1152]. noncompetitive [1586]. nonconcentric [1619]. nonconforming [1088]. nonconvex [809]. nondestructive [1841]. nondifferentiable [42, 313, 415]. nonequivariant [1601]. nonexistence [1166]. nonhomogeneous [1155, 1631]. nonidentical [1710]. Nonlinear [831, 566, 1121, 361, 573, 1230, 30, 1928, 1098, 424, 1577, 17, 71, 63, 939, 1139, 106, 272, 457, 103, 124, 743, 641, 473, 749, 808, 951, 21, 889, 652, 685, 362, 442, 1227, 524, 668, 734, 1382, 767, 437, 1291, 1106, 1197, 1182, 1217, 839, 651, 416, 154, 1387, 1128, 946, 765, 1511, 644, 942, 562, 1316, 1474, 1490, 1417, 1475, 585, 386, 381, 1185, 705, 879, 221, 287, 646, 987, 732, 964, 160, 312, 694, 928, 647, 876, 69, 787, 120, 1262, 1259, 1218, 677, 1018, 868, 1385, 910, 1266, 358, 593, 779, 1390, 1061, 1292]. nonlinear [1228]. Nonlinear-cobweb [1230]. Nonlinearities [830, 702, 632]. Nonlinearly [592, 657, 1517]. nonlocal [1465, 1416]. nonmonotone [779]. nonmonotonic [960]. Nonnegative [1246, 1397]. nonorthogonal [1752]. nonoscillatory [593]. Nonresonance [1151]. nonrigid [1611]. nonselective [1247]. nonsmooth [1395, 1609, 835, 1531, 1742, 766, 1167]. nonstandard [1868]. Nonsymmetric [1764, 1631]. nontensile [1907]. nonuniform [347, 1559]. Nordsieck [1887]. norm [29, 1392]. Normal [544, 655, 1359]. norms [1383]. North [747, 1446]. nose [949]. note [63, 1704, 26, 15, 228, 329, 360, 415, 1638, 786, 114, 100, 811, 1591, 1760, 1375, 1354, 586, 1752, 600, 1135, 1422, 1615, 1632, 1233, 510, 49, 1316, 1413, 249, 1746, 434, 127, 1338, 1487, 1870, 341, 1495, 816, 1489, 784]. Notes [471, 343]. notion [1178]. Novel [1909, 1730, 1674, 719]. nozzle [704]. NT [1539]. NT-step [1539]. null [1343]. number [1153, 1057, 302, 512, 150, 433, 268, 1715, 848, 488, 510, 1723, 1075, 16]. numbers [1225]. Numerical [1527, 223, 706, 1557, 363, 1283, 1179, 1663, 1436, 1876, 1125, 703, 1811, 984, 867, 622, 715, 1103, 1907, 1922, 1741, 192, 185, 862, 1579, 1418, 165, 614, 954, 971, 311, 50, 1641, 565, 1497, 1755, 1560, 1545, 1624, 919, 1487, 487, 1400, 256, 736, 676, 804, 1177, 1659, 747]. numerically [1054]. nutrient [1763, 1809]. nutrient-limited [1809]. Nyström [355].

oar [1457]. objective [713, 1788, 1895, 835, 29, 1626, 1103, 1348]. Oblique [1642, 372]. Observability [956, 493]. Observability-enhanced [956]. Observables [472]. observation [333]. Observer [1169]. Observer-based [1169]. obstacle [1875, 1504, 562, 1019]. obstacles [951]. obstruction [1640]. occurring [735, 741]. ODEs [1887]. off [737]. offer [1568]. offline [1705, 1820]. Ogden [276]. oil [601, 1207]. oilspill [1085]. Oldroyd [1698]. Oldroyd-B [1698]. One [711, 886, 1842, 1242, 111, 100, 1858, 1100, 596, 83, 1209, 1868, 513, 153, 1002, 1445]. One-dimensional [711, 1842, 111, 100, 1858, 1100, 83, 153, 1445]. One-field [886]. one-phase [1868]. one-predator [1209]. one-speed [596]. one-step [513]. online [1705, 1725, 1820]. only [956]. onset [1096]. onto [585]. open [1886, 1826, 693, 342, 155, 653, 1002]. open-channel [653, 1002]. openings [239]. operation [450, 1200]. operations [104]. operator [449, 1132, 333, 405, 784]. operators [414, 521, 78, 136, 730, 1152, 251, 1274, 875, 1926, 1174]. optical [209, 457]. optical-fibre [209]. optics [457]. optima [174, 212, 213]. Optimal [110, 28, 208, 704, 1472, 1705, 1286, 396, 1136, 1488, 1609, 1540, 1476, 858, 841, 1790, 1498, 1625, 1524, 1703, 538, 1060, 744, 1835, 878, 1558, 907, 1869, 1680, 1494, 1694, 473, 62, 526, 554, 681, 587, 536, 1858, 1007, 1673, 576, 1159, 1832, 869, 1600, 765, 1592, 888, 1239, 1473, 1479, 1745, 1180, 1624, 1689, 79, 427, 105, 1634, 338, 52, 879, 1587, 1277, 1464, 935, 676, 98, 132, 419, 445, 527, 592, 657, 941, 1103, 809, 908, 322, 1721, 794, 840, 945, 1690]. optimal-growth [427]. Optimality [677, 349, 42, 1307, 946, 330, 549, 99, 945]. Optimisation [420, 1340, 1178, 1216, 525, 1005, 1228]. optimization [1892, 421, 416, 1576, 750, 1539, 250, 1633, 1789, 1851, 965, 1830, 1531, 221, 1742, 1599, 828, 1481, 276, 1348, 766, 960]. optimum [161, 261]. Option [1779, 1859, 1673, 1676, 1585, 1862, 1647, 1727, 1864]. options [1677, 1472, 786, 1813, 1810, 1753, 1497, 1678, 1772, 1863, 1865, 1740]. orbits [853, 805]. order [1161, 1350, 1297, 1300, 889, 414, 846, 1579, 1, 1215, 1227, 758, 1884, 1666, 355, 349, 1733, 165, 360, 344, 1291, 1804, 1106, 724, 1351, 1419, 1469, 1720, 1151, 351, 946, 806, 1630, 765, 823, 957, 1641, 1131, 408, 1474, 1585, 595, 1397, 1468, 1125, 297, 1824, 1734, 1901, 321, 1042, 513, 1109, 1236, 1259, 1499, 98, 868, 1202, 1506, 1865, 593, 1781, 1690, 1221]. order-bounded [414]. orders [1795, 730, 1311]. Ordinary [407, 1365, 1050, 730, 200, 210, 222]. organic [1780]. organisation [468]. organization [1617]. oriented [602]. origins [1446]. Ornstein [1732]. orthochronous [574]. Orthogonal [1360, 1448, 1, 88, 1593, 1240]. orthonormalisation [479]. orthotropic [684]. Oscillation [1226, 678, 1267, 1221, 351, 430, 1061]. Oscillations [611, 830, 408, 566, 443, 603, 1639, 816, 1805]. oscillator [594, 795, 853, 56, 147, 211]. oscillators [164, 734]. Oscillatory [839, 805, 148, 1292]. Ostrowski [1424, 1402]. ot [74]. other [179]. Ou-Iang [1423]. out-of-plane [1760]. outflow [1543, 1894]. outflows [1608]. output [107, 1207]. overdetermined [1406]. overstable [312]. oxidation [1200, 1257]. oxide [1444]. oxygen [1076, 513].

packet [1702, 177]. Painlevé [1118, 1130, 1127, 1120]. pair [602, 1798]. pairs [1858]. palindromic [1421]. pantograph [1846]. paper [471]. Papua [1505]. para [164]. para-Bose [164]. parabolic [633, 352, 612, 1609, 799, 255, 1239, 105, 1060, 879, 98, 1202, 917]. parade [1770]. paradigm [918]. paradox [702, 1310, 1844]. Parallel [1718, 307, 1719, 637]. Parallelisation [1321]. parallelization [1462]. parameter [1104, 1297, 633, 820, 1229, 1245, 720, 112, 1862, 193, 621, 527]. parameterization [1540, 941]. parameters [1139, 362, 382, 1315, 1476, 1843, 1252, 1755, 241, 403, 438, 546]. parametric [1302, 1036, 908]. parametrically [805]. parametrically-excited [805]. parametrisation [1297]. parametrization [1624]. parasites [520, 591]. Pareto [1648]. Parisian [1675]. Parrondo [1310]. Parseval [397]. Part [1004, 1034, 582, 583, 584, 273, 274, 376, 1809, 200, 1275, 1276, 99, 217, 218]. Partial [695, 969, 272, 633, 1189, 1598, 765, 903, 957, 1478, 366, 105, 407, 964, 935, 1713, 98, 1202, 1124, 1061]. partially [120, 222]. particle [279, 1907]. particles [1080, 1606]. particles-modelling [1080]. partly [431]. passage [1849, 869]. past [1504, 1334, 1398, 1526, 260, 1574]. patch [1191]. patches [1751]. path [1136, 1689, 543]. path/trajectory [543]. pathlines [329]. pathogen [1187]. paths [536, 275]. Patient [1368]. Patient-dependent [1368]. pattern [1697, 711, 487]. Pauli [522]. PDE [1802, 1094]. PDEs [1042]. Pearce [1353]. peg [843]. penalty [425]. pendulum [582, 583, 584, 673, 805]. Penetration [1655, 1075]. penguin [1770]. pension [1625]. peptide [1658]. Performance [1696, 1215, 473, 1717, 555, 558]. perfusion [1765]. period [339, 318, 1453]. Periodic [458, 1191, 1292, 1298, 1751, 428, 537, 909, 203, 656, 309, 374, 395, 1849, 97, 1063, 1456, 1275, 1276, 1329, 1164, 1724, 1489, 1280]. periodically [582, 583, 584, 673]. periodicity [1140]. peristaltic [1192]. permanence [1033]. permanent [1072]. permeable [848]. Persistence [567, 1234, 1086, 683]. pertinent [665]. Perturbation [1421, 458, 1404, 304, 1229, 695, 970, 69]. Perturbations [734, 773, 1909, 58]. perturbed [1825, 1048, 1329, 1177, 1399]. Petri [555]. Petrov [303, 622]. pH [1175]. Phase [1214, 924, 769, 1868, 886, 487, 132, 1149]. Phase-retardation [1214]. phases [1442]. phenomena [1186, 1200, 1257]. phenomenon [1844]. Phillip [1770]. Photofocusing [1761]. pick [1820]. pick-up-in-store [1820]. picking [1666]. PID [1580]. Piecewise [623, 727, 234, 52, 592, 187]. Piecewise-constant [623]. Pinning [1720]. pipe [1728, 970, 243, 1480]. pipeline [110]. Pitaevskii [1797]. planar [1269, 1875, 888, 298, 736]. Plane [1275, 1276, 638, 268, 1662, 1760, 113, 1050, 1220, 596, 57, 800, 37, 49, 3, 266, 716, 687]. plane-symmetric [266]. Planform [674]. planing [150, 217, 218]. planing-surface [217, 218]. planner [543]. planning [2, 1689]. plaque [1557]. plasmas [1000]. plastic [641, 1282, 1359]. plasticity [1445]. plate [726, 19, 1214, 1088, 603, 1398, 1587, 1231, 1458]. plates [359, 332, 246, 1093]. plug [20]. plumes [1793, 1552, 653]. Poincaré [574]. Point [1269, 1466, 1389, 1287, 1527, 1352, 1500, 597, 811, 1543, 1528, 387, 1365, 1135, 1632, 1686, 91, 557, 1539, 523, 250, 1397, 1490, 383, 1138, 570, 1268, 987, 1289, 1333, 989, 1228]. Point-joint [1269]. points [706, 416, 1103]. Pointwise [272, 758, 1833]. Poiseuille [1496, 925]. Poisoning [167]. Poisoning-induced [167]. Poisson [607, 751, 1743, 1000, 1648]. poker [1771]. Pol [862]. pole [1104]. poles [806]. policy [1141, 1600, 1464, 672]. pollution [1885]. poloidal [535, 1275, 1276]. poloidal-toroidal [1275, 1276]. Pólya [1544]. polygonal [327, 70, 93]. polyhedra [901]. polyhedral [91]. polyhex [1420]. polymers [1441]. Polynomial [440, 758, 1302, 565, 187]. polynomials [1188, 1, 1388, 1351, 1595, 604, 1407, 891, 1593, 1020, 341, 901, 675, 1448]. polytropic [1416]. pond [687]. Pontrayagin [62]. Pontryagin [1239]. poor [1877]. Popoviciu [1099]. population [1107, 429, 1324, 1094, 1185, 1001, 1267]. populations [1382, 575, 1826]. Porous [621, 173, 1801, 686, 812, 912, 897, 1762, 328, 363, 1422, 1615, 253, 519, 970, 881, 1787, 1526, 867, 847, 892, 1922]. portfolio [1704, 1625, 907, 1869]. posed [420, 760, 877, 1833, 1239, 1475]. position [66]. Positive [1218, 1825, 999, 573, 38, 240, 1182, 761, 1456, 1043, 703, 1545, 1412]. positive-definite [1545]. positive-type [703]. possessing [698, 922, 373]. Potential [483, 359, 179, 1181, 260]. potentials [1603, 251, 70, 93]. Potts [1297, 1486]. powder [1080]. power [702, 70, 93]. power-law [702]. powered [1717]. Practical [1107, 1439]. Prandtl [1027, 1334]. pre [1533]. pre-stressed [1533]. precision [1215]. preconditioned [1509, 989]. Preconditioning [823]. Predator [1738, 838, 1688, 1884, 1209, 1607, 1515, 1538, 567, 1234, 1191]. predator-prey [1884]. predict [1765]. prediction [636]. predictive [1851, 1631]. predictors [1788]. Preece [1330]. Preface [1435, 1013, 938, 974, 1012, 1115]. preforms [209]. prescribed [77]. presence [889, 930, 897, 1106, 724, 854, 628, 1166, 364, 275, 279, 721, 881]. presentation [1196]. preserving [758, 1831]. pressure [329, 1281, 684, 4, 1208]. prevention [1525]. prey [838, 1688, 1884, 1209, 1607, 1515, 1738, 567, 1234, 1191]. Price [1795, 1704, 1472, 786, 1725]. prices [1677]. Pricing [1810, 1865, 1740, 1111, 1779, 1908, 1859, 1927, 1813, 1860, 1753, 1862, 1622, 1674, 1727, 1863, 1864, 1618]. primal [1500, 1372, 1833, 1219]. primal-dual [1500, 1372]. principal [815]. principle [866, 62, 757, 1743, 1239, 1290, 794, 840]. principles [77, 106, 615, 612, 760, 275, 635, 682]. prisms [70, 93]. probabilistic [1212, 1264, 166]. probabilities [319, 1464]. probability [735, 900, 959, 1175]. problem [1176, 428, 272, 1016, 1303, 743, 480, 737, 45, 359, 141, 726, 738, 685, 768, 704, 1038, 352, 1162, 191, 126, 146, 115, 1508, 619, 785, 313, 1291, 1804, 547, 587, 1375, 628, 113, 900, 877, 665, 539, 488, 1778, 750, 1895, 806, 869, 1036, 511, 765, 330, 1592, 562, 1335, 1509, 1731, 211, 1474, 1397, 1479, 516, 843, 444, 1043, 29, 631, 481, 459, 1487, 476, 301, 69, 1262, 20, 798, 548, 893, 1412, 1333, 1127, 163, 204, 241, 403, 546, 867, 419, 1103, 1201, 1493, 263, 333, 794, 731, 461, 955, 1680]. Problems [1614, 1161, 1186, 77, 106, 537, 1395, 124, 1886, 1879, 641, 1406, 323, 846, 1032, 67, 1227, 856, 880, 912, 89, 1666, 1421, 184, 420, 614, 179, 822, 554, 439, 1409, 1151, 1340, 1488, 760, 1365, 304, 1725, 1372, 1679, 1576, 1307, 1833, 946, 123, 1595, 1229, 172, 253, 1786, 1473, 1490, 1745, 1789, 383, 1475, 210, 970, 1517, 1824, 1624, 1181, 549, 246, 3, 1689, 79, 381, 538, 1634, 338, 1818, 128, 879, 1277, 1274, 321, 183, 716, 694, 1216, 647, 787, 1578, 676, 804, 1218, 1626, 132, 445, 527, 592, 657, 941, 1769, 809, 1452, 1819]. problems [1266, 405, 493, 322, 902, 1721, 578, 766, 945, 1390, 960, 1690, 479]. procedure [366, 1462]. procedures [383]. process [738, 1544, 1732, 1616, 1586, 1780, 1464, 1454, 513, 783, 1869]. processes [1795, 1886, 933, 1354, 1108, 1726, 1575, 636, 792]. processing [1667, 161, 261]. processors [27]. product [932, 469, 780]. production [1811, 1794, 460]. products [1723]. Professor [1913, 1838, 1069, 1067]. profile [800]. profiles [1057, 643]. program [120]. programme [21]. programming [1287, 1395, 2, 361, 426, 713, 1788, 349, 42, 313, 477, 658, 697, 1833, 91, 166, 666, 946, 869, 942, 1147, 1213, 1745, 386, 835, 29, 481, 459, 570, 571, 732, 133, 677, 1626, 578, 943]. programs [1643, 425, 99, 1256, 961]. progressing [177]. progressive [1442]. projected [1576]. projectile [445]. projection [188, 1241, 1786, 1898, 1253, 1385, 1714]. projective [1133]. prolate [173]. Proof [1486]. Propagation [846, 140, 572, 1126, 1117, 921, 308, 36]. Proper [1360, 1178]. Properties [463, 36, 1697, 1059, 119, 756, 351, 827, 244, 1901, 1192, 140, 1428, 186, 1323]. property [1118]. proportional [956, 1498, 1464, 1680]. proposed [192]. providing [604]. Province [1505]. proximal [1633]. proximal-gradient [1633]. proximity [1389, 1539]. pseudo [641, 386, 610, 1499, 1926, 1724]. pseudo-convexity [386]. pseudo-Hermitian [1926]. pseudo-plastic [641]. pseudo-sliding [610]. pseudo-spectral [1499]. pseudomonotone [1433]. pseudoplastic [1186]. pseudospectral [1831]. psi [1364]. pulmonary [1765]. Pulsatile [725]. pulsation [1805]. pulse [1606]. pump [1408]. punch [1155, 113]. punching [246]. punctured [65]. pure [139, 1602]. pursuit [992]. put [1673, 1497].

qSV [1533]. qSV-waves [1533]. quadratic [473, 822, 823, 377, 229, 776]. Quadrature [1025, 1528, 1039, 1029, 1545, 256]. quadratures [1367, 1099]. qualification [658]. qualifications [525, 943]. quality [1929]. quantitative [1857]. quantization [278, 201, 229]. quantized [1123]. quantum [522, 36]. Quasi [1537, 1773, 1148, 863, 491, 402, 933, 1409, 1742, 1042, 1056, 917, 792, 1581, 1553]. quasi-complementarity [1409]. quasi-linear [1042]. Quasi-Monte [1537, 1581, 1553]. quasi-Newton [1742, 917]. quasi-static [1773, 402]. quasi-stationarity [933, 792]. quasi-stationary [863]. quasi-steady [1056]. quasi-steady-state [491]. quasi-trapezoid [1148]. quasilinear [799, 255, 105, 1060]. Quasistationarity [999]. Questions [273, 376]. queue [931]. queue-length [931]. Queueing [1254, 1347, 863, 923, 1141, 553, 561, 558]. queues [1821]. quiescent [689]. Quintic [1400].

R [1231]. Rachford [1614]. Radau [1099]. Radial [686, 1879, 1863]. Radiating [506]. radiation [139, 155]. radiative [387]. radio [1214, 1717]. radius [1638]. radius-dependent [1638]. rainfall [1701]. Ramaswami [1264]. random [1564, 1891, 319, 1798, 152, 375, 83, 1159, 1832, 1598, 1410, 606, 1745, 249, 1820, 1448, 1120]. randomness [930]. range [114, 445]. Rank [773, 1242]. Rank-1 [773]. rapid [219, 750, 330, 1095]. Rapidly [242]. Rapidly-convergent [242]. Rapp [220]. rare [1320]. rate [652, 347, 1212, 1264, 123, 1774, 1137, 1750, 907]. rates [1108, 1094, 1405, 459]. ratio [426, 1607, 261, 1219, 1191]. ratio-dependent [1607, 1191]. rational [636]. ratios [521, 1356]. Ray [246, 40, 64]. Rayleigh [1035, 1543, 1785, 1431, 16, 1803]. re [1731]. re-entrant [1731]. reactant [1077]. Reaction [817, 1739, 685, 1906, 273, 274, 376, 1418, 1591, 651, 1898, 1316, 759, 1071, 1510, 1200, 1257, 1487, 1639, 816, 620, 1018, 715, 1577, 1465]. reactions [254, 1842]. reactive [307, 262, 504]. reactor [167, 1073, 1200, 1288, 1257, 1586, 878]. real [1696]. realisations [1159]. realizations [1654]. rebalancing [299]. reciprocal [125]. Reciprocity [1773, 521, 1143, 1578, 1379]. Reconsidering [1438]. reconstruction [1516]. record [432]. recovery [1693]. rectangles [1419, 1469]. rectangular [359, 1642, 267, 857]. rectilinear [478]. recurrences [793]. recursive [1302, 1414]. redistribution [646, 868]. reduce [1551]. reduced [88]. reducing [1360]. Reduction [595, 1188, 1136, 1734, 775]. Reductions [964]. reefs [1568]. reference [424]. refinement [1004, 1034, 588, 1241]. Refinements [1047, 1105]. Reflected [1454, 1451]. Reflection [812, 897, 372, 1154, 585, 466, 1356, 544]. reflection-transmission [544]. Reflection/refraction [1154]. refraction [1154]. regime [1860, 1861, 1740, 1869]. regime-switching [1861, 1740, 1869]. regimes [265]. region [173, 14, 20, 1167, 577, 960]. regions [1584, 478, 136, 857, 688]. registration [1611, 1830]. regression [1788, 1036]. regular [1539, 1657, 1481]. regular-shaped [1657]. regularisation [420]. regularised [499]. Regularity [1514, 402, 1097, 627, 1250]. regularization [760, 1475]. reinsurance [1498, 1524, 1703, 1464, 1680]. related [67, 1421, 554, 1220, 1335, 1721]. relating [796]. relation [232]. relations [1773, 846, 179]. Relationship [834, 1754]. relationships [586, 377]. Relative [808, 91, 8, 1294, 1313]. relativistic [90]. relativity [139, 279]. relaxation [1693, 1434]. relaxed [1385]. release [43]. Reliability [1519, 735, 476]. remainder [1349]. Remarks [694, 491, 522, 8, 175, 1201]. remodelling [1925]. removal [1630, 1780]. rendezvous [1851]. renewable [1717]. repairable [1519]. reparable [1342]. replacement [1336]. replicator [1825]. representation [940, 107, 417, 1274]. representations [15, 201, 1657]. reptation [1016]. repulsion [83]. repulsive [158]. research [532, 530]. reservoirs [601, 1207]. residual [1833, 1392]. resistance [1505, 530]. resistive [1606]. resolution [12]. resonance [668, 51, 722]. resonances [807]. resonant [594]. resonators [239]. resource [1304]. resource-consumer [1304]. respect [1638, 1314]. response [800, 178, 1905, 371, 736, 204, 1234, 747]. responses [1925, 320]. restricted [126, 146, 132]. result [1406, 988, 107]. results [1741, 1886, 921, 402, 1388, 1421, 174, 1100, 1142, 1456, 919, 1659]. retailers [1705, 1820]. retardation [1214]. retarded [1060]. retina [1749, 1883]. retinal [1805]. retrials [1347]. retrieval [227]. return [1672]. reverse [995]. reverting [1810, 1862, 1618]. review [1875, 1868, 1654]. revised [1242]. revisited [1559, 942, 401, 30, 943, 432]. Reynolds [302, 512, 268, 848]. rheological [1925]. rho [1863]. RI [134]. RI-invariant [134]. Riccati [669, 353, 952]. Ricci [1050]. rich [504]. Riemann [469]. Riemann-integrable [469]. Riesz [1017, 1418]. rigid [1244, 1155, 113]. rings [1907]. rising [300, 899]. risk [1900, 1498, 1524, 1464, 1680]. risk-neutralized [1900]. Ritz [1785]. RLW [1283, 1400]. Robbins [1509]. robin [1617, 988]. Robinson [980, 984]. robot [1083, 888]. robotic [543]. Robust [1315, 1341, 1263, 1169, 1851, 1262, 1430]. Robustness [1361, 1314]. rod [832]. role [1602, 1905]. Rolling [1244]. rolls [312]. root [952, 501]. roots [921]. Rossby [452]. Rössler [1918]. rotating [173, 1243, 1278, 293, 848, 971, 312]. rotation [55, 1797]. rotational [32, 160, 1163]. rotor [1920]. rough [1813, 1359]. round [1617]. routing [863, 561]. row [119]. rows [134]. rubber [55, 544]. rubber-like [544]. rugs [271]. Ruin [1464]. rule [118, 43]. rules [1016]. Rumours [1206, 1255]. run [1900]. Runge [698, 194, 1439, 355, 1899].

Saddle [1287, 250, 1490]. saddle-point [1490]. safe [1917]. safety [1729]. salesman [1666]. Sal’nikov [1639]. sampled [1622, 1674]. sandpile [1335]. satisfying [1313]. Savageau [1193]. scalar [1161, 1696, 1546, 279, 266]. scalar-tensor [266]. Scale [125, 1564, 887, 1321, 142, 1789, 1290, 1734, 1534, 1440]. scale-free [1564]. scaled [1242, 1663]. scales [1326]. scaling [1026, 550, 1228]. scatterer [906, 1773]. Scattering [821, 898, 1478, 770, 288, 906, 789, 726, 880, 1038, 1642, 1132, 1019, 929, 1723, 1610, 1327, 896, 140, 1787]. schedule [931]. schedules [1361]. Scheduling [1719, 1434, 141, 1718, 1293]. scheme [1797, 1689, 736, 1262]. schemes [347, 1598, 1625, 1831]. Schoenberg [119]. Scholes [1453, 1864]. Schrödinger [988, 1902, 1132, 1511, 588, 251, 287]. Schwarzschild [983]. science [1653]. sciences [1905]. scientific [1439]. scientist [1068]. Scoring [1015]. screen [54]. sea [1437]. sealed [1312]. seamount [1279]. search [681, 418, 1742, 989]. searching [514, 520, 591]. seashell [1766]. seasonal [1494, 1730]. Seasonality [1160]. Second [866, 349, 722, 957, 1188, 1579, 1215, 188, 165, 360, 1291, 1351, 946, 1630, 595, 1397, 1901, 1782, 189, 1887, 321, 1042, 1109, 1236, 1259, 1127, 1411, 98, 1202, 1865]. second-grade [1782]. Second-order [957, 1579, 1215, 1291, 1351, 946, 1630, 1397, 1042, 1109, 1236, 1259, 1202, 1865]. section [302, 512, 725, 1463, 637, 448]. sector [427]. Segal [278]. segmented [1313]. Segur [1127]. SEIR [1328]. select [1788]. selection [146, 1625, 527]. selections [1146]. Selective [1556, 829, 1654]. Self [1080, 1805, 468, 1132, 1539, 1928, 1481]. self-adjoint [1132]. Self-excited [1805]. Self-heating [1080]. self-modulation [1928]. self-organisation [468]. self-regular [1539, 1481]. Semi [1739, 78, 348, 425, 1813, 1252, 1387, 1213, 1417, 69]. Semi-analytical [1739, 1813]. semi-definite [1213]. semi-infinite [425, 1213, 69]. semi-Markovian [1252, 1387, 1417]. semi-simple [78, 348]. semidefinite [1539]. semidiscrete [1609]. semilinear [1804, 757, 1180]. semivariograms [764]. sense [1178]. sensing [1606]. Sensitivity [553, 536, 695, 1294]. sensitized [1876]. sensors [391]. separable [1159, 823]. separation [707, 1778, 1129, 982]. sequences [216]. Sequential [1176, 571]. serial [865]. Series [1463, 104, 1769, 1003, 437, 18, 38, 240, 458, 1730, 1595, 397, 1518, 970, 398, 1901, 256, 1156, 748, 1308, 1403]. Serre [1684]. servers [1808, 1719]. services [554]. set [958, 349, 1433, 756, 1372, 91, 250, 1468, 1434, 1868, 481, 175, 634]. set-valued [1433, 175]. sets [990, 746, 796, 819, 809]. setting [1537, 1553, 1581]. several [1139, 1388, 483, 1323, 1577, 1325]. SH [638, 1154]. SH-waves [1154]. shaft [682]. shallow [616, 6, 991, 1282, 1811, 1794, 155, 687]. shallow-water [6]. Shannon [994, 1105]. shape [332, 704, 277, 800, 736, 577]. shaped [1657, 1698]. Sharp [1466, 1268, 1038, 922, 1424, 1402]. Shaw [328, 363, 1868, 487]. shear [1496, 1761, 1205, 724, 1760, 61, 292, 37, 49, 1928]. shear-index [1205]. shear-thinning [1760]. sheared [293]. sheet [800, 466, 1782]. sheets [1839, 230]. shelf [51, 393, 747]. shell [1526, 977]. shell-crossing [977]. shells [686, 566, 958, 616, 1282]. Shepp [990, 746, 819]. Sherman [1462]. shield [533]. shifted [1545]. shim [1095]. Ship [404, 530, 1173]. shock [372, 171, 1871, 544, 1703, 919]. shock-waves [171]. shooting [845]. short [490, 489, 1363, 463]. short-crested [463]. short-time [1363]. short-wave [490]. shout [1772]. Side [51]. sided [845, 1182]. Sigmoidal [1308, 1403, 967]. Sigmoidal-type [1403]. sign [1298]. signal [665, 27, 161]. signals [950, 1159, 161, 261, 748, 1648]. significance [530]. signs [1577]. Similarity [799, 40, 643, 221, 1149]. Simple [1693, 381, 373, 1104, 1004, 1602, 141, 1085, 78, 348, 440, 1519, 1396, 1299, 148, 1622, 816, 1018, 1565, 1467]. simplified [760, 1809]. Simpson [1301]. simulate [1751]. Simulation [1676, 1527, 1701, 764, 1557, 1771, 1854]. simulations [1907]. Sinc [1021]. single [17, 71, 612, 1720, 1434, 1920, 1299]. single-bubble [1299]. Singular [1186, 1229, 195, 303, 190, 1190, 1871, 1039, 903, 957, 1316, 1378, 1534, 1098, 69, 1218, 187, 186, 1781, 1430]. singularities [854, 1116, 669, 434]. Singularity [185, 977]. singularly [1177]. sink [1801, 433, 597, 689, 811, 364, 488, 600, 1135, 1422, 1632, 1686, 317, 847, 892]. SIR [1687, 1565]. SIRS [1567]. sites [1729, 1265]. situ [754]. Sivashinsky [1802]. six [730]. size [1382, 1160, 575, 1789, 1687, 1538, 865, 1664]. size-structured [1382]. sizes [1407]. Skorohod [1727]. slab [602, 654, 115, 1559, 1078, 262]. slaughter [865]. slaughtered [1175]. sleep [1531]. sleeve [848, 971]. slender [392, 929, 7, 1550]. slider [1208, 153]. sliding [1479, 610, 1414, 230]. slip [1443, 1560, 1235, 1619, 1659]. slit [1549]. slope [218]. sloping [364]. slotted [971]. Slow [1850]. Slow-burning [1850]. Slowly [564, 173, 512, 38, 494]. Slowly-varying [564]. sludge [1780]. sluice [831]. Small [848, 148, 1139, 1389, 239, 92, 129, 724, 1314, 1849, 308, 1316, 603, 1687, 1075, 1640, 467]. small-time [1316]. small-update [1389]. Smooth [217, 707, 1307, 592]. smoothed [1907]. smoother [501]. Smoothing [766, 1643, 482, 500, 842]. Smoothly [1045, 949]. Smoothness [500, 1811]. Sobolev [1386, 1593]. social [1892]. soft [896]. soil [913]. solar [1876]. solid [845, 1080, 1559, 1154]. solidification [769]. solids [1079, 372, 3]. solitary [452, 1125, 1338, 870]. soliton [1129, 1126]. solitons [1123]. soluble [75]. solute [1922]. Solution [726, 280, 588, 121, 176, 151, 168, 1018, 63, 470, 272, 192, 185, 743, 45, 141, 21, 652, 1117, 273, 1038, 188, 6, 165, 184, 1351, 1302, 439, 1885, 533, 100, 954, 1166, 1363, 971, 1559, 1595, 50, 565, 255, 1509, 1675, 1316, 1474, 353, 1213, 1463, 1635, 1518, 703, 1545, 1624, 1185, 202, 266, 1400, 1042, 120, 1333, 1103, 1769, 809, 187, 1452, 64, 1724, 1445, 892, 1309, 1684]. Solutions [339, 1679, 1298, 1496, 1739, 1087, 1107, 615, 1121, 846, 1906, 203, 318, 1051, 274, 376, 1227, 524, 656, 1418, 360, 1804, 547, 1419, 1469, 614, 1182, 179, 458, 1608, 1918, 1151, 1340, 1375, 1525, 1673, 839, 799, 1220, 1166, 154, 488, 351, 761, 535, 669, 1190, 1511, 981, 97, 1397, 1456, 393, 499, 1876, 1510, 1043, 1514, 1846, 3, 1075, 1818, 1277, 221, 13, 80, 287, 646, 987, 964, 852, 647, 876, 1495, 1259, 1246, 968, 1218, 1412, 1018, 504, 1845, 710, 186, 910, 40, 1074, 1149, 245, 593, 1465, 1191, 1292, 1489, 1232, 1280, 558]. Solvability [939, 1339, 1089, 1161, 1517]. solvable [337]. solvers [849]. Solving [1242, 1634, 1626, 1482, 969, 1879, 126, 1197, 793, 1833, 1592, 1789, 1689, 498, 1499, 1411, 784, 1379]. Some [139, 641, 1, 491, 1388, 522, 1142, 643, 604, 8, 1105, 175, 13, 625, 1274, 1330, 1594, 1201, 710, 855, 1037, 515, 282, 470, 1057, 1107, 1406, 299, 231, 846, 1032, 274, 1080, 437, 439, 1340, 304, 959, 1459, 644, 1152, 393, 25, 84, 1755, 1088, 1047, 1689, 1195, 964, 1089, 1909, 256, 430, 1659]. sonoluminescence [1299]. sor [1455]. Source [876, 433, 1543, 1894, 364, 488, 510, 688, 987, 317, 911, 910, 64, 1577, 1416]. sources [1377, 1323]. Space [1415, 1426, 1032, 278, 87, 1155, 1418, 1715, 983, 643, 1159, 1832, 994, 523, 1537, 1553, 1581, 101, 1534, 901, 494, 574, 1165, 890]. space-fractional [1418]. Space-like [1415]. spacecraft [1851]. spaced [359]. spaces [125, 344, 1433, 1409, 1314, 1154, 127, 1089, 1533]. spacetimes [980]. spacies [159]. spanning [1722]. Sparse [451, 1321, 1818, 358]. Spatial [1565, 1544, 1566, 1765, 1844, 1594]. spatially [1849, 645]. spatially-distributed [645]. Special [1115, 1913, 274, 1435, 111, 1676, 377, 1838, 1626, 1124, 1857]. species [428, 612, 252, 320, 605, 1257, 1247, 1191]. specification [954]. specified [764]. spectra [1538]. Spectral [12, 903, 1241, 636, 1004, 1034, 278, 490, 1351, 806, 984, 1499, 515]. spectral-Galerkin [1351]. Spectrally [1864]. spectrum [1665, 1693, 232, 788]. Speech [1069]. speed [596, 1225, 744]. SPH [1401]. sphere [223, 439, 1619, 1074, 296]. spheres [1746]. Spherical [377, 1655, 1891, 1712, 1894, 300, 1204, 1619, 1095, 1526]. spheroids [173]. spillway [68]. spin [1122, 36]. spinor [1426]. Splash [949]. splashless [404]. spline [344, 1006, 500, 1400]. splines [1024, 1283, 823, 1872, 482]. split [1723]. Splitting [383]. spontaneous [1486, 1755]. spot [476, 487, 204]. spray [1916]. spray-on [1916]. spread [1298, 235, 887, 1477, 1727, 1505]. spreading [1817]. spreading-vanishing [1817]. sputtering [832]. square [27, 952, 501]. square-root [952, 501]. squares [126, 146, 114, 1036, 383, 30, 358]. squeeze [148]. squeezing [102]. Sraffa [343]. stabilisation [1263]. Stability [1087, 194, 268, 724, 142, 138, 1417, 1515, 1534, 1304, 918, 1342, 1146, 1058, 1044, 1107, 930, 31, 1638, 1217, 1340, 1375, 320, 374, 395, 51, 416, 1314, 135, 1252, 1387, 942, 8, 1086, 97, 759, 1468, 1580, 1663, 664, 683, 1290, 1341, 1378, 1073, 1140, 1901, 1185, 879, 1538, 1909, 798, 1093, 1329, 1266, 1234, 1232]. stabilization [1898, 1378, 1916]. Stable [479, 449, 514, 925]. stage [355, 1895, 1531]. stagnant [688]. stagnation [1527]. standard [1537, 1553, 1581]. state [491, 1315, 1340, 1108, 1372, 1679, 1239, 595, 112, 1075, 1534, 40, 64, 1074]. states [1741, 164, 158, 273, 274, 376, 167]. Static [266, 1773, 125, 402, 1580, 32]. stationarity [933, 792]. Stationary [1886, 293, 863, 416, 981, 80, 498, 5, 1232]. statistic [261]. statistical [477, 1294, 774]. Steady [273, 274, 376, 1334, 491, 167, 1279, 575, 1632, 1559, 1075, 1578, 40, 64, 1074, 1056]. steady-state [1075, 64]. Stefan [547]. Steiner [1835]. stenosed [1698]. step [1539, 1474, 1789, 251, 1208, 513]. stepping [1824]. stern [777, 1173, 1092]. Stieltjes [89, 922]. Stiff [200, 210, 222, 953]. stiffness [1026]. stirred [1200, 1257]. Stochastic [520, 1252, 1851, 794, 1354, 1477, 1859, 1927, 1900, 555, 1387, 1810, 1592, 50, 1790, 1826, 1929, 1454, 1909, 968, 1863, 1865, 1671]. stock [1861]. Stokes [15, 491, 1785, 300, 1752, 1276, 723, 849, 882, 1250]. storage [1657, 1917, 104]. store [1820]. stored [1071]. straddling [392]. straight [637, 243]. strain [199, 148, 1571, 1445]. strain-gradient [1445]. straining [1608]. Strategic [1821, 1808]. strategies [514, 681, 858, 744, 137]. strategy [1494, 1827, 135, 576, 1745, 503]. stratification [870]. stratified [925, 854, 293, 484, 829]. stream [949]. Streamlined [1818]. streamlines [329]. streams [829]. streets [365]. stress [1496, 26, 183, 1526, 276]. stress-deformation [276]. stressed [1533]. stresses [566, 1533]. stretched [246]. stretching [1782]. Strict [944]. string [978]. Strong [384, 385, 942, 386, 571, 33, 1871, 172, 1901, 852]. Strongly [928, 449, 1405, 694]. structural [1557]. structure [90, 134, 1220, 1479, 610, 1738, 1426, 662]. structured [21, 1382, 1264, 1185, 120]. structures [1133]. Struve [297]. student [2]. studies [862, 777, 487, 715]. study [764, 1523, 1771, 27, 1560, 637, 1877, 865, 867, 622, 1505, 1711, 842, 1177, 1740, 1379]. studying [1566]. Sturm [1104, 537, 128, 1218]. style [1753, 1861]. Sub [1335]. Sub-supersolutions [1335]. subcritical [1386]. subdifferentiability [944]. Subgradient [940]. Subharmonic [853, 1280]. subject [954, 1559, 1586]. submarine [1642]. submerged [898, 433, 1875, 364, 488, 1632, 510, 929, 603, 770, 1327, 688, 290, 155, 317, 911]. submicron [1606]. Suboptimality [233]. subpopulations [683]. subquadratic [1280]. subset [1788, 146]. Subsonic [260, 462]. subspaces [665, 773]. substrate [1586]. subtraction [185]. Successive [1412, 951]. suction [725]. suction/injection [725]. Sudden [631]. Sufficient [42, 1307, 549, 443]. suitable [418]. sum [737, 1892, 990, 27, 746, 796, 819]. summable [1089]. summation [1355]. Summing [1003, 38, 18]. sums [240, 1245, 1364, 1588]. sup [882]. superadditivity [1059]. superalgebras [440]. Supercritical [68, 951, 1615]. superexponential [742]. superfast [459]. superfluid [308]. supergravity [976]. superintegrable [1486]. superlinear [1055]. superposition [1128]. supersolutions [1335]. Supplementary [923]. supporting [1905]. Surface [290, 1516, 391, 1801, 77, 831, 1527, 642, 889, 930, 1244, 880, 912, 897, 1038, 433, 597, 689, 628, 394, 364, 600, 1686, 1871, 308, 510, 434, 688, 1338, 1782, 310, 721, 1668, 1289, 868, 217, 218, 317, 710, 911, 289]. surface-tension [689]. Surface-wave [290]. surfaces [300, 1443, 5, 1056, 1359]. survey [1029]. suspensions [504]. SVM [1711]. swaps [1732, 1860, 1900, 1622, 1674, 1618]. sweep [487]. Swift [1771]. Swimming [7, 1761, 1760]. swirling [1638]. switched [1384, 1341, 1558, 1263]. switching [28, 1860, 1679, 1387, 556, 1417, 1861, 1195, 676, 890, 1740, 1869, 558]. Symm [675]. Symmetric [713, 1525, 939, 415, 614, 1543, 365, 516, 386, 13, 80, 266, 1415]. Symmetries [301, 964, 1654]. Symmetry [1406, 1278, 1122, 706, 685, 147]. symmetry-breaking [706]. symplectic [1226]. synchronization [1720]. synchronized [1821]. Synthesis [958, 288]. system [384, 862, 1688, 668, 923, 1141, 385, 553, 219, 111, 119, 337, 374, 395, 788, 1519, 1363, 1062, 1607, 765, 994, 1316, 1515, 1000, 634, 610, 312, 494, 9, 281, 1124, 1164, 794, 1234, 1263, 1342, 917, 1191, 1292, 1416, 558]. Systems [830, 1347, 1242, 1665, 1044, 1107, 473, 663, 749, 808, 1570, 44, 1226, 1339, 838, 89, 233, 1048, 767, 107, 613, 1567, 1798, 573, 564, 1302, 626, 1591, 756, 1315, 1007, 1488, 320, 1169, 1133, 1540, 1476, 57, 539, 1252, 1387, 227, 1343, 1190, 1233, 820, 1511, 142, 1229, 605, 696, 1898, 888, 595, 138, 1384, 1290, 1341, 1417, 1631, 1378, 200, 210, 222, 1254, 1545, 1734, 1076, 1534, 175, 1060, 918, 373, 545, 36, 498, 301, 966, 98, 419, 875, 1329, 1695, 1385, 1574, 776, 1558, 1391, 1323, 779, 840, 662, 1465, 1781, 502, 1280, 1430]. Szekeres [1333].

T. [531]. table [913]. tale [1119]. Taming [1116]. tank [1200, 1257]. Tanner [1884]. Tasmanian [1570]. Taupo [1440]. Taylor [437, 1771, 1543, 1349, 1901, 748, 1803]. technique [1895, 765, 1624]. techniques [764, 1049, 479]. technology [1111, 1653]. telecommunications [607, 751]. Temperature [1431, 832, 643, 1516, 546]. temperature-dependent [546]. temperatures [262]. Temporal [1765, 798, 281, 621]. tension [912, 897, 689, 628, 684, 1686, 1338, 721, 1668, 217, 218]. tensor [87, 78, 266]. term [859, 883, 1349, 703, 1367, 1587, 545, 979]. termination [1672]. terms [38, 240, 1577]. test [1392, 1076]. testing [209, 1841, 1413]. Texas [1771]. th [513]. their [1121, 871, 1423, 66, 698, 1245, 287, 58, 1359]. theorem [378, 195, 1217, 162, 723, 1330]. theorems [738, 521, 278, 1410, 678, 105, 1110, 1109, 797, 1061]. theoretical [150, 1594]. theory [1111, 1186, 1176, 743, 307, 738, 62, 208, 278, 735, 1205, 564, 1603, 179, 587, 741, 665, 1629, 511, 472, 991, 250, 211, 298, 390, 434, 1047, 41, 266, 494, 140, 74, 1835, 241, 281, 403, 546, 230, 217, 218, 1265, 1404, 1173]. therapy [877]. Thermal [262, 254, 14, 645, 1436, 438, 1074]. thermally [925]. thermocapillary [1516]. thermocline [889, 930, 1106, 291]. Thermodynamic [9]. Thermodynamics [1362]. thermoelastic [115, 1100, 1356]. thermons [1070]. theta [242]. thickness [642, 652]. Thin [467, 898, 889, 930, 1106, 533, 1782, 736, 1550]. thinning [1760]. third [1350, 1300, 1227, 355, 461]. third-order [1350, 1300, 1227]. Thompson [1826]. Three [1603, 1885, 1397, 1640, 221, 737, 75, 318, 355, 1528, 1354, 991, 275, 1883, 901, 1312, 1840, 1232, 1455, 747]. three-axial [737]. Three-dimensional [1603, 1885, 75, 901, 1840, 1232, 1455, 747]. Three-layer [1640, 1883]. three-point [1528, 1397]. three-stage [355]. three-variable [1312]. thresholds [1858]. Thrifty [1760]. tight [1713]. Time [834, 1496, 1694, 1825, 737, 909, 44, 838, 1884, 234, 233, 999, 1451, 822, 983, 587, 536, 159, 1169, 1540, 1108, 1476, 1363, 1532, 1730, 1251, 1849, 1616, 957, 1731, 56, 1316, 1474, 138, 1479, 1625, 1405, 1417, 952, 1378, 1514, 1824, 79, 338, 1534, 1862, 501, 574, 74, 966, 1575, 636, 677, 1426, 132, 419, 445, 1103, 1864, 322, 1558, 1326, 1191, 792]. time-consistent [1625]. time-delay [1169]. time-delayed [838]. time-dependent [1496, 56, 1862, 636]. time-fractional [1731, 1514, 1864]. time-fuel [1479]. time-lag [132, 419, 322]. time-periodic [909]. time-varying [138, 1417]. timer [1865]. times [371, 1878]. timescales [1575]. Timoshenko [1137]. tissue [1763, 1809]. tissues [1925]. Toeplitz [1325, 1399]. Tolman [1374]. tomography [1253]. tone [119, 134]. topographies [393]. topography [293, 1794]. topography-driven [1794]. Topological [1083, 428, 1433, 1409]. Topology [1559, 631, 101, 908]. toroidal [205, 535, 1275, 1276]. toroidal-poloidal [535]. torque [582]. torsion [720]. total [1407, 1336]. tournament [263]. tournaments [1617]. Tower [843]. trace [1386, 251]. tracers [298]. Traces [1174]. trachoma [1877]. track [1694, 858, 1778, 1434]. Tracking [1384]. traction [3]. trade [737]. trade-off [737]. trading [1858]. traffic [1254, 1854, 41]. trailing [467]. trailing-edge [467]. train [1694, 576, 858, 1778, 744]. trains [1434, 288]. Trajectory [81, 543]. transcendents [1120]. Transcritical [1504]. transfer [1243, 451, 387, 1204, 14, 1458, 719]. transform [969, 818, 803, 297, 932, 635, 844, 941, 1714]. transformation [231, 1128, 1592, 25, 84, 1545, 1130]. transformations [983, 644, 595, 53, 682]. transforms [742, 922, 1377, 1322]. Transient [371]. Transition [1662, 594, 668, 319, 1623, 1728, 1108]. Translations [53]. Transmission [1649, 1303, 856, 544, 1854, 1187, 1551, 1877]. transonic [260]. transport [1441, 596, 519, 40, 64, 1440, 1922]. transportation [480, 737]. transversely [1155, 3]. trapezoid [1148]. trapezoidal [1366]. trapped [1668]. trapping [587]. traps [1849]. Trautman [980, 984]. Travelling [769, 1324, 1510, 1639, 504, 1465, 845, 1666, 1591]. treatment [1650, 1770, 25, 84]. tree [523, 1835]. trees [1136, 1722, 1721]. trench [1642]. tribute [1759]. tridiagonal [1399]. trigonometric [1438]. triple [25, 84, 398]. tritrophic [1754]. trivariate [901]. tropical [747]. Truck [1729]. truncated [1595]. trust [1167, 960]. Trying [1320]. Tsuno [1392]. tube [302, 512, 899, 102]. tuberculosis [1505]. tubes [55, 725]. tumour [1570]. Turán [1367]. turbulence [1623, 1662, 1728]. turbulent [1623]. twelve [134]. twelve-tone [134]. Two [239, 67, 433, 1138, 487, 692, 1466, 845, 359, 1255, 654, 832, 264, 1906, 921, 764, 362, 382, 1557, 107, 1556, 328, 363, 1182, 990, 887, 1646, 337, 1525, 159, 252, 320, 190, 1365, 1281, 1843, 154, 1422, 1778, 1895, 1719, 1209, 1607, 991, 827, 605, 1154, 1775, 519, 899, 1498, 383, 1327, 1078, 807, 688, 919, 427, 1534, 32, 746, 819, 1268, 1247, 721, 1571, 886, 1428, 1668, 1495, 342, 1787, 1411, 1312, 780, 1149, 1002, 1174, 1698, 1191, 1292]. two- [1312]. two-body [1525]. Two-dimensional [239, 692, 328, 363, 887, 1281, 899, 1078, 1411]. two-fluid [654, 832]. two-input [107]. two-layer [1556, 1422, 807, 919, 1668, 1495]. two-layered [827, 721, 1698]. two-output [107]. two-patch [1191]. Two-phase [487, 886, 1149]. Two-point [1138, 1466, 1365, 383, 1268]. two-prey [1209]. two-sector [427]. two-sided [845, 1182]. two-species [320, 605, 1191]. two-stage [1895]. two-strain [1571]. two-time-scale [1534]. two-train [1778]. two-way [1557]. type [1879, 1032, 1423, 767, 38, 859, 1528, 839, 757, 696, 1424, 1402, 1180, 995, 1138, 703, 1515, 1846, 1593, 41, 506, 625, 1089, 852, 1099, 1021, 1391, 1234, 430, 1489, 1403]. types [468, 964].

Uhlenbeck [1732]. ultimate [22, 1175]. ultrasonic [1841]. ultraspherical [1351]. unbounded [1108]. uncertain [1169, 1378, 865, 1263]. Uncertainty [635, 1571]. unconstrained [1576, 1742]. uncoupling [539, 366]. underground [1835]. understands [1771]. undulatory [1760]. unfair [1310]. unified [880, 1196, 321]. Uniform [1027, 128, 594, 1902, 1775, 683, 7]. unilateral [431]. uniquely [66]. Uniqueness [274, 738, 376, 1804, 626, 452, 1846, 571, 1232]. unitary [88]. unmanned [1580]. unstable [363, 1074]. Unsteady [512, 342, 1289, 1458, 1698, 100, 991, 1560, 1782, 102]. unsymmetric [262]. up-and-in [1675]. update [1389, 1598]. upper [1432, 1166]. uptake [513]. upwards [692]. upwind [1690]. urn [1544]. Usage [1392]. usd [1678]. use [969]. used [1876]. Using [793, 1360, 1188, 1795, 1153, 1584, 216, 789, 1, 1080, 1557, 1351, 425, 1283, 823, 991, 1241, 1898, 1635, 1678, 1755, 1088, 1434, 390, 1377, 1531, 1634, 1401, 513, 1835, 104, 1698]. utilising [1095]. utility [1498, 970]. Uzawa [1490].

vacation [931]. vacations [1821]. vaccination [1494, 1569]. vacuum [980, 981, 80]. validation [193]. valleys [638]. valuation [1900, 1453, 1309]. value [1161, 106, 272, 641, 1406, 738, 685, 1291, 1804, 179, 439, 1808, 1365, 877, 1595, 1229, 1786, 1374, 1474, 1397, 1498, 383, 210, 1043, 1425, 321, 787, 1262, 1218, 1412, 1201, 405, 493, 902, 461, 479]. valued [1433, 175]. values [442, 1425]. vanilla [1678]. vanishing [1817]. vapour [228]. variable [923, 165, 1008, 595, 1479, 1463, 1338, 865, 610, 1103, 1312, 1448, 662]. variable-structure [662]. variables [1168, 1388, 1726, 166, 606, 1129]. Variance [834, 1732, 1827, 1900, 1625, 1703, 1622, 1618]. variates [1245]. variation [1383, 750, 1630, 330, 1235, 1165]. variation-of-constants [1165]. Variational [89, 682, 939, 106, 1584, 615, 1024, 1162, 33, 1152, 1335, 549, 879, 1385]. variations [6]. various [1215]. vary [494]. varying [302, 512, 725, 564, 786, 858, 827, 138, 1417]. Vector [1044, 1433, 1409, 698, 535, 1178, 1275, 1276, 1216]. vehicle [1580]. vehicular [41]. velocities [32]. velocity [7, 1587, 1235]. velocity-slip [1235]. vendor [1745]. vendor-managed [1745]. venous [1805]. ventilation [682]. version [1392]. versus [1741]. Vertical [584, 1801, 821, 898, 726, 812, 880, 897, 888, 1478, 1610, 603, 1327, 466, 881, 896, 1787, 1458]. vertically [293, 1154, 692]. vessels [831, 1765, 1764]. VI [625]. via [1677, 1466, 1287, 1107, 615, 278, 1851, 133, 1348, 833]. vibrations [1035]. viscoelastic [958, 1762, 113]. viscosity [1235]. viscous [1793, 1556, 1646, 1552, 625, 627, 1495, 102, 484, 16]. visits [1365]. visual [924]. Voigt [1142, 1196]. Volatility [1900, 1704, 1859, 1927, 1810, 1678, 1861, 1865, 1618]. volcanic [1396, 1446, 1440]. Volterra [194, 663, 749, 808, 1048, 252, 374, 395, 1128, 605, 696, 1515, 968]. Volume [94, 143, 196, 236, 283, 324, 368, 409, 454, 495, 540, 579, 659, 699, 872, 914, 1009, 1064, 1112, 1170, 1222, 1380, 1381, 1012, 10, 11, 23, 24, 34, 35, 47, 48, 59, 60, 72, 73, 85, 86, 95, 96, 108, 109, 116, 117, 130, 131, 156, 157, 169, 170, 144, 145, 181, 182, 197, 198, 206, 207, 214, 215, 224, 225, 237, 238, 247, 248, 258, 259, 269, 270, 284, 285, 294, 295, 305, 306, 314, 315, 325, 326, 335, 336, 345, 346, 356, 357]. volume [369, 370, 379, 380, 388, 389, 399, 400, 410, 411, 422, 423, 435, 436, 446, 447, 455, 456, 464, 465, 474, 475, 485, 486, 496, 497, 507, 508, 517, 518, 528, 529, 541, 542, 551, 552, 559, 560, 568, 569, 580, 581, 589, 590, 598, 599, 608, 609, 617, 618, 629, 630, 639, 640, 648, 649, 660, 661, 670, 671, 679, 680, 690, 691, 700, 701, 708, 709, 717, 718, 728, 729, 739, 740, 752, 753, 762, 763, 771, 772, 781, 782, 790, 791]. volume [801, 802, 813, 814, 825, 826, 836, 837, 850, 851, 860, 861, 873, 874, 884, 885, 894, 895, 904, 905, 915, 916, 926, 927, 936, 937, 947, 948, 962, 963, 972, 973, 985, 986, 997, 998, 1010, 1011, 1030, 1031, 1040, 1041, 1052, 1053, 1065, 1066, 1081, 1082, 1090, 1091, 1101, 1102, 1113, 1114, 1134, 1144, 1145, 1157, 1158, 1171, 1172, 1183, 1184, 1198, 1199, 1210, 1211, 1223, 1224, 1237, 1238, 1248, 1249, 1260, 1261, 1271, 1272, 1284, 1285, 1295, 1296, 1270, 1305, 1306]. volume [1318, 1319, 1331, 1332, 1345, 1346, 1317, 1357, 1358, 1370, 1371, 1393, 1394, 1369, 1449, 1450, 1460, 1461, 1470, 1471, 1483, 1484, 1491, 1492, 1501, 1502, 1512, 1513, 1521, 1522, 1529, 1530, 1535, 1536, 1541, 1542, 1547, 1548, 1554, 1555, 1561, 1562, 1572, 1573, 1582, 1583, 1589, 1590, 1596, 1597, 1604, 1605, 1612, 1613, 1620, 1621, 1627, 1628, 1636, 1637, 1644, 1645, 1651, 1652, 1660, 1661, 1669, 1670, 1682, 1683, 1691, 1692, 1699, 1700, 1706, 1707, 1735, 1736, 1747, 1748, 1756, 1757, 1767, 1768]. volume [1776, 1777, 1783, 1784, 1791, 1792, 1799, 1800, 1806, 1807, 1814, 1815, 1822, 1823, 1828, 1829, 1836, 1837, 1847, 1848, 1855, 1856, 1866, 1867, 1873, 1874, 1881, 1882, 1888, 1889, 1896, 1897, 1903, 1904, 1911, 1912, 1923, 1924, 1930, 1931, 1544, 1819, 1690]. volumetric [1701]. Vortex [1398, 1279, 365, 298, 466]. vortices [909, 1685, 928]. vorticity [1092, 1045, 1794, 1250]. Vries [1641, 1831].

w [1698]. w-shaped [1698]. Wairakei [1447]. Waiting [834]. Waiting-Time [834]. Wake [1838, 1069]. walk [152, 375, 83, 249]. walks [1598]. Wall [1916, 821, 898, 812, 897, 899, 1192]. walls [533]. Wang [1413]. warehouse [1666]. warfare [135]. warrants [1908]. warring [503]. wastewaters [1650]. Water [1327, 287, 1787, 821, 898, 726, 930, 286, 913, 812, 880, 897, 1038, 6, 490, 1106, 1798, 1603, 822, 177, 394, 1902, 707, 991, 929, 1775, 601, 1207, 434, 603, 770, 1929, 463, 585, 712, 714, 466, 1811, 1338, 310, 896, 1658, 5, 155, 911, 1440, 1005]. Wave [1084, 921, 90, 912, 1642, 6, 66, 490, 1702, 1798, 925, 1603, 1072, 1601, 628, 51, 177, 292, 393, 1610, 1327, 1510, 1125, 1225, 290, 721, 1787, 288, 504, 530, 493, 1577, 1465]. wave-free [1603]. wave-maker [628]. waveguides [1741]. wavelength [909]. Wavelet [818, 950, 1322, 1233, 994, 1689, 1634]. wavelets [1174]. wavemaking [777]. Waves [492, 845, 821, 898, 769, 789, 726, 889, 930, 1079, 286, 92, 129, 334, 812, 880, 897, 1038, 1733, 638, 1106, 724, 1603, 372, 293, 171, 1591, 1077, 51, 489, 1902, 57, 291, 308, 1092, 452, 1154, 929, 1775, 1324, 1478, 434, 603, 770, 463, 585, 1225, 466, 1338, 1928, 1639, 287, 310, 881, 896, 870, 342, 1533, 1649, 289, 817, 864, 296]. way [1557]. weak [1804, 1151, 1468, 1517]. weakly [1741, 582, 583, 584, 673, 190, 903, 957, 187, 186]. wedge [585]. weibull [1407, 1245]. weibull-distributed [1407]. weight [1016]. weighted [1537, 1553, 1581, 1366]. welding [476]. Welfare [672, 828, 1892]. well [1239, 1033]. well-known [1033]. wells [712]. West [747]. Western [1505]. wheelset [311]. which [494, 1359]. Wick [1727]. Wiener [1017, 1420, 450]. Wigner [844]. wind [1775, 624]. wing [1057, 1921]. Withdrawal [847, 1556, 811, 1135, 1422, 991, 1289, 829, 892]. without [832, 658, 696, 1180, 1207, 1356]. WKB [1127]. Wolbachia [1551]. work [531]. woven [1839].

XXI [144, 145, 143].

yield [1309].

Zakharov [490]. Zealand [1446, 1905, 1569, 1440]. Zernike [1448]. zero [1, 97, 13]. zero-finding [1]. zeros [244, 1593]. Ziebur [429]. zigzag [1420]. zinc [1444]. zonal [1712, 833]. zone [1440]. zones [468].

References

Brent:1975:SHO

[1] Richard P. Brent. Some high-order zero-finding methods using almost orthogonal polynomials. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):1–29, June 1975. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/some-highorder-zerofinding-
methods-using-almost-orthogonal-polynomials/3A3E44B4410EABE7AFED6DDB8236FECA.

Campbell:1975:PMS

[2] L. H. Campbell. A programming model for student enrolment planning. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):30–39, June 1975. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/programming-model-for-
student-enrolment-planning/45473546B51ED497AD406732AA86AA55.

Morland:1975:ESP

[3] L. W. Morland. Existence of solutions of plane traction problems for inextensible transversely isotropic elastic solids. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):40–54, June 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-of-solutions-of-plane-traction-problems-for-
inextensible-transversely-isotropic-elastic-solids/AA652553D37C9C5880C03223924AFDF3.

Patterson:1975:MBP

[4] J. C. Patterson and J. M. Fitz-Gerald. The measurement of blood pressure. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):55–65, June 1975. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/measurement-of-blood-pressure/29B5FEF6FE54B7098D51C68C2938F67A.

Tuck:1975:AFF

[5] E. O. Tuck. On air flow over free surfaces of stationary water. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):66–80, June 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-air-flow-over-free-surfaces-of-stationary-water/
E3E4D2ACF24A36C98F31F7C7472D5A36.

Clements:1975:ASL

[6] D. L. Clements and C. Rogers. Analytic solution of the linearized shallow-water wave equations for certain continuous depth variations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):81–94, June 1975. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/analytic-solution-of-the-linearized-
shallowwater-wave-equations-for-certain-continuous-depth-variations/
84F0A58A17B02D0FD982AEA09CBCE4E8.

Newman:1975:SSF

[7] J. N. Newman. Swimming of slender fish in a non-uniform velocity field. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):95–111, June 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/swimming-of-slender-fish-in-a-nonuniform-velocity-field/
6FAD07A3CECF7F1FAC67E90B68B610C1.

Kloeden:1975:SRR

[8] Peter E. Kloeden. Some remarks on relative stability. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):112–115, June 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-remarks-on-relative-stability/ACA7BD61084AD55235DCA43571F7325E.

Smith:1975:TLS

[9] E. R. Smith and J. W. Perram. Thermodynamic limit for a system with dipole–dipole interactions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):116–128, June 1975. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/thermodynamic-limit-for-a-system-
with-dipoledipole-interactions/75E5B878D562A6976C6208608A9D5280.

Anonymous:1975:AVIa

[10] Anonymous. ANZ volume 19 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):f1–f2, June 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-19-issue-1-cover-and-front-matter/EAF0FABE6CB6A47E1940FCC6AE1E5BF0.

Anonymous:1975:AVIb

[11] Anonymous. ANZ volume 19 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(1):b1–b2, June 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-19-issue-1-cover-and-back-matter/486887B32056A4F5DCDC935F10EEA746.

Green:1975:SRI

[12] H. S. Green. Spectral resolution of the identity for matrices of elements of a Lie algebra. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):129–139, December 1975. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/spectral-resolution-of-the-
identity-for-matrices-of-elements-of-a-lie-algebra/C4E64AE062E86E31278BBA13F33521CE.

Patel:1975:SAS

[13] L. K. Patel and V. M. Trivedi. Some axially symmetric zero mass meson solutions of Einstein’s equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2): 140–145, December 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-axially-symmetric-zero-mass-meson-solutions-of-
einsteins-equations/CF967A4D5221658476A94C1F2D10891F.

Jones:1975:HTT

[14] A. S. Jones. Heat transfer in the thermal entrance region of a flat duct. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):146–160, December 1975. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/heat-transfer-in-the-thermal-entrance-region-
of-a-flat-duct/4004EF3E4591A523197C4E8D4E718E36.

Blake:1975:NIR

[15] J. R. Blake. A note on integral representations in Stokes flow. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):161–164, December 1975. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/note-on-integral-representations-in-stokes-
flow/4122BF455DFCB0FD5BB515245233BA37.

VanderBorght:1975:EVD

[16] R. Van der Borght. The effect of viscous dissipation on nonlinear convection at high Rayleigh number. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19 (2):165–172, December 1975. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/effect-of-viscous-dissipation-on-nonlinear-
convection-at-high-rayleigh-number/0CF3799D8172F584096D5B741E6D94F1.

Abbott:1975:FLC

[17] James P. Abbott and Richard P. Brent. Fast local convergence with single and multistep methods for nonlinear equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19 (2):173–199, December 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fast-local-convergence-with-single-and-multistep-methods-
for-nonlinear-equations/DFE20C8FE10034CDEC4D249FC8BD9E5B. See

correction [71].

Drummond:1975:FSD

[18] J. E. Drummond. A formula for summing divergent series. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):200–214, December 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/formula-for-summing-divergent-series/AA97F0381EA7D1FEF0A4283EFB1CD326.

Bhargava:1975:CCM

[19] R. D. Bhargava and Ram Narayan. Concentric circular misfit and a circular arc crack in an infinite isotropic elastic plate. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):215–228, December 1975. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/concentric-circular-misfit-
and-a-circular-arc-crack-in-an-infinite-isotropic-elastic-plate/
AE28CB39ED2A83F98E1DF1DB0B543234.

Steven:1975:CPE

[20] G. P. Steven. Contact problem of an elastic plug in an elastic region. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):229–241, December 1975. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/contact-problem-of-an-
elastic-plug-in-an-elastic-region/4C6B2E49AA5B2BCABC0FF49A584F381E.

Berry:1975:SSN

[21] L. T. M. Berry. On the solution of a structured nonlinear programme. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):242–248, December 1975. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-the-solution-of-a-structured-nonlinear-programme/
13EC1C96AAA47206147EA27D69B2B789.

Kloeden:1975:ADE

[22] P. E. Kloeden. Aggregation-decomposition and equi-ultimate boundedness. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):249–258, December 1975. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/aggregationdecomposition-and-equiultimate-
boundedness/DCD3853780FCA54F50F18C1B56320526.

Anonymous:1975:AVIc

[23] Anonymous. ANZ volume 19 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):f1–f2, December 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-19-issue-2-cover-and-front-matter/114822AC54C7C92FB457A1E3B6959221.

Anonymous:1975:AVId

[24] Anonymous. ANZ volume 19 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(2):b1–b2, December 1975. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-19-issue-2-cover-and-back-matter/D996167D76BEC3BA3EF234A6035A0462.

Love:1976:TCT

[25] E. R. Love and D. L. Clements. A transformation of Cooke’s treatment of some triple integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):259–288, June 1976.

CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/transformation-of-
cookes-treatment-of-some-triple-integral-equations/6DE390DA6B6D8D65C8D5CBD3C8CFD53A.

Basu:1976:NSC

[26] Animesh Basu. A note on stress concentration around an elliptic hole in micropolar elasticity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):289–293, June 1976. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
stress-concentration-around-an-elliptic-hole-in-micropolar-
elasticity/C2970B00D4DCF9EE059523EB5027B613.

Keats:1976:GSS

[27] R. G. Keats and V. K-K. Yu. A generalisation of the study of sum and square law signal processors with multiple clipped inputs. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):294–315, June 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalisation-of-the-study-of-sum-and-square-law-signal-
processors-with-multiple-clipped-inputs/B2C6A4602A0798CEC3CFB4C28D6AD5B2.

Blatt:1976:OCC

[28] John M. Blatt. Optimal control with a cost of switching control. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):316–332, June 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-control-with-a-cost-of-switching-control/85799AECC9C0072C63047C0403BBCF50.

Mond:1976:PPN

[29] Bertram Mond and Murray Schechter. A programming problem with an L_{p} norm in the objective function. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):333–342, June 1976. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/programming-problem-with-an-lp-norm-in-the-
objective-function/172EC170F1A38C557EC9CE423F147E8C.

Osborne:1976:NLS

[30] M. R. Osborne. Nonlinear least squares — the Levenberg algorithm revisited *. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):343–357, June 1976. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-

journal/article/nonlinear-least-squares-the-levenberg-algorithm-
revisited/BD5B9EA8DB2F092020CC173354C4C86E.

Braddock:1976:SDD

[31] R. D. Braddock and P. Van Den Driessche. On the stability of differential–difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):358–370, June 1976. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-stability-of-differentialdifference-equations/EC053938154F4962A70A318CE1D53641.

Paria:1976:ESM

[32] G. Paria and A. K. Sharma. Effect of static magnetic field on the helical flow of incompressible cholesteric liquid crystal between two coxial circular cylinders having rotational and axial velocities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):371–380, June 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/effect-of-static-magnetic-field-on-the-helical-flow-
of-incompressible-cholesteric-liquid-crystal-between-two-coxial-
circular-cylinders-having-rotational-and-axial-velocities/A3A27AE17889F06A31D0B512512DC003.

Gregory:1976:VDE

[33] L. J. Gregory and A. H. Klotz. On the variational derivation of Einstein’s strong field equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):381–386, June 1976. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-variational-derivation-of-einsteins-strong-field-equations/
01C83C9705CAFCB11053AD6C922D3A85.

Anonymous:1976:AVIa

[34] Anonymous. ANZ volume 19 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):f1–f2, June 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-19-issue-3-cover-and-front-matter/9718602478D6E8612771476D115F2CF9.

Anonymous:1976:AVIb

[35] Anonymous. ANZ volume 19 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(3):b1–b2, June 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-19-issue-3-cover-and-back-matter/405D0A3F4160983ABDAA829985C0F986.

Robinson:1976:PPQ

[36] Derek W. Robinson. Properties of propagation of quantum spin systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):387–399, December 1976. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/properties-of-propagation-of-quantum-spin-systems/
AE9D7982D7D1316B8D2C51B045B672E6.

Knowles:1976:FAP

[37] James K. Knowles. On finite anti-plane shear for imcompressible elastic materials. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):400–415, December 1976. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
finite-antiplane-shear-for-imcompressible-elastic-materials/
22C6D106458573E66B0EAE509E491882.

Drummond:1976:SCT

[38] J. E. Drummond. Summing a common type of slowly convergent series of positive terms. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):416–421, December 1976. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/summing-a-common-type-
of-slowly-convergent-series-of-positive-terms/49CA10C2EE7B9FF8E64A095BB123BE90.

Sloan:1976:CDK

[39] Ian H. Sloan. Convergence of degenerate-kernel methods. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):422–431, December 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/convergence-of-degeneratekernel-methods/EA35AB6EB596F6E8B84E045E2ECFDFA1.

Webb:1976:SSS

[40] G. M. Webb. Similarity solutions of the steady state cosmic-ray equation of transport. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):432–451, December 1976. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/similarity-solutions-of-the-
steady-state-cosmicray-equation-of-transport/53AACB9657AD68C5796C25949E74B454.

Pao:1976:BTE

[41] C. V. Pao. A Boltzmann-type equation in the kinetic theory of vehicular traffic. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):452–461, December 1976. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/boltzmanntype-equation-in-the-kinetic-
theory-of-vehicular-traffic/D40D30CC4BC7D5C6492F1361DFB15023.

Craven:1976:SFJ

[42] B. D. Craven and B. Mond. Sufficient Fritz John optimality conditions for nondifferentiable convex programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4): 462–468, December 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/sufficient-fritz-john-optimality-conditions-for-nondifferentiable-
convex-programming/86E6EFBD5F7D59D79ADF839367C23D08.

Yeo:1976:DGR

[43] Geoffrey Yeo. A dam with general release rule. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):469–477, December 1976. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/dam-
with-general-release-rule/BF37691FF4B9348B4DAE1F5FC3AFB622.

Blatt:1976:CST

[44] John M. Blatt. Control systems with time lags. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19 (4):478–492, December 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/control-systems-with-time-lags/912C3B1E692C34BC46D9AD6927D197BB.

Barton:1976:ASE

[45] N. G. Barton. On the asymptotic solution of an elliptic interior layer problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):493–512, December 1976. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-asymptotic-solution-
of-an-elliptic-interior-layer-problem/A8FE65DEAA36CE7E022ACC1A99888C5B.

Anonymous:1976:C

[46] Anonymous. Contents. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):513–514, December 1976. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/contents/CD7E7EE1B79078EF1176C753145187C8.

Anonymous:1976:AVIc

[47] Anonymous. ANZ volume 19 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):f1–f2, December 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-19-issue-4-cover-and-front-matter/E453E25258C096D5492BF1E9DB85A8E2.

Anonymous:1976:AVId

[48] Anonymous. ANZ volume 19 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 19(4):b1–b2, December 1976. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-19-issue-4-cover-and-back-matter/1572331AAEA81331F3F3A8934648458B.

Knowles:1977:NAP

[49] James K. Knowles. A note on anti-plane shear for compressible materials in finite elastostatics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):1–7, June 1977. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/note-on-antiplane-
shear-for-compressible-materials-in-finite-elastostatics/7BD1D60789D3E4EBDD86475F4298D7BF.

Kloeden:1977:NSS

[50] P. E. Kloeden and R. A. Pearson. The numerical solution of stochastic differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):8–12, June 1977. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/numerical-solution-of-
stochastic-differential-equations/124E9E04D8CCB636687BDD6B0D8DC2D7.

Grimshaw:1977:SCS

[51] R. Grimshaw. The stability of continental shelf waves I. Side band instability and long wave resonance. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20 (1):13–30, June 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stability-of-continental-shelf-waves-i-side-band-instability-
and-long-wave-resonance/B8406AFEDD71F28F6F6004A2B62BA6AC.

Noussair:1977:EPC

[52] E. S. Noussair. On the existence of piecewise continuous optimal controls. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):31–37, June 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-the-existence-of-piecewise-continuous-optimal-controls/
F922E6CE4E0114EB3809D54118068F41.

Smrz:1977:TGT

[53] P. K. Smrz. Translations as gauge transformations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):38–45, June 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/translations-as-gauge-transformations/F7DABD9C70EAFCB9209230C15A2C2CAC.

Macaskill:1977:EAI

[54] C. Macaskill and E. O. Tuck. Evaluation of the acoustic impedance of a screen. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):46–61, June 1977. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/evaluation-of-the-acoustic-
impedance-of-a-screen/DEF743AB7A552B9938361B60154BE5DF.

Chadwick:1977:DRC

[55] P. Chadwick, C. F. M. Creasy, and V. G. Hart. The deformation of rubber cylinders and tubes by rotation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):62–96, June 1977. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/deformation-of-rubber-
cylinders-and-tubes-by-rotation/63C3F93A7C71F1C6F404E05A74E1A6D2.

Leach:1977:DMD

[56] P. G. L. Leach. On a direct method for the determination of an exact invariant for the time-dependent harmonic oscillator. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):97–105, June 1977. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-a-direct-method-for-the-
determination-of-an-exact-invariant-for-the-timedependent-harmonic-
oscillator/4B7038E8ED5D1668CA0B68200C65244A.

Hayes:1977:EFP

[57] M. Hayes. Energy flux for plane waves in linear conservative systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):106–113, June 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/energy-flux-for-plane-waves-in-linear-conservative-
systems/96FB3C5377F18D722A7B438F1B4F1A1A.

Szekeres:1977:HNC

[58] Peter Szekeres and John R. Rankin. Homogeneous Newtonian cosmologies and their perturbations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):114–128, June 1977. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/homogeneous-newtonian-
cosmologies-and-their-perturbations/7FC2062EF3336C4219111DF21EE0D109.

Anonymous:1977:AVIa

[59] Anonymous. ANZ volume 20 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):f1–f2, June 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-20-issue-1-cover-and-front-matter/524A6C35165B3D696A1AEB4C8E63B210.

Anonymous:1977:AVIb

[60] Anonymous. ANZ volume 20 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(1):b1–b2, June 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-20-issue-1-cover-and-back-matter/8946A2F14AC7B7ABEB8C9668E994D131.

Hill:1977:GSD

[61] James M. Hill. Generalized shear deformations for isotropic incompressible hyperelastic materials. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):129–141, December 1977. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/generalized-shear-
deformations-for-isotropic-incompressible-hyperelastic-materials/
D5A8D4194AAC920924790017A8C94D9A.

Blatt:1977:EDP

[62] J. M. Blatt and J. D. Gray. An elementary derivation of Pontrayagin’s maximum principle of optimal control theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):142–156, December 1977. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
elementary-derivation-of-pontrayagins-maximum-principle-of-
optimal-control-theory/6FA53C7CB078C1FB2D200768F3B2D4FB.

Abbott:1977:NCM

[63] James P. Abbott and Richard P. Brent. A note on continuation methods for the solution of nonlinear equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):157–164, December 1977. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
continuation-methods-for-the-solution-of-nonlinear-equations/
31599904BC03B0BFD4086B073B79D642.

Webb:1977:MSS

[64] G. M. Webb. A monoenergetic-source solution of the steadystate cosmic-ray equation of transport. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):165–

174, December 1977.  CODEN JAMMDU.  ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/monoenergeticsource-solution-of-the-steadystate-cosmicray-
equation-of-transport/076C7E54E4A433A52D9290BE5B85D586.

Nakai:1977:GFC

[65] Mitsuru Nakai and Leo Sario. Green’s function of the clamped punctured disk. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):175–181, December 1977. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/greens-function-of-the-clamped-punctured-
disk/8B070AF4D88963A44E5DBB685C7BA656.

Corbett:1977:WFU

[66] J. V. Corbett and C. A. Hurst. Are wave functions uniquely determined by their position and momentum distributions? Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):182–201, December 1977. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/are-
wave-functions-uniquely-determined-by-their-position-and-momentum-
distributions/228E4A34D0B3C63C54B1A01006278C42.

Brisley:1977:TRE

[67] W. Brisley. Two related estimation problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):202–210, December 1977. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/two-
related-estimation-problems/50F318CC7985084BCD44AB32E01EF595.

Collings:1977:SFI

[68] I. L. Collings and R. Grimshaw. Supercritical flow of an ideal fluid over a spillway. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):211–225, December 1977. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/supercritical-flow-of-
an-ideal-fluid-over-a-spillway/27D9C87F09E8BE436CDA9905F275F691.

Shepherd:1977:NSP

[69] J. J. Shepherd. A nonlinear singular perturbation problem on a semiinfinite interval. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):226–240, December 1977. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/nonlinear-singular-perturbation-
problem-on-a-semiinfinite-interval/F19ABD6E742B4B0765B07F9C23C71E16.

Philip:1977:IPL

[70] J. R. Philip. Inverse power law potentials about polygonal prisms and in polygonal cavities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):241–253, December 1977. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/inverse-power-law-
potentials-about-polygonal-prisms-and-in-polygonal-cavities/
164AFE8107FFA1CCA045F890C7B61BDF.

Abbott:1977:CFL

[71] James P. Abbott and Richard P. Brent. [Correction:] Fast local convergence with single and multistep methods for nonlinear equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):254, December 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fast-local-convergence-with-single-and-multistep-methods-
for-nonlinear-equations/B725437DCD6E2A41074C6AFDE201FB12. See

[17].

Anonymous:1977:AVIc

[72] Anonymous. ANZ volume 20 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):f1–f2, December 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-20-issue-2-cover-and-front-matter/3B2BAA508B30A223586B1654080F03C7.

Anonymous:1977:AVId

[73] Anonymous. ANZ volume 20 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(2):b1–b2, December 1977. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-20-issue-2-cover-and-back-matter/2ED2591F810D811B8EE43815FC630CE9.

Scott:1978:DAO

[74] C. H. Scott and T. R. Jefferson. A duality approach ot discrete time control theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):257–264, June 1978. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/duality-
approach-ot-discrete-time-control-theory/A9319FA75B033E7F69AEB967899B6EA8.

Barton:1978:IDS

[75] N. G. Barton. The initial dispersion of soluble matter in threedimensional flow. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):265–279, June 1978. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/initial-dispersion-of-
soluble-matter-in-threedimensional-flow/E3414EBAF8A83CFC7F55AB3EE769D7EA.

Diamond:1978:DLF

[76] Phil Diamond. Discrete Liapunov functions with Δ^{2}V > 0. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):280–284, June 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/discrete-liapunov-functions-with-2v0/13CBD1ED308ACC3F437971613FAE480D.

Anderson:1978:MMP

[77] N. Anderson and A. M. Arthurs. Maximum and minimum principles for capillary surface problems with prescribed contact angle. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):285–289, June 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/maximum-and-minimum-principles-for-capillary-surface-
problems-with-prescribed-contact-angle/E96239B0A0B34417AA34ADC2DA22BA02.

Gould:1978:TOC

[78] M. D. Gould. On tensor operators and characteristic identities for semi-simple Lie algebras. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):290–314, June 1978. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
tensor-operators-and-characteristic-identities-for-semisimple-
lie-algebras/09EAF891A8D41FE95AE8640A450DB1BE.

Murray:1978:CAC

[79] J. M. Murray and K. L. Teo. On a computational algorithm for a class of optimal control problems involving discrete time delayed arguments. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):315–343, June 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-a-computational-algorithm-for-a-class-of-optimal-
control-problems-involving-discrete-time-delayed-arguments/
B19BA3BF1A97DB0444A1B31D4E7CF9C1.

Patel:1978:CAS

[80] M. D. Patel. A class of axially symmetric stationary exact solutions of Einstein’s vacuum field equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20 (3):344–351, June 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/class-of-axially-symmetric-stationary-exact-solutions-
of-einsteins-vacuum-field-equations/370BC297F432D714E5B9F4514630B889.

Jittorntrum:1978:TAE

[81] Krisorn Jittorntrum and M. R. Osborne. Trajectory analysis and extrapolation in barrier function methods. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20 (3):352–369, June 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/trajectory-analysis-and-extrapolation-in-barrier-function-
methods/FF444EB5B9E6BF1863E8578E7D7DEB5A.

Fisher:1978:ABC

[82] M. E. Fisher and P. E. Kloeden. Asymptotic behaviour of a class of discontinuous difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20 (3):370–374, June 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-behaviour-of-a-class-of-discontinuous-difference-
equations/8AAD698BAEB1F49ACA67B833498770E1.

Hines:1978:ODR

[83] D. F. Hines and C. J. Thompson. A one-dimensional random walk with repulsion. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):375–380, June 1978. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/

journals/anziam-journal/article/onedimensional-random-walk-
with-repulsion/75396BBE77113C8997FD7FDCAE616C63.

Love:1978:TCT

[84] E. R. Love and D. L. Clements. A transformation of Cooke’s treatment of some triple integral equations +. Journal of the Australian Mathematical Society: Series B, Applied Mathematics,

20(3):381–383, June 1978.  CODEN JAMMDU.  ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/transformation-of-cookes-treatment-of-some-
triple-integral-equations/3F7A41FAA9E43FCE2FA1EFF223DAC026.

Anonymous:1978:AVIa

[85] Anonymous. ANZ volume 20 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):f1–f2, June 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-20-issue-3-cover-and-front-matter/9677B33500A25330C10F0A16B521FA0F.

Anonymous:1978:AVIb

[86] Anonymous. ANZ volume 20 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(3):b1–b2, June 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-20-issue-3-cover-and-back-matter/1AF3EFF9273D904E9E58EE70ECFADADD.

Buchen:1978:EGT

[87] Peter W. Buchen. The elastodynamic Green’s tensor for the 2D halfspace. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4):385–400, December 1978. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/elastodynamic-greens-tensor-for-the-
2d-halfspace/64B71BF348A86A32C31A7C8BBEB684A5.

Gould:1978:CIR

[88] M. D. Gould. The characteristic identities and reduced matrix elements of the unitary and orthogonal groups. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4): 401–433, December 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/characteristic-identities-and-reduced-matrix-elements-
of-the-unitary-and-orthogonal-groups/0AFB8F0AADBC97D568EF52D28B3FA1D8.

Chan:1978:VCP

[89] W. L. Chan and S. K. Ng. Variational control problems for linear differential systems with Stieltjes boundary conditions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20 (4):434–445, December 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/variational-control-problems-for-linear-differential-
systems-with-stieltjes-boundary-conditions/F7D49B17E7CCC11FBB5BB7ED98F9089B.

Cant:1978:ASR

[90] A. Cant and C. A. Hurst. The algebraic structure of relativistic wave equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4):446–486, December 1978. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/algebraic-structure-of-relativistic-
wave-equations/998BAD8E12B3C37B403F9AFD92110264.

Jefferson:1978:FPR

[91] T. R. Jefferson and C. H. Scott. On finding a point in the relative interior of a polyhedral set and degeneracy in geometric programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4):487–494, December 1978. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/on-finding-a-point-in-
the-relative-interior-of-a-polyhedral-set-and-degeneracy-in-
geometric-programming/283DDF06CA24D4F571155225A60ED59F.

Buchwald:1978:DEW

[92] V. T. Buchwald. The diffraction of elastic waves by small cylindrical cavities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4):495–507, December 1978. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/diffraction-of-elastic-waves-by-small-
cylindrical-cavities/B26BC7ED3CA05596351902AD177E7259.

Philip:1978:IPL

[93] J. R. Philip. Inverse power law potentials about polygonal prisms and in polygonal cavities +. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4):508, December 1978. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/inverse-power-law-
potentials-about-polygonal-prisms-and-in-polygonal-cavities/
8330DE351AD04EC38BA73B7640267435.

Anonymous:1978:CV

[94] Anonymous. Contents of volume 20. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4):509–

510, December 1978.  CODEN JAMMDU.  ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/contents-of-volume-20/25DEC5049D25135DE42C73EB79D6D867.

Anonymous:1978:AVIc

[95] Anonymous. ANZ volume 20 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4):f1–f2, December 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-20-issue-4-cover-and-front-matter/E5AD8B14301A3376DC6F761C888E897F.

Anonymous:1978:AVId

[96] Anonymous. ANZ volume 20 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 20(4):b1–b2, December 1978. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-20-issue-4-cover-and-back-matter/767ED7EE4196ED6364884FAFDE5776FC.

Landman:1979:BSP

[97] K. A. Landman. Bifurcation and stability of periodic solutions from a zero eigenvalue. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):2–20, April 1979. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/bifurcation-and-stability-of-periodic-
solutions-from-a-zero-eigenvalue/83E27652AE139146A398F5BBE1E7BF70.

Teo:1979:EOC

[98] K. L. Teo. Existence of optimal controls for systems governed by second order linear parabolic partial delay-differential equations with first boundary conditions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):21–36, April 1979. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/existence-of-optimal-controls-
for-systems-governed-by-second-order-linear-parabolic-partial-
delaydifferential-equations-with-first-boundary-conditions/
B30D21FA3BE3F96F528EE6C72661672A.

Scott:1979:COC

[99] C. H. Scott and T. R. Jefferson. Characterizations of optimality for continuous convex mathematical programs. Part I. Linear constraints. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):37–44, April 1979. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/characterizations-of-optimality-for-
continuous-convex-mathematical-programs-part-i-linear-constraints/
988D27BDC951568178A6120773EA2FCC.

Forbes:1979:NSO

[100] L. K. Forbes. A note on the solution of the one-dimensional unsteady equations of arterial blood flow by the method of characteristics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):45–52, April 1979. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/note-on-the-solution-of-the-onedimensional-
unsteady-equations-of-arterial-blood-flow-by-the-method-of-
characteristics/6B48D374149F3FC2970B9BB1A31DC0CD.

Nanda:1979:TMS

[101] Sribatsa Nanda. A-topology for Minkowski space. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):53–64, April 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/atopology-for-minkowski-space/E08956A790E839C4CE108103DA430978.

Skalak:1979:USV

[102] F. M. Skalak and C. Y. Wang. On the unsteady squeezing of a viscous fluid from a tube. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):65–74, April 1979. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-unsteady-squeezing-of-a-viscous-fluid-from-a-tube/06BC682A4691C70951B1509D5863F0CF.

Arthurs:1979:MHC

[103] A. M. Arthurs and V. G. Hart. The method of the hypercircle for a class of nonlinear equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21 (1):75–83, April 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/method-of-the-hypercircle-for-a-class-of-nonlinear-
equations/911019F22720F6D24D6FB0287D073CB0.

Stokes:1979:SOU

[104] A. N. Stokes. Series operations using minimum storage. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):84–89, April 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/series-operations-using-minimum-storage/9B4F6FBEE1DD077E6647AEF7C0A76FE6.

Nababan:1979:ETO

[105] S. Nababan and E. S. Noussair. Existence theorems for optimal control of quasilinear parabolic partial differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):90–101, April 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-theorems-for-optimal-control-of-quasilinear-
parabolic-partial-differential-equations/560D1213569267E77D43196811D44C35.

Anderson:1979:CVP

[106] N. Anderson and A. M. Arthurs. Complementary variational principles for a class of nonlinear boundary value problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):102–112, April 1979. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/complementary-variational-principles-for-a-
class-of-nonlinear-boundary-value-problems/93DFCDDAEF4DA17AFD33D9D232862406.

Clements:1979:RRT

[107] David J. Clements. A representation result for two-input two-output decentralized control systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):113–127, April 1979. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/representation-result-
for-twoinput-twooutput-decentralized-control-systems/9A153671506820F0CDDE755EE2997363.

Anonymous:1979:AVIa

[108] Anonymous. ANZ volume 21 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):f1–f3, April 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-21-issue-1-cover-and-front-matter/3F84774A9A92D9A9B1D1E6D8BD854890.

Anonymous:1979:AVIb

[109] Anonymous. ANZ volume 21 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(1):b1–b2, April 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-21-issue-1-cover-and-back-matter/57FDA7C8C766592049D0C403EA7935BF.

Bhaskaran:1979:ODA

[110] Sita Bhaskaran and Franz J. M. Salzborn. Optimal diameter assignment for gas pipeline networks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(2):129–144, August 1979. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
diameter-assignment-for-gas-pipeline-networks/BDD468A0C365C3940B85492A3AF0F40B.

Evans:1979:AKE

[111] J. W. Evans. Aspects of the kinetic equations for a special one-dimensional system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(2):145–175, August 1979. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/aspects-
of-the-kinetic-equations-for-a-special-onedimensional-system/
C6C06D64ABF738A3E814612CE4DDC282.

Moore:1979:MAP

[112] J. B. Moore and G. Ledwich. Multivariable adaptive parameter and state estimators with convergence analysis. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(2):176–197, August 1979. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/multivariable-adaptive-parameter-and-state-
estimators-with-convergence-analysis/FE6E377791F58ED6C21E63DF2E55F567.

Golden:1979:PRP

[113] J. M. Golden. The problem of a rigid punch moving on a viscoelastic half-plane with inertial effects approximately included. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(2):198–229, August 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/problem-of-a-rigid-punch-moving-on-a-viscoelastic-halfplane-
with-inertial-effects-approximately-included/3EC1D4BB40AD7538274FBCC25F1EBAD5.

Evans:1979:NLS

[114] R. J. Evans and A. Cantoni. A note on least-squares approximation with range constraints. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(2):230–242, August 1979. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
leastsquares-approximation-with-range-constraints/BCD63EA0D3BB20FEF5669F09DB6F7371.

Clements:1979:TCP

[115] D. L. Clements and T. R. Tauchert. A thermoelastic crack problem for an anisotropic slab. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(2):243–255, August 1979. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/thermoelastic-crack-problem-
for-an-anisotropic-slab/B35AB7D44566FAD63067D5D69E89433F.

Anonymous:1979:AVIc

[116] Anonymous. ANZ volume 21 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(2):f1–f2, August 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-21-issue-2-cover-and-front-matter/018722C8CAE959D2729CCCFC458C63F9.

Anonymous:1979:AVId

[117] Anonymous. ANZ volume 21 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(2):b1–b2, August 1979. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-21-issue-2-cover-and-back-matter/1774ED61A66025FC6220774441EB16ED.

Loxton:1980:KRA

[118] J. H. Loxton and J. W. Sanders. The kernel of a rule of approximate integration. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):257–267, January 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/kernel-of-a-rule-of-approximate-integration/
7DFC2523A6E0A15DBE2F411262605F42.

Fill:1980:IPS

[119] James A. Fill and Alan J. Izenman. Invariance properties of Schoenberg’s tone row system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):268–282, January 1980. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/invariance-properties-
of-schoenbergs-tone-row-system/E2F7E71E4E1B59440E92B1C79B59E043.

Simms:1980:SPS

[120] P. D. Simms. On the solution of a partially structured nonlinear program. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):283–292, January 1980. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-

journal/article/on-the-solution-of-a-partially-structured-nonlinear-
program/7FC2193898354C7996A6D5649F85F776.

Love:1980:SHL

[121] J. D. Love. Solution of homogeneous linear difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):293–296, January 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/solution-of-homogeneous-linear-difference-equations/
56F94F3D10BC4275586A879CBD0D12FE.

Rosenblat:1980:AB

[122] S. Rosenblat. Asymmetric bifurcation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21 (3):297–304, January 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymmetric-bifurcation/FE1FBDB559BE753B3A0ADA7143028399.

Jittorntrum:1980:MBF

[123] Krisorn Jittorntrum and M. R. Osborne. A modified barrier function method with improved rate of convergence for degenerate problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):305–329, January 1980. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/modified-barrier-function-method-with-improved-
rate-of-convergence-for-degenerate-problems/F11B25ABDACA4DF551D23454483B82A3.

Arthurs:1980:HEN

[124] A. M. Arthurs. Hypercircle estimates for nonlinear problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):330–337, January 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/hypercircle-estimates-for-nonlinear-
problems/67EF5F34FFB1EE8AB72235A7FC987217.

Buchdahl:1980:SCL

[125] H. A. Buchdahl. Scale covariant Lagrangians and spaces reciprocal to static Einstein spaces. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):338–344, January 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/scale-covariant-lagrangians-
and-spaces-reciprocal-to-static-einstein-spaces/C1E9DD368FE5286F7ABD9271A51A7EEE.

Clark:1980:ASR

[126] David Clark. An algorithm for solving the restricted least squares problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):345–356, January 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-algorithm-for-solving-the-restricted-
least-squares-problem/18C1DF8E6F5030CF75AD6E8868B5046B.

Nanda:1980:NMM

[127] Sribatsa Nanda. A note on maximal and minimal causal spaces. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):357–364, January 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-maximal-and-minimal-causal-spaces/253F0A5409A4B0BBDCB8DE1471AE2039.

Paine:1980:UEE

[128] John Paine and Frank de Hoog. Uniform estimation of the eigenvalues of Sturm–Liouville problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):365–383, January 1980. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/uniform-estimation-of-
the-eigenvalues-of-sturmliouville-problems/F8D5EE89083432AE643639F159A23086.

Buchwald:1980:DEW

[129] V. T. Buchwald. The diffraction of elastic waves by small cylindrical cavities +. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):384, January 1980. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/diffraction-of-elastic-waves-by-small-cylindrical-
cavities/DFADD56665E55DDBD8E94A20B8C5BBC3.

Anonymous:1980:AVIa

[130] Anonymous. ANZ volume 21 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):f1–f2, January 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-21-issue-3-cover-and-front-matter/22444B84F9DCF2485CFDA1049304BBF6.

Anonymous:1980:AVIb

[131] Anonymous. ANZ volume 21 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(3):b1–b2, January 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-21-issue-3-cover-and-back-matter/883BE77FA9FA864B5F65DDEE26F718A2.

Teo:1980:CAT

[132] K. L. Teo and B. D. Craven. On a computational algorithm for timelag optimal control problems with restricted phase coordinates. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):385–397, April 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-a-computational-algorithm-for-timelag-optimal-control-
problems-with-restricted-phase-coordinates/C9B67BA530E7F650CB53D0CFA5E0D6F0.

Scott:1980:FPD

[133] C. H. Scott and T. R. Jefferson. Fractional programming duality via geometric programming duality. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21 (4):398–401, April 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fractional-programming-duality-via-geometric-programming-
duality/89AB5CF4DEB1A5EA9403E22F838A20A5.

Fill:1980:SRI

[134] James A. Fill and Alan J. Izenman. The structure of RI-invariant twelvetone rows. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):402–417, April 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/structure-of-riinvariant-twelevetone-
rows/3958B61CF24EFF2F68A3FE5C0F8FC6CE.

Hathaway:1980:SBW

[135] Neville J. Hathaway, Ian W. Murray, and Johan A. Rickard. The stability of Bishop’s warfare strategy. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21 (4):418–432, April 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stability-of-bishops-warfare-strategy/B6C1E749E9DE542635F2643C0B81C202.

Hennagin:1980:ERE

[136] Stephen C. Hennagin and Peter Linz. Exclusion regions for eigenvalues of linear operators. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):433–439, April 1980. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/exclusion-regions-for-
eigenvalues-of-linear-operators/B420831B0F93C93FCA42E14C25042E14.

Tuck:1980:DSB

[137] E. O. Tuck. Doubling strategies for backgammon-like games. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):440–451, April 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/doubling-strategies-for-backgammonlike-games/2CF4D8E3EC0471A1B08CD76F2C0017FD.

Lewis:1980:SNL

[138] R. M. Lewis. Stability of non-linear systems with time-varying delays. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):452–463, April 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/stability-of-nonlinear-systems-with-
timevarying-delays/932FA768F2C6D0D7D78E9F55E90365A3.

Akabari:1980:SPR

[139] R. P. Akabari, U. K. Dave, and L. K. Patel. Some pure radiation filds in general relativity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):464–473, April 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/some-pure-radiation-filds-in-general-
relativity/D2D2408398E8F783014FF15AE0D33C4F.

Robinson:1980:PPS

[140] Derek W. Robinson. Propagation properties in scattering theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):474–485, April 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/propagation-properties-in-scattering-theory/A66CCF26AD8FDD143B079FEA095B34C7.

Beecham:1980:SPS

[141] A. F. Beecham and A. C. Hurley. A scheduling problem with a simple graphical solution. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):486–495, April 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/scheduling-problem-with-a-simple-
graphical-solution/6FA421C73BB7BA64B02C2C85D874C935.

Kloeden:1980:SLS

[142] P. E. Kloeden. Stability of large-scale systems with lagged interconnections. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):496–509, April 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/stability-of-largescale-systems-with-
lagged-interconnections/3DB79489E6286C8FD67FB8653DF3C4CF.

Anonymous:1980:CVX

[143] Anonymous. Contents of volume XXI. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21 (4):510–511, April 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/contents-of-volume-xxi/F9BADA6AE7F51561E077C07343436F18.

Anonymous:1980:AVXa

[144] Anonymous. ANZ volume XXI issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):f1–f2, April 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-xxi-issue-4-cover-and-front-matter/DB53245A7203857F8C32C6B880CBC0D3.

Anonymous:1980:AVXb

[145] Anonymous. ANZ volume XXI issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 21(4):b1–b2, April 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-xxi-issue-4-cover-and-back-matter/22901E43AD56B1A583280AAFBE67D017.

Clark:1980:ISS

[146] D. I. Clark and M. R. Osborne. On the implementation of a subset selection algorithm for the restricted least squares problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1): 2–11, July 1980. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-implementation-of-a-subset-selection-algorithm-for-the-
restricted-least-squares-problem/F7BBC01C6E6017EB1471F4ECED8B33A5.

Leach:1980:CSG

[147] P. G. L. Leach. The complete symmetry group of a forced harmonic oscillator. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):12–21, July 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/complete-symmetry-group-of-a-forced-
harmonic-oscillator/28D8D6FE5C971E0E2D8ADD7F6E1F6123.

Phan-Thien:1980:SSO

[148] N. Phan-Thien. Small strain oscillatory squeeze film flow of simple fluids. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):22–27, July 1980. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/small-strain-oscillatory-squeeze-film-flow-
of-simple-fluids/F7A6BE04587DFDC39D850B8AAE836EC2.

Dolan:1980:CMG

[149] P. Dolan, P. Choudhury, and J. L. Safko. A ‘constant of the motion’ for the geodesic deviation equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):28–33, July 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/constant-of-the-motion-for-
the-geodesic-deviation-equation/B418DDA433A8E962DA6F5574B05FD8B8.

Casling:1980:TIA

[150] E. M. Casling. A theoretical investigation of an asymmetric planing hull at infinite Froude number. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22 (1):34–52, July 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/theoretical-investigation-of-an-asymmetric-planing-
hull-at-infinite-froude-number/37D613A699088D93E7710561E324C138.

Singh:1980:SGHa

[151] V. N. Singh. Solution of a general homogeneous linear difference equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):53–57, July 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/solution-of-a-general-
homogeneous-linear-difference-equation/FD508CCAFE731B332B4C8177F8A67DE4.

Hill:1980:DRW

[152] James M. Hill. A discrete random walk model for diffusion in media with double diffusivity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):58–74, July 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/discrete-random-walk-model-
for-diffusion-in-media-with-double-diffusivity/4513172C80D7F1A5B068AF8DF4E36744.

Shepherd:1980:ODG

[153] J. J. Shepherd. The one-dimensional gas-lubricated slider bearing. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):75–97, July 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/onedimensional-gaslubricated-slider-
bearing/B8ECED35A961A16728C9C11BB1B8AB51.

Hart:1980:EST

[154] V. G. Hart. Exact solutions of two nonlinear equations and hypercircle estimates. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):98–103, July 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/exact-solutions-of-two-nonlinear-equations-
and-hypercircle-estimates/EDA961638E274785C9550D3620E2A2E8.

Tuck:1980:ESB

[155] E. O. Tuck. The effect of a submerged barrier on the natural freqencies and radiation damping of a shallow basin connected to open water. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):104–128, July 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/effect-of-a-submerged-
barrier-on-the-natural-freqencies-and-radiation-damping-of-
a-shallow-basin-connected-to-open-water/D5BE70A31BD120F545A7212561C7CD83.

Anonymous:1980:AVIc

[156] Anonymous. ANZ volume 22 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):f1–f3, July 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-22-issue-1-cover-and-front-matter/AA6632AB6312386305699313BC7C1FCD.

Anonymous:1980:AVId

[157] Anonymous. ANZ volume 22 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(1):b1–b2, July 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-22-issue-1-cover-and-back-matter/727DA318DE66F6249A5B4173E9B4B96A.

Bratteli:1980:ESB

[158] Ola Bratteli and Derek W. Robinson. Equilibrium states of a Bose gas with repulsive interactions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):129–147, October 1980. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/equilibrium-states-of-
a-bose-gas-with-repulsive-interactions/067F5EFEDFDFEDC7455CE722BAD88BB0.

Gopalsamy:1980:LCT

[159] K. Gopalsamy and B. D. Aggarwala. Limit cycles in two spacies competition with time delays. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):148–160, October 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/limit-cycles-in-two-spacies-
competition-with-time-delays/333EAC34FC7561961629AEDEB510CE7D.

Riahi:1980:NHC

[160] N. Riahi. On a nonlinear hydromagnetic convection under a rotational constraint. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):161–174, October 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-a-nonlinear-hydromagnetic-convection-
under-a-rotational-constraint/99557E3D3C740E2B662029F71FB37B2D.

Keats:1980:OPC

[161] R. G. Keats and Joan Cooper. The optimum processing of clipped signals: an approach based on minimum signal distortion. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):175–184, October 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/optimum-processing-of-
clipped-signals-an-approach-based-on-minimum-signal-distortion/
EEA7BF5FFF326644421F14D0B20FD21A.

Nieuwenhuis:1980:ALA

[162] J. W. Nieuwenhuis. About a local approximation theorem and an inverse function theorem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):185–192, October 1980. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/about-a-
local-approximation-theorem-and-an-inverse-function-theorem/
2AA5C7AF32ED5ACF5B952334410C8837.

Tam:1980:IID

[163] K. K. Tam. On the influence of the initial data in a combustion problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):193–209, October 1980. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-the-influence-of-the-initial-data-in-a-combustion-
problem/40C213CCF79CB38A80E98F395575B277.

Biswas:1980:CSP

[164] S. N. Biswas and T. S. Santhanam. Coherent states of para-Bose oscillators. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):210–217, October 1980. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/coherent-states-of-parabose-oscillators/
892F1160B92E6ECBDC1D034BA3A13BBE.

Clements:1980:BIE

[165] D. L. Clements. A boundary integral equation method for the numerical solution of a second order elliptic equation with variable coefficients. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):218–228, October 1980. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/boundary-integral-equation-
method-for-the-numerical-solution-of-a-second-order-elliptic-
equation-with-variable-coefficients/89D6F88443D18FE55D982993E8B26B3F.

Jefferson:1980:GPP

[166] T. R. Jefferson and C. H. Scott. Geometric programming with probabilistic decision variables. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):229–236, October 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/geometric-programming-with-
probabilistic-decision-variables/077A720001525B05768982DC316EC824.

Do:1980:PIE

[167] D. D. Do and R. H. Weiland. Poisoning-induced exchange of steady states in a catalytic chemical reactor. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22 (2):237–253, October 1980. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/poisoninginduced-exchange-of-steady-states-
in-a-catalytic-chemical-reactor/0FA6BA27CC913A21DA4041577D6D0820.

Singh:1980:SGHb

[168] V. N. Singh. Solution of a general homogeneous linear difference equation +. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):254, October 1980. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/solution-of-a-general-
homogeneous-linear-difference-equation/7E7DE67898C684B46FFED07D16C5A77D.

Anonymous:1980:AVIe

[169] Anonymous. ANZ volume 22 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):f1–f2, October 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-22-issue-2-cover-and-front-matter/A70146E7F0169630B64F9793D180F79A.

Anonymous:1980:AVIf

[170] Anonymous. ANZ volume 22 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(2):b1–b2, October 1980. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-22-issue-2-cover-and-back-matter/7F45F04FDEB9652C4F058899A3DF20B0.

Forbes:1981:ESW

[171] L. K. Forbes. On the evolution of shock-waves in mathematical models of the aorta. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):257–269, January 1981. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-evolution-of-shockwaves-in-mathematical-models-of-the-aorta/
2EC07C8656BE9EED9A80A3D45F390E45.

Kulkarani:1981:EBS

[172] R. P. Kulkarani and B. V. Limaye. On error bounds in strong approximations for eigenvalue problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):270–283, January 1981. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
error-bounds-in-strong-approximations-for-eigenvalue-problems/
17DAD81CEB9B9DFECC82906DAF85343B.

Agarwal:1981:FPR

[173] N. K. Agarwal and Sunil Datta. Flow in the porous region between slowly rotating prolate spheroids. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):284–290, January 1981. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/flow-in-
the-porous-region-between-slowly-rotating-prolate-spheroids/
A075DE2F54B81045F9E048C48CF99728.

Gates:1981:EOEa

[174] D. J. Gates and M. Westcott. Extended optima and equilibria for continuous games. I. General results. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22 (3):291–307, January 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/extended-optima-and-equilibria-for-continuous-games-
i-general-results/E9441AAFC8F683FE0A6EC009E44F48B2.

Nieuwenhuis:1981:SRS

[175] J. W. Nieuwenhuis. Some remarks on set-valued dynamical systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):308–313, January 1981. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/some-remarks-on-setvalued-dynamical-
systems/F1566902691460AA4666D5D788C87097.

Rizvi:1981:SGL

[176] S. A. H. Rizvi. Solution of a general liner difference equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):314–317, January 1981. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/solution-of-a-general-
liner-difference-equation/0589DA6B316FFDCB1EF43A67B5279A6D.

Grimshaw:1981:MFG

[177] R. Grimshaw. Mean flows generated by a progressing water wave packet. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):318–347, January 1981. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/mean-flows-generated-by-a-progressing-water-
wave-packert/360A7569F064BC9EA4B9F3006751BCE3.

McNabb:1981:ARC

[178] A. McNabb. The asymptotic response of a calorimeter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):348–352, January 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-response-of-a-calorimeter/A02B26811B80F07EA8D2BE0EC282EA38.

England:1981:LIO

[179] A. H. England. Love’s integral and other relations between solutions to mixed boundary-value problems in potential theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):353–367, January 1981. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/loves-integral-and-other-relations-between-
solutions-to-mixed-boundaryvalue-problems-in-potential-theory/
1C392E857208A099FECE6B686A7B6499.

Stuart:1981:NCI

[180] S. N. Stuart. Non-classical integrals of Bessel functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):368–378, January 1981. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/nonclassical-integrals-of-bessel-functions/
3DF8C2C82C3F6C46560F15B156793DEF.

Anonymous:1981:AVIa

[181] Anonymous. ANZ volume 22 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):f1–f2, January 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-22-issue-3-cover-and-front-matter/5125E384B3468FFAA33A6DB0A72E3298.

Anonymous:1981:AVIb

[182] Anonymous. ANZ volume 22 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(3):b1–b2, January 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-22-issue-3-cover-and-back-matter/32EE0BB1AF5321B83E1692FF1DCDC276.

Rizzo:1981:BIE

[183] F. J. Rizzo and D. J. Shippy. The boundary integral equation method with application to certain stress concentration problems in elasticity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):381–393, April 1981. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/boundary-integral-equation-
method-with-application-to-certain-stress-concentration-problems-
in-elasticity/47128EE167DBF5A572C964E08B91C5A0.

Clements:1981:BIE

[184] David L. Clements and Oscar A. C. Jones. The boundary integral equation method for the solution of a class of problems in anisotropic elasticity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):394–407, April 1981. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/boundary-integral-
equation-method-for-the-solution-of-a-class-of-problems-in-
anisotropic-elasticity/85DABD6C6F22C3C98850B86E63D9BF71.

Anselone:1981:SSN

[185] P. M. Anselone. Singularity subtraction in the numerical solution of integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):408–418, April 1981. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/singularity-subtraction-in-
the-numerical-solution-of-integral-equations/C42E161C2F215CFC2278C29E8BA88C9D.

Vainikko:1981:PSW

[186] G. Vainikko and A. Pedas. The properties of solutions of weakly singular integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):419–430, April 1981. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/properties-of-solutions-of-
weakly-singular-integral-equations/F4ACE59ED7546E32414C867376CE9B97.

Vainikko:1981:PPA

[187] G. Vainikko and P. Uba. A piecewise polynomial approximation to the solution of an integral equation with weakly singular kernel. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):431–438, April 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/piecewise-polynomial-approximation-to-the-solution-
of-an-integral-equation-with-weakly-singular-kernel/E3A3FEFB8EE4FFE86A587B358AA23B22.

Chatelin:1981:IPS

[188] Françoise Chatelin and Rachid Lebbar. The iterated projection solution for the Fredholm integral equation of second kind. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):439–451, April 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/iterated-projection-solution-for-the-fredholm-integral-
equation-of-second-kind/03829C11C7B5718E0AAEA5F6667E69AA.

Qun:1981:DCE

[189] Lin Qun. Deferred corrections for equations of the second kind. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):452–455, April 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/deferred-corrections-for-equations-of-the-second-kind/
512CA5D3FA13E014A967CE397BAF651C.

Graham:1981:CMT

[190] Ivan G. Graham. Collocation methods for two dimensional weakly singular integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):456–473, April 1981. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/collocation-methods-for-two-
dimensional-weakly-singular-integral-equations/153FBE6AC2757196B8DEA771E0688847.

Chu:1981:DCI

[191] K. W. Chu and A. Spence. Deferred correction for the integral equation eigenvalue problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):474–487, April 1981. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/deferred-correction-for-the-
integral-equation-eigenvalue-problem/0373BBC1B7D323A0651A200A27F0FDDE.

Anderssen:1981:FCM

[192] R. S. Anderssen and P. M. Prenter. A formal comparison of methods proposed for the numerical solution of first kind integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):488–500, April 1981. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/formal-comparison-
of-methods-proposed-for-the-numerical-solution-of-first-kind-
integral-equations/13BA34FED784A5C09AD840B55BE370BF.

OBrien:1981:EGC

[193] D. M. O’Brien and J. N. Holt. The extension of generalized crossvalidation to a multi-parameter class of estimators. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22 (4):501–514, April 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/extension-of-generalized-crossvalidation-to-a-multiparameter-
class-of-estimators/C3420D29FA122395B414D5ADEA0CA657.

Baker:1981:SAR

[194] Christopher T. H. Baker and Joan C. Wilkinson. Stability analysis of Runge–Kutta methods applied to a basic Volterra integral equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):515–538, April 1981. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/stability-analysis-of-
rungekutta-methods-applied-to-a-basic-volterra-integral-equation/
8505C2B72DCFA420DE9F5F8410F81D8E.

Elliott:1981:CTS

[195] David Elliott. A convergence theorem for singular integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):539–552, April 1981. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/convergence-theorem-for-singular-integral-equations/
9D990C4FAE8B37AD6C10BBA5248CCAB5.

Anonymous:1981:CV

[196] Anonymous. Contents of volume 22. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):553–554, April 1981. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
22/D1B4D066548F423A102128E35E066546.

Anonymous:1981:AVIc

[197] Anonymous. ANZ volume 22 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):f1–f3, April 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-22-issue-4-cover-and-front-matter/E51E209D375E86F915292CDB05CE5C93.

Anonymous:1981:AVId

[198] Anonymous. ANZ volume 22 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 22(4):b1–b2, April 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-22-issue-4-cover-and-back-matter/AA03FE859E5872115CC966BF3A20613B.

OBrien:1981:GMB

[199] R. W. O’Brien. The gliding motion of a bacterium: Flexibacter strain BH3. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):2–16, July 1981. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/gliding-motion-of-
a-bacterium-flexibacter-strain-bh3/A8370B14C53EF5B55CF65A9DAE7553C1.

Mahony:1981:SSOa

[200] J. J. Mahony and J. J. Shepherd. Stiff systems of ordinary differential equations. Part 1. Completely stiff, homogeneous systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):17–51, July 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stiff-systems-of-ordinary-differential-equations-part-
1-completely-stiff-homogeneous-systems/95A9BE38222837FB74F1733F3B63A782.

Kleeman:1981:RGM

[201] R. Kleeman. On representations of a generalized method of field quantization. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):52–63, July 1981. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-representations-of-a-generalized-
method-of-field-quantization/43A57234EF8813989E4DD33521D1B549.

Potts:1981:ESD

[202] Renfrey B. Potts. Exact solution of a difference approximation to Duffing’s equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):64–77, July 1981. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/exact-solution-of-
a-difference-approximation-to-duffings-equation/E87ED77D8AC3427640D19171F13F09F1.

Brown:1981:EPS

[203] A. Brown. Equations for periodic solutions of a logistic difference equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):78–94, July 1981. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/equations-for-periodic-solutions-of-
a-logistic-difference-equation/F1E41F534D91E463E0F5901B83F5C7B8.

Tam:1981:RHS

[204] K. K. Tam and M. T. Kiang. The response to a hot spot in a combustion problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):95–102, July 1981. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/response-to-a-hot-
spot-in-a-combustion-problem/248678BBEDF3B6918B413088047ADA51.

Caldwell:1981:EIT

[205] James Caldwell. Effects of iron on a toroidal conductor. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):103–114, July 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/effects-of-iron-on-a-toroidal-conductor/0BCEEDBB2154932070E2959DBCD11429.

Anonymous:1981:AVIe

[206] Anonymous. ANZ volume 23 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):f1–f3, July 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-23-issue-1-cover-and-front-matter/F8189DC0E096FC1028FAB6B6808C630C.

Anonymous:1981:AVIf

[207] Anonymous. ANZ volume 23 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(1):b1–b2, July 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-23-issue-1-cover-and-back-matter/5E463B4EB1ECE4331D8684A7E88D0F97.

Blatt:1981:OCT

[208] John M. Blatt. Optimal control theory with general constraints. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(2):115–126, October 1981. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/optimal-control-theory-with-general-
constraints/4F66C296F7C552409D7729806F0A5C12.

Anderssen:1981:NDT

[209] R. S. Anderssen and R. B. Calligaro. Non-destructive testing of opticalfibre preforms. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(2):127–135, October 1981. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/nondestructive-testing-of-opticalfibre-
preforms/B4B6119A6D0BDA68255E0D0DAB0F5B08.

Mahony:1981:SSOb

[210] J. J. Mahony and J. J. Shepherd. Stiff systems of ordinary differential equations: II. Boundary-value problems for completely stiff systems.

Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(2):136–172, October 1981. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/stiff-systems-of-ordinary-differential-equations/
555836EF64FBB47929FFA460A93235AC.

Leach:1981:ALT

[211] P. G. L. Leach. Applications of the Lie theory of extended groups in Hamiltonian mechanics: the oscillator and the Kepler problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(2):173–186, October 1981. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/applications-of-the-lie-theory-of-extended-
groups-in-hamiltonian-mechanics-the-oscillator-and-the-kepler-
problem/8D6E4EC87483207853F1F0765A5518A4.

Gates:1981:EOEb

[212] D. J. Gates and M. Westcott. Extended optima and equilibria for continuous games: II. A class of economic models. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23 (2):187–209, October 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/extended-optima-and-equilibria-for-continuous-games/
FA301D3493C62B9AE5663A3ECB1ADD74.

Gates:1981:EOEc

[213] D. J. Gates and M. Westcott. Extended optima and equilibria for continuous games: III. Comparison with bargaining experiments. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(2):210–227, October 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/extended-optima-and-equilibria-for-continuous-games/
1FAA37F9B4A5B87B9A0490DB4A3FEC5E.

Anonymous:1981:AVIg

[214] Anonymous. ANZ volume 23 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(2):f1–f2, October 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-23-issue-2-cover-and-front-matter/CC253A432878DE98D7C7729F86C0EEA2.

Anonymous:1981:AVIh

[215] Anonymous. ANZ volume 23 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(2):b1–b2, October 1981. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-23-issue-2-cover-and-back-matter/4DAC0F85162C3CAAD8D4C7E64B6C242F.

Barber:1982:ESU

[216] Michael N. Barber and C. J. Hamer. Extrapolation of sequences using a generalized epsilon-algorithm. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(3):229–240, January 1982. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/extrapolation-of-sequences-
using-a-generalized-epsilonalgorithm/F0C7DF562C6D8B78877D80ABD3FB3A3B.

Tuck:1982:LPSa

[217] E. O. Tuck. Linearized planing-surface theory with surface tension. Part I: Smooth detachment. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(3):241–258, January 1982. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/linearized-planingsurface-theory-
with-surface-tension-part-i-smooth-detachment/F773E81C8906C4CCDC937169AECB5628.

Tuck:1982:LPSb

[218] E. O. Tuck. Linearized planing-surface theory with surface tension. Part II: Detachment with discontinuous slope. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23 (3):259–277, January 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/linearized-planingsurface-theory-with-surface-tension-
part-ii-detachment-with-discontinuous-slope/B2247C78F6A8E6E0ECEB55687BD8DE87.

Do:1982:DRE

[219] D. D. Do and R. H. Weiland. Dynamics of rapid extinction in a lumped system with Arrhenius chemistry. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(3):278–290, January 1982. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/dynamics-of-rapid-
extinction-in-a-lumped-system-with-arrhenius-chemistry/51426E3D34024BD98A8FB8BFE5C0E25C.

Pearce:1982:GRF

[220] C. E. M. Pearce. A generalization of Rapp’s formula. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(3):291–296, January 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalization-of-rapps-formula/D1F4364AC9CB696CDCFC132D20C9CEDD.

Parlange:1982:TDS

[221] J.-Y. Parlange, R. D. Braddock, G. Sander, and F. Stagnitti. Three dimensional similarity solutions of the nonlinear diffusion equation from optimization and first integrals. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(3):297–309, January 1982. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/three-dimensional-
similarity-solutions-of-the-nonlinear-diffusion-equation-from-
optimization-and-first-integrals/DFAF0D3527E1BCFE5CF175766EC471BC.

Mahony:1982:SSO

[222] J. J. Mahony and J. J. Shepherd. Stiff systems of ordinary differential equations. III. Partially stiff systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23 (3):310–331, January 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stiff-systems-of-ordinary-differential-equations-iii-
partially-stiff-systems/9E1B889921F3F76EC1B4B62634998DDD.

Atkinson:1982:NIS

[223] Kendall Atkinson. Numerical integration on the sphere. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(3):332–347, January 1982. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/numerical-integration-on-the-sphere/
1E811EF7FC701F86F8F530C020F65AC8.

Anonymous:1982:AVIa

[224] Anonymous. ANZ volume 23 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(3):f1–f2, January 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-23-issue-3-cover-and-front-matter/AE763B3DE30EEAF9E57C4629A0734FBA.

Anonymous:1982:AVIb

[225] Anonymous. ANZ volume 23 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(3):b1–b2, January 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-23-issue-3-cover-and-back-matter/3EFAA3AD990E43437EB9C7A117579C60.

Potts:1982:BDE

[226] Renfrey B. Potts. Best difference equation approximation to Duffing’s equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):349–356, April 1982. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/best-difference-equation-approximation-
to-duffings-equation/8E7F7BCFD40CC5A5E143678ABAD7846F.

Howlett:1982:IRF

[227] P. G. Howlett. Input retrieval in finite dimensional linear systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):357–382, April 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/input-retrieval-in-finite-dimensional-linear-systems/
EB52227AB7FA624E135291183ADE1D13.

Blake:1982:NID

[228] J. R. Blake and P. Cerone. A note on the impulse due to a vapour bubble near a boundary. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):383–393, April 1982. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/note-on-the-impulse-due-to-a-vapour-
bubble-near-a-boundary/567E5541C575B0011B218ACD0F15DA97.

Vaisman:1982:QQM

[229] Izu Vaisman. On the quantization of quadratic momenta. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4): 394–402, April 1982. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-quantization-of-quadratic-momenta/D9D2561CDA27C7ACCED97EDB27565EDE.

Tuck:1982:ITS

[230] E. O. Tuck. An inviscid theory for sliding flexible sheets. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4): 403–415, April 1982. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
inviscid-theory-for-sliding-flexible-sheets/44657C3BFC1808E44AD52514797F2763.

Bracken:1982:TLS

[231] A. J. Bracken. A transformation to linearity of some non-linear differential-difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):416–419, April 1982.

CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/transformation-to-
linearity-of-some-nonlinear-differentialdifference-equations/
F4EE49EFB588E0AF82EFC93AC6674246.

Corbett:1982:CRZ

[232] J. V. Corbett. The commutation relation i[Y, Z] = 2Y and the absolutely continuous spectrum of Y. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23 (4):420–427, April 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/commutation-relation-iy-z-2y-and-the-absolutely-continuous-
spectrum-of-y/9B448601FE2A28B213A7C5D8DDC8F4ED.

Chan:1982:SLC

[233] W. L. Chan. Suboptimality in linear control systems with time delay. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):428–437, April 1982. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/suboptimality-in-linear-control-systems-with-
time-delay/3948943F3FFA192B785969588CC4AB86.

Cerone:1982:ABD

[234] P. Cerone and K. P. Tognetti. The asymptotic behaviour due to a piecewise time dependent net maternity function. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23 (4):438–450, April 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-behaviour-due-to-a-piecewise-time-dependent-
net-maternity-function/677EECC3693E180C1CE8F095C4D55A8C.

Anderson:1982:MSG

[235] D. H. Anderson, E. A. Catchpole, N. J. de Mestre, and T. Parkes. Modelling the spread of grass fires. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):451–466, April 1982. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/modelling-the-spread-of-
grass-fires/CDDEEB47686374AA03E15C9FBF825A38.

Anonymous:1982:CV

[236] Anonymous. Contents of volume 23. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):467–468, April 1982.

CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
23/8E8C3D7E0A3AC9EECA1C6C7236E7B319.

Anonymous:1982:AVIc

[237] Anonymous. ANZ volume 23 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):f1–f2, April 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-23-issue-4-cover-and-front-matter/A0AD77CC15CDB67422A33523A89C0158.

Anonymous:1982:AVId

[238] Anonymous. ANZ volume 23 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 23(4):b1–b2, April 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-23-issue-4-cover-and-back-matter/496BDF3B39032AF8B76109E076369B0B.

Bigg:1982:TDR

[239] G. R. Bigg and E. O. Tuck. Two-dimensional resonators with small openings. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):2–27, July 1982. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/twodimensional-resonators-with-small-openings/
9E583BF8323B533E1CA66D525AF34ABB.

Drummond:1982:ISA

[240] J. E. Drummond. On interpreting the sums of asymptotic series of positive terms. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):28–39, July 1982. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
interpreting-the-sums-of-asymptotic-series-of-positive-terms/
24D702C23138EB94463648E33D88D6F2.

Tam:1982:CCP

[241] K. K. Tam. Computation of critical parameters for a problem in combustion theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):40–46, July 1982. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/computation-of-critical-parameters-for-
a-problem-in-combustion-theory/91123051E1DFA5D47C8F129E18DCCDFC.

Fenton:1982:RCM

[242] J. D. Fenton and R. S. Gardiner-Garden. Rapidly-convergent methods for evaluating elliptic integrals and theta and elliptic functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):47–58, July 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/rapidlyconvergent-methods-for-evaluating-elliptic-integrals-
and-theta-and-elliptic-functions/2D993C9A7C9EB1D4B61B856E22B45A34.

Sinha:1982:EFS

[243] P. C. Sinha and Meena Aggarwal. Entry flow in a straight circular pipe. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):59–66, July 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/entry-flow-in-a-straight-circular-pipe/652A5C0AEB7E47CAE90BBB733D6666F1.

McCann:1982:MPZ

[244] Roger C. McCann and E. R. Love. Monotonicity properties of the zeros of Bessel functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):67–85, July 1982. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/monotonicity-properties-of-
the-zeros-of-bessel-functions/6709D8C3E024EDAC348CCA8ADCA850CC.

Winley:1982:ASC

[245] Graham Winley and Keith Tognetti. Approximate solutions for coupled moment equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):86–97, July 1982. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/approximate-solutions-
for-coupled-moment-equations/443E2C7AA1A8E23806A7142D6FBCCD0A.

Moodie:1982:RAP

[246] T. Bryant Moodie, R. J. Tait, and D. W. Barclay. Ray analysis and punching problems for stretched elastic plates. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1): 98–119, July 1982. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/ray-
analysis-and-punching-problems-for-stretched-elastic-plates/
8192FFE7E8A8356E722DC9001F0F6ADD.

Anonymous:1982:AVIe

[247] Anonymous. ANZ volume 24 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):f1–f3, July 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-24-issue-1-cover-and-front-matter/B5B1ABE32E4C558A1777D758A7F29406.

Anonymous:1982:AVIf

[248] Anonymous. ANZ volume 24 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):b1–b2, July 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-24-issue-1-cover-and-back-matter/3D58473ADBED054F1E4F689829B204A7.

Liyanage:1982:NRW

[249] L. H. Liyanage, J. M. Hill, and C. M. Gulati. A note on the random walk model arising in double diffusion. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(2):121–129, October 1982. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
the-random-walk-model-arising-in-double-diffusion/AC2B5654BD8CC30F795880ADAC65FBDB.

Lai:1982:SPD

[250] Hang-Chin Lai and Shu-Shih Yang. Saddle point and duality in the optimization theory of convex set functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (2):130–137, October 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/saddle-point-and-duality-in-the-optimization-theory-
of-convex-set-functions/04F2BB33A22DC35D49B7879712805ADB.

OBrien:1982:TFS

[251] D. M. O’Brien. A trace formula for Schrödinger operators with step potentials. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(2):138–159, October 1982. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/trace-formula-for-
schrodinger-operators-with-step-potentials/AC660575C69199D50732AF0672C316AC.

Gopalsamy:1982:EET

[252] K. Gopalsamy. Exchange of equilibria in two species Lotka–Volterra competition models. Journal of the Australian Mathematical Society:

Series B, Applied Mathematics, 24(2):160–170, October 1982. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/exchange-of-equilibria-in-two-
species-lotkavolterra-competition-models/FD601F434A8FE6A81CA353E6892B2645.

Lacey:1982:MBP

[253] A. A. Lacey. Moving boundary problems in the flow of liquid through porous media. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(2):171–193, October 1982. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/moving-boundary-problems-in-the-flow-
of-liquid-through-porous-media/4BAC8404B278BD17372871A681B7815D.

Ayeni:1982:ECT

[254] R. O. Ayeni. On the explosion of chain-thermal reactions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(2):194–202, October 1982. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-the-explosion-of-chainthermal-reactions/
067A9C28AEF407DDC06D45B154C5CD16.

Lavery:1982:ISQ

[255] John E. Lavery. Iterative solution of quasilinear parabolic equations by parabolic equations with constant coefficients. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (2):203–222, October 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/iterative-solution-of-quasilinear-parabolic-equations-
by-parabolic-equations-with-constant-coefficients/334B06FB56A2BCF18FC443C8C6E0A0D5.

Sidi:1982:CFS

[256] Avram Sidi. Converging factors for some asymptotic moment series that arise in numerical quadrature. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (2):223–233, October 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/converging-factors-for-some-asymptotic-moment-series-
that-arise-in-numerical-quadrature/64756EACD87BDB6A0B4AEC6A74302DED.

Karmakar:1982:DHC

[257] Sudhangshu B. Karmakar. Detection of Hamiltonian circuits in a directed graph. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(2):234–242, October 1982. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/detection-of-hamiltonian-circuits-in-
a-directed-graph/DB2DA61F672269069A9C4A5C86A51DE7.

Anonymous:1982:AVIg

[258] Anonymous. ANZ volume 24 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(2):f1–f2, October 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-24-issue-2-cover-and-front-matter/E7FB7BA837890C6D8FBC8F38F08BD611.

Anonymous:1982:AVIh

[259] Anonymous. ANZ volume 24 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(2):b1–b2, October 1982. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-24-issue-2-cover-and-back-matter/8D34D2EA79970CC2ADDDF4B1C3834092.

VanDyke:1983:SPF

[260] M. D. Van Dyke and A. J. Guttmann. Subsonic potential flow past a circle and the transonic controversy. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (3):243–261, January 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/subsonic-potential-flow-past-a-circle-and-the-transonic-
controversy/6864BD8F2276DA7E93108D4AD1CBC998.

Keats:1983:OPC

[261] R. G. Keats, Winifred Frost, and Annette Dobson. The optimum processing of clipped signals: an approach based an a likelihood ratio statistic. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(3):262–278, January 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/optimum-processing-of-
clipped-signals-an-approach-based-an-a-likelihood-ratio-statistic/
DA4145E08F849D95731636497A686682.

Tam:1983:TIR

[262] K. K. Tam and P. B. Chapman. Thermal ignition in a reactive slab with unsymmetric boundary temperatures. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(3):279–288, January 1983. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/thermal-

ignition-in-a-reactive-slab-with-unsymmetric-boundary-temperatures/
3132409693099DA262C54DA3746BC559.

Wallis:1983:TP

[263] W. D. Wallis. A tournament problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (3):289–291, January 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/tournament-problem/F33341D5E68B6A0293B28A2246472705.

Braddock:1983:TLD

[264] R. D. Braddock and P. van den Driessche. On a two lag differential delay equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(3):292–317, January 1983. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
a-two-lag-differential-delay-equation/97BA0DEE41B7AB1A85CB7544BE7E711B.

Dutton:1983:EHC

[265] John A. Dutton and Peter E. Kloeden. The existence of Hadley convective regimes of atmospheric motion. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (3):318–338, January 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-of-hadley-convective-regimes-of-atmospheric-
motion/366927A9FE6EBD3F65C7E24C9FE8E6CF.

Reddy:1983:SPS

[266] D. R. K. Reddy and V. U. M. Rao. Static plane-symmetric solution of a scalar-tensor theory of gravitation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (3):339–342, January 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/static-planesymmetric-solution-of-a-scalartensor-theory-
of-gravitation/1A8649563AB2E7D65FA93CA050B681D3.

Gottlieb:1983:EEA

[267] H. P. W. Gottlieb. The effect of an enclosed air cavity on a rectangular drum. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(3):343–349, January 1983. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/effect-of-an-enclosed-air-cavity-on-
a-rectangular-drum/14270B1889A060C13148B70ED7ED86AB.

Davis:1983:SPC

[268] G. B. Davis and A. G. Morris. Stability of plane Couette flow for high Reynolds number. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(3):350–357, January 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/stability-of-plane-couette-
flow-for-high-reynolds-number/A3DFEE9ABAD3B13E013C54E985C03994.

Anonymous:1983:AVIa

[269] Anonymous. ANZ volume 24 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(3):f1–f2, January 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-24-issue-3-cover-and-front-matter/4EA958612A71FCAAC214613FC01CFADD.

Anonymous:1983:AVIb

[270] Anonymous. ANZ volume 24 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(3):b1–b2, January 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-24-issue-3-cover-and-back-matter/32BB25AB8C80FB0AD0835417C08D5DF9.

Weir:1983:ICR

[271] G. J. Weir. Insulating circular rugs. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):359–365, April 1983. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/insulating-circular-
rugs/E250F3D78453B18068B9DC1F923C92A9.

Anderson:1983:PBS

[272] N. Anderson and A. M. Arthurs. Pointwise bounds for a solution of a nonlinear boundary value problem in partial differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):366–373, April 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/pointwise-bounds-for-a-solution-of-a-nonlinear-boundary-
value-problem-in-partial-differential-equations/6ACA42A879EFB55D854268375EA15991.

Burnell:1983:SSRa

[273] J. G. Burnell, A. A. Lacey, and G. C. Wake. Steady states of the reaction–diffusion equations. Part 1: Questions of existence and continuity of solution branches. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (4):374–391, April 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/steady-states-of-the-reactiondiffusion-equations-part-
1-questions-of-existence-and-continuity-of-solution-branches/
5342A765DA22B0E3F04DCB62CF3A1A9F.

Burnell:1983:SSRb

[274] J. G. Burnell, A. A. Lacey, and G. C. Wake. Steady states of the reaction–diffusion equations. Part II: Uniqueness of solutions and some special cases. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):392–416, April 1983. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/steady-states-of-the-reactiondiffusion-
equations-part-ii-uniqueness-of-solutions-and-some-special-
cases/B34BD3EDD9778A1BBF172BA93DFBE64F.

Lee:1983:MPD

[275] Alexander I. Lee and James M. Hill. On maximum principles for diffusion in the presence of three diffusion paths. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):417–423, April 1983. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-maximum-principles-for-diffusion-in-the-
presence-of-three-diffusion-paths/032C455ACF6C34C4E78E043380946658.

Twizell:1983:NLO

[276] E. H. Twizell and R. W. Ogden. Non-linear optimization of the material constants in Ogden’s stress-deformation function for incompressinle isotropic elastic materials. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24 (4):424–434, April 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinear-optimization-of-the-material-constants-in-
ogdens-stressdeformation-function-for-incompressinle-isotropic-
elastic-materials/58A088EB0EB2D66BF3551C8069FBBB63.

Gottlieb:1983:HSA

[277] H. P. W. Gottlieb. Hearing the shape of an annular drum. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):435–438, April 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/hearing-the-shape-of-an-annular-drum/310B747E6FC44E3694E9991DC331852D.

Broadbridge:1983:ETS

[278] P. Broadbridge. Existence theorems for Segal quantization via spectral theory in Krein space. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):439–460, April 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/existence-theorems-for-segal-
quantization-via-spectral-theory-in-krein-space/59E5EC758511C66B7981D9DAAF3AEA1E.

Reddy:1983:FCP

[279] D. R. K. Reddy and V. U. M. Rao. Field of a charged particle in the presence of scalar meson fields in general relativity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):461–465, April 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/field-of-a-charged-particle-in-the-presence-of-scalar-
meson-fields-in-general-relativity/373751387037646DA0387FF199F24878.

Hoa:1983:SGI

[280] H. Van Hoa. Solution of general inhomogeneous linear difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):466–472, April 1983. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/solution-of-general-inhomogeneous-linear-
difference-equations/6032676AE8ADBA883511D51A7D2D70F2.

Tam:1983:TES

[281] K. K. Tam. The temporal evolution of a system in combustion theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):473–483, April 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/temporal-evolution-of-
a-system-in-combustion-theory/424F0F2492D055258CD4ECADAD90D64A.

deMestre:1983:SEE

[282] Neville de Mestre and Trevor Pakes. Some elementary exact channel flows. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):484–491, April 1983. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/some-elementary-exact-channel-flows/
ABEB9C66654A07325CC2C625E6AA955A.

Anonymous:1983:CV

[283] Anonymous. Contents of volume 24. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):492–493, April 1983. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
24/F99F4260D21CAFCF8F1CEEFC6E79EDAF.

Anonymous:1983:AVIc

[284] Anonymous. ANZ volume 24 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):f1–f2, April 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-24-issue-4-cover-and-front-matter/A73B079975D64AB04F653876E381F88F.

Anonymous:1983:AVId

[285] Anonymous. ANZ volume 24 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(4):b1–b2, April 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-24-issue-4-cover-and-back-matter/2AB2D52C793131B2F579E67C39D00D82.

Bryant:1983:CGW

[286] P. J. Bryant. Cyclic gravity waves in deep water. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):2–15, July 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/cyclic-gravity-waves-in-deep-water/5B2630E12F17F431FCC40730E49758A4.

Peregrine:1983:WWN

[287] D. H. Peregrine. Water waves, nonlinear Schrödinger equations and their solutions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):16–43, July 1983. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/water-waves-nonlinear-
schrodinger-equations-and-their-solutions/D87F5416C657F3B5C35AE96DF9F73DD0.

Sobey:1983:SAS

[288] R. J. Sobey and E. J. Colman. Scattering analysis and synthesis of wave trains. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):44–63, July 1983. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/scattering-analysis-and-synthesis-of-wave-trains/
8812C64EFBDD60C7B5842F88B74A1985.

Viera:1983:GSW

[289] F. Viera and V. T. Buchwald. The generation of surface waves by an intense cyclone. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):64–83, July 1983. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/generation-of-surface-
waves-by-an-intense-cyclone/F6F5C24463F4035F3FF51B3C3CE80456.

Miles:1983:SWI

[290] John W. Miles. Surface-wave interaction with a deeply submerged circular duct. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):84–93, July 1983. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/surfacewave-interaction-with-a-deeply-
submerged-circular-duct/2D467CA6C0FCA3A348F78C4E0A537AA9.

Hung:1983:IIW

[291] N. T. Hung and S. A. Maslowe. Interaction of internal waves in a continuous thermocline model. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):94–109, July 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/interaction-of-internal-waves-
in-a-continuous-thermocline-model/7253974E0B0866ACEF87CB6076EB2BDB.

Hughes:1983:ACS

[292] Roger L. Hughes. The anti-cyclonic shear wave: a new geophysical wave. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):110–126, July 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anticyclonic-shear-wave-a-new-geophysical-wave/171B755BC6100D131158978AAE6B388F.

Fandry:1983:SWF

[293] C. B. Fandry, R. L. Hughes, and L. M. Leslie. Stationary waves forced by topography in a vertically sheared, stratified, rotating fulid. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):127–144, July 1983. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/stationary-waves-forced-by-topography-in-a-
vertically-sheared-stratified-rotating-fulid/FC59C3274C943EBC7A7F7B17E1622B86.

Anonymous:1983:AVIe matics, 25(1):f1–f3, July 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-25-issue-1-cover-and-front-matter/83281093813066B8066306842958D1C0.

Anonymous:1983:AVIf

[295] Anonymous. ANZ volume 25 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(1):b1–b2, July 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-25-issue-1-cover-and-back-matter/E4038BB339A4B8E03439980DC6EB6028.

Wood:1983:EII

[296] W. W. Wood. Energy intensity of inertial waves in a sphere. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):145–160, October 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/energy-intensity-of-inertial-waves-in-a-sphere/01985B00588E203F6A66A57D9CD5C002.

McKellar:1983:IST

[297] B. H. J. McKellar, M. A. Box, and E. R. Love. Inversion of the Struve transform of half integer order. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25 (2):161–174, October 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/inversion-of-the-struve-transform-of-half-integer-order/
9E3204A3960C96ABD17A6B1BBD339279.

Leipnik:1983:CTM

[298] R. B. Leipnik. On Charwat’s theory of motion of tracers in planar vortex flows. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):175–189, October 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/on-charwats-theory-of-
motion-of-tracers-in-planar-vortex-flows/6BBB19E5D59488982A926DFD395DC3B1.

Barry:1983:DFA

[299] J. M. Barry, J. H. Jenkinson, and J. P. Pollard. A discrete Fourier analysis of coarse mesh rebalancing and some associated iterative methods. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):190–216, October 1983. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/discrete-fourier-analysis-of-coarse-
mesh-rebalancing-and-some-associated-iterative-methods/BE23E8C44ED628C311FA8ABD42E57A76.

Harper:1983:ASF

[300] J. F. Harper. Axisymmertic Stokes flow images in spherical free surfaces with applications to rising bubbles. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25 (2):217–231, October 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/axisymmertic-stokes-flow-images-in-spherical-free-surfaces-
with-applications-to-rising-bubbles/0B2F857F4913B18861035D185BA9A9A4.

Sarlet:1983:SIP

[301] W. Sarlet. Symmetries and the inverse problem of Lagrangian dynamics for linear systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):232–243, October 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/symmetries-and-the-inverse-
problem-of-lagrangian-dynamics-for-linear-systems/88707A84BBF9369AC9A0E42DD8224337.

Bestman:1983:LRN

[302] A. R. Bestman. Low Reynolds number flow in a heated tube of varying section. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):244–260, October 1983. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/low-reynolds-number-flow-in-a-heated-
tube-of-varying-section/62D2A9DF8EA9B975C5F7AED1E3660D52.

Elliott:1983:GPM

[303] David Elliott. A Galerkin–Petrov method for singular integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):261–275, October 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/galerkinpetrov-method-
for-singular-integral-equations/C00480B9B4B0ABD115057503C4BE9456.

Gupta:1983:BGP

[304] J. R. Gupta, S. K. Sood, R. G. Shandil, M. B. Banerjee, and K. Banerjee. Bounds for the growth of a perturbation in some double-diffusive convection problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):276–285, October 1983. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/bounds-for-the-growth-of-a-
perturbation-in-some-doublediffusive-convection-problems/67DBF123E5B44A4501AC624693C8D639.

Anonymous:1983:AVIg

[305] Anonymous. ANZ volume 25 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):f1–f2, October 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-25-issue-2-cover-and-front-matter/44D7BB26588E545E91486605ACB1F9B4.

Anonymous:1983:AVIh

[306] Anonymous. ANZ volume 25 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(2):b1–b2, October 1983. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-25-issue-2-cover-and-back-matter/381645A90AB63C68E92E53360D8CD1D3.

Barton:1984:ATD

[307] N. G. Barton. An asymptotic theory for dispersion of reactive contaminants in parallel flow. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):287–310, January 1984. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
asymptotic-theory-for-dispersion-of-reactive-contaminants-in-
parallel-flow/F8830C26500C185DD694D4A186A89434.

Johnson:1984:PSA

[308] R. S. Johnson. The propagation of small amplitude long waves on the surface of superfluid helium. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):311–348, January 1984. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/propagation-of-small-amplitude-
long-waves-on-the-surface-of-superfluid-helium/81B3BA13DCE2BFB0AE83034A0D29CA31.

Gopalsamy:1984:HDP

[309] K. Gopalsamy. Harmless delays in a periodic ecosystem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):349–365, January 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/harmless-delays-in-a-periodic-ecosystem/F917AE9C0B1A028D9F9384203F27AA89.

Rhodes-Robinson:1984:GWW

[310] P. F. Rhodes-Robinson. On the generation of water waves at an inertial surface. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):366–383, January 1984. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-the-generation-of-water-waves-at-
an-inertial-surface/E10D96A6BCFE9DDEFAEBEA241274E2A6.

Huilgol:1984:ANA

[311] R. R. Huilgol, R. Janus, M. A. Lohe, and T. W. Sag. On the application of a numerical algorithm for Hopf bifuraction to the hunting of a wheelset. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):384–405, January 1984. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/on-the-application-of-
a-numerical-algorithm-for-hopf-bifuraction-to-the-hunting-of-
a-wheelset/494A2316802B7922AF1499B965D1702A.

Riahi:1984:NOC

[312] N. Riahi. On nonlinear overstable convection rolls in a rotating system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):406–418, January 1984. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-nonlinear-overstable-convection-rolls-in-
a-rotating-system/356FE84F4C66EB4F8A4B3B4B92DD1246.

Craven:1984:CPA

[313] B. D. Craven, B. Mond, and J. Parida. On a complementarity problem associated with nondifferentiable programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):419–430, January 1984. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
a-complementarity-problem-associated-with-nondifferentiable-
programming/0E09B5A6B2B03B171326EE2279BF597B.

Anonymous:1984:AVIa

[314] Anonymous. ANZ volume 25 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):f1–f2, January 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-25-issue-3-cover-and-front-matter/890E3BBC3F8D607B72FB6E6E0B256115.

Anonymous:1984:AVIb

[315] Anonymous. ANZ volume 25 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(3):b1–b2, January 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-25-issue-3-cover-and-back-matter/ECDB0C8F2A62440638C0E07084453DD0.

Anderssen:1984:LFF

[316] R. S. Anderssen and D. R. Jackett. Linear functionals of foliage angle density. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(4):431–442, April 1984. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/linear-functionals-of-foliage-angle-
density/5E8B30FACA1F2DECC73F3C763C8B1FD4.

Tuck:1984:CLF

[317] E. O. Tuck and J.-M. Vanden Broeck. A cusp-like free-surface flow due to a submerged source or sink. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25 (4):443–450, April 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/cusplike-freesurface-flow-due-to-a-submerged-source-
or-sink/1C3EF055F8D2E1EF206B99BB76BF3262.

Brown:1984:SPT

[318] A. Brown. solutions of period three for a non-linear difference equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(4):451–462, April 1984. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/solutions-of-period-three-for-a-nonlinear-difference-
equation/881E306B1E07413B3603FCF828C17DAC.

Chesson:1984:IDR

[319] Peter L. Chesson. Infinitely divisible random transition probabilities with application to dependent Markov chains. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25 (4):463–472, April 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/infinitely-divisible-random-transition-probabilities-
with-application-to-dependent-markov-chains/144A9368DE2D4A76C37A9515A7354ED5.

Gopalsamy:1984:DRS

[320] K. Gopalsamy. Delayed responses and stability in two-species systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(4):473–500, April 1984. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/delayed-responses-and-
stability-in-twospecies-systems/98A1790E37E816F9D5DE2E4CF1C7A592.

Rezayat:1984:UBI

[321] M. Rezayat, F. J. Rizzo, and D. J. Shippy. A unified boundary integral equation method for a class of second order elliptic boundary value problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(4):501–517, April 1984. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/unified-
boundary-integral-equation-method-for-a-class-of-second-order-
elliptic-boundary-value-problems/71B5E0176091EE706620A5A179FB101F.

Wong:1984:CGM

[322] K. H. Wong and K. L. Teo. A conditional gradient method for a class of time-lag optimal control problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25 (4):518–537, April 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/conditional-gradient-method-for-a-class-of-timelag-
optimal-control-problems/D44EA6D35A4A5E84C1522396ECF7263F.

Bass:1984:MPI

[323] L. Bass, A. J. Bracken, and R. Vyborny. Minimisation problems for implicit functionals defined by differential equations of liver kinetics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(4):538–562, April 1984. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/minimisation-problems-
for-implicit-functionals-defined-by-differential-equations-
of-liver-kinetics/FFD29A03683C77937758102C252811A2.

Anonymous:1984:CV

[324] Anonymous. Contents of volume 25. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(4):563–564, April 1984. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
25/61BA5E8BAAC782D30CD9DA7F1530A05E.

Anonymous:1984:AVIc

[325] Anonymous. ANZ volume 25 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25(4):f1–f2, April 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-25-issue-4-cover-and-front-matter/ACCCCB345736AC6A1ABD22234443D0F6.

Anonymous:1984:AVId

[326] Anonymous. ANZ volume 25 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 25 (4):b1, April 1984. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-25-issue-4-cover-and-back-matter/4A968453390CEE2D7535D754B2CFB806.

Chandler:1984:GMB

[327] G. A. Chandler. Galerkin’s method for boundary integral equations on polygonal domains. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):1–13, July 1984. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/galerkins-method-for-boundary-integral-
equations-on-polygonal-domains/386F7B7A2AC2C0E3B7593719446E3474.

Davidson:1984:IEI

[328] M. R. Davidson. An integral equation for immiscible fluid displacement in a two-dimensional porous medium or Hele–Shaw cell. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):14–30, July 1984. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-integral-equation-for-immiscible-
fluid-isplacement-in-a-twodimensional-porous-medium-or-heleshaw-
cell/601511113924494666BFA60D6C06932C.

Cerone:1984:NIS

[329] P. Cerone and J. R. Blake. A note on the instantaneous streamlines, pathlines and pressure contours for a cavitation bubble near a boundary. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):31–44, July 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-the-instantaneous-streamlines-pathlines-and-
pressure-contours-for-a-cavitation-bubble-near-a-boundary/4D30706DC31EFDFDBF658A12133B1F65.

Kibalczyc:1984:NOC

[330] K. Kibalczyc and S. Walczak. Necessary optimality conditions for a problem with costs of rapid variation of control. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):45–55, July 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/necessary-optimality-conditions-for-a-problem-with-
costs-of-rapid-variation-of-control/503520DCF031C795AEF8E05639171ECC.

Gair:1984:DMN

[331] F. C. Gair. On the design of mortgages and the need for indexation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):56–76, July 1984. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-design-of-mortgages-
and-the-need-for-indexation/371A63EC15B7F0960AF0AF5589A906A8.

Bucco:1984:BAP

[332] D. Bucco and J. Mazumdar. Buckling analysis of plates of arbitrary shape. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):77–91, July 1984. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/buckling-analysis-
of-plates-of-arbitrary-shape/4E0727D0DB77AE3E4323D712E8F292E7.

Werner:1984:OPB

[333] K.-D. Werner. An observation problem for the Bessel differential operator. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):92–107, July 1984. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/an-observation-problem-for-
the-bessel-differential-operator/B28320B9B463375D6304DE889E2E85E3.

Buchwald:1984:DLE

[334] V. T. Buchwald and T. Tran Cong. The diffraction of long elastic waves by elliptic cylindrical cavities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26 (1):108–118, July 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/diffraction-of-long-elastic-waves-by-elliptic-cylindrical-
cavities/909C3BC3F6E7AF3CA07EB13A49F89063.

Anonymous:1984:AVIe

[335] Anonymous. ANZ volume 26 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):f1–f3, July 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-26-issue-1-cover-and-front-matter/97897063582FCF2CC57FEA1C8DBD5F14.

Anonymous:1984:AVIf

[336] Anonymous. ANZ volume 26 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(1):b1–b2, July 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-26-issue-1-cover-and-back-matter/A81592E585D7FDC5AF108F22AA8BF103.

Forrester:1984:EST

[337] Peter J. Forrester. An exactly solvable two component classical Coulomb system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):119–128, October 1984. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
exactly-solvable-two-component-classical-coulomb-system/3E47E973E72B178564BE367D5E3DD02D.

Nedeljkovic:1984:NAD

[338] Nikola B. Nedeljković. New algorithms for discrete-time optimal control problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):129–145, October 1984. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/new-
algorithms-for-discretetime-optimal-control-problems/9A9B2E01973B2D5FD9701E72AE968E6D.

Brown:1984:SPF

[339] A. Brown. Solutions of period four for a non-linear difference equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):146–164, October 1984. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/solutions-of-period-four-for-a-nonlinear-difference-
equation/06D016D5CE3A8B307F60471BBE48DF39.

Hill:1984:ILB

[340] James M. Hill and Jeffrey N. Dewynne. Improved lower bounds for the motion of moving boundaries. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):165–175, October 1984. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/improved-lower-bounds-
for-the-motion-of-moving-boundaries/459B271013FC437CCEDC58DA6F0D8341.

Qureshi:1984:NHP

[341] M. I. Qureshi and M. A. Pathan. A note on hypergeometric polynomials. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):176–182, October 1984. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/note-on-hypergeometric-polynomials/7FE7BE6AEB9151EADFF1FD16B8D64848.

Sharp:1984:UWO

[342] P. W. Sharp. Unsteady waves on an open two layer fluid. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):183–199, October 1984. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/unsteady-waves-on-an-open-two-layer-
fluid/C60C0F75640022AA6DFC07B560EBFB7D.

Woods:1984:NSF

[343] J. E. Woods. Notes on Sraffa’s fixed capital model. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26 (2):200–232, October 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/notes-on-sraffas-fixed-capital-model/AB2EF4319D4746A0B37F08E4837B0470.

Dahmen:1984:AOC

[344] Wolfgang Dahmen and Charles A. Micchelli. On the approximation order from certain multivariate spline spaces. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):233–246, October 1984. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-approximation-order-from-certain-multivariate-spline-spaces/
BB43AAC1F658BA07732CA32E3D648F70.

Anonymous:1984:AVIg

[345] Anonymous. ANZ volume 26 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):f1–f2, October 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-26-issue-2-cover-and-front-matter/4905209494605B6D1F619016A6D4FC8B.

Anonymous:1984:AVIh

[346] Anonymous. ANZ volume 26 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(2):b1–b2, October 1984. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-26-issue-2-cover-and-back-matter/FF1077086BAB2BBBECE772240D828924.

DeHoog:1985:RCF

[347] Frank De Hoog and David Jackett. On the rate of convergence of finite difference schemes on nonuniform grids. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):247–256, January 1985. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-rate-of-convergence-of-finite-difference-schemes-on-nonuniform-
grids/2C707C4CECAE7E4A5359863ACDF8E9B6.

Gould:1985:CIS

[348] M. D. Gould. Characteristic identities for semi-simple Lie algebras. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):257–283, January 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/characteristic-identities-for-semisimple-lie-algebras/
8718BD7E01B55329B1DEE408849304F6.

Chou:1985:SOO

[349] J. H. Chou, Wei-Shen Hsia, and Tan-Yu Lee. Second order optimality conditions for mathematical programming with set functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):284–292, January 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/second-order-optimality-conditions-for-mathematical-
prograramming-with-set-functions/D0A5EACB08A33F394E1C1B9432C4DC2C.

Gottlieb:1985:ELN

[350] H. P. W. Gottlieb. Eigenvalues of the Laplacian with Neumann boundary conditions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):293–309, January 1985. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/eigenvalues-of-the-laplacian-with-neumann-
boundary-conditions/E8834C7E11E70E07C095E9EBF10612E8.

Hooker:1985:GOP

[351] John W. Hooker and William T. Patula. Growth and oscillation properties of solutions of a fourth order linear difference equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):310–328, January 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/growth-and-oscillation-properties-of-solutions-of-a-
fourth-order-linear-difference-equation/5642F6EC1A82441BB5447916688847D6.

Chen:1985:LMF

[352] C. M. Chen and V. Thomée. The lumped mass finite element method for a parabolic problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):329–354, January 1985. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/lumped-mass-finite-element-
method-for-a-parabolic-problem/0D2F96A1EC3713DECBBB9A55F54BF666.

Leipnik:1985:CFS

[353] R. B. Leipnik. A canonical form and solution for the matrix Riccati differential equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):355–361, January 1985. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/canonical-form-and-solution-
for-the-matrix-riccati-differential-equation/7FABE4CB1A6E019D37A03855FC3106E5.

Craven:1985:GFA

[354] B. D. Craven. Generalized functions for applications. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26 (3):362–374, January 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalized-functions-for-applications/531B0E408273D4E4332645BF66C4BC41.

Chawla:1985:FTS

[355] M. M. Chawla and S. R. Sharma. Families of three-stage third order Runge–Kutta–Nyström methods for y′′ = f(x, y, y′). Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):375–386, January 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/families-of-threestage-third-order-rungekuttanystrom-
methods-for-y-f-x-y-y/86E75518DD20899BB60208DFA310734B.

Anonymous:1985:AVIa

[356] Anonymous. ANZ volume 26 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):f1–f2, January 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-26-issue-3-cover-and-front-matter/44386773CD8596F760CA287F613E6ABA.

Anonymous:1985:AVIb

[357] Anonymous. ANZ volume 26 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(3):b1–b2, January 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-26-issue-3-cover-and-back-matter/FED5045E879A376772955D64CB9E7025.

Wright:1985:ILM

[358] S. J. Wright and J. N. Holt. An inexact Levenberg–Marquardt method for large sparse nonlinear least squares. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):387–403, April 1985. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
inexact-levenbergmarquardt-method-for-large-sparse-nonlinear-
least-squres/C92147BBF93B355F317369800FF8CF6A.

Barton:1985:EET

[359] N. G. Barton. End effects in a two dimensional potential problem for closely spaced rectangular plates. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):404–414, April 1985. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/end-
effects-in-a-two-dimensional-potential-problem-for-closely-
spaced-rectangular-plates/2FA7BD2B238229113CDD87DBAE084763.

Clements:1985:NBI

[360] D. L. Clements, M. Haselgrove, and D. M. Barnett. A note on the boundary integral equation method for the solutions of second order elliptic equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):415–421, April 1985. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
the-boundary-integral-equation-method-for-the-solutions-of-
second-order-elliptic-equations/18101728DF0E62337E61581673CCFE02.

Chandra:1985:NPD

[361] S. Chandra, B. D. Craven, and B. Mond. Nonlinear programming duality and matrix game equivalence. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):422–429, April 1985. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/nonlinear-programming-
duality-and-matrix-game-equivalence/748862E883A13AE62A76F96134B5B9F8.

Brown:1985:NDE

[362] A. Brown. A nonlinear difference equation with two parameters. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):430–451, April 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinear-difference-equation-with-two-parameters/BE0F30D6A61F4C2900347FEF82EBCFE1.

Davidson:1985:NCU

[363] M. R. Davidson. Numerical calculation of unstable immiscible fluid displacement in a two-dimensional porous medium or Hele–Shaw cell. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):452–469, April 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-calculation-of-unstable-immiscible-fluid-
displacement-in-a-twodimensional-porous-medium-or-heleshaw-
cell/3FC4F23B5632E01D348CE29FDF660EFF.

Hocking:1985:CLF

[364] G. C. Hocking. Cusp-like free-surface flows due to a submerged source or sink in the presence of a flat or sloping bottom. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):470–486, April 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/cusplike-freesurface-flows-due-to-a-submerged-source-
or-sink-in-the-presence-of-a-flat-or-sloping-bottom/D36F17FEABD168F5CC9DC9194D3F8034.

Keady:1985:AES

[365] G. Keady. Asymptotic estimates for symmetric vortex streets. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):487–502, April 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-estimates-for-symmetric-vortex-streets/6E236B56D06A90DFE2B0D1195C276D07.

McNabb:1985:UPC

[366] A. McNabb. An uncoupling procedure for a class of coupled linear partial differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):503–516, April 1985. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
uncoupling-procedure-for-a-class-of-coupled-linear-partial-
differential-equations/59CA7C5F1AEF80A5F08A1EDDA032663C.

Scott:1985:DCM

[367] C. H. Scott, T. R. Jefferson, and E. Sirri. On duality for convex minimization with nested maxima. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):517–522, April 1985. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/on-duality-for-convex-
minimization-with-nested-maxima/1388E70648AA6964B41D9F7FE0D502E3.

Anonymous:1985:CV

[368] Anonymous. Contents of volume 26. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):523–524, April 1985. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
26/BBA8A615ABD0016035E4CEC14FA8B698.

Anonymous:1985:AVIc

[369] Anonymous. ANZ volume 26 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):f1–f2, April 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-26-issue-4-cover-and-front-matter/6193479DC5ED1CF16CE529AF80FC013C.

Anonymous:1985:AVId

[370] Anonymous. ANZ volume 26 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 26(4):b1–b2, April 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-26-issue-4-cover-and-back-matter/DF28EBD3D6AAD89ACC5ECD15D535F4C3.

OBrien:1985:TER

[371] D. M. O’Brien and R. S. Smith. Transient electromagnetic response of a layered conducting medium at asymptotically late times. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1):1–30, July 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/transient-electromagnetic-response-of-a-layered-conducting-
medium-at-asymptotically-late-times/E8906FE37093B58056BD2C73A9F6A371.

Duszczyk:1985:ROS

[372] B. Duszczyk, S. Kosinski, and Z. Wesolowski. Reflection of oblique shock waves in incompressible elastic solids. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27 (1):31–47, July 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reflection-of-oblique-shock-waves-in-incompressible-
elastic-solids/5397718A8410D7CF22F9A0B0271F7A84.

Roberts:1985:SED

[373] A. J. Roberts. Simple examples of the derivation of amplitude equations for systems of equations possessing bifurcations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1):48–65, July 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/simple-examples-of-the-derivation-of-amplitude-equations-
for-systems-of-equations-possessing-bifurcations/AD8A58174E9FDFC1F3AFD34C9A9C61BC.

Gopalsamy:1985:GAS

[374] K. Gopalsamy. Global asymptotic stability in a periodic Lotka– Volterra system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1):66–72, July 1985. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/global-asymptotic-
stability-in-a-periodic-lotkavolterra-system/8921220141ABF69559181D5869BA0067.

Hill:1985:GRW

[375] James M. Hill and Barry D. Hughes. On the general random walk formulation for diffusion in media with diffusivities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1): 73–87, July 1985. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-general-random-walk-formulation-for-diffusion-in-media-
with-diffusivities/73222C5D2B639F578A09F04BCA03E4D8.

Burnell:1985:SSR

[376] J. G. Burnell, A. A. Lacey, and G. C. Wake. Steady states of the reaction– diffusion equations. Part III: Questions of multiplicity and uniqueness of solutions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1):88–110, July 1985. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/steady-states-of-the-reactiondiffusion-
equations-part-iii-questions-of-multiplicity-and-uiqueness-
of-solutions/C1F938D26ED91B5014EE95FECC95489C.

Mehlum:1985:SCQ

[377] Even Mehlum and Jet Wimp. Spherical curves and quadratic relationships for special functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1):111–124, July 1985. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/spherical-curves-and-quadratic-
relationships-for-special-functions/2D539990857E007B7CBFD499AF3AD95B.

Bakulev:1985:GMT

[378] A. P. Bakulev, N. N. Bogolubov, and A. M. Kurbatov. The generalized Mayer theorem in the approximating Hamiltonian method. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1):125–130, July 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalized-mayer-theorem-in-the-approximating-hamiltonian-
method/DA83B34656A519C60223A3BFA10C128E.

Anonymous:1985:AVIe

[379] Anonymous. ANZ volume 27 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1):f1–f3, July 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-27-issue-1-cover-and-front-matter/384B68D1B39E763DCBC8AF610B74A85A.

Anonymous:1985:AVIf

[380] Anonymous. ANZ volume 27 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(1):b1–b2, July 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-27-issue-1-cover-and-back-matter/338DD6959BF56032650171A7931504F6.

Murray:1985:SND

[381] J. M. Murray. Simple nonlinear dual control problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(2):131–144, October 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/simple-nonlinear-dual-control-problems/81B76EA505B4AC67FF938D44C3485F6A.

Brown:1985:NLD

[382] A. Brown. A non-linear difference equation with two parameters. II. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(2):145–166, October 1985. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/nonlinear-difference-equation-with-two-parameters-
ii/DC311F7A867D41EC8E80B6D875DA7AE1.

Locker:1985:SLS

[383] John Locker and P. M. Prenter. Splitting least squares and collocation procedures for two-point boundary value problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(2):167–193, October 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/splitting-least-squares-and-collocation-procedures-
for-twopoint-boundary-value-problems/593574F7607790334F38E67383B76E67.

Anderson:1985:SSE

[384] B. D. O. Anderson, W. A. Coppel, and D. J. Cullen. Strong system equivalence (I). Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(2):194–222, October 1985. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/strong-system-equivalence-i/C7FCBE939005F4F99AF059CD6808EA8D.

Coppel:1985:SSE

[385] W. A. Coppel and D. J. Cullen. Strong system equivalence (II). Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(2):223–237, October 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/strong-system-equivalence-ii/05DE4BD685071F028FEBD8818002F9AA.

Mishra:1985:SPC

[386] M. S. Mishra, S. Nanda, and D. Acharya. Strong pseudo-convexity and symmetric duality in nonlinear programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27 (2):238–244, October 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/strong-pseudoconvexity-and-symmetric-duality-in-nonlinear-
programming/80A9AE208AEB3AE152FA83A65BC1D1A0.

Grant:1985:CPB

[387] I. F. Grant and B. H. J. McKellar. Critical point behaviour of the diffusion length for radiative transfer. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27 (2):245–257, October 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/critical-point-behaviour-of-the-diffusion-length-for-
radiative-transfer/574863E53AF9F7C7BEB092FA3D4E3FD7.

Anonymous:1985:AVIg

[388] Anonymous. ANZ volume 27 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(2):f1–f2, October 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-27-issue-2-cover-and-front-matter/8C0AFBBC664BBAC42A6F7FFC653629C6.

Anonymous:1985:AVIh

[389] Anonymous. ANZ volume 27 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(2):b1–b2, October 1985. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-27-issue-2-cover-and-back-matter/8B34DB21C89DA97AAA2A184A1BBF2C4F.

Mahony:1986:ALN

[390] J. J. Mahony and J. Norbury. Asymptotic location of nodal lines using geodesic theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):259–280, January 1986. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/asymptotic-location-of-nodal-
lines-using-geodesic-theory/FC229D785298075256B4EA5919760638.

Weir:1986:SMH

[391] G. J. Weir. Surface mounted heat flux sensors. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):281–294, January 1986. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/surface-
mounted-heat-flux-sensors/956F948B296EE6DAEE378140656ABAE0.

Fulford:1986:FDA

[392] G. R. Fulford and J. R. Blake. Force distribution along a slender body straddling an interface. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):295–315, January 1986. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/force-distribution-along-a-
slender-body-straddling-an-interface/1E181CAC5368959BE80E068662A0DA53.

Louis:1986:EEW

[393] J. P. Louis and D. J. Clarke. Exact edge wave solutions for some generalised exponential shelf topographies. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27 (3):316–326, January 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/exact-edge-wave-solutions-for-some-generalised-exponential-
shelf-topographies/255091238714B54E598C679C810E9834.

Grundy:1986:AMA

[394] I. H. Grundy. Airfoils moving in air close to a dynamic water surface. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):327–345, January 1986. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/airfoils-moving-in-air-close-to-a-dynamic-water-
surface/73BB2C9836F49058133BAAED17708861.

Gopalsamy:1986:GAS

[395] K. Gopalsamy. Global asymptotic stability in an almost-periodic Lotka–Volterra system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):346–360, January 1986. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/global-asymptotic-
stability-in-an-almostperiodic-lotkavolterra-system/7058291180D9555A823FCF81B23AFD7D.

Fink:1986:OCL

[396] A. M. Fink. Optimal control in liver kinetics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):361–369, January 1986. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
control-in-liver-kinetics/D56ABAE5543C9FFC51A6353A16695A68.

Lekner:1986:PIJ

[397] John Lekner. Parseval’s integral and the Jacobi expansions in series of Bessel functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):370–375, January 1986. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/parsevals-integral-and-the-
jacobi-expansions-in-series-of-bessel-fuinctions/CB59963A40ADE3D8FB915C41427273D5.

Mohd:1986:EFG

[398] Ch. Wali Mohd and M. I. Qureshi. Expansion formulae for general triple hypergeometric series. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):376–385, January 1986. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/expansion-formulae-for-general-
triple-hypergeometric-series/D24A596F416FE990211A315269DA5538.

Anonymous:1986:AVIa

[399] Anonymous. ANZ volume 27 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):f1–f2, January 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-27-issue-3-cover-and-front-matter/5ECD106E89163A427304E9336C5669E5.

Anonymous:1986:AVIb

[400] Anonymous. ANZ volume 27 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(3):b1–b2, January 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-27-issue-3-cover-and-back-matter/53571170BF1626677DB8CC30D749CB7C.

Miller:1986:FDE

[401] John Boris Miller. The foliage density equation revisited. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):387–401, April 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/foliage-density-equation-revisited/74E55245618680744D083412C8623A96.

Carey:1986:ERR

[402] A. L. Carey and D. M. O’Brien. Existence and regularity results for Maxwell’s equations in the quasi-static limit. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27 (4):402–415, April 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-and-regularity-results-for-maxwells-equations-
in-the-quasistatic-limit/55602D512504D50B14AD577F9B4CA389.

Tam:1986:CDD

[403] K. K. Tam. Criticality dependence on data and parameters for a problem in combustion theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):416–441, April 1986. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/criticality-dependence-on-data-
and-parameters-for-a-problem-in-combustion-theory/8A7AE66F669F7D3A30C1ECCAA91299B7.

Madurasinghe:1986:SBC

[404] M. A. D. Madurasinghe and E. O. Tuck. Ship bows with continuous and splashless flow attachment. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27 (4):442–452, April 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/ship-bows-with-contiunous-and-splashless-flow-attachment/
9EA1F078AD7015E9D2D003259AEBC02E.

Werner:1986:BVC

[405] K.-D. Werner. Boundary value control problems involving the Bessel differential operator. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):453–472, April 1986. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/boundary-value-control-problems-
involving-the-bessel-differential-operator/8BA91B40AE412C5922513AF8A25F36AA.

Woods:1986:DAL

[406] J. E. Woods. On the dynamic analysis of a Leontief model with fixed capital. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):473–487, April 1986. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-dynamic-analysis-of-
a-leontief-model-with-fixed-capitial/62254CA92B4ACB40A6297EDB2920AD93.

Potts:1986:OPD

[407] Renfrey B. Potts. Ordinary and partial difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):488–501, April 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/ordinary-and-partial-difference-equations/923B6B8C37D5E568B461148F692E045E.

Ladas:1986:OHO

[408] G. Ladas and Y. G. Sficas. Oscillations of higher-order neutral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):502–511, April 1986. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/oscilations-of-higherorder-
neutral-equations/CA757124553E47404667F0B3FD2AF7F7.

Anonymous:1986:CV

[409] Anonymous. Contents of volume 27. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):512–513, April 1986. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
27/16C244E07C0CCF6CE6AAD4423BC1CD12.

Anonymous:1986:AVIc

[410] Anonymous. ANZ volume 27 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):f1–f2, April 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-27-issue-4-cover-and-front-matter/FA4EB56FBFAC5A1344EEF9874118563E.

Anonymous:1986:AVId

[411] Anonymous. ANZ volume 27 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 27(4):b1–b2, April 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-27-issue-4-cover-and-back-matter/C5EF895948E394B2C09355C6EB782B4B.

Ben-Israel:1986:WI

[412] A. Ben-Israel and B. Mond. What is invexity? Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):1–9, July 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/what-is-invexity/82C3938EDF3585B2AC5C27093E23814F.

Egudo:1986:DGC

[413] R. R. Egudo and B. Mond. Duality with generalized convexity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):10–21, July 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/duality-with-generalized-convexity/B306D4750BE4EEF42254D240749AF383.

Borwein:1986:GDO

[414] Jonathan M. Borwein. Generic differentiability of order-bounded convex operators. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):22–29, July 1986. CODEN JAMMDU. ISSN

0334-2700. URL http://docserver.carma.newcastle.edu.au/1607/;
 http://journals.cambridge.org/action/displayAbstract?fromPage=
online&aid=3973752;   https://www.cambridge.org/core/journals/
anziam-journal/article/generic-differentiability-of-orderbounded-
convex-oparators/EBC10922AB270DAAF5784308C7E3157A.

Craven:1986:NNS

[415] B. D. Craven. A note on nondifferentiable symmetric duality. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):30–35, July 1986. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/note-on-nondifferentiable-symmetric-
duality/7C7CEF545D690951756D9CBB3E1C66C7.

Guddat:1986:SSP

[416] J. Guddat, H. Th. Jongen, and J. Rueckmann. On stability and stationary points in nonlinear optimization. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):36–56, July 1986. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
stability-and-stationary-points-in-nonlinear-optimization/663E094816B996223FA28694CC488317.

Osborne:1986:CRG

[417] M. R. Osborne, S. A. Pruess, and R. S. Womersley. Concise representation of generalised gradients. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):57–74, July 1986. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/concise-
representation-of-generalised-gradients/14E3D0718C4E347C7F95255A9B13EBCC.

Griewank:1986:GCB

[418] Anderas Griewank. The ‘global’ convergence of Broyden-like methods with suitable line search. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):75–92, July 1986. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/global-convergence-of-
broydenlike-methods-with-suitable-line-search/0CADF26A8423FB200A5F224D0C7B2E85.

Teo:1986:OCP

[419] K. L. Teo, K. H. Wong, and Z. S. Wu. An optimal control problem involving a class of linear time-lag systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1): 93–113, July 1986. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
optimal-control-problem-involving-a-class-of-linear-timelag-
systems/F9A011A869CF695DF13F5B4E25B4FDCA.

Davies:1986:ORI

[420] A. R. Davies and R. S. Anderssen. Optimisation in the regularisation illposed problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):114–133, July 1986. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/optimisation-in-the-regularisation-illposed-
problems/42B1FA9051E8B6724C38131BEE6CDDC0.

Duffin:1986:ADC

[421] R. J. Duffin, D. F. Karney, and E. Z. Prisman. Apex duality for constrained optimization. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):134–146, July 1986. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/apex-duality-for-constrained-
optimization/C0CD0BCABDEF21D2E733E024A54796BD.

Anonymous:1986:AVIe

[422] Anonymous. ANZ volume 28 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):f1–f3, July 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-28-issue-1-cover-and-front-matter/FF662E8D5A52B3F6857144450372EA63.

Anonymous:1986:AVIf

[423] Anonymous. ANZ volume 28 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(1):b1–b2, July 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-28-issue-1-cover-and-back-matter/7877DE8140B62B3F8A5F0620ED730F80.

Skowronski:1986:NMR

[424] J. M. Skowronski. Nonlinear model reference adaptive control. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):147–157, October 1986. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/nonlinear-model-reference-adaptive-control/
907D15536F159D7868AFA8FED9742E5D.

Gustafson:1986:ISI

[425] Sven-Åke Gustafson. Investigating semi-infinite programs using penalty functions and Lagrangian methods. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (2):158–169, October 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/investigating-semiinfinite-programs-using-penalty-functions-
and-lagrangian-methods/666507ACFF977407A6FE165F8E1368FB.

Chandra:1986:GFP

[426] S. Chandra, B. D. Craven, and B. Mond. Generalized fractional programming durability: a ratio game approach. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (2):170–180, October 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalized-fractional-programming-duablity-a-ratio-
game-approach/49329E0C2A74B5B56778EFF87BE98882.

Mustard:1986:GMN

[427] David Mustard. Green manure and nitrogenous fertilizer — a two-sector optimal-growth model. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):181–201, October 1986. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/green-manure-and-nitrogenous-
fertilizera-twosector-optimalgrowth-model/6F32AF924BED779BD7E4F672498FA887.

Alvarez:1986:ATD

[428] Carlos Alvarez and Alan C. Lazer. An application of topological degree to the periodic competing species problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):202–219, October 1986. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
application-of-topological-degree-to-the-periodic-competing-
species-problem/C02B3A10E73E2A62B9E13742DF40AA1F.

Brown:1986:ZME

[429] A. Brown. Ziebur’s matrix equation for population growth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):220–228, October 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/zieburs-matrix-equation-for-population-growth/D9A59F0B0A69F5495DC5696880917166.

Zahariev:1986:SOC

[430] A. I. Zahariev and D. D. Bainov. On some oscillation criteria for a class of neutral type functional differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):229–239, October 1986. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/on-some-
oscillation-criteria-for-a-class-of-neutral-type-functional-
differential-equations/07989CD0E546B58AFB37349DF438553E.

Pathan:1986:PBP

[431] M. A. Pathan and Yasmeen. On partly bilateral and partly unilateral generating functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):240–245, October 1986. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
partly-bilateral-and-partly-unilateral-generating-functions/
594F1C505C580E0E0E80813BBE51EBAA.

deMestre:1986:LJR

[432] Neville de Mestre. The long jump record revisited. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (2):246–259, October 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/long-jump-record-revisited/FF7FFD97AF483E9F6B6EB8E157AEB5AD.

Collings:1986:TIF

[433] I. L. Collings. Two infinite-Froude-number cusped free-surface flows due to a submerged line source or sink. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):260–270, October 1986. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/two-
infinitefroudenumber-cusped-freesurface-flows-due-to-a-submerged-
line-source-or-sink/9C3980928C27B2A248AACF288B9C48AF.

Mandal:1986:NST

[434] B. N. Mandal and Krishna Kundu. A note on the singularities in the theory of water waves with an inertial surface. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):271–278, October 1986. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/note-on-the-singularities-in-the-theory-of-
water-waves-with-an-inertial-surface/925BFD503A4C8CE1D555C3F1C330EF45.

Anonymous:1986:AVIg

[435] Anonymous. ANZ volume 28 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):f1–f2, October 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-28-issue-2-cover-and-front-matter/C990A865CEA8C5EE1553C649D2DA18D4.

Anonymous:1986:AVIh

[436] Anonymous. ANZ volume 28 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(2):b1–b2, October 1986. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-28-issue-2-cover-and-back-matter/3AB9204599AEB08A06F2075DAA26530E.

Coleman:1987:HBI

[437] C. J. Coleman. A hybrid boundary integral/Taylor series approach to some nonlinear equations from fluid mechanics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (3):279–286, January 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/hybrid-boundary-integraltaylor-series-approach-to-some-
nonlinear-equations-from-fluid-mechanics/E5CAC15FF47EABE321B3B3D63F211DB6.

Tam:1987:ITE

[438] K. K. Tam. Initiation of thermal explosion by intense light: criticality dependence on data and parameters. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (3):287–295, January 1987. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/initiation-of-thermal-explosion-by-intense-
light-criticality-dependence-on-data-and-parameters/56922A134BCFAC7D164073E5E6C2C98E.

Fabrikant:1987:CFS

[439] V. I. Fabrikant. Closed form solution to some mixed boundary value problems for a charged sphere. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (3):296–309, January 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/closed-form-solution-to-some-mixed-boundary-value-problems-
for-a-charged-sphere/79882B0395F3225F34884E51F5260BF7.

Gould:1987:PIS

[440] M. D. Gould. Polynomial identities for simple Lie superalgebras. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(3):310–327, January 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/polynomial-identities-for-simple-lie-superalgebras/
851E9385026EFB62770DC2988C6A45C7.

Karmakar:1987:HMD

[441] Sudhangshu B. Karmakar. A heuristic method for the determination of a Hamiltonian circuit in a graph. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (3):328–339, January 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/heuristic-method-for-the-determination-of-a-hamiltonian-
circuit-in-a-graph/FEF1F92476BB94C023494181BD11D6B6.

Brown:1987:CVN

[442] A. Brown. Critical values for a nonlinear difference equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(3):340–361, January 1987. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/critical-values-for-a-nonlinear-difference-
equation/6589573E72CC37F706E106B680667304.

Kulenovic:1987:NSC

[443] M. R. S. Kulenović, G. Ladas, and A. Meimaridou. Necessary and sufficient condition for oscillations of neutral differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(3):362–375, January 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/necessary-and-sufficient-condition-for-oscillations-
of-neutral-differential-equations/63CFF8D2A2239404669FA366472BBE8F.

Matula:1987:EP

[444] Joanna Matula. On an extremum problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(3):376–392, January 1987. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
an-extremum-problem/802F090ED3ECDA0F036C7546B369218F.

Teo:1987:CMF

[445] K. L. Teo, G. Jepps, E. J. Moore, and S. Hayes. A computational method for free time optimal control problems, with application to maximizing the range of an aircraft-like projectile. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (3):393–413, January 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/computational-method-for-free-time-optimal-control-
problems-with-application-to-maximizing-the-range-of-an-aircraftlike-
projectile/FD62328A05280C297CC02E3CB0F0C686.

Anonymous:1987:AVIa

[446] Anonymous. ANZ volume 28 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(3):f1–f2, January 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-28-issue-3-cover-and-front-matter/37A1E1C9DDFE747DDDF24FDA8595268C.

Anonymous:1987:AVIb

[447] Anonymous. ANZ volume 28 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(3):b1–b2, January 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-28-issue-3-cover-and-back-matter/0DE702455AC83B338AB72B4B19722E00.

deHoog:1987:FSA

[448] Frank de Hoog and Ian H. Sloan. The finite-section approximation for integral equations on the half-line. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28 (4):415–434, April 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/finitesection-approximation-for-integral-equations-
on-the-halfline/3F94D10D101AD61DAA1CAB69C2C88165.

Ahues:1987:CSS

[449] Mario Ahues. A class of strongly stable operator approximations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(4):435–442, April 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/class-of-strongly-stable-operator-approximations/FD89949C9237CE84F28B19E2489E1A84.

Green:1987:CWH

[450] Michael Green and Brian D. O. Anderson. On the continuity of the Wiener–Hopf factorization operation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(4):443–461, April 1987. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-continuity-of-the-wienerhopf-factorization-operation/885FDC8B1A49066AB9FC0D8C3421D678.

Batchelor:1987:SMF

[451] Murray T. Batchelor. Sparse matrix factorizations of transfer matrices. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(4):462–475, April 1987. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/sparse-matrix-factorizations-of-transfer-matrices/
1CE98AE8E44B9522A5C6F801666B3F31.

Kloeden:1987:USR

[452] P. E. Kloeden. On the uniqueness of solitary Rossby waves. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(4):476–485, April 1987. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-the-uniqueness-of-solitary-rossby-waves/
D57DC5B1F18C062183823869BD0BF4DC.

Dangelmayr:1987:DBN

[453] Gerhard Dangelmayr. Degenerate bifurcations near a double eigenvalue in the Brusselator. Journal of the Australian Mathematical Society:

Series B, Applied Mathematics, 28(4):486–535, April 1987. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/degenerate-bifurcations-near-
a-double-eigenvalue-in-the-brusselator/9135AA27613B4F6B2A8A6E50AFD1FA25.

Anonymous:1987:CV

[454] Anonymous. Contents of volume 28. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(4):536–537, April 1987. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
28/C623FB0C39FB5BC11ECF14C7EA5D0FDF.

Anonymous:1987:AVIc

[455] Anonymous. ANZ volume 28 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(4):f1–f2, April 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-28-issue-4-cover-and-front-matter/6A6E8B8EA309A789062183A99EABF305.

Anonymous:1987:AVId

[456] Anonymous. ANZ volume 28 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 28(4):b1–b2, April 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-28-issue-4-cover-and-back-matter/0AAA68862C96C43317FEC85C0F196B6B.

Ankiewicz:1987:COF

[457] A. Ankiewicz and C. Pask. Chaos in optics: field fluctuations for a nonlinear optical fibre loop closed by a coupler. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(1):1–20, July 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/chaos-in-optics-field-fluctuations-for-a-nonlinear-
optical-fibre-loop-closed-by-a-coupler/EBC59A37F19A96E1CA2F0D740C8B76CD.

Forbes:1987:PSH

[458] Lawrence K. Forbes. Periodic solutions of high accuracy to the forced Duffing equation: Perturbation series in the forcing amplitude. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(1):21–38, July 1987. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/periodic-solutions-of-high-accuracy-to-the-
forced-duffing-equation-perturbation-series-in-the-forcing-
amplitude/D023754EE668CC42323346943EA32148.

Osborne:1987:DBF

[459] M. R. Osborne. Dual barrier functions with superfast rates of convergence for the linear programming problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(1):39–58, July 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/dual-barrier-functions-with-superfast-rates-of-convergence-
for-the-linear-programming-problem/42070E481762F569321A4BE0C44FC26B.

Woods:1987:AJP

[460] J. E. Woods. On the analysis of joint production. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(1): 59–78, July 1987. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-analysis-of-joint-production/883188A68F1028047FA85201ABD2F84E.

Zayed:1987:ELT

[461] E. M. E. Zayed. Eigenvalues of the Laplacian for the third boundary value problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(1):79–87, July 1987. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/eigenvalues-of-the-laplacian-for-the-
third-boundary-value-problem/E95126446147E281E5CB5251AA843404.

Chow:1987:FEE

[462] S. S. Chow and G. F. Carey. Finite element error estimates for subsonic flow. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(1):88–102, July 1987. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/finite-element-error-estimates-for-subsonic-
flow/6FE91F7D9695C812A451CD7234EC97CD.

Marchant:1987:PSC

[463] T. R. Marchant and A. J. Roberts. Properties of short-crested waves in water of finite depth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(1):103–125, July 1987. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/properties-of-shortcrested-
waves-in-water-of-finite-depth/D0FE5C942E90F79986D19E68A5D16925.

Anonymous:1987:AVIe matics, 29(1):f1–f3, July 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-29-issue-1-cover-and-front-matter/C1D3EC78F2AA9444838E171BA78C933B.

Anonymous:1987:AVIf

[465] Anonymous. ANZ volume 29 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(1):b1–b2, July 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-29-issue-1-cover-and-back-matter/1E9F32C5A85748DE0EE5318ACDF9FD67.

McKee:1987:RWW

[466] W. D. McKee and F. Tesoriero. Reflection from a water waves from a vertical vortex sheet in water of finite depth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29 (2):127–141, October 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reflection-from-a-water-waves-from-a-vertical-vortex-
sheet-in-water-of-finite-depth/FF206C0D871926C7E7F804738B53BF8B.

Tuck:1987:TAS

[467] E. O. Tuck. Thin airfoils with small trailing-edge flaps at arbitrary angles. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(2):142–155, October 1987. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/thin-airfoils-with-small-trailingedge-flaps-
at-arbitrary-angles/C9AF00874D5CFD4B4DFAB7F86C322842.

Bass:1987:IDE

[468] L. Bass, A. J. Bracken, K. Holmåker, and B. R. F. Jefferies. Integrodifferential equations for the self-organisation of liver zones by competitive exclusion of cell-types. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(2):156–194, October 1987. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/integrodifferential-equations-
for-the-selforganisation-of-liver-zones-by-competitive-exclusion-
of-celltypes/61634E905A00B8BEC48824EAD85E4914.

Rabinowitz:1987:IPI

[469] Philip Rabinowitz and William E. Smith. Interpolatory product integration for Riemann-integrable functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29 (2):195–202, October 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/interpolatory-product-integration-for-riemannintegrable-
functions/26E08AD97686BC0DDCC6ECCA9C8CDD0F.

Ahner:1987:SSI

[470] John F. Ahner and John S. Lowndes. The solution of some integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(2):203–215, October 1987. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/solution-of-some-integral-equations/
EC506869E10F5D7F22397472766DA7A9.

Lavoie:1987:NPJ

[471] J. L. Lavoie. Notes on a paper by J. B. Miller. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29 (2):216–220, October 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/notes-on-a-paper-by-j-b-miller/33ED402D46EC866D313DFD749E3CBAD8.

Kleeman:1987:OMF

[472] R. Kleeman. Observables in modular field theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29 (2):221–248, October 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/observables-in-modular-field-thoey/D843867E522DBA806C8C8B7A3A3F4467.

Balachandran:1987:EOC

[473] K. Balachandran and D. Somasundaram. Existence of optimal control for nonlinear systems with quadratic performance. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(2):249–255, October 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-of-optimal-control-for-nonlinear-systems-
with-quadratic-performance/C49FBC032AB2F4CC9A52EF0A949F8903.

Anonymous:1987:AVIg

[474] Anonymous. ANZ volume 29 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(2):f1–f2, October 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-29-issue-2-cover-and-front-matter/991D624B0377FA37C8759B4F5C2660BD.

Anonymous:1987:AVIh

[475] Anonymous. ANZ volume 29 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(2):b1–b2, October 1987. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-29-issue-2-cover-and-back-matter/352E52A34571137A22A2370C1DF43263.

Rumsewicz:1988:SWR

[476] M. Rumsewicz and P. Taylor. A spot welding reliability problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):257–265, January 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/spot-welding-reliability-problem/5086D9651856DE06E8295E995AAC5047.

Drynan:1988:DSD

[477] Ross G. Drynan. Determining statistical dominance by linear programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):266–269, January 1988. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/determining-statistical-dominance-by-
linear-programming/2625CCF4554A6BF3C3F7CA5FE9D187F3.

Gottlieb:1988:ELR

[478] H. P. W. Gottlieb. Eigenvalues of the Laplacian for rectilinear regions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):270–281, January 1988. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/eigenvalues-of-the-laplacian-for-rectilinear-
regions/EBB6AFA3EFB866C6F115E5F7FB1C97C5.

vanLoon:1988:SCO

[479] P. M. van Loon and R. M. M. Mattheij. Stable continuous orthonormalisation techniques for linear boundary value problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):282–295, January 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stable-continuous-orthonormalisation-techniques-for-
linear-boundary-value-problems/FDF02F375276C36407DF8470A9C4FA4F.

Bandopadhyaya:1988:IFM

[480] Lakshmisree Bandopadhyaya and M. C. Puri. Impaired flow multiindex transportation problem with axial constraints. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29 (3):296–309, January 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/impaired-flow-multiindex-transportation-problem-with-
axial-constraints/4FF9ABFEC937DD07CC1C5113F52618E5.

Nagahisa:1988:NCM

[481] Youji Nagahisa. Necessary conditions for a mathematical programming problem with set and cone constraints. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29 (3):310–321, January 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/necessary-conditions-for-a-mathematical-programming-
problem-with-set-and-cone-constraints/F4065BE07308365FB41DE1CE14714F4F.

Osborne:1988:AGS

[482] M. R. Osborne and Tania Prvan. On algorithms for generalised smoothing splines. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):322–341, January 1988. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/on-algorithms-for-
generalised-smoothing-splines/5D5895D8B8AF9EB3B57E297F8884E5C6.

Fabrikant:1988:PSA

[483] V. I. Fabrikant. Potential of several arbitrarily located disks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):342–351, January 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/potential-of-several-arbitrarily-located-disks/8632250CDFAFE41E0C17CF71FEC69D94.

Smyth:1988:MFI

[484] Noel Smyth. Mean flow induced by the viscous critical layer in a stratified fluid. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):352–373, January 1988. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/mean-flow-induced-
by-the-viscous-critical-layer-in-a-stratified-fluid/013DAB1BEDA363DC0512FAA696CCE0CB.

Anonymous:1988:AVIa

[485] Anonymous. ANZ volume 29 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):f1–f2, January 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-29-issue-3-cover-and-front-matter/8242181801C27F896262B50462CEBE2C.

Anonymous:1988:AVIb

[486] Anonymous. ANZ volume 29 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(3):b1–b2, January 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-29-issue-3-cover-and-back-matter/9ED9B54E69BA030857E037886F457ABB.

Sahwartiz:1988:TPF

[487] Leonard W. Sahwartiz and Anthony J. Degregoria. Two-phase flow in Hele–Shaw cells: numerical studies of sweep efficiency in a five-spot pattern. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):375–400, April 1988. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/twophase-flow-in-heleshaw-cells-numberical-
studies-of-sweep-efficiency-in-a-fivespot-pattern/00F98DC5EDD9CFF2D226739C86BE7267.

Hocking:1988:IFN

[488] G. C. Hocking. Infinite Froude number solutions to the problem of a submerged source or sink. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):401–409, April 1988. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/infinite-froude-number-solutions-
to-the-problem-of-a-submerged-source-or-sink/74D4B6B947AF07C0D1B30E37B3AA1BC3.

Grimshaw:1988:MSG

[489] R. Grimshaw. The modulation of short gravity waves by long waves or currents. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):410–429, April 1988. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/modulation-of-short-gravity-waves-by-
long-waves-or-currents/23BF5F36294285A8AC32912DAEC0C393.

Craik:1988:ISW

[490] A. D. D. Craik. Interaction of a short-wave field with a dominant long wave in deep water: derivation form Zakharov’s spectral formulation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):430–439, April 1988. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/interction-of-a-shortwave-field-with-a-dominant-
long-wave-in-deep-water-derivation-form-zakharovs-spectral-
formulation/B92A0D1BF3060239558A869FEA0228A5.

Cannon:1988:SRQ

[491] John R. Cannon and George H. Knightly. Some remarks on a quasisteady-state approximation of the Navier–Stokes equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):440–447, April 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-remarks-on-a-quasisteadystate-approximation-of-
the-navierstokes-equation/17E741DF84A07F2654046D0FE14B5ECF.

Davidson:1988:WBF

[492] Rodney F. Davidson. Waves below first cutoff in a duct. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):448–460, April 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/waves-below-first-cutoff-in-a-duct/0D1D14CBE7D212667E01FAA4DFC4CC90.

Werner:1988:BVC

[493] K.-D. Werner. Boundary value controllability and observability problems for the wave and heat equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29 (4):461–479, April 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/boundary-value-controllability-and-observability-problems-
for-the-wave-and-heat-equation/DFD69417BF4802FAB0B6579266E0146C.

Roberts:1988:ACM

[494] A. J. Roberts. The application of centre-manifold theory to the evolution of system which vary slowly in space. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29 (4):480–500, April 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/application-of-centremanifold-theory-to-the-evolution-
of-system-which-vary-slowly-in-space/547F20B033ADA277B499EFED84217262.

Anonymous:1988:CV

[495] Anonymous. Contents of volume 29. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):501–502, April 1988. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
29/E6EEF8ABDADCAF2BAA2DC77CF40A1506.

Anonymous:1988:AVIc

[496] Anonymous. ANZ volume 29 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):f1–f2, April 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-29-issue-4-cover-and-front-matter/6C0E73E1BB089EB97FEF3AE4AF195362.

Anonymous:1988:AVId

[497] Anonymous. ANZ volume 29 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 29(4):b1–b2, April 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-29-issue-4-cover-and-back-matter/96355F4CA00D952348E867F31FECE640.

Robinson:1988:MAS

[498] G. K. Robinson. A method of accelerating stationary iterative methods for solving linear systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):1–23, July 1988. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/method-of-accelerating-stationary-
iterative-methods-for-solving-linear-systems/78C20EC969E2B81B3267C1CBC06B8605.

Lukas:1988:ARS

[499] M. A. Lukas. Assessing regularised solutions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):24–42, July 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/assessing-regularised-solutions/59BFBB8B9D1EAB928FC51F328D24006F.

Osborne:1988:SCG

[500] M. R. Osborne and Tania Prvan. Smoothness and conditioning in generalised smoothing spline calculations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):43–56, July 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/smoothness-and-conditioning-in-generalised-smoothing-
spline-calculations/8007071E415D587D04A90B79E9C13C12.

Prvan:1988:SRF

[501] Tania Prvan and M. R. Osborne. A square-root fixed-interval discretetime smoother. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):57–68, July 1988. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/squareroot-fixedinterval-discretetime-
smoother/C7EE227195795557A8B46BE66AD64C2E.

Zhaolin:1988:CGD

[502] Chen Zhaolin, Hong Huimin, and Zhang Jifeng. Controllability of generalised dynamical systems with constrained control. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):69–78, July 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/controllability-of-generalised-dynamical-systems-with-
constrained-control/64E133CA1E60BBD629D3327C7A5CDCE8.

Rickard:1988:FSW

[503] John Rickard and Neville Hathaway. A flexible strategy for warring duopolists. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):79–88, July 1988. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/flexible-strategy-for-
warring-duopolists/2B447AA18CBFEC108F53F79C59BFEDE4.

Tam:1988:TWS

[504] K. K. Tam. Travelling wave solutions for rich flames of reactive suspensions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):89–100, July 1988. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/travelling-wave-solutions-for-rich-flames-of-
reactive-suspensions/C68362AB7BCE1E3BB3E56F1229B8473C.

Bracken:1988:GDI

[505] A. J. Bracken, H. S. Green, and L. Bass. Groups defined on images in fluid diffusion. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):101–119, July 1988. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/groups-defined-on-
images-in-fluid-diffusion/E6CE55EEF2E2F092E50285D02B10D2A8.

Patel:1988:RDT

[506] L. K. Patel, R. P. Akabari, and U. K. Dave. Radiating Demianskitype metrics and the Einstein–Maxwell fields. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30 (1):120–126, July 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/radiating-demianskitype-metrics-and-the-einsteinmaxwell-
fields/F1D99A2D568CCBA33C2FC29A74BB1C7C.

Anonymous:1988:AVIe

[507] Anonymous. ANZ volume 30 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):f1–f3, July 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-30-issue-1-cover-and-front-matter/505B10C368929CE7A1F46F2BCC77F92E.

Anonymous:1988:AVIf

[508] Anonymous. ANZ volume 30 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(1):b1–b2, July 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-30-issue-1-cover-and-back-matter/95084679A24E6D36100CA8439ABC0174.

Blake:1988:KIA

[509] J. R. Blake. The Kelvin impulse: application to cavitation bubble dynamics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):127–146, October 1988. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/kelvin-impulse-application-
to-cavitation-bubble-dynamics/B4A12BB1C0DFC5C03EE2FB09F9D28274.

King:1988:NFS

[510] A. C. King and M. I. G. Bloor. A note on the free surface induced by a submerged source at infinite Froude number. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):147–156, October 1988. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
the-free-surface-induced-by-a-submerged-source-at-infinite-
froude-number/44A0D4F627585960C71E867EA8988582.

Kachoyan:1988:FDP

[511] B. J. Kachoyan and P. J. Blennerhassett. On the finite Dean problem: linear theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):157–178, October 1988. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-finite-dean-problem-
linear-theory/73F94C14C7F30F8E7F1FB2C2779EB116.

Bestman:1988:ULR

[512] A. R. Bestman. Unsteady low Reynolds number flow in a heated tube of slowly varying section. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):179–202, October 1988. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/unsteady-low-reynolds-
number-flow-in-a-heated-tube-of-slowly-varying-section/E45A239F3C4A3965A238DE16A780C3D7.

Sharan:1988:MMP

[513] Maithili Sharan, M. P. Singh, and Balbir Singh. A mathematical model for the process of gas exchange in lung capillaries using n th order one-step kinetics of oxygen uptake by haemoglobin. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):203–213, October 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/mathematical-model-for-the-process-of-gas-exchange-
in-lung-capillaries-using-nth-order-onestep-kinetics-of-oxygen-
uptake-by-haemoglobin/BB8C5E2B528BA6A31414DC5EFD82D0E6.

Breen:1988:DSS

[514] P. Breen and W. Henderson. Dominated strategies in searching for evolutionary stable strategies. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):214–219, October 1988. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/dominated-strategies-in-searching-
for-evolutionary-stable-strategies/E8E873A6A0957367F64A29E62137A4C1.

Zayed:1988:SAS

[515] E. M. E. Zayed. Some asymptotic spectral formulae for the eigenvalues of the Laplacian. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):220–229, October 1988. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/some-asymptotic-spectral-formulae-for-
the-eigenvalues-of-the-laplacian/11081E8892DF6C354138CC8F559124A3.

Lin:1988:AHM

[516] Wen-Wei Lin and Gerhard Lutzer. An application of the homotopy method to the generalised symmetric eigenvalue problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):230–249, October 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/an-application-of-the-homotopy-method-to-the-generalised-
symmetric-eigenvalue-problem/87DB98AB4CDE0CD50D5364E9B1A02FB0.

Anonymous:1988:AVIg

[517] Anonymous. ANZ volume 30 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):f1–f2, October 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-30-issue-2-cover-and-front-matter/79CF45D57D5AF0BF42CAD04F457601E7.

Anonymous:1988:AVIh

[518] Anonymous. ANZ volume 30 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(2):b1–b2, October 1988. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-30-issue-2-cover-and-back-matter/896CBF0AB250ADF97EE4984CEB082C5B.

Landman:1989:BIM

[519] K. A. Landman. A boundary integral method for contaminant transport in two adjacent porous media. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30 (3):251–267, January 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/boundary-integral-method-for-contaminant-transport-
in-two-adjacent-porous-media/DB088C7B8C32EC895D7BEA670A98C166.

Diamond:1989:SMI

[520] Phil Diamond. Stochastic models for interference between searching insect parasites. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):268–277, January 1989. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/stochastic-models-for-
interference-between-searching-insect-parasites/6EF62D561A9E2AEAFE0250642D61EAC5.

Bracken:1989:FRB

[521] A. J. Bracken, K. Holmåker, L. V. Maloney, and L. Bass. Flux ratios for biological membranes and reciprocity theorems for linear operators. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):278–297, January 1989. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/flux-ratios-for-biological-
membranes-and-reciprocity-theorems-for-linear-operators/75089D3E72A9AEF4ED61122214E9C0D9.

Friedman:1989:SRP

[522] Charles. N. Friedman. Some remarks on Pauli data in quantum mechanics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):298–303, January 1989. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/some-remarks-on-pauli-data-
in-quantum-mechanics/587EF502DB7C6132144A99961930D777.

Kulshrestha:1989:DBD

[523] D. K. Kulshrestha. Duality between distant point and median of a tree network space. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):304–312, January 1989. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/duality-
between-distant-point-and-median-of-a-tree-network-space/7C646330A15B3DB40E51ACAEC16B30E8.

Chan:1989:ESN

[524] W. L. Chan and Yu-Kun Zheng. Exact solutions of nonlinear evolution equations of the AKNS class. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30 (3):313–325, January 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/exact-solutions-of-nonlinear-evolution-equations-of-
the-akns-class/7A5B4593808E6F0A6E9E497B37B1F925.

VanRooyen:1989:CQI

[525] M. Van Rooyen, M. Sears, and S. Zlobec. Constraint qualifications in input optimisation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):326–342, January 1989. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/constraint-qualifications-in-
input-optimisation/2694671D3356B0951812885D1BA38B15.

Craven:1989:BCO

[526] B. D. Craven.  Boundary conditions optimal control.  Jour-

nal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):343–349, January 1989. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/boundary-conditions-optimal-control/
6483EFDFF46376F4C3DDD3F0DB998162.

Teo:1989:CMC

[527] K. L. Teo and C. J. Goh. A computational method for combined optimal parameter selection and optimal control problems with general constraints. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):350–364, January 1989. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/computational-method-for-combined-optimal-
parameter-selection-and-optimal-control-problems-with-general-
constraints/22E801483A40CBE764BFCDA14A67A3F2.

Anonymous:1989:AVIa

[528] Anonymous. ANZ volume 30 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):f1–f2, January 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-30-issue-3-cover-and-front-matter/1E227AA429A20655348A97B9952C28C4.

Anonymous:1989:AVIb

[529] Anonymous. ANZ volume 30 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(3):b1–b2, January 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-30-issue-3-cover-and-back-matter/B3EE1ACA7558686B0F52E664C688903B.

Tuck:1989:WRF

[530] E. O. Tuck. The wave resistance formula of J. H. Michell (1898) and its significance to recent research in ship hydrodynamics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):365–377, April 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/wave-resistance-formula-of-jh-michell-1898-and-its-
significance-to-recent-research-in-ship-hydrodynamics/6D0B69CE2AE6BDC1D06BA675F1C4DEDD.

Grimshaw:1989:AIM

[531] R. Grimshaw. An analysis of the impact of T. M. Cherry’s work on asymptotic expansions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):378–388, April 1989. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
analysis-of-the-impact-of-t-m-cherrys-work-on-asymptotic-expansions/
CE2E24EA5F3474FFBED0008A2FA148A4.

Hurley:1989:MRA

[532] D. G. Hurley. Mathematical research at the Aeronautical Research Laboratories 1939–1960. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):389–413, April 1989. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/mathematical-research-
at-the-aeronautical-research-laboratories-19391960/8AB3EA1068890B188B8DA24220B4DECD.

Fitz-Gerald:1989:ASI

[533] G. F. Fitz-Gerald and N. A. McDonald. An asymptotic solution of an integro-differential equation arising in magnetic coupling through thin shield walls. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):414–423, April 1989. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
asymptotic-solution-of-an-integrodifferential-equation-arising-
in-magnetic-coupling-through-thin-shield-walls/70D5523230FD0030CC4325658BAD2845.

Hall:1989:FDE

[534] A. J. Hall and G. C. Wake. A functional differential equation arising in modelling of cell growth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30 (4):424–435, April 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/functional-differential-equation-arising-in-modelling-
of-cell-growth/95D60D37F79C3B73E17418D368EF9A50.

Ivers:1989:GTP

[535] D. J. Ivers. On generalised toroidal-poloidal solutions of vector field equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):436–449, April 1989. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-generalised-toroidalpoloidal-solutions-
of-vector-field-equations/C8151CF8060F614159E0A1DF77424D1D.

Flytzanis:1989:ESA

[536] E. Flytzanis and Nikolaos S. Papageorgiou. On the existence and sensitivity analysis of optimal capital accumulation paths in continuous time, infinite horizon models. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):450–459, April 1989. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-existence-and-sensitivity-analysis-of-optimal-catital-accumulation-
paths-in-continuous-time-infinite-horizon-models/7CB81E6E8653BE85D181D455B5EEF545.

Andrew:1989:CFD

[537] Alan L. Andrew. Correction of finite difference eigenvalues of periodic Sturm–Liouville problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):460–469, April 1989. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/correction-of-finite-difference-
eigenvalues-of-periodic-sturmliouville-problems/942FF24F2A3B577649B12D7CD17EFDDE.

Murray:1989:OCP

[538] J. M. Murray. Optimal control problems with elastic collisions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):470–482, April 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-control-problems-with-elastic-collisions/9A61E7DF52B75CEB501F3604F06E27C6.

Hill:1989:PUS

[539] James M. Hill and Alex McNabb. On the problem of uncoupling systems of linear differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):483–501, April 1989. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/on-the-problem-of-
uncoupling-systems-of-linear-differential-equations/CF83C3D582A6F2AE98E045337659D3E2.

Anonymous:1989:CV

[540] Anonymous. Contents of volume 30. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):502–504, April 1989. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
30/D102C0325AEDD9892B0D577DE9276D85.

Anonymous:1989:AVIc

[541] Anonymous. ANZ volume 30 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):f1–f3, April 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-30-issue-4-cover-and-front-matter/A17BDB02A2AFBD20C16EED9E474174E3.

Anonymous:1989:AVId

[542] Anonymous. ANZ volume 30 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 30(4):b1–b2, April 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-30-issue-4-cover-and-back-matter/5702BBA38B66808D507B5A1AF3CD6C80.

Tan:1989:DPT

[543] H. H. Tan and R. B. Potts. A discrete path/trajectory planner for robotic arms. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):1–28, July 1989. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/discrete-pathtrajectory-planner-for-
robotic-arms/8B49152B1CE915A7DDBEDD5C2EFCB66E.

Kosinski:1989:NSR

[544] S. Kosinski and B. Duszczyk. Normal shock reflection-transmission in rubber-like elastic material. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):29–47, July 1989. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/normal-shock-reflectiontransmission-
in-rubberlike-elastic-material/60CFF821D6CD4104E010F9B496C18CC1.

Roberts:1989:AIC

[545] A. J. Roberts. Appropriate initial conditions for asymptotic descriptions of the long term evolution of dynamical systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):48–75, July 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/appropriate-initial-conditions-for-asymptotic-descriptions-
of-the-long-term-evolution-of-dynamical-systems/DD2AC177D07E6E819ABA18E250E12351.

Tam:1989:CDD

[546] K. K. Tam. Criticality dependence on data and parameters for a problem in combustion theory, with temperature-dependent conductivity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):76–80, July 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/criticality-dependence-on-data-and-parameters-for-a-
problem-in-combustion-theory-with-temperaturedependent-conductivity/
05D24061FC0C135CB2B4FDED8091EFD5.

Dewynne:1989:ABS

[547] J. N. Dewynne, S. D. Howison, J. R. Ockendon, and Weiqing Xie. Asymptotic behavior of solutions to the Stefan problem with a kinetic condition at the free boundary. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):81–96, July 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-behavior-of-solutions-to-the-stefan-problem-
with-a-kinetic-condition-at-the-free-boundary/CE48D3F428EEC0434179FB0A3B71F983.

Sun:1989:EMF

[548] Min Sun. An evolutionary monotone follower problem in [0, 1]. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):97–107, July 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/an-evolutionary-monotone-follower-problem-in-0-1/49FB7A35EAF4F581F3AC10074BF2D018.

Mond:1989:SOC

[549] B. Mond and I. Husain. Sufficient optimality criteria and duality for variational problems with generalised invexity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):108–121, July 1989. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/sufficient-optimality-criteria-and-duality-
for-variational-problems-with-generalised-invexity/2A9A399584F38FDC465C67C297CFC13A.

Coope:1989:CDI

[550] Ian D. Coope. A conjugate direction implementation of the BFGS algorithm with automatic scaling. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):122–134, July 1989. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/conjugate-direction-implementation-
of-the-bfgs-algorithm-with-automatic-scaling/D4A2CD5F01D63B2668B72C2AB5A9721D.

Anonymous:1989:AVIe

[551] Anonymous. ANZ volume 31 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):f1–f3, July 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-31-issue-1-cover-and-front-matter/945747B857CE8A1395C17E1DDD574F3F.

Anonymous:1989:AVIf

[552] Anonymous. ANZ volume 31 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(1):b1–b2, July 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-31-issue-1-cover-and-back-matter/B98CCA3EE35AF6414670F9CEDC7D904E.

Coyle:1989:SBG

[553] Andrew Coyle. Sensitivity bounds on a GI/M/n/n queueing system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(2):135–149, October 1989. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/sensitivity-bounds-on-
a-gimnn-queueing-system/0322B50218085E3F808ED049F60BC018.

Evans:1989:MMR

[554] Suzanne P. Evans. A mathematical model and related problems of optimal management and design in a broadband integrated services network. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(2):150–175, October 1989. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/mathematical-model-and-related-problems-of-
optimal-management-and-design-in-a-broadband-integrated-services-
network/5B5162C718E9546D9304CBB0B53E77A6.

Henderson:1989:NLP

[555] W. Henderson, D. Lucic, and P. G. Taylor. A net level performance analysis of stochastic Petri nets. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (2):176–187, October 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/net-level-performance-analysis-of-stochastic-petri-
nets/C4096A71CB0E385A8DCC954429D09ED9.

Jajszczyk:1989:NMN

[556] Andrzej Jajszczyk. On nonblocking multiconnection networks composed of digital switching matrices. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(2):188–203, October 1989. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
nonblocking-multiconnection-networks-composed-of-digital-switching-
matrices/CE783916389E9442EB70AE180F62DC5F.

Kelly:1989:FPM

[557] F. P. Kelly. Fixed point models of loss networks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (2):204–218, October 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fixed-point-models-of-loss-networks/FD5AA6764623CC6E4836E7893DFA1524.

Zukerman:1989:AMG

[558] Moshe Zukerman. Applications of matrix-geometric solutions for queueing performance evaluation of a hybrid switching system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(2):219–239, October 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/applications-of-matrixgeometric-solutions-for-queueing-
performance-evaluation-of-a-hybrid-switching-system/A3F84D2999FE686025774027CE6DC5E2.

Anonymous:1989:AVIg

[559] Anonymous. ANZ volume 31 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(2):f1–f3, October 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-31-issue-2-cover-and-front-matter/D2784DF2FB8DA42DBC58FA000A00A30A.

Anonymous:1989:AVIh

[560] Anonymous. ANZ volume 31 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(2):b1–b2, October 1989. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-31-issue-2-cover-and-back-matter/68A9EC4A66BBD398276366545113E7BA.

VanDijk:1990:AEB

[561] Nico M. Van Dijk. Analytic error bounds for approximations of queueing networks with an application to alternate routing. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (3):241–258, January 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/analytic-error-bounds-for-approximations-of-queueing-
networks-with-an-application-to-alternate-routing/E3C83A58CE67CB1211BCA18B74736DF4.

Korman:1990:MIN

[562] Philip Korman, Anthony W. Leung, and Srdjan Stojanovic. Monotone iterations for nonlinear obstacle problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (3):259–276, January 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/monotone-iterations-for-nonlinear-obstacle-problem/
A8DD64001C0B5858B41386DBF197E41D.

Carey:1990:DFM

[563] A. L. Carey and K. McNamara. Degenerate forms of Maxwell’s equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(3):277–300, January 1990. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/degenerate-forms-of-maxwells-equations/7916FD844AFF071C3D0B8C28D12E9AEC.

Denier:1990:SVB

[564] J. P. Denier and R. Grimshaw. Slowly-varying bifurcation theory in dissipative systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(3):301–318, January 1990. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/slowlyvarying-bifurcation-theory-
in-dissipative-systems/E45966FF0AD064A82DAE1DD5535A8995.

Kumar:1990:NSH

[565] Sunil Kumar. The numerical solution of Hammerstein equations by a method based on polynomial collocation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (3):319–329, January 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-solution-of-hammerstein-equations-by-a-method-
based-on-polynomial-collocation/5CF8D72488988E285F55A076D565B0F4.

Bera:1990:NOB

[566] Rasajit Kumar Bera. Nonlinear oscillations and buckling of anisotropic cylindrical shells under large initial stresses. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (3):330–346, January 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinear-oscillations-and-buckling-of-anisotropic-
cylindrical-shells-under-large-initial-stresses/4C4F18852217442115EC7AB7F744BD60.

So:1990:PEP

[567] Joseph W.-H. So. Persistence and extinction in a predator–prey model consisting of nine prey genotypes. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (3):347–365, January 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/persistence-and-extinction-in-a-predatorprey-model-
consisting-of-nine-prey-genotypes/86DE1838C50CFA97495F0E607B3FC7C8.

Anonymous:1990:AVIa

[568] Anonymous. ANZ volume 31 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(3):f1–f2, January 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-31-issue-3-cover-and-front-matter/8482557D0610D98E12C16FBE56E9AB1E.

Anonymous:1990:AVIb

[569] Anonymous. ANZ volume 31 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(3):b1–b2, January 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-31-issue-3-cover-and-back-matter/1BDB9BA4A71C87CC7EF5827799A3CC8C.

Osborne:1990:IPM

[570] M. R. Osborne. An interior point method for linear programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):367–378, April 1990. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-interior-point-method-for-linear-
programming/6AD8643810A8EB46D5D48DC30C3BED23.

Osborne:1990:SUS

[571] M. R. Osborne and R. S. Womersley. Strong uniqueness in sequential linear programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):379–384, April 1990. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/strong-uniqueness-
in-sequential-linear-programming/1C2D26C02894B50EEACAAC08791C8070.

Smyth:1990:PFF

[572] N. F. Smyth. Propagation of flame fronts. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (4):385–396, April 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/propagation-of-flame-fronts/D3D32BAB971060B67E1374CDAB0ACC11.

Denier:1990:NIP

[573] J. P. Denier and R. H. J. Grimshaw. Nonlinear interaction of positive and negative energy modes in Hamiltonian systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):397–424, April 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinear-interaction-of-positive-and-negative-energy-
modes-in-hamiltonian-systems/897C918B71C92C5564AC4E9CC8213D57.

Schutz:1990:IMM

[574] John W. Schutz. The isotropy mappings of Minkowski space–time generate the orthochronous Poincaré group. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31 (4):425–433, April 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/isotropy-mappings-of-minkowski-spacetime-generate-the-
orthochronous-poincare-group/EC250840B92D3C9F05F7CE3E02E520F9.

Hall:1990:FDE

[575] A. J. Hall and G. C. Wake. Functional differential equations determining steady size distributions for populations of cells growing exponentially. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):434–453, April 1990. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/functional-differential-equations-determining-
steady-size-distributions-for-populations-of-cells-growing-
exponentially/3F0716E7CACA63DC67319D1769C607D3.

Howlett:1990:OSC

[576] Phil Howlett. An optimal strategy for the control of a train. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):454–471, April 1990. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-optimal-strategy-for-the-control-
of-a-train/7D2EE7A1B1846024DE51141B0534CCF0.

Zayed:1990:HSA

[577] E. M. E. Zayed. On hearing the shape of an arbitrary doubly-connected region in R^{2}. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):472–483, April 1990. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-hearing-the-shape-of-an-
arbitrary-doublyconnected-region-in-r2/13D944F2D08F62333BCA95D39CBD557B.

Yadav:1990:DFM

[578] Shri Ram Yadav and R. N. Mukherjee. Duality for fractional minimax programming problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):484–492, April 1990. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/duality-
for-fractional-minimax-programming-problems/B2FF1B5A7C31EBBCDF7271689708B898.

Anonymous:1990:CV

[579] Anonymous. Contents of volume 31. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):493–494, April 1990.

CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
31/6BB92D07F9D78B5B64E5B38A5710D245.

Anonymous:1990:AVIc

[580] Anonymous. ANZ volume 31 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):f1–f2, April 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-31-issue-4-cover-and-front-matter/1196BD33DB57DB47A6177395575A9E2B.

Anonymous:1990:AVId

[581] Anonymous. ANZ volume 31 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 31(4):b1–b2, April 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-31-issue-4-cover-and-back-matter/02EF005D96B7A388C48EC1FF42AA2BB1.

Bryant:1990:PFWa

[582] Peter J. Bryant and John W. Miles. On a periodically forced, weakly damped pendulum. Part 1: Applied torque. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(1): 1–22, July 1990. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
a-periodically-forced-weakly-damped-pendulum-part-1-applied-
torque/40A8C15F7910866D37CAB1E36F09A9B2.

Bryant:1990:PFWb

[583] Peter J. Bryant and John W. Miles. On a periodically forced, weakly damped pendulum. Part 2: Horizontal forcing. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(1): 23–41, July 1990. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
a-periodically-forced-weakly-damped-pendulum-part-2-horizontal-
forcing/2513E758D2A331BCF1A8837E90A61B4F.

Bryant:1990:PFWc

[584] Peter J. Bryant and John W. Miles. On a periodically forced, weakly damped pendulum. Part 3: Vertical forcing. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(1): 42–60, July 1990. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
a-periodically-forced-weakly-damped-pendulum-part-3-vertical-
forcing/C279AA663F76676EBA3C7D527A79E633.

Marchant:1990:RND

[585] T. R. Marchant and A. J. Roberts. Reflection of nonlinear deepwater waves incident onto a wedge of arbitrary angle. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(1):61–96, July 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reflection-of-nonlinear-deepwater-waves-incident-onto-
a-wedge-of-arbitrary-angle/8C92AA4A893B868BC609C5ABEC585116.

Giorgi:1990:NRB

[586] Giorgio Giorgi. A note on the relationships between convexity and invexity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(1):97–99, July 1990. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/note-on-the-relationships-between-convexity-
and-invexity/472FBE45BEA6296D61A3A79465DCDA1F.

Fisher:1990:MTT

[587] M. E. Fisher, J. L. Noakes, and K. L. Teo. A minimum trapping time problem in optimal control theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(1):100–114, July 1990. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/minimum-
trapping-time-problem-in-optimal-control-theory/A643D94DBA65ACFEDF791AD3C940A356.

Kulkarni:1990:SSE

[588] Rekha P. Kulkarni and Balmohan V. Limaye. Solution of a Schrödinger equation by iterative refinement. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32 (1):115–131, July 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/solution-of-a-schrodinger-equation-by-iterative-refinement/
F755AB5FEE85A9935B795B6A196A026E.

Anonymous:1990:AVIe

[589] Anonymous. ANZ volume 32 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(1):f1–f3, July 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-32-issue-1-cover-and-front-matter/083C06AA6D7525C50E7DD8400FD2673C.

Anonymous:1990:AVIf

[590] Anonymous. ANZ volume 32 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(1):b1–b2, July 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-32-issue-1-cover-and-back-matter/8A2FFB3DB816349783BE0EC265940E11.

Pollett:1990:MIB

[591] P. K. Pollett. On a model for interference between searching insect parasites. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(2):133–150, October 1990. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-a-model-for-interference-between-
searching-insect-parasites/989CE1AD9ABFFBBB6AD9850E031F192F.

Teo:1990:NCO

[592] K. L. Teo, K. K. Leong, and G. J. Goh. Nonlinearly constrained optimal control problems involving piecewise smooth controls. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(2):151–179, October 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinearly-constrained-optimal-control-problems-involving-
piecewise-smooth-controls/A0726FC2CF515FC462920583DD16052E.

Wudu:1990:ENS

[593] Lu Wudu. Existence of nonoscillatory solutions of first order nonlinear neutral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(2):180–192, October 1990. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/existence-of-nonoscillatory-
solutions-of-first-order-nonlinear-neutral-equations/C2F17081C4E91912C0BDB3F4BF951AB8.

Chapman:1990:UDO

[594] P. B. Chapman. A uniform description of an oscillator’s resonant transition. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(2):193–206, October 1990. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/uniform-description-of-an-oscillators-
resonant-transition/F6CF4D47E13817F7C557548E7DCD5EE9.

Leipnik:1990:RSO

[595] R. B. Leipnik. Reduction of second order linear dynamical systems, with large dissipation, by state variable transformations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(2):207–222, October 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reduction-of-second-order-linear-dynamical-systems-
with-large-dissipation-by-state-variable-transformations/1D5083DDA921ACF532E06FB587EE1C88.

Haidar:1990:EOS

[596] Nassar H. S. Haidar. On eigensolutions of the one-speed neutron transport equation in plane geometry. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(2):223–230, October 1990. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/on-eigensolutions-
of-the-onespeed-neutron-transport-equation-in-plane-geometry/
E8E9FBF2057B3DABA8B9E6D778F9FA36.

Forbes:1990:FCP

[597] Lawrence K. Forbes and Graeme C. Hocking. Flow caused by a point sink in a fluid having a free surface. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32 (2):231–249, October 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/flow-caused-by-a-point-sink-in-a-fluid-having-a-free-
surface/8179E14991E4600B33BAAA35460CD93D.

Anonymous:1990:AVIg

[598] Anonymous. ANZ volume 32 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(2):f1–f2, October 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-32-issue-2-cover-and-front-matter/5C493DB19919B2DF323FC5FDA1757261.

Anonymous:1990:AVIh

[599] Anonymous. ANZ volume 32 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(2):b1–b2, October 1990. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-32-issue-2-cover-and-back-matter/D4962C475BE5134B7E34991A68F425AE.

Hocking:1991:NFI

[600] G. C. Hocking and L. K. Forbes. A note on the flow induced by a line sink beneath a free surface. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):251–260, January 1991. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
the-flow-induced-by-a-line-sink-beneath-a-free-surface/74E7DCAC0852CAA457D8C088E4A9C816.

Lucas:1991:BIM

[601] S. K. Lucas, J. R. Blake, and A. Kucera. A boundary-integral method applied to water coning in oil reservoirs. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32 (3):261–283, January 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/boundaryintegral-method-applied-to-water-coning-in-
oil-reservoirs/20D7C284AF69124D3367B3D79A62154E.

Ang:1991:PAO

[602] W. T. Ang. A pair of arbitrarily-oriented coplanar cracks in an anisotropic elastic slab. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):284–295, January 1991. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/pair-of-
arbitrarilyoriented-coplanar-cracks-in-an-anisotropic-elastic-
slab/218D1F38D300CF397B9D9DFE05B0833C.

Mandal:1991:WDS

[603] B. N. Mandal. On waves due to small oscillations of a vertical plate submerged in deep water. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):296–303, January 1991. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/on-waves-due-to-small-
oscillations-of-a-vertical-plate-submerged-in-deep-water/87F08737FDBEECE30A91B87E6910CB9F.

Klemm:1991:SII

[604] Anthony D. Klemm and Sigurd Y. Larsen. Some integrals involving Legendre polynomials providing combinatorial identities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):304–310, January 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-integrals-involving-legendre-polynomials-providing-
combinatorial-identities/4004FC9622EDBA67B490DB392F34A059.

Kuang:1991:NDT

[605] Y. Kuang. On neutral-delay two-species Lotka–Volterra competitive systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):311–326, January 1991. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-neutraldelay-twospecies-
lotkavolterra-competitive-systems/4D9132DE0297B57FFACF1B1D1AD555DD.

Leipnik:1991:LRV

[606] Roy B. Leipnik. On lognormal random variables: I — the characteristic function. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):327–347, January 1991. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
lognormal-random-variables-ithe-characteristic-function/F1563B5AD8918EF2CD51092F82EB0B73.

Brown:1991:PAT

[607] T. C. Brown and P. K. Pollett. Poisson approximations for telecommunications networks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):348–364, January 1991. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/poisson-approximations-for-
telecommunications-networks/E666D6315B77850CBA4F3EEC5F24C9BF.

Anonymous:1991:AVIa

[608] Anonymous. ANZ volume 32 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):f1–f2, January 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-32-issue-3-cover-and-front-matter/87FC1E86E379130162D35D21AC437274.

Anonymous:1991:AVIb

[609] Anonymous. ANZ volume 32 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(3):b1–b2, January 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-32-issue-3-cover-and-back-matter/0E6971E3EDFE3F606354C22BFD5E27FD.

Potts:1991:DVS

[610] R. B. Potts and X. Yu. Discrete variable structure system with pseudo-sliding mode. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(4):365–376, April 1991. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/discrete-variable-
structure-system-with-pseudosliding-mode/2F7F6A1B2F443A6E42D038ACAF0D9EA7.

Gopalsamy:1991:ODD

[611] K. Gopalsamy and G. Ladas. Oscillations of delay differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(4):377–381, April 1991. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/oscillations-of-delay-
differential-equations/8D7C81CB4421F610C121AAFA2A1D0303.

Erbe:1991:CPI

[612] L. H. Erbe, H. I. Freedman, X. Z. Liu, and J. H. Wu. Comparison principles for impulsive parabolic equations with applications to models of single species growth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(4):382–400, April 1991. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/comparison-principles-
for-impulsive-parabolic-equations-with-applications-to-models-
of-single-species-growth/7C47E034462A0BA31C58AB1967CF5C60.

Cox:1991:CMF

[613] S. M. Cox and A. J. Roberts. Centre manifolds of forced dynamical systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(4):401–436, April 1991. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/centre-manifolds-of-forced-
dynamical-systems/DC47E8C5F53E2A1882252CAF6FEFA4A6.

Dzeng:1991:HCM

[614] D. C. Dzeng and W. W. Lin. Homotopy continuation method for the numerical solutions of generalised symmetric eigenvalue problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(4):437–456, April 1991. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/homotopy-continuation-
method-for-the-numerical-solutions-of-generalised-symmetric-
eigenvalue-problems/1DBCA0E704755A34164E5694F083292C.

Bhutani:1991:CNE

[615] O. P. Bhutani and K. Vijayakumar. On certain new and exact solutions of the Emden–Fowler equation and Emden equation via invariant variational principles and group invariance. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(4):457–468, April 1991. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
certain-new-and-exact-solutions-of-the-emdenfowler-equation-
and-emden-equation-via-invariant-variational-principles-and-
group-invariance/241C5350F06CC73C4F767FA7AACCF05A.

Bucco:1991:ASS

[616] D. Bucco and J. Mazumdar. Analysis of shallow shells by a combination of finite elements and contour methods. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32 (4):469–491, April 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/analysis-of-shallow-shells-by-a-combination-of-finite-
elements-and-contour-methods/64AD27C9759EDDFD7E2854060CFD9419.

Anonymous:1991:AVIc

[617] Anonymous. ANZ volume 32 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(4):f1–f2, April 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-32-issue-4-cover-and-front-matter/13A22E34282F06CB94758DD8D013BC63.

Anonymous:1991:AVId

[618] Anonymous. ANZ volume 32 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 32(4):b1–b2, April 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-32-issue-4-cover-and-back-matter/EB962BAD15044BB1B42F177F697117AA.

Coleman:1991:MHP

[619] C. J. Coleman. On the microwave hotspot problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1): 1–8, July 1991. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-microwave-hotspot-problem/B1AD7A29AF490135CACF5A793F525F6C.

So:1991:IMR

[620] Joseph Wai Hung So. Inertial manifold for a reaction diffusion equation model of competition in a chemostat. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):9–15, July 1991. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/inertial-manifold-
for-a-reaction--diffusion-equation-model-of-competition-in-
a-chemostat/62A8A7D6651C05B993FDA61059770D9A.

Tam:1991:PMC

[621] K. K. Tam. Porous medium combustion: ignition, temporal evolution, and parameter dependence. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):16–26, July 1991. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/porous-medium-combustion-ignition-
temporal-evolution-and-parameter-dependence/2B2B3415929E6D7A32300E3923D11ECA.

Tang:1991:NSF

[622] S. Tang and R. O. Weber. Numerical study of Fisher’s equation by a Petrov–Galerkin finite element method. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):27–38, July 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-study-of-fishers-equation-by-a-petrovgalerkin-
finite-element-method/8FB618FBC0CF771D577F9E4326C7FBEA.

Graham:1991:PCC

[623] I. G. Graham and Y. Yan. Piecewise-constant collocation for first-kind boundary integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):39–64, July 1991. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/piecewiseconstant-
collocation-for-firstkind-boundary-integral-equations/D19985E0123CF5AF8B502F92E1AB11A3.

deMestre:1991:MAW

[624] Neville de Mestre. A mathematical analysis of wind effects on a long-jumper. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):65–76, July 1991. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/mathematical-analysis-
of-wind-effects-on-a-longjumper/832F203D6E7C071FBD583E5D2204C850.

Patel:1991:SBT

[625] L. K. Patel and Sharda S. Koppar. Some Bianchi type VI_{0} viscous fluid cosmological models. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):77–84, July 1991. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/some-bianchi-type-vi0-
viscous-fluid-cosmological-models/1DBBEBDDAF7F9EBE56C2AAD770C8C441.

Erbe:1991:CUG

[626] Lynn H. Erbe and Zhongchao Liang. Continuation and uniqueness for generalised Emden–Fowler systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33 (1):85–93, July 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/continuation-and-uniqueness-for-generalised-emdenfowler-
systems/C385B0D3AE1CD0FDE1B7E2E0656E2D35.

Salvi:1991:EVI

[627] Rodolfo Salvi. The equations of viscous incompressible non-homogeneous fluids: on the existence and regularity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33 (1):94–110, July 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/equations-of-viscous-incompressible-nonhomogeneous-
fluids-on-the-existence-and-regularity/BA4E107F59CEB64C16AA10672CEEC4DE.

Ghosh:1991:CWM

[628] N. K. Ghosh. A cylindrical wave-maker problem in a liquid of finite depth with an inertial surface in the presence of surface tension. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):111–121, July 1991. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/cylindrical-wavemaker-
problem-in-a-liquid-of-finite-depth-with-an-inertial-surface-
in-the-presence-of-surface-tension/D2E5DF50BCDF96ECBEDB9CC35B9E1E4A.

Anonymous:1991:AVIe

[629] Anonymous. ANZ volume 33 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):f1–f2, July 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-33-issue-1-cover-and-front-matter/2D616C07A8F9EC0BA9937CB667E102D6.

Anonymous:1991:AVIf

[630] Anonymous. ANZ volume 33 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(1):b1–b2, July 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-33-issue-1-cover-and-back-matter/B8B5528A8F07EBBB528DC5E3EEFDF8BC.

Morris:1991:SCA

[631] Sidney A. Morris and Peter J. Nyikos. Sudden cardiac arrest and a problem in topology. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(2):123–132, October 1991. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/sudden-cardiac-arrest-
and-a-problem-in-topology/71E98C4360A9A6737896EE075833B386.

He:1991:HBI

[632] Xiangjian He. Hopf bifurcation at infinity with discontinuous nonlinearities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(2):133–148, October 1991. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/hopf-bifurcation-at-infinity-with-discontinuous-
nonlinearities/35337CC5677223180050B5EB7EB320A8.

Cannon:1991:DCP

[633] J. R. Cannon, Yanping Lin, and Shingmin Wang. Determination of a control parameter in a parabolic partial differential equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(2):149–163, October 1991. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/determination-of-a-control-
parameter-in-a-parabolic-partial-differential-equation/7C126ED3C30FAEB92EAE1995C3545F28.

Peng:1991:DCS

[634] Shige Peng and Jiongmin Yong. Determination of a controllable set for a controlled dynamic system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33 (2):164–179, October 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/determination-of-a-controllable-set-for-a-controlled-
dynamic-system/86215DDB78DB3391774FCCB75B9D93BA.

Mustard:1991:UPI

[635] David Mustard. Uncertainty principles invariant under the fractional Fourier transform. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(2):180–191, October 1991. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/uncertainty-principles-invariant-
under-the-fractional-fourier-transform/3D318CC1E70DF6646EF4C0E3BF5AC5A9.

Spencer:1991:SFP

[636] N. M. Spencer and V. V. Anh. Spectral factorisation and prediction of multivariate processes with time-dependent rational spectral density matrices. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(2):192–210, October 1991. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/spectral-factorisation-and-
prediction-of-multivariate-processes-with-timedependent-rational-
spectral-density-matrices/D6C3AAB5C091592F66792A4CB3EB636D.

Mazumdar:1991:MSF

[637] J. Mazumdar and R. N. Dubey. A method for the study of fully developed parallel flow in straight ducts of arbitrary cross section. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(2):211–239, October 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/method-for-the-study-of-fully-developed-parallel-flow-
in-straight-ducts-of-arbitrary-cross-section/E1D7DAD0B71633EA60BD4FE2F5809001.

Clements:1991:GMA

[638] David L. Clements and Ashley Larsson. Ground motion on alluvial valleys under incident plane SH waves. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33 (2):240–253, October 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/ground-motion-on-alluvial-valleys-under-incident-plane-
sh-waves/DB1E24FC08B4570FD6FCBE2FF6076729.

Anonymous:1991:AVIg

[639] Anonymous. ANZ volume 33 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(2):f1–f2, October 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-33-issue-2-cover-and-front-matter/0BBD5A6755E48EAF19B3489C95AFECBF.

Anonymous:1991:AVIh

[640] Anonymous. ANZ volume 33 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(2):b1–b2, October 1991. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-33-issue-2-cover-and-back-matter/D79EABD855D23F2B6DA695DFAFC1EC53.

Atkinson:1992:SBV

[641] C. Atkinson and C. R. Champion. Some boundary-value problems for nonlinear (N) diffusion and pseudo-plastic flow. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33 (3):255–268, January 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-boundaryvalue-problems-for-nonlinear-n-diffusion-
and-pseudoplastic-flow/13F3E18A5E4FCA9760BCCC003286F5CC.

Barton:1992:CSD

[642] N. G. Barton, C.-H. Li, and S. J. Spencer. Control of a surface of discontinuity in continuous thickness. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(3):269–289, January 1992. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/control-
of-a-surface-of-discontinuity-in-continuous-thickness/327494B8956A29D14465ED6BF250DABC.

Hill:1992:SST

[643] James M. Hill and Adrian H. Pincombe. Some similarity temperature profiles for the microwave heating of a half-space. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(3):290–320, January 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-similarity-temperature-profiles-for-the-microwave-
heating-of-a-halfspace/50DF6E4E29BF6D3324B4C22AEA33FB63.

King:1992:LTB

[644] J. R. King. Local transformations between some nonlinear diffusion equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(3):321–349, January 1992. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/local-transformations-
between-some-nonlinear-diffusion-equations/BB89FF016B953D7F9C33727179F1F646.

Lacey:1992:CIC

[645] A. A. Lacey and G. C. Wake. Critical initial conditions for spatiallydistributed thermal explosions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(3):350–362, January 1992. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/critical-initial-conditions-
for-spatiallydistributed-thermal-explosions/5C0954A85348E6AAB4E056428BF6599E.

Philip:1992:ESR

[646] J. R. Philip. Exact solutions for redistribution by nonlinear convection– diffusion. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(3):363–383, January 1992. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/exact-solutions-for-redistribution-by-
nonlinear-convectiondiffusion/5799164638E739563F658D9A1F07498E.

Sander:1992:ESN

[647] G. C. Sander. Exact solutions to nonlinear diffusion-convection problems on finite domains. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(3):384–401, January 1992. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/exact-solutions-to-nonlinear-
diffusionconvection-problems-on-finite-domains/C8DDCB6512A6AA819F95D061FA01CD17.

Anonymous:1992:AVIa

[648] Anonymous. ANZ volume 33 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(3):f1–f2, January 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-33-issue-3-cover-and-front-matter/5336BC7FB4E72BDC1F952CD1E83FDB9B.

Anonymous:1992:AVIb

[649] Anonymous. ANZ volume 33 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(3):b1–b2, January 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-33-issue-3-cover-and-back-matter/8F42AA29F78F2A77A03F3874E4B30786.

Smyth:1992:ECH

[650] N. F. Smyth. The effect of conductivity on hotspots. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):403–413, April 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/effect-of-conductivity-on-hotspots/AB37E0707C91E1160962DB411FE8E971.

Grundy:1992:AEN

[651] R. E. Grundy. The asymptotics of extinction in nonlinear diffusion reaction equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):414–429, April 1992. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/asymptotics-of-extinction-in-nonlinear-
diffusion-reaction-equations/229F93D7436691699578175DB32CFAE7.

Broadbridge:1992:ENS

[652] P. Broadbridge and P. J. Banksa. Exact nonlinear solution for constant-rate expression from material of finite thickness. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):430–450, April 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/exact-nonlinear-solution-for-constantrate-expression-
from-material-of-finite-thickness/247C17F4FE19F47328B20B4C83ECE553.

Wicks:1992:IBP

[653] P. J. Wicks. Interaction of buoyant plumes in open-channel flow. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):451–473, April 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/interaction-of-buoyant-plumes-in-openchannel-flow/B7DD508CCEBD245CB90457BCD10F405A.

Bera:1992:CIS

[654] R. K. Bera and A. Chakrabarti. Cooling of an infinite slab in a two-fluid medium. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):474–485, April 1992. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/cooling-
of-an-infinite-slab-in-a-twofluid-medium/D57C0C34E7F4C912E14DF8D438E90087.

Thompson:1992:NFE

[655] Gerard Thompson. Normal forms for elements of o(p, q) and Hamiltonians with integrals linear in momenta. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33 (4):486–507, April 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/normal-forms-for-elements-of-op-q-and-hamiltonians-
with-integrals-linear-in-momenta/4010040621C5A52B469AED2AEA8D989D.

Chen:1992:EPS

[656] Yongshao Chen. The existence of periodic solutions for a class of neutral differential difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):508–516, April 1992. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/existence-of-periodic-

solutions-for-a-class-of-neutral-differential-difference-equations/
203134870600998F25AF19CB8D7E7BFC.

Teo:1992:NCO

[657] K. L. Teo and K. H. Wong. Nonlinearly constrained optimal control problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):517–530, April 1992. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/nonlinearly-constrained-optimal-control-
problems/E824575888C212D169EFE778AA6CD462.

Egudo:1992:DCQ

[658] R. R. Egudo, T. Weir, and B. Mond. Duality without constraint qualification for multiobjective programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):531–544, April 1992. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/duality-
without-constraint-qualification-for-multiobjective-programming/
2A41FE8669D7C189A88D88BE8F32CFCA.

Anonymous:1992:CV

[659] Anonymous. Contents of volume 33. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):545–546, April 1992. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
33/B097F16E0628E2E9A4A234DDC3909356.

Anonymous:1992:AVIc

[660] Anonymous. ANZ volume 33 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):f1–f3, April 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-33-issue-4-cover-and-front-matter/08E5C88F62095AF63827A651C6C4D6FE.

Anonymous:1992:AVId

[661] Anonymous. ANZ volume 33 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 33(4):b1–b2, April 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-33-issue-4-cover-and-back-matter/441878FB11BFDDFCEE0E69D856E4A0FA.

Yu:1992:CCV

[662] Xinghuo Yu and Renfrey B. Potts. Computer-controlled variablestructure systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):1–17, July 1992. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/computercontrolled-variablestructure-
systems/9ABEF1CC4CB7D2F72D32E4827240316E.

Balachandran:1992:CNV

[663] K. Balachandran. Controllability of neutral Volterra integrodifferential systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):18–25, July 1992. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/controllability-of-neutral-volterra-
integrodifferential-systems/42F9A73A9EA57884D1EB15EFA0BA5161.

Lin:1992:GSD

[664] Xiaodong Lin. On the global stability of a delay epidemic model. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):26–34, July 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-the-global-stability-of-a-delay-epidemic-model/07E927FAD1529445AB033E0A62CE6552.

Halpern:1992:EPC

[665] P. H. Halpern, R. N. Mohapatra, P. J. O’Hara, and R. S. Rodriguez. An extremal problem concerning finite dimensional subspaces of C[a, b] pertinent in signal theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):35–42, July 1992. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
extremal-problem-concerning-finite-dimensional-subspaces-of-
ca-b-pertinent-in-signal-theory/F7D73E0516AB7053A76FF3BBDE6CA516.

Jeyakumar:1992:GCM

[666] V. Jeyakumar and B. Mond. On generalised convex mathematical programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):43–53, July 1992. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/on-generalised-convex-
mathematical-programming/A864783D512F49DBC13891230FE445A1.

Roberts:1992:BCA

[667] A. J. Roberts. Boundary conditions for approximate differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):54–80, July 1992. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/boundary-conditions-for-approximate-
differential-equations/572D1F1F3D5DE009BC99A0BAFED90D65.

Chapman:1992:TTR

[668] P. B. Chapman. The transition through resonance of a nonlinear nonautonomous system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):81–111, July 1992. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/transition-through-resonance-
of-a-nonlinear-nonautonomous-system/9AF62419D334875A425BF340BD619686.

Jank:1992:ADS

[669] Gerhard Jank. Asymptotic distribution of singularities of solutions of matrix-Riccati differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34 (1):112–131, July 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-distribution-of-singularities-of-solutions-
of-matrixriccati-differential-equations/0FC969E366D95084A53A42FC0A7BDA82.

Anonymous:1992:AVIe

[670] Anonymous. ANZ volume 34 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):f1–f2, July 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-34-issue-1-cover-and-front-matter/A69A2250812EC607A71E85298F422832.

Anonymous:1992:AVIf

[671] Anonymous. ANZ volume 34 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(1):b1–b2, July 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-34-issue-1-cover-and-back-matter/A0CCD727629E560CC59ABB53017925C7.

Rickard:1992:WPM

[672] John Rickard, Allen Russell, and Christine Martini. Welfare policy and multi-national monopolies. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(2):133–152, October 1992. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/welfare-
policy-and-multinational-monopolies/E46B9F77F97DC64A392E2C7294EF5DA7.

Bryant:1992:CBP

[673] Peter J. Bryant. Chaotic breakdown of a periodically forced, weakly damped pendulum. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(2):153–173, October 1992. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/chaotic-breakdown-
of-a-periodically-forced-weakly-damped-pendulum/535128DF20069D9525D5303D7AC9E9F5.

Roberts:1992:PEC

[674] A. J. Roberts. Planform evolution in convection — an embedded centre manifold. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(2):174–198, October 1992. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/planform-evolution-in-
convectionan-embedded-centre-manifold/3F516CDBA06D4EB10C45452F529B407E.

Sloan:1992:CCP

[675] I. H. Sloan and E. P. Stephan. Collocation with Chebyshev polynomials for Symm’s integral equation on an interval. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(2):199–211, October 1992. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/collocation-with-chebyshev-polynomials-for-
symms-integral-equation-on-an-interval/4E1EF11E1C7F5280CDB46DC81DD39B5B.

Stewart:1992:NAO

[676] David E. Stewart. A numerical algorithm for optimal control problems with switching costs. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(2):212–228, October 1992. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/numerical-algorithm-for-optimal-
control-problems-with-switching-costs/8018728FF71499A9F1DEB2364FF471F9.

Suneja:1992:ODC

[677] S. Suneja, C. Singh, and R. N. Kaul. Optimality and duality in continuous-time nonlinear fractional programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34 (2):229–244, October 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimality-and-duality-in-continuoustime-nonlinear-
fractional-programming/B02EAACE6D72F0CE8C9797D9C4130FC1.

Lalli:1992:OCT

[678] B. S. Lalli and B. G. Zhang. Oscillation and comparison theorems for certain neutral difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34 (2):245–256, October 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/oscillation-and-comparison-theorems-for-certain-neutral-
difference-equations/2E573078D8AEA8538BB799B28F7ADEE6.

Anonymous:1992:AVIg

[679] Anonymous. ANZ volume 34 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(2):f1–f2, October 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-34-issue-2-cover-and-front-matter/31C1C633F6A87C1D54FB44CC2C6F58C9.

Anonymous:1992:AVIh

[680] Anonymous. ANZ volume 34 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(2):b1–b2, October 1992. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-34-issue-2-cover-and-back-matter/467B06000429426442C17B466F2D591A.

Fink:1993:LAI

[681] A. M. Fink and L. Bass. The likely antecedents of improbable events: optimal search strategies. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):257–273, January 1993. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/likely-antecedents-of-
improbable-events-optimal-search-strategies/70332921F622F364F97BE07703685F64.

Watts:1993:VPD

[682] A. M. Watts. Variational principles, duality, Legendre transformations and mine shaft ventilation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):274–281, January 1993. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/variational-principles-
duality-legendre-transformations-and-mine-shaft-ventilation/
7FBB13CC7F482C492074D676301F4DF6.

Lin:1993:GSE

[683] Xiaodong Lin and Joseph W.-H. So. Global stability of the endemic equilibrium and uniform persistence in epidemic models with subpopulations.

Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):282–295, January 1993. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/global-stability-of-the-endemic-equilibrium-
and-uniform-persistence-in-epidemic-models-with-subpopulations/
FC91ADDFEA08F27560779ACAAAD61C8A.

Hart:1993:JDO

[684] V. G. Hart and Jingyu Shi. Joined dissimilar orthotropic elastic cylindrical membranes under internal pressure and longitudinal tension. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):296–317, January 1993. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/joined-dissimilar-
orthotropic-elastic-cylindrical-membranes-under-internal-pressure-
and-longitudinal-tension/CAC5D3642C01854262E57FF101BE06C0.

Broadbridge:1993:NRD

[685] Philip Broadbridge and Colin Rogers. On a nonlinear reaction–diffusion boundary-value problem: application of a Lie–Bäcklund symmetry. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):318–332, January 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-a-nonlinear-reactiondiffusion-boundaryvalue-problem-
application-of-a-liebacklund-symmetry/97E753884B1DC03F66F081ECDF0159E5.

Barry:1993:RFT

[686] S. I. Barry and G. K. Aldis. Radial flow through deformable porous shells. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):333–354, January 1993. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/radial-flow-through-deformable-porous-shells/
6D8C8A94495B6E6D4334B9ED961CD030.

Weir:1993:LEA

[687] G. J. Weir. Linearised evaporation about a shallow half-plane pond. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):355–367, January 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/linearised-evaporation-about-a-shallow-halfplane-pond/
60040B2209C3C5032C72562A4DD45221.

Mekias:1993:FSF

[688] Hocine Mekias and Jean-Marc Vanden-Broeck. Free-surface flow due to a source submerged in a fluid of infinite depth with two stagnant regions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):368–376, January 1993. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/freesurface-flow-due-to-a-source-submerged-
in-a-fluid-of-infinite-depth-with-two-stagnant-regions/32FC403360147E28FCC6BA0BACF61DF6.

Forbes:1993:FIL

[689] Lawrence K. Forbes and Graeme C. Hocking. Flow induced by a line sink in a quiescent fluid with surface-tension effects. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):377–391, January 1993. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/flow-induced-by-a-line-sink-in-a-quiescent-
fluid-with-surfacetension-effects/3814791FCA5BEB92B3254A2FB2FF30AC.

Anonymous:1993:AVIa

[690] Anonymous. ANZ volume 34 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):f1–f2, January 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-34-issue-3-cover-and-front-matter/CFDED4E63807F67188D0FFADE5074F95.

Anonymous:1993:AVIb

[691] Anonymous. ANZ volume 34 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(3):b1–b2, January 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-34-issue-3-cover-and-back-matter/F76F01235547CBB4E7CC2F94CFFBA80C.

Vanden-Broeck:1993:TDJ

[692] Jean-Marc Vanden-Broeck. Two-dimensional jet aimed vertically upwards. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):393–400, April 1993. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/twodimensional-jet-aimed-vertically-
upwards/FC3E4C490893697912E8858100787C1B.

Prossdorf:1993:DML

[693] S. Prössdorf, J. Saranen, and I. H. Sloan. A discrete method for the logarithmic-kernel integral equation on an open arc. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):401–418, April 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/discrete-method-for-the-logarithmickernel-integral-
equation-on-an-open-arc/AA62BEA2640A15BEEE1045043E3EEA0C.

Ruotsalainen:1993:RBE

[694] Keijo Ruotsalainen. Remarks on the boundary element method for strongly nonlinear problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):419–438, April 1993. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/remarks-
on-the-boundary-element-method-for-strongly-nonlinear-problems/
644367C75B52879535F58005B4C6B4E3.

Leipnik:1993:PDE

[695] Roy B. Leipnik. Partial differential equations for eigenvalues: sensitivity and perturbation analysis. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):439–470, April 1993. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/partial-
differential-equations-for-eigenvalues-sensitivity-and-perturbation-
analysis/BEC5301EB452678D361D0EDD2E4C2922.

Kuang:1993:CLV

[696] Yang Kuang and Hal L. Smith. Convergence in Lotka–Volterra type diffusive delay systems without dominating instantaneous negative feedbacks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):471–494, April 1993. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/convergence-in-lotkavolterra-type-diffusive-
delay-systems-without-dominating-instantaneous-negative-feedbacks/
D0E9816D202C04D46695B9BF4C106479.

Helmke:1993:IFL

[697] U. Helmke. Isospectral flows and linear programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):495–510, April 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/isospectral-flows-and-linear-programming/49CA891AFC1301B2D4A06B365350FA9E.

Gorringe:1993:FIT

[698] V. M. Gorringe and P. G. L. Leach. The first integrals and their Lie algebra of the most general autonomous Hamiltonian of the form H = T + V possessing a Laplace–Runge–Lenz vector. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):511–522, April 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/first-integrals-and-their-lie-algebra-of-the-most-general-
autonomous-hamiltonian-of-the-form-h-t-v-possessing-a-laplacerungelenz-
vector/ACAB2C23E0A73AD09FF8CF24E1E0992C.

Anonymous:1993:CV

[699] Anonymous. Contents of volume 34. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):523–524, April 1993. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
34/9F7DEA1E8E2253ADA2829D618583CE20.

Anonymous:1993:AVIc

[700] Anonymous. ANZ volume 34 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):f1–f2, April 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-34-issue-4-cover-and-front-matter/B3D697CE94EC758D7EF1B28962F526E1.

Anonymous:1993:AVId

[701] Anonymous. ANZ volume 34 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 34(4):b1–b2, April 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-34-issue-4-cover-and-back-matter/E4CEE7E2BBA58F5D6F40062938E570C8.

Calvert:1993:BPP

[702] Bruce Calvert and Grant Keady. Braess’s paradox and powerlaw nonlinearities in networks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(1):1–22, July 1993. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/braesss-
paradox-and-powerlaw-nonlinearities-in-networks/E797B9F96D467A0047286DDA9DD0ABCC.

McLean:1993:NSE

[703] W. McLean and V. Thomée. Numerical solution of an evolution equation with a positive-type memory term. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(1):23–70, July 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-solution-of-an-evolution-equation-with-a-
positivetype-memory-term/1753922379E643F136A1874AC0181EF7.

Butt:1993:OSD

[704] R. Butt. Optimal shape design for a nozzle problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(1):71– 86, July 1993. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
shape-design-for-a-nozzle-problem/DC70E8F0F8537902CE78D6C95AAF337F.

Pani:1993:FEM

[705] Amiya K. Pani. A finite element method for a diffusion equation with constrained energy and nonlinear boundary conditions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(1):87–102, July 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/finite-element-method-for-a-diffusion-equation-with-
constrained-energy-and-nonlinear-boundary-conditions/5064329213EBDCC2F1BBC2FC08564963.

Attili:1993:NCS

[706] B. S. Attili. Numerical computation of symmetry-breaking bifurcation points. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(1):103–113, July 1993. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/numerical-computation-
of-symmetrybreaking-bifurcation-points/C96059E3FA07530F2532DD8A73EBF7B8.

Hocking:1993:BFS

[707] G. C. Hocking. Bow flows with smooth separation in water of finite depth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(1):114–126, July 1993. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/bow-flows-with-smooth-separation-in-
water-of-finite-depth/90E6B29E972BB5E2701572A0C53A236B.

Anonymous:1993:AVIe

[708] Anonymous. ANZ volume 35 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(1):f1–f2, July 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-35-issue-1-cover-and-front-matter/09562C97526A26BFE5CAB449139E2393.

Anonymous:1993:AVIf

[709] Anonymous. ANZ volume 35 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(1):b1–b2, July 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-35-issue-1-cover-and-back-matter/8B1BB161158A01EC83675D85183586F6.

Tuck:1993:SAS

[710] E. O. Tuck. Some accurate solutions of the lifting surface integral equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(2):127–144, October 1993. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/some-accurate-solutions-of-
the-lifting-surface-integral-equation/FBDC56A1E7525527499DD01DFAAB8552.

Forbes:1993:ODP

[711] Lawrence K. Forbes. One-dimensional pattern formation in a model of burning. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(2):145–173, October 1993. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/onedimensional-pattern-formation-in-
a-model-of-burning/F60C9D986C99B3F57CCAE2B3F45B347C.

McCarthy:1993:GWC

[712] J. F. McCarthy. Gas and water cresting towards horizontal wells. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(2):174–197, October 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/gas-and-water-cresting-towards-horizontal-wells/E6031E6314B16A0B5CC895B545CF9C88.

Chandra:1993:SDM

[713] S. Chandra and M. V. Durga Prasad. Symmetric duality in multiobjective programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(2):198–206, October 1993. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/symmetric-duality-
in-multiobjective-programming/BA1ABD46B6A4E6BA9B34077F504CB28E.

McCarthy:1993:IMW

[714] J. F. McCarthy. Improved model of water cresting. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35 (2):207–222, October 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/improved-model-of-water-cresting/F6AAE4FD09C788E273B12804F01F5AE7.

Tang:1993:NSD

[715] S. Tang, S. Qin, and R. O. Weber. Numerical studies on 2-dimensional reaction–diffusion equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(2):223–243, October 1993. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/numerical-studies-
on-2dimensional-reactiondiffusion-equations/4E9DCE0CA7AC6107BAFDA1A48CBDE2C1.

Roberts:1993:LEA

[716] Stephen Roberts. A line element algorithm for curve flow problems in the plane. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(2):244–261, October 1993. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/line-element-algorithm-for-
curve-flow-problems-in-the-plane/FCEB1281DD0332FC8372DBDFAAFA4A3A.

Anonymous:1993:AVIg

[717] Anonymous. ANZ volume 35 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(2):f1–f2, October 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-35-issue-2-cover-and-front-matter/CBED36D17A2C442A68B50989371232E4.

Anonymous:1993:AVIh

[718] Anonymous. ANZ volume 35 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(2):b1–b2, October 1993. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-35-issue-2-cover-and-back-matter/B610721544CDADCA6F95E352036C7434.

Wu:1994:NFE

[719] Yong-Hong Wu, James M. Hill, and Paul J. Flint. A novel finite element method for heat transfer in the continuous caster. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):263–288, January 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/novel-finite-element-method-for-heat-transfer-in-the-
continuous-caster/896BEA24180DE13D30245EA6C436F724.

McNabb:1994:DTP

[720] Alex McNabb and Grant Keady. Diffusion and the torsion parameter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):289–301, January 1994. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/diffusion-and-the-torsion-parameter/
BDFDEFDBFF68E3560D743C5D19031F82.

Rhodes-Robinson:1994:WMT

[721] P. F. Rhodes-Robinson. On wave motion in a two-layered liquid of infinite depth in the presence of surface and interfacial tension. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):302–322, January 1994. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/on-wave-motion-in-
a-twolayered-liquid-of-infinite-depth-in-the-presence-of-surface-
and-interfacial-tension/84B2078EB2FFE6ED87F4542D9E126125.

Khosla:1994:SHR

[722] H. K. Khosla and R. K. Chhabra. Second harmonic resonance in magnetohydrodynamic jet. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):323–334, January 1994. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/second-harmonic-resonance-in-
magnetohydrodynamic-jet/9448E341780F5FC47D9C142B499437BF.

Palaniappan:1994:TFS

[723] D. Palaniappan, S. D. Nigam, and T. Amaranath. A theorem for a fluid Stokes flow. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):335–347, January 1994. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/theorem-for-a-fluid-stokes-
flow/2C4C5B87E3DB0FF9AE9D0F4FE85E07E5.

Dhar:1994:SAF

[724] A. K. Dhar and K. P. Das. Stability analysis from fourth order evolution equation for small but finite amplitude interfacial waves in the presence of a basic current shear. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35 (3):348–365, January 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stability-analysis-from-fourth-order-evolution-equation-

for-small-but-finite-amplitude-interfacial-waves-in-the-presence-
of-a-basic-current-shear/BE5833BB641F96D3AAC4AD1F448F7A59.

Chandra:1994:PFC

[725] Peeyush Chandra and J. S. V. R. Krishna Prasad. Pulsatile flow in circular tubes of varying cross-section with suction/injection. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):366–381, January 1994. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/pulsatile-flow-in-circular-tubes-of-
varying-crosssection-with-suctioninjection/29005A8DBF65CD888DF3B2F5836C143A.

Bharathi:1994:SPS

[726] L. Vijaya Bharathi, A. Chakrabarti, B. N. Mandal, and S. Banerjee. Solution of the problem of scattering of water waves by a nearly vertical plate. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):382–395, January 1994. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/solution-of-the-problem-of-
scattering-of-water-waves-by-a-nearly-vertical-plate/96B48D8C6CE14D409BDA57524CFE77E0.

Graham:1994:PCC

[727] I. G. Graham and Y. Yan. Piecewise constant collocation for first-kind boundary integral equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):396–398, January 1994. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/piecewise-constant-collocation-
for-firstkind-boundary-integral-equations/31A18F3D2E93D3242A7240E2BB5FB274.

Anonymous:1994:AVIa

[728] Anonymous. ANZ volume 35 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):f1–f2, January 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-35-issue-3-cover-and-front-matter/D32EF75851BD31E5B38F2EC772F8F8B4.

Anonymous:1994:AVIb

[729] Anonymous. ANZ volume 35 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(3):b1–b2, January 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-35-issue-3-cover-and-back-matter/C19CC3B51A15024276EC3ED177B2E6C1.

Latham:1994:CAC

[730] Geoff A. Latham. The computational approach to commuting ordinary differential operators of orders six and nine. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35 (4):399–419, April 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/computational-approach-to-commuting-ordinary-differential-
operators-of-orders-six-and-nine/309DA9CD24D31232307615166730D932.

Yao:1994:GCP

[731] Jen-Chih Yao. On the generalized complementarity problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):420–428, April 1994. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-the-generalized-complementarity-problem/
6405031F04432CE647E8704A7307FD5B.

Preda:1994:NPM

[732] Vasile Preda. On nonlinear programming and matrix game equivalence. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):429–438, April 1994. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-nonlinear-programming-and-matrix-
game-equivalence/641B2DED13E7A07340878006506D853C.

Chaudhry:1994:ILB

[733] M. Aslam Chaudhry. On an integral of Lommel and Bessel functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):439–444, April 1994. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-an-integral-of-lommel-and-bessel-
functions/5CFB7B68A25F416B02A900A2A98C3124.

Chapman:1994:PNA

[734] P. B. Chapman. Perturbations of nonlinear autonomous oscillators. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):445–468, April 1994. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/perturbations-of-nonlinear-
autonomous-oscillators/BE3747FF6A15CF856C97A78F9A501B3A.

Chaudhry:1994:FLE

[735] M. Aslam Chaudhry. On a family of logarithmic and exponential integrals occurring in probability and reliability theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):469–478, April 1994. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-a-family-of-logarithmic-and-exponential-
integrals-occurring-in-probability-and-reliability-theory/413059A8EFCE2FAD3DAE9798423811F5.

Siew:1994:NSE

[736] P. F. Siew. A numerical scheme for the electromagnetic response in thin conductors of arbitrary planar shape. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35 (4):479–497, April 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-scheme-for-the-electromagnetic-response-in-
thin-conductors-of-arbitrary-planar-shape/32B45DEFFBC364A436B87EAF4592BBFE.

Bandopadhyaya:1994:CTT

[737] Lakshmisree Bandopadhyaya. Cost-time trade-off in three-axial sums’ transportation problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):498–505, April 1994. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/costtime-tradeoff-
in-threeaxial-sums-transportation-problem/06624B0604BE724735E124C60109FFF7.

Bich:1994:BVP

[738] Dao Huy Bich. A boundary value problem of elastoplastic deformation process theory: Existence and uniqueness theorems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):506–524, April 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/boundary-value-problem-of-elastoplastic-deformation-
process-theory-existence-and-uniqueness-theorems/2D4EC010C0B3F5EB30BE55D314C6181B.

Anonymous:1994:AVIc

[739] Anonymous. ANZ volume 35 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):f1–f2, April 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-35-issue-4-cover-and-front-matter/532F2A2C32E4EA73607ADB6EA760FC19.

Anonymous:1994:AVId

[740] Anonymous. ANZ volume 35 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 35(4):b1–b2, April 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-35-issue-4-cover-and-back-matter/65D83B33D434A3A496E163161930979D.

Grimshaw:1994:ADE

[741] R. Grimshaw and J. Gan. Analysis of a differential equation occurring in the theory of flame fronts. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):1–16, July 1994. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/analysis-of-a-differential-
equation-occurring-in-the-theory-of-flame-fronts/EF341F7CFD411F304C97462A4A234A2A.

Deakin:1994:LTS

[742] Michael A. B. Deakin. Laplace transforms for superexponential functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):17–25, July 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/laplace-transforms-for-superexponential-functions/647F8DEC809111E951761ABF637C1A83.

Arthurs:1994:SNP

[743] A. M. Arthurs and J. Clegg. On the solution of a nonlinear problem in cell membrane theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):26–37, July 1994. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-solution-of-a-nonlinear-problem-in-cell-membrane-theory/
710E823773C156BD615328692DD46C5D.

Pudney:1994:ODS

[744] Peter Pudney and Phil Howlett. Optimal driving strategies for a train journey with speed limits. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):38–49, July 1994. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/optimal-driving-strategies-
for-a-train-journey-with-speed-limits/C7B9AB6FF47A8F1B70FEF511A23971CB.

Craven:1994:CDA

[745] B. D. Craven. Convergence of discrete approximations for constrained minimization. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):50–59, July 1994. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/convergence-of-discrete-approximations-
for-constrained-minimization/EAE8420AEE3090097DBC3688BA541F27.

Pearce:1994:TLB

[746] C. E. M. Pearce and J. E. Pecarić. On two lemmas of Brown and Shepp having application to sum sets and fractals. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1): 60–63, July 1994. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
two-lemmas-of-brown-and-shepp-having-application-to-sum-sets-
and-fractals/D1AEE14C21DC08F602E24F8185570750.

Zhu:1994:TDN

[747] Songping Zhu and Jörg Imberger. A three-dimensional numerical model of the response of the Australian North West Shelf to tropical cyclones. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):64–100, July 1994. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/threedimensional-numerical-model-of-
the-response-of-the-australian-north-west-shelf-to-tropical-
cyclones/154C7AB11443A8EDFFDA0F2300BEDFE2.

Veron:1994:TSB

[748] M. A. Hernandez Veron. The Taylor series for bandlimited signals. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):101–106, July 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/taylor-series-for-bandlimited-signals/79A682437322C7D067A3B55F560057F2.

Balachandran:1994:CNN

[749] K. Balachandran and P. Balasubramaniam. Controllability of nonlinear neutral Volterra integrodifferential systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):107–116, July 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/controllability-of-nonlinear-neutral-volterra-integrodifferential-
systems/F7D68BB580D85E80FBB1811BA19D9936.

Idczak:1994:OPC

[750] D. Idczak, K. Kibalczyc, and S. Walczak. On an optimization problem with cost of rapid variation of control. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):117–131, July 1994. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
an-optimization-problem-with-cost-of-rapid-variation-of-control/
29CCD0B0B442EBD2209A4DD77D67BDA6.

Brown:1994:PAT

[751] T. C. Brown and P. K. Pollett. Poisson approximations for telecommunications networks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):132, July 1994. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/poisson-approximations-for-
telecommunications-networks/1CB5F9690850737FD1424EB2EE2378D1.

Anonymous:1994:AVIe

[752] Anonymous. ANZ volume 36 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):f1–f2, July 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-36-issue-1-cover-and-front-matter/61427C673E881ADF432115476C50EFD0.

Anonymous:1994:AVIf

[753] Anonymous. ANZ volume 36 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(1):b1–b2, July 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-36-issue-1-cover-and-back-matter/1BF0FBF0F9040C31C75FE5AAA173ADDD.

Forbes:1994:FFA

[754] Lawrence K. Forbes, Anthony M. Watts, and Graeme A. Chandler. Flow fields associated with in situ mineral leaching. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(2):133–151, October 1994. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/flow-fields-associated-with-in-situ-
mineral-leaching/65329A58A9FD77EFE6FA94F197E6E314.

Chapman:1994:ALL

[755] P. B. Chapman. Aspects of laser Lorenz dynamics. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (2):152–174, October 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/aspects-of-laser-lorenz-dynamics/F691BFF3612F207E95DF42C192414790.

Forte:1994:CPA

[756] B. Forte, M. Lo Schiavo, and E. R. Vrscay. Continuity properties of attractors for iterated fuzzy set systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (2):175–193, October 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/continuity-properties-of-attractors-for-iterated-fuzzy-
set-systems/65F2BFD7CC25A139BD63E22E53AB6128.

He:1994:MPT

[757] Feiyue He. The maximum principle for a type of hereditary semilinear differential equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(2):194–212, October 1994. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/maximum-
principle-for-a-type-of-hereditary-semilinear-differential-
equation/4F4FAF5002CD847CBD2B944103FE683D.

Cao:1994:PEH

[758] Jia-Ding Cao and Heinz H. Gonska. Pointwise estimates for higher order convexity preserving polynomial approximation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (2):213–233, October 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/pointwise-estimates-for-higher-order-convexity-preserving-
polynomial-approximation/187F32D9E343EA0DD2E5DC66C74CEDE9.

Li:1994:CSA

[759] Nian Li, Joseph Steiner, and Shimin Tang. Convergence and stability analysis of an explicit finite difference method for 2-dimensional reaction– diffusion equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(2):234–241, October 1994. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/convergence-and-stability-analysis-
of-an-explicit-finite-difference-method-for-2dimensional-reactiondiffusion-
equations/896BF4CD0CA94C3CAC1EEEEA8B82EF9A.

George:1994:CDP

[760] Santhosh George and M. Thamban Nair. A class of discrepancy principles for the simplified regularization of ill-posed problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(2):242–248, October 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/class-of-discrepancy-principles-for-the-simplified-
regularization-of-illposed-problems/D0FE21CABE0F95627D0B25ADE10669E6.

Huang:1994:EPS

[761] Yin Xi Huang. Existence of positive solutions for a class of the p-Laplace equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(2):249–264, October 1994. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/existence-of-positive-solutions-for-
a-class-of-the-plaplace-equations/0081B7892A530548F459C0B6716AC7EB.

Anonymous:1994:AVIg

[762] Anonymous. ANZ volume 36 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(2):f1–f2, October 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-36-issue-2-cover-and-front-matter/7E9E39E4B35B2EC9F9F2BE6DC6DBAAF8.

Anonymous:1994:AVIh

[763] Anonymous. ANZ volume 36 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(2):b1–b2, October 1994. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-36-issue-2-cover-and-back-matter/071886D20501367E7B93FAB428DF3157.

Brooker:1995:CSS

[764] P. I. Brooker and M. A. Stewart. A comparative study of simulation techniques for two dimensional data honouring specified exponential semivariograms. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(3):249–260, January 1995. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/comparative-study-
of-simulation-techniques-for-two-dimensional-data-honouring-
specified-exponential-semivariograms/E8100378CC8ACC7951A6FD4B806AF166.

Kazemi:1995:GTO

[765] Mohammad A. Kazemi. A gradient technique for an optimal control problem governed by a system of nonlinear first order partial differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(3):261–273, January 1995. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/gradient-technique-for-an-optimal-control-

problem-governed-by-a-system-of-nonlinear-first-order-partial-
differential-equations/0CBB870A62677775D62E90BB110C71F2.

Yang:1995:SAN

[766] X. Q. Yang. Smoothing approximations to nonsmooth optimization problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(3):274–285, January 1995. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/smoothing-approximations-to-nonsmooth-
optimization-problems/875E89FF2F2EED9B086051E434C323A8.

Chukwu:1995:CNI

[767] E. N. Chukwu. Control in W_{2}^{(1)} of nonlinear interconnected systems of neutral type. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(3):286–312, January 1995. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/control-
in-w21-of-nonlinear-interconnected-systems-of-neutral-type/
6ACE7BAE2786E02E128C62623E577DD6.

Bu:1995:CPC

[768] Charles Bu. On the Cauchy problem for the 1+2 complex Ginzburg– Landau equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(3):313–324, January 1995. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/on-the-cauchy-problem-
for-the-12-complex-ginzburglandau-equation/0AE928409B20E4C5C9CF146669D63F02.

Barber:1995:TWP

[769] Michael N. Barber and David Singleton. Travelling waves in phase field models of solidification. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(3):325–371, January 1995. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/travelling-waves-in-phase-field-
models-of-solidification/7A5DF8D7DB14C07CC309319E2DFE032E.

Mandal:1995:SWW

[770] B. N. Mandal and Sudeshna Banerjea. Scattering of water waves by a submerged nearly circular cylinder. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (3):372–380, January 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/

article/scattering-of-water-waves-by-a-submerged-nearly-circular-
cylinder/F54315E21F48F8B9EFFBA4B60FAC9C4F.

Anonymous:1995:AVIa

[771] Anonymous. ANZ volume 36 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(3):f1–f2, January 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-36-issue-3-cover-and-front-matter/5A96F3BBF0DF78F0C021D44CB202728E.

Anonymous:1995:AVIb

[772] Anonymous. ANZ volume 36 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(3):b1–b2, January 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-36-issue-3-cover-and-back-matter/C674A65A254EFF6A88B070FD90E1C31D.

Lenard:1995:RPL

[773] Christopher T. Lenard. Rank-1 perturbations and the Lanczos method, inverse iteration, and Krylov subspaces. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (4):381–388, April 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/rank1-perturbations-and-the-lanczos-method-inverse-
iteration-and-krylov-subspaces/16AE94BFA5A1CFFC32DBB94942B8920A.

Nonnenmacher:1995:NFJ

[774] D. J. F. Nonnenmacher and R. Zagst. A new form of Jensen’s inequality and its application to statistical experiments. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(4):389–398, April 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/new-form-of-jensens-inequality-and-its-application-
to-statistical-experiments/F0B95E607E49A1D5E9D878EC5687E2BB.

Park:1995:DRB

[775] Yunbeom Park and U Jin Choi. Degree reduction of Bézier curves and its error analysis. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(4):399–413, April 1995. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/degree-reduction-of-bezier-
curves-and-its-error-analysis/61992BF695CBD49BB26A48857FE294EA.

Wong:1995:CMC

[776] K. H. Wong, N. Lock, and K. Kaji. A computational method for a class of jump linear quadratic systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (4):414–423, April 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/computational-method-for-a-class-of-jump-linear-quadratic-
systems/330EE2773119540990F043D19ADD3355.

Farrow:1995:FSS

[777] D. E. Farrow and E. O. Tuck. Further studies of stern wavemaking. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(4):424–437, April 1995. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/further-studies-of-stern-wavemaking/
A2B9DD18E478291C4BC5D568AC4C5145.

Philip:1995:FDL

[778] J. R. Philip. Fast diffusion with loss at infinity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (4):438–459, April 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fast-diffusion-with-loss-at-infinity/8C9B79177038523C6075EEF5BDEA0527.

Xiao:1995:NIN

[779] Yi Xiao and Eric King-Wah Chu. A nonmonotone inexact Newton algorithm for nonlinear systems of equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (4):460–492, April 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonmonotone-inexact-newton-algorithm-for-nonlinear-
systems-of-equations/271038059EA05C2B9A533B273D4066B6.

Valvi:1995:HPT

[780] F. N. Valvi. The Hadamard product of two Brownian matrices: Analytic inverse and determinant. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36 (4):493–497, April 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/hadamard-product-of-two-brownian-matrices-analytic-
inverse-and-determinant/533BFBA1CB5772E97A3C0C6E97E90913.

Anonymous:1995:AVIc

[781] Anonymous. ANZ volume 36 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(4):f1–f2, April 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-36-issue-4-cover-and-front-matter/5BAFA6AD510002AA20C74464327C88CA.

Anonymous:1995:AVId

[782] Anonymous. ANZ volume 36 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 36(4):b1–b2, April 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-36-issue-4-cover-and-back-matter/FB9E654753FC3E1CEE6955C1451D8593.

Whale:1995:MIM

[783] J. Whale, N. Fowkes, G. Hocking, and D. Hill. A model of the injection moulding process. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1):1–15, July 1995. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/model-of-the-injection-moulding-process/
C8717198B5A80D41B42C83F2AD3E737A.

Zheng:1995:NCH

[784] Shiming Zheng and Desmond Robbie. A note on the convergence of Halley’s method for solving operator equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1):16– 25, July 1995. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
the-convergence-of-halleys-method-for-solving-operator-equations/
1D6347D4054AFD111904DCAADC7D3632.

Connolly:1995:GMF

[785] T. J. Connolly, K. A. Landman, and L. R. White. On Gerchberg’s method for the Fourier inverse problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1):26–44, July 1995. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
gerchbergs-method-for-the-fourier-inverse-problem/BC62E5AC5B24C15704053FFD6927842B.

Dewynne:1995:NAO

[786] J. N. Dewynne and P. Wilmott. A note on American options with varying exercise price. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1):45–57, July 1995. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
american-options-with-varying-exercise-price/302F0338893A79EE6054876C13046084.

Shridharan:1995:GIM

[787] Radha Shridharan and Ravi P. Agarwal. General iterative methods for nonlinear boundary value problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1):58–85, July 1995. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/general-iterative-
methods-for-nonlinear-boundary-value-problems/9BEDCED64D5883C036AB852F2BEB4E0D.

Guo:1995:ABS

[788] Bao Zhu Guo. Asymptotic behaviour of the spectrum of a direct feedback control system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1):86–98, July 1995. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/asymptotic-behaviour-of-the-spectrum-
of-a-direct-feedback-control-system/EBC89F894A83DFB9DD7EDC6443B81F66.

Bencheikh:1995:LFS

[789] L. Bencheikh. Low frequency scattering of elastic waves by a cavity using a matched asymptotic expansion method. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1): 99–120, July 1995. CODEN JAMMDU. ISSN 0334-2700. URL https://

www.cambridge.org/core/journals/anziam-journal/article/low-
frequency-scattering-of-elastic-waves-by-a-cavity-using-a-matched-
asymptotic-expansion-method/35749A3C97C128413EF2FFD400F0E581.

Anonymous:1995:AVIe

[790] Anonymous. ANZ volume 37 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1):f1–f2, July 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-37-issue-1-cover-and-front-matter/3052AA2865464C2EB58D35448B97079A.

Anonymous:1995:AVIf

[791] Anonymous. ANZ volume 37 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(1):b1–b2, July 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-37-issue-1-cover-and-back-matter/D9A1738ADC116A17C75494710989418A.

vanDoorn:1995:GEQ

[792] Erik A. van Doorn and Pauline Schrijner. Geometric ergodicity and quasi-stationarity in discrete-time birth–death processes. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):121–144, October 1995. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/geomatric-ergodicity-and-quasistationarity-
in-discretetime-birthdeath-processes/9A6E800B22F8DB888B554A6085967394.

Dube:1995:UFG

[793] Simant Dube. Using fractal geometry for solving divide-and-conquer recurrences. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):145–171, October 1995. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/using-fractal-geometry-for-
solving-divideandconquer-recurrences/1FF704BD0ED82C26BA52B7FA1ACDF77F.

Xu:1995:SMP

[794] Wensheng Xu. Stochastic maximum principle for optimal control problem of forward and backward system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37 (2):172–185, October 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stochastic-maximum-principle-for-optimal-control-problem-
of-forward-and-backward-system/58A7F9D899473DE945341E2E79BD81BB.

Christie:1995:CAO

[795] J. R. Christie and K. Gopalsamy. Chaos in an anharmonic oscillator. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):186–207, October 1995. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/chaos-in-an-anharmonic-oscillator/2080CF70A81D7921C4613C6C4FBB8BE6.

Pearce:1995:IRS

[796] C. E. M. Pearce and J. E. Pecarić. On an inequality relating to sum sets. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):208–211, October 1995. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-an-inequality-relating-to-sum-sets/637CBC92CD7DEC648B12FA37F8107C60.

Winch:1995:DAT

[797] D. E. Winch and P. H. Roberts. Derivatives of addition theorems for Legendre functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):212–234, October 1995. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/derivatives-of-addition-theorems-
for-legendre-functions/495655C12A4D5D19D3AEC6E238879295.

Stott:1995:TSD

[798] Jillian A. K. Stott and James P. Denier. The temporal stability of a developing jet: a model problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):235–252, October 1995. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/temporal-stability-of-
a-developing-jet-a-model-problem/DA2806C930446DDEF67BE2F99F6A61F3.

Guo:1995:SSQ

[799] Jong-Sheng Guo. Similarity solutions for a quasilinear parabolic equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):253–266, October 1995. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/similarity-solutions-for-a-
quasilinear-parabolic-equation/07450CC7FC709B4547942487B1014007.

Hurley:1995:ERD

[800] D. G. Hurley and P. F. Siew. The EM response due to a plane sheet of arbitrary shape and conductivity profile. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):267–278, October 1995. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/em-response-due-to-a-plane-sheet-of-
arbitrary-shape-and-conductivity-profile/0ED49979FE5C4DB468681D55AC7FDB68.

Anonymous:1995:AVIg

[801] Anonymous. ANZ volume 37 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):f1–f2, October 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-37-issue-2-cover-and-front-matter/7293ACC6B78BBE40F5D924998BBA27B3.

Anonymous:1995:AVIh

[802] Anonymous. ANZ volume 37 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(2):b1–b2, October 1995. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-37-issue-2-cover-and-back-matter/FD2CD92EE63A14095BB9BD3675E5D016.

Deakin:1996:LTE

[803] Michael A. B. Deakin. The Laplace transform of exp(e^{t}). Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):279–287, January 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/laplace-transform-of-expet/02BA63C8D603231FBD0BD209A8A1C516.

Stewart:1996:NMF

[804] David E. Stewart. A numerical method for friction problems with multiple contacts. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):288–308, January 1996. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/numerical-method-for-friction-problems-
with-multiple-contacts/753612E603D420040178AA58C08442C1.

Clifford:1996:LOO

[805] M. J. Clifford and S. R. Bishop. Locating oscillatory orbits of the parametrically-excited pendulum. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):309–319, January 1996. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/locating-
oscillatory-orbits-of-the-parametricallyexcited-pendulum/D592814128A674CB228688443B5F6480.

Jiang:1996:MSP

[806] Zhuhan Jiang. Matrix spectral problem with multiple-order jumps and poles. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):320–333, January 1996. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/matrix-spectral-problem-with-multipleorder-
jumps-and-poles/8CFA6477EA9E64CEBDCD0F75ECA1D58A.

McKee:1996:BRT

[807] W. D. McKee. Bragg resonances in a two-layer fluid. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):334–345, January 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/bragg-resonances-in-a-twolayer-fluid/40783376EC3CC8EF181C80F8D72C061C.

Balachandran:1996:RCN

[808] K. Balachandran and J. P. Dauer. Relative controllability of nonlinear neutral Volterra integrodifferential systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37 (3):346–353, January 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/relative-controllability-of-nonlinear-neutral-volterra-
integrodifferential-systems/36D84695D1C7ED17915716CC0A553BC3.

Tuan:1996:SSN

[809] H. D. Tuan. On solution sets of nonconvex Darboux problems and applications to optimal control with endpoint constraints. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):354–391, January 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-solution-sets-of-nonconvex-darboux-problems-and-
applications-to-optimal-control-with-endpoint-constraints/345E6326891D1EE59C17E61983D0E8E9.

Chaudhry:1996:EGI

[810] M. Aslam Chaudhry and S. M. Zubair. On an extension of generalized incomplete Gamma functions with applications. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):392–405, January 1996. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
an-extension-of-generalized-incomplete-gamma-functions-with-
applications/06E99485F893A1301A12A15F6B3A2B8B.

Forbes:1996:NWT

[811] Lawrence K. Forbes, Graeme C. Hocking, and Graeme A. Chandler. A note on withdrawal through a point sink in fluid of finite depth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):406–416, January 1996. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/note-on-withdrawal-through-a-point-sink-in-
fluid-of-finite-depth/BB1037B4442C6FEFC9500BD6BFE98085.

Chakrabarti:1996:RWW

[812] A. Chakrabarti and T. Sahoo. Reflection of water waves by a nearly vertical porous wall. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):417–429, January 1996. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/reflection-of-water-waves-by-
a-nearly-vertical-porous-wall/900F5408261626AC70A9C08983AD06A0.

Anonymous:1996:AVIa

[813] Anonymous. ANZ volume 37 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):f1–f2, January 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-37-issue-3-cover-and-front-matter/B1891F75F2B0EAB0FD96BEDD95F7BDF2.

Anonymous:1996:AVIb

[814] Anonymous. ANZ volume 37 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(3):b1–b2, January 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-37-issue-3-cover-and-back-matter/81D1B205ABAF4019BF5742480E317995.

Mahony:1996:GAP

[815] R. E. Mahony, U. Helmke, and J. B. Moore. Gradient algorithms for principal component analysis. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37 (4):430–450, April 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/gradient-algorithms-for-principal-component-analysis/
4A68F08C7E1D10784520D78BCA5D9518.

Sexton:1996:NOS

[816] M. Jane Sexton and Lawrence K. Forbes. A note on oscillations in a simple model of a chemical reaction. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):451–457, April 1996. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
oscillations-in-a-simple-model-of-a-chemical-reaction/0701F09929972F3D0EA1B567C2DB15F7.

Watt:1996:RWN

[817] S. D. Watt and R. O. Weber. Reaction waves and non-constant diffusivities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):458–473, April 1996. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/reaction-waves-and-nonconstant-diffusivities/
5F8AC80272D295FDC09FF601A7E5BDDD.

Cabrelli:1996:WTD

[818] Carlos A. Cabrelli and Ursula M. Molter. Wavelet transform of the dilation equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):474–489, April 1996. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/wavelet-transform-
of-the-dilation-equation/7A1BB248B8F6482BB32CAECF673937E0.

Pearce:1996:TLB

[819] C. E. M. Pearce and J. Pecarić. On two lemmas of Brown and Shepp having application to sum sets and fractals, II. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):490–494, April 1996. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
two-lemmas-of-brown-and-shepp-having-application-to-sum-sets-
and-fractals-ii/4FA534F8FDB688A774EF4A792BAA1D2A.

Kazemi:1996:NCC

[820] Mohammad A. Kazemi. Necessary conditions for constrained distributed parameter systems with deviating argument. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):495–511, April 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/necessary-conditions-for-constrained-distributed-parameter-
systems-with-deviating-argument/FBBA28A366B9F250B27B6CA9F4F80F94.

Banerjea:1996:SWW

[821] Sudeshna Banerjea. Scattering of water waves by a vertical wall with gaps. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):512–529, April 1996. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/scattering-of-water-waves-by-a-vertical-
wall-with-gaps/B937437AE324C59ABF1DFE63C5AC8E4B.

Ernst:1996:CTQ

[822] A. T. Ernst. Continuous-time quadratic cost flow problems with applications to water distribution networks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37 (4):530–548, April 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/continuoustime-quadratic-cost-flow-problems-with-applications-
to-water-distribution-networks/13FF0C2B7D47076B97658D93FD1DDC1E.

Kim:1996:PCM

[823] Sang Dong Kim. Preconditioning collocation method using quadratic splines with applications to 2nd-order separable elliptic equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):549–570, April 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/preconditioning-collocation-method-using-quadratic-

splines-with-applications-to-2ndorder-separable-elliptic-equations/
8018B2B65C1D003E394B5ACD6BF6FA91.

Philip:1996:FDL

[824] J. R. Philip. Fast diffusion with loss at infinity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):571, April 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fast-diffusion-with-loss-at-infinity/F3C85F50E04C08740A18C8080AFD97FE.

Anonymous:1996:AVIc

[825] Anonymous. ANZ volume 37 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):f1–f2, April 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-37-issue-4-cover-and-front-matter/4679EFDD21762F84055D4E6620420D3E.

Anonymous:1996:AVId

[826] Anonymous. ANZ volume 37 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 37(4):b1–b2, April 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-37-issue-4-cover-and-back-matter/34C23C76DE23D6D8AF99EE952C7F0878.

Kolyshkin:1996:DCL

[827] A. A. Kolyshkin and Rémi Vaillancourt. Double conductor line above a two-layered cylinder with varying properties. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):1–15, July 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/double-conductor-line-above-a-twolayered-cylinder-with-
varying-properties/F8AB2800D0B628CCB9C87AD270F3CC49.

Russell:1996:WOM

[828] A. M. Russell, C. A. Martini, and J. A. Rickard. Welfare optimization and multinational monopolies. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):16–25, July 1996. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/welfare-
optimization-and-multinational-monopolies/A916DE8EB184D126AB8FD914E0A9E5B0.

Yih:1996:SWS

[829] Chia-Shun Yih and Songping Zhu. Selective withdrawal from stratified streams. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):26–40, July 1996. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/selective-withdrawal-from-stratified-
streams/E2F57A5F747084009E343BA18E4064C9.

He:1996:OIS

[830] Xiangjian He. Oscillations of interconnected systems with C^{0} nonlinearities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):41–62, July 1996. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/oscillations-of-interconnected-systems-with-
c0-nonlinearities/B4B86A50065B04CFFE89FFF645E5DBD2.

Asavanant:1996:NFS

[831] J. Asavanant and J.-M. Vanden-Broeck. Nonlinear free-surface flows emerging from vessels and flows under a sluice gate. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):63–86, July 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinear-freesurface-flows-emerging-from-vessels-and-
flows-under-a-sluice-gate/4C839B01B43D0EA3C9F4D1239F2BB3B4.

Bera:1996:STC

[832] R. K. Bera and A. Chakrabarti. The sputtering temperature of a cooling cylindrical rod without and with an insulated core in a two-fluid medium. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):87–100, July 1996. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/sputtering-temperature-of-a-cooling-
cylindrical-rod-without-and-with-an-insulated-core-in-a-twofluid-
medium/9B2B2F2A23BE6D60FB1648121670C0B9.

Watt:1996:CZM

[833] S. D. Watt and A. J. Roberts. The construction of zonal models of dispersion in channels via matched centre manifolds. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):101–125, July 1996. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/construction-of-zonal-models-of-dispersion-
in-channels-via-matched-centre-manifolds/4E021E5C5937844D0CAC72B30B8E0AFE.

Zhou:1996:VMR

[834] S. Zhou and X. Cai. Variance minimization — relationship between completion-time variance and waiting-time variance. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):126–139, July 1996. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/variance-minimization-relationship-between-
completiontime-variance-and-waitingtime-variance/D8DEFE8075D5A74EF5434AB92107A66E.

Mishra:1996:GCM

[835] S. K. Mishra and R. N. Mukherjee. On generalised convex multi-objective nonsmooth programming. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):140–148, July 1996. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/on-generalised-convex-
multiobjective-nonsmooth-programming/770CFF28C5FBF0B0E42B7B83BC5700B5.

Anonymous:1996:AVIe

[836] Anonymous. ANZ volume 38 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):f1–f2, July 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-38-issue-1-cover-and-front-matter/34D36991FBB80FA291FF280A23EC70A7.

Anonymous:1996:AVIf

[837] Anonymous. ANZ volume 38 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(1):b1–b2, July 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-38-issue-1-cover-and-back-matter/4FC34E790C1D0B870D0864B2F708F6BE.

Cao:1996:GAT

[838] Yulin Cao and H. I. Freedman. Global attractivity in time-delayed predator–prey systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):149–162, October 1996. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/global-attractivity-in-timedelayed-
predatorprey-systems/B4A50F39E5DF5B2E809B8D0EE31E0434.

Graef:1996:OAB

[839] John R. Graef, Agnes Miciano, Paul W. Spikes, P. Sundaram, and E. Thandapani. Oscillatory and asymptotic behavior of solutions of nonlinear neutral-type difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38 (2):163–171, October 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/oscillatory-and-asymptotic-behavior-of-solutions-of-
nonlinear-neutraltype-difference-equations/018F7DBDD7C024D0CC70C122A9E2BEA1.

Xu:1996:MPO

[840] Wensheng Xu. A maximum principle for optimal control for a class of controlled systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):172–181, October 1996. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/maximum-principle-for-
optimal-control-for-a-class-of-controlled-systems/5F2AC53416CEA3F1EF6CCA202E799A35.

Jennings:1996:OCC

[841] L. S. Jennings, K. H. Wong, and K. L. Teo. Optimal control computation to account for eccentric movement. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):182–193, October 1996. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/optimal-control-computation-to-account-for-
eccentric-movement/890859E11EFC702FFE995647F61C981E.

Yang:1996:CSS

[842] X. Q. Yang. A comparative study of smoothing approximations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):194–200, October 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/comparative-study-of-smoothing-approximations/32A5FFED40F6E87B4E47C6FA88F4AADB.

Majumdar:1996:GMP

[843] A. A. K. Majumdar. Generalized multi-peg Tower of Hanoi problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):201–208, October 1996. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/generalized-multipeg-tower-of-hanoi-
problem/694840EC7758F88854744576A287118E.

Mustard:1996:FFT

[844] David Mustard. The fractional Fourier transform and the Wigner distribution. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):209–219, October 1996. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/fractional-fourier-transform-and-the-
wigner-distribution/D5460A0A5FFF8DFBE2B8A9855D40A42E.

Andonowati:1996:TSS

[845] Andonowati. A two-sided shooting method in computation of travelling combustion waves of a solid material. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38 (2):220–228, October 1996. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/twosided-shooting-method-in-computation-of-
travelling-combustion-waves-of-a-solid-material/A87D7E4787DCEE7710CDCD8E2C47B2B9.

Bragg:1996:PRS

[846] L. R. Bragg. Propagation relations for solutions of some higher order Cauchy problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):229–239, October 1996. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/propagation-relations-
for-solutions-of-some-higher-order-cauchy-problems/B5ACD64EC645617438D3CE6C166368A5.

Zhang:1996:WLF

[847] H. Zhang and G. C. Hocking. Withdrawal of layered fluid through a line sink in a porous medium. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):240–254, October 1996. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/withdrawal-of-layered-
fluid-through-a-line-sink-in-a-porous-medium/29202E4FE07EB95320558BFB7C5CF7DB.

Hird:1996:SRN

[848] L. D. Hird and P. F. Siew. Small Reynolds number flow between eccentric rotating cylinders with a permeable sleeve. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38 (2):255–273, October 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/small-reynolds-number-flow-between-eccentric-rotating-
cylinders-with-a-permeable-sleeve/504BA08DA26B095DC92988745905B72D.

Shin:1996:FSF

[849] Dongho Shin and John C. Strikwerda. Fast solvers for finite difference approximations for the Stokes and Navier–Stokes equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):274–290, October 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fast-solvers-for-finite-difference-approximations-for-
the-stokes-and-navierstokes-equations/1F9256C162CC6BEBFCF62F7C695852F7.

Anonymous:1996:AVIg

[850] Anonymous. ANZ volume 38 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):f1–f2, October 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-38-issue-2-cover-and-front-matter/2931FFD446F15B035CE1907B2C28C9E8.

Anonymous:1996:AVIh

[851] Anonymous. ANZ volume 38 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(2):b1–b2, October 1996. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-38-issue-2-cover-and-back-matter/ED13323FC09DB024003560B414E3DD69.

Rojas-Medar:1997:GSS

[852] Marko A. Rojas-Medar and José Luiz Boldrini. Global strong solutions of equations of magnetohydrodynamic type. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38 (3):291–306, January 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/global-strong-solutions-of-equations-of-magnetohydrodynamic-
type/DF09CEE6F018650796F793334303AA82.

Christie:1997:SOA

[853] J. R. Christie, K. Gopalsamy, and M. P. Panizza. Subharmonic orbits in an anharmonic oscillator. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(3):307–315, January 1997. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/subharmonic-orbits-in-
an-anharmonic-oscillator/61818583CC75EB85BDB0FB9589AC48CE.

Dostovalova:1997:FSS

[854] Anna S. Dostovalova and Sergey T. Simakov. Formation of singularities in a stratified fluid in the presence of a critical level. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(3):316–324, January 1997. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/formation-of-singularities-in-a-stratified-
fluid-in-the-presence-of-a-critical-level/BE21AC0E526574D921C901362792BCCD.

Varosanec:1997:SII

[855] Sanja Varosanec and Josip Pecarić. Some integral inequalities, with application to bounds for moments of a distribution. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38 (3):325–335, January 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-integral-inequalities-with-application-to-bounds-
for-moments-of-a-distribution/F7BE392ADE488A043428AB3BA67B8E2A.

Carstensen:1997:ABE

[856] Carsten Carstensen and Ernst P. Stephan. Adaptive boundaryelement methods for transmission problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38 (3):336–367, January 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/adaptive-boundaryelement-methods-for-transmission-problems/
ABA7D2D5C8864CCB22B80189FF064B84.

Leyk:1997:CCL

[857] Zbigniew Leyk. A C^{0}-collocation-like method for elliptic equations on rectangular regions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(3):368–387, January 1997. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/c0collocationlike-
method-for-elliptic-equations-on-rectangular-regions/14E9269482A16F98DF2D1F7599645747.

Howlett:1997:ODS

[858] P. G. Howlett and J. Cheng. Optimal driving strategies for a train on a track with continuously varying gradient. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(3):388–410, January 1997. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
driving-strategies-for-a-train-on-a-track-with-continuously-
varying-gradient/A8904A79571D4BB33C1E6C56AB73CEE9.

El-Gabali:1997:EMT

[859] M. El-Gabali. Efficient multiple-term approximations for the generalised elliptic-type integrals. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(3):411–426, January 1997. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/efficient-multipleterm-approximations-
for-the-generalised-elliptictype-integrals/E4A876F6BE1D096AA903D070033F3644.

Anonymous:1997:AVIa

[860] Anonymous. ANZ volume 38 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(3):f1–f2, January 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-38-issue-3-cover-and-front-matter/C7616B0625AB4034EB2B7F747F7644AE.

Anonymous:1997:AVIb

[861] Anonymous. ANZ volume 38 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(3):b1–b2, January 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-38-issue-3-cover-and-back-matter/49128726B56CF8E4621915C2232BEE66.

Barnes:1997:ANS

[862] Belinda Barnes and Roger Grimshaw. Analytical and numerical studies of the Bonhoeffer van der Pol system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38 (4):427–453, April 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/analytical-and-numerical-studies-of-the-bonhoeffer-
van-der-pol-system/B9B99987D5AE10CD84BBB78718BD86CB.

Boucherie:1997:QSD

[863] Richard J. Boucherie. On the quasi-stationary distribution for queueing networks with defective routing. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):454–463, April 1997. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-quasistationary-distribution-for-queueing-networks-with-
defective-routing/FFF46B186EEC80A06EF8B6B5D2F3E092.

Weber:1997:CW

[864] R. O. Weber and S. D. Watt. Combustion waves. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):464–476, April 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/combustion-waves/CF2B9AE85758D59049BF0F2FA6FF3942.

Pleasants:1997:AHW

[865] A. B. Pleasants, G. C. Wake, and A. L. Rae. The allometric hypothesis when the size variable is uncertain: issues in the study of carcass composition by serial slaughter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):477–488, April 1997. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/allometric-hypothesis-
when-the-size-variable-is-uncertain-issues-in-the-study-of-
carcass-composition-by-serial-slaughter/841683996DBC614DEE5D681F3FE39E51.

Bainov:1997:SML

[866] D. D. Bainov and I. M. Stamova. Second method of Lyapunov and comparison principle for impulsive differential-difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):489–505, April 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/second-method-of-lyapunov-and-comparison-principle-
for-impulsive-differentialdifference-equations/E09439032FB387829A92D64EBA7850F6.

Tam:1997:NSP

[867] K. K. Tam and Andonowati. Numerical study of a problem in the combustion of a porous medium. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38 (4):506–517, April 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-study-of-a-problem-in-the-combustion-of-a-
porous-medium/5F41DBC3D0922618222BAC5AF82368C9.

Tritscher:1997:IFO

[868] Peter Tritscher. An integrable fourth-order nonlinear evolution equation applied to surface redistribution due to capillarity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4): 518–541, April 1997. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
integrable-fourthorder-nonlinear-evolution-equation-applied-
to-surface-redistribution-due-to-capillarity/897CEE0DF92EC7B6B3F45A92B23B7EF0.

Jianyong:1997:MDP

[869] Liu Jianyong and Liu Ke. Markov decision programming — the moment optimal problem for the first-passage model. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):542–562, April 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/markov-decision-programmingthe-moment-optimal-problem-
for-the-firstpassage-model/ABA3B8BB437243A934DD06B58AF53FB5.

Sha:1997:ISW

[870] Huyun Sha and J.-M. Vanden-Broeck. Internal solitary waves with stratification in density. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):563–580, April 1997. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/internal-solitary-
waves-with-stratification-in-density/069B54051776167764BCB2D2CD6BC1B6.

Chaudhry:1997:CBG

[871] M. Aslam Chaudhry and S. M. Zubair. On a connection between the generalized incomplete gamma functions and their extensions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):581–589, April 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-a-connection-between-the-generalized-incomplete-
gamma-functions-and-their-extensions/52B0C2E2C258260F3036FF3EF56F63A8.

Anonymous:1997:CV

[872] Anonymous. Contents of volume 38. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):590–592, April 1997. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
38/4CAB85702EEA4DE8A80B0A81D6A7A222.

Anonymous:1997:AVIc

[873] Anonymous. ANZ volume 38 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):f1–f2, April 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-38-issue-4-cover-and-front-matter/4A214E2F8454C6084F1615971C189D47.

Anonymous:1997:AVId

[874] Anonymous. ANZ volume 38 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 38(4):b1–b2, April 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-38-issue-4-cover-and-back-matter/6617980C6A9AB34BF0FE5C99960059C7.

Torokhti:1997:CAN

[875] A. P. Torokhti and P. G. Howlett. On the constructive approximation of non-linear operators in the modelling of dynamical systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39 (1):1–27, July 1997. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-constructive-approximation-of-nonlinear-operators-in-the-
modelling-of-dynamical-systems/58C0338E9FBF4B882AD63756E4C4F8BB.

Sander:1997:SSN

[876] G. C. Sander, R. D. Braddock, I. F. Cunning, J. Norbury, and S. W. Weeks. Source solutions for the nonlinear diffusion-convection equation.

Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):28–45, July 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/source-solutions-for-the-nonlinear-diffusionconvection-
equation/AD1E396DE40894F03B1C370814653B13.

Haidar:1997:BVP

[877] Nassar H. S. Haidar. On a boundary-value problem posed by cancer therapy with neutron beams. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):46–60, July 1997. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-a-boundaryvalue-problem-
posed-by-cancer-therapy-with-neutron-beams/8FC1C9D6A9862EAE5A506376AE9B980A.

Wong:1997:OCC

[878] K. H. Wong and N. Lock. Optimal control of a chemical reactor. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):61–76, July 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-control-of-a-chemical-reactor/CD63B177CE47EEDF8BF3088AD8EC26B2.

Papageorgiou:1997:VSC

[879] Nikolaos S. Papageorgiou. On the variational stability of a class of nonlinear parabolic optimal control problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1): 77–92, July 1997. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-variational-stability-of-a-class-of-nonlinear-parabolic-
optimal-control-problems/F5DC7D6CC1EDCC164038AA8BD548B620.

Chakrabarti:1997:UAP

[880] A. Chakrabarti, Sudeshna Banerjea, B. N. Mandal, and T. Sahoo. A unified approach to problems of scattering of surface water waves by vertical barriers. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):93–103, July 1997. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/unified-
approach-to-problems-of-scattering-of-surface-water-waves-by-
vertical-barriers/EEC2E8AC7CCB2B9BCC8C386FBAA39FE3.

Rhodes-Robinson:1997:WPV

[881] P. F. Rhodes-Robinson. On waves in the presence of vertical porous boundaries. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):104–120, July 1997. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
waves-in-the-presence-of-vertical-porous-boundaries/B7DDC078D8E121AEBC06ADC4C10776E6.

Shin:1997:ISC

[882] Dongho Shin and John C. Strikwerda. Inf-sup conditions for finitedifference approximations of the Stokes equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):121–134, July 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/infsup-conditions-for-finitedifference-approximations-
of-the-stokes-equations/1E7C512E108E2CE90B0B8ABA0BACDE0E.

El-Gabali:1997:MTA

[883] Magdi A. El-Gabali. Multiple-term approximations for Appell’s F_{1} function. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):135–148, July 1997. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/multipleterm-approximations-for-appells-f1-
function/8E218109899D68FD0DC15B6E9D61E8BD.

Anonymous:1997:AVIe

[884] Anonymous. ANZ volume 39 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):f1–f2, July 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-39-issue-1-cover-and-front-matter/40D06C9F7CDB70BB0CE654E77AF03D11.

Anonymous:1997:AVIf

[885] Anonymous. ANZ volume 39 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):b1–b2, July 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-39-issue-1-cover-and-back-matter/91C17010400158E932275F12BC01A609.

Rudman:1997:OFE

[886] Murray Rudman. One-field equations for two-phase flows. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):149–170, October 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/onefield-equations-for-twophase-flows/B84479989E83EE272A1EB4D3E81CEC00.

Forbes:1997:TDM

[887] Lawrence K. Forbes. A two-dimensional model for large-scale bushfire spread. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):171–194, October 1997. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/twodimensional-model-for-largescale-
bushfire-spread/7DAC7ACCBA18C225B0646110B8DEB74E.

Lee:1997:OCM

[888] H. W. J. Lee, K. L. Teo, and L. S. Jennings. On optimal control of multi-link vertical planar robot arms systems moving under the effect of gravity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):195–213, October 1997. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/on-optimal-control-of-
multilink-vertical-planar-robot-arms-systems-moving-under-the-
effect-of-gravity/AA2A337B9FF951D28775F2E0D31CEEAC.

Bhattacharyya:1997:FON

[889] Sudebi Bhattacharyya and K. P. Das. Fourth-order nonlinear evolution equations for surface gravity waves in the presence of a thin thermocline. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):214–229, October 1997. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/fourthorder-nonlinear-
evolution-equations-for-surface-gravity-waves-in-the-presence-
of-a-thin-thermocline/3DD62E13EDB2A2FB28237FC9981834D7.

Yung:1997:DGS

[890] Siu Pang Yung. Differential game with switching controls on Hilbert space. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):230–256, October 1997. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/differential-game-
with-switching-controls-on-hilbert-space/D17C249541DF1661C337F1B46A2961F4.

Mohammad:1997:BBG

[891] Ch. Wali Mohammad. Bilinear and bilateral generating functions of generalized polynomials. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):257–270, October 1997. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/bilinear-and-bilateral-generating-
functions-of-generalized-polynomials/1CD69EBEBBA41DFDCFFD8F20AB59E339.

Zhang:1997:ASC

[892] H. Zhang, G. C. Hocking, and D. A. Barry. An analytical solution for critical withdrawal of layered fluid through a line sink in a porous medium. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):271–279, October 1997. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-analytical-solution-for-critical-
withdrawal-of-layered-fluid-through-a-line-sink-in-a-porous-
medium/F3B080191F9189BB2303D767BA6AE1F2.

Sun:1997:CIM

[893] Wenyu Sun. On the convergence of an iterative method for the minimax problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):280–292, October 1997. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/on-the-convergence-of-
an-iterative-method-for-the-minimax-problem/5AA4DA9C799A0DDD8BD7B7B766B315FE.

Anonymous:1997:AVIg

[894] Anonymous. ANZ volume 39 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):f1–f2, October 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-39-issue-2-cover-and-front-matter/2EB51372EB9B5979F0229A2F67D720E6.

Anonymous:1997:AVIh

[895] Anonymous. ANZ volume 39 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(2):b1–b2, October 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-39-issue-2-cover-and-back-matter/1500974840BD4A5372F321AEA98175C2.

Rhodes-Robinson:1998:SWN

[896] P. F. Rhodes-Robinson. On the scattering of waves by nearly hard or soft incomplete vertical barriers in water of infinite depth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(3):293–307, January 1998. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-scattering-of-waves-
by-nearly-hard-or-soft-incomplete-vertical-barriers-in-water-
of-infinite-depth/5E706BFB46799EE7BEE73CC39C2E8DAE.

Chakrabarti:1998:RWW

[897] A. Chakrabarti and T. Sahoo. Reflection of water waves in the presence of surface tension by a nearly vertical porous wall. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(3):308–317, January 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reflection-of-water-waves-in-the-presence-of-surface-
tension-by-a-nearly-vertical-porous-wall/41B51ACFCBDCFD1989535CD292E08B6D.

Banerjea:1998:SWW

[898] Sudeshna Banerjea and B. N. Mandal. Scattering of water waves by a submerged thin vertical wall with a gap. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39 (3):318–331, January 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/scattering-of-water-waves-by-a-submerged-thin-vertical-
wall-with-a-gap/A16682D93639E88F0855B3A76E055A9F.

Lee:1998:BRI

[899] J. Lee and J.-M. Vanden-Broeck. Bubbles rising in an inclined twodimensional tube and jets falling along a wall. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(3):332–349, January 1998. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/bubbles-
rising-in-an-inclined-twodimensional-tube-and-jets-falling-
along-a-wall/778677A45B8ED3C90F8C21352099C448.

Guljas:1998:IPD

[900] Boris Guljas, C. E. M. Pearce, and Josip Pecarić. An inequality for probability density functions arising from a distinguishability problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(3):350–354, January 1998. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/an-inequality-for-
probability-density-functions-arising-from-a-distinguishability-
problem/7D1F0D4067AF6375DFF760C210D292CF.

Rathod:1998:ITP

[901] H. T. Rathod and H. S. Govinda Rao. Integration of trivariate polynomials over linear polyhedra in Euclidean three-dimensional space. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(3):355–385, January 1998. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/integration-of-trivariate-polynomials-over-
linear-polyhedra-in-euclidean-threedimensional-space/2EC1DD2A423191EE3F2935B6E34DA970.

Wong:1998:ECP

[902] Patricia J. Y. Wong and Ravi P. Agarwal. Eigenvalue characterization for (n · p) boundary-value problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39 (3):386–407, January 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/eigenvalue-characterization-for-n-p-boundaryvalue-problems/
9BBF9972E805EC63DED7259DB6187B4E.

Kim:1998:SCM

[903] Chang Ho Kim and U Jin Choi. Spectral collocation methods for a partial integro-differential equation with a weakly singular kernel. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(3):408–430, January 1998. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/spectral-collocation-methods-for-a-partial-
integrodifferential-equation-with-a-weakly-singular-kernel/
D149C5390E08FA8A55F9D924FA2552B9.

Anonymous:1998:AVIa

[904] Anonymous. ANZ volume 39 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(3):f1–f2, January 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-39-issue-3-cover-and-front-matter/6E7E06C13718053B7103433D3A164FE7.

Anonymous:1998:AVIb

[905] Anonymous. ANZ volume 39 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(3):b1–b2, January 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-39-issue-3-cover-and-back-matter/8D90D9D18A5F37DA005A63DE8D2825E0.

Athanasiadis:1998:ASA

[906] Christodoulos Athanasiadis. On the acoustic scattering amplitude for a multi-layered scatterer. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):431–448, April 1998. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-acoustic-scattering-
amplitude-for-a-multilayered-scatterer/5902820F95465C835C39FA625600BDAD.

Xu:1998:OCP

[907] Wensheng Xu and Shuping Chen. Optimal consumption/portfolio choice with borrowing rate higher than deposit rate. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):449–462, April 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-consumptionportfolio-choice-with-borrowing-
rate-higher-than-deposit-rate/E3F805BBEA0B31833701FA3925D5DC51.

Weber:1998:TPO

[908] Gerhard-W. Weber. On the topology of parametric optimal control. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):463–497, April 1998. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-the-topology-of-parametric-optimal-control/
9793C352255171884699EB1BF90B7AC7.

Bassom:1998:LWV

[909] Andrew P. Bassom and P. J. Blennerhassett. Long wavelength vortices in time-periodic flows. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):498–512, April 1998. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/long-wavelength-vortices-in-
timeperiodic-flows/3165245076673C41610106351C6E6950.

Vijayakumar:1998:IES

[910] K. Vijayakumar. On the integrability and exact solutions of the nonlinear diffusion equation with a nonlinear source. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):513–527, April 1998. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-integrability-and-exact-solutions-of-the-nonlinear-diffusion-
equation-with-a-nonlinear-source/48CC071EC96612F8F30E3DC7D49A3698.

Vanden-Broeck:1998:MFS

[911] J.-M. Vanden-Broeck. A model for the free-surface flow due to a submerged source in water of infinite depth. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39 (4):528–538, April 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/model-for-the-freesurface-flow-due-to-a-submerged-source-
in-water-of-infinite-depth/72BDE2E6AD0527E0E27BF70825C15339.

Chakrabarti:1998:EST

[912] A. Chakrabarti and T. Sahoo. The effect of surface tension in porous wave maker problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):539–556, April 1998. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/effect-of-surface-tension-in-porous-
wave-maker-problems/F84C7F370612D67F2F96018B1B9F62E7.

Buchwald:1998:LES

[913] V. T. Buchwald and F. Viera. Linearised evaporation from a soil of finite depth above a water table. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39 (4):557–576, April 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/linearised-evaporation-from-a-soil-of-finite-depth-
above-a-water-table/4FF2BA6DBAD8D7494AAF4CEFA6B2AEF6.

Anonymous:1998:CV

[914] Anonymous. Contents of volume 39. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):577–579, April 1998. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/contents-of-volume-
39/7BCAC0E81E856F9030B41ED297A09AC9.

Anonymous:1998:AVIc

[915] Anonymous. ANZ volume 39 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):f1–f2, April 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-39-issue-4-cover-and-front-matter/F17ACC45F05EE04C5998DC94998519D8.

Anonymous:1998:AVId

[916] Anonymous. ANZ volume 39 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(4):b1–b2, April 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-39-issue-4-cover-and-back-matter/3F3C756F4073B131CB5DCFF4EDA3C0C0.

Yu:1998:QNA

[917] Wenhuan Yu. A quasi-Newton approach to identification of a parabolic system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):1–22, July 1998. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/quasinewton-approach-to-identification-
of-a-parabolic-system/F6FB37049DE434C6099C5D46E8B2634C.

Rixon:1998:SPB

[918] Andrew J. Rixon, Craig R. Johnson, and Alans S. Jones. Stability in paradigm biological systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):23–34, July 1998. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/stability-in-paradigm-
biological-systems/60DDF7CC01B265233CED47D707B97762.

Montgomery:1998:ANR

[919] P. J. Montgomery and T. B. Moodie. Analytical and numerical results for flow and shock formation in two-layer gravity currents. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):35–58, July 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/analytical-and-numerical-results-for-flow-and-shock-
formation-in-twolayer-gravity-currents/6038B40E6F11AA6E09FFE131C310A027.

Poznanski:1998:ELC

[920] Roman R. Poznanski. Electrophysiology of a leaky cable model for coupled neurons. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):59–71, July 1998. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/electrophysiology-
of-a-leaky-cable-model-for-coupled-neurons/73A770B0020ABE7C2FE10CACAF7932AE.

Brock:1998:ARR

[921] L. M. Brock. Analytic results for roots of two irrational functions in elastic wave propagation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):72–79, July 1998. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/analytic-results-for-roots-
of-two-irrational-functions-in-elastic-wave-propagation/C421DDEF06385187BDF077B0EA32745C.

Guljas:1998:JID

[922] B. Guljas, C. E. M. Pearce, and J. Pecarić. Jensen’s inequality for distributions possessing higher moments, with application to sharp bounds for Laplace–Stieltjes transforms. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):80–85, July 1998. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/jensens-
inequality-for-distributions-possessing-higher-moments-with-
application-to-sharp-bounds-for-laplacestieltjes-transforms/
F798C70FB7E31CCC0B5BD795EA4BFD66.

Choi:1998:SVM

[923] Bong Dae Choi, Gang Uk Hwang, and Dong Hwan Han. Supplementary variable method applied to the MAP/G/1 queueing system. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):86–96, July 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/supplementary-variable-method-applied-to-the-mapg1-
queueing-system/960475251FF98F4357B928C09A78968F.

Aw:1998:ALE

[924] Y. K. Aw, Robyn Owens, and John Ross. An analysis of local energy and phase congruency models in visual feature detection. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1): 97–122, July 1998. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
analysis-of-local-energy-and-phase-congruency-models-in-visual-
feature-detection/41E65143ADFAD3917898901A685CF067.

Denier:1998:NSW

[925] James P. Denier and Andrew P. Bassom. Neutrally stable wave motions in thermally stratified Poiseuille–Couette flow. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):123–144, July 1998. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/neutrally-stable-wave-motions-in-thermally-
stratified-poiseuillecouette-flow/7D261FCE57E53E60F204DBCA4F5BD214.

Anonymous:1998:AVIe

[926] Anonymous. ANZ volume 40 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):f1–f2, July 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-40-issue-1-cover-and-front-matter/B7C0CA9279DD56E9D4CFFF2DBD9C4B78.

Anonymous:1998:AVIf

[927] Anonymous. ANZ volume 40 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(1):b1–b2, July 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-40-issue-1-cover-and-back-matter/0F2DF5F42CA1D87E4F314AD4C28E3C03.

Russell:1998:SNV

[928] Craig L. Russell, P. J. Blennerhassett, and P. J. Stiles. Strongly nonlinear vortices in magnetized ferrofluids. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (2):146–170, October 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/strongly-nonlinear-vortices-in-magnetized-ferrofluids/
63D8E57AE36FE20C5B2BBE18D1CEBDE1.

Kundu:1998:SWW

[929] P. K. Kundu and N. K. Saha. On the scattering of water waves by a submerged slender barrier. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(2):171–189, October 1998. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-scattering-of-water-
waves-by-a-submerged-slender-barrier/915950DF8A2BB7A6D6191771456BE2B9.

Bhattacharyya:1998:ERS

[930] Sudebi Bhattacharyya and K. P. Das. The effect of randomness on the stability of deep water surface gravity waves in the presence of a thin thermocline. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(2):190–206, October 1998. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/effect-of-randomness-on-the-stability-
of-deep-water-surface-gravity-waves-in-the-presence-of-a-thin-
thermocline/E39E88369DE2B7AA166A4DA31C76962E.

Shin:1998:BGV

[931] Yang Woo Shin and Chareles E. M. Pearce. The BMAP/G/ 1 vacation queue with queue-length dependent vacation schedule. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(2):207–221, October 1998. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/bmapg1-vacation-queue-with-
queuelength-dependent-vacation-schedule/621D6B7A7567E05F1139918CA3112A69.

Miller:1998:MTP

[932] Allen R. Miller and H. M. Srivastava. On the Mellin transform of a product of hypergeometric functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(2):222–237, October 1998. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-mellin-transform-of-a-product-of-hypergeometric-functions/
8FAD1BFB7BBD51CF2C125EBA878B0AC0.

Chan:1998:LDQ

[933] Terence Chan. Large deviations and quasi-stationarity for densitydependent birth–death processes. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(2):238–256, October 1998. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/large-deviations-and-
quasistationarity-for-densitydependent-birthdeath-processes/
90398B27C94418A53A0E39562D10A95E.

Mustard:1998:FC

[934] David Mustard. Fractional convolution. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (2):257–265, October 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fractional-convolution/5F77E68A692F334232326450D3040334.

Rehbock:1998:COC

[935] V. Rehbock, S. Wang, and K. L. Teo. Computing optimal control with a hyperbolic partial differential equation. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (2):266–287, October 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/computing-optimal-control-with-a-hyperbolic-partial-
differential-equation/6B469E3C614ABFA653379952FEA5CC28.

Anonymous:1998:AVIg

[936] Anonymous. ANZ volume 40 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(2):f1–f3, October 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-40-issue-2-cover-and-front-matter/4D5425D182A36D8D37482EA0BBD2B29D.

Anonymous:1998:AVIh

[937] Anonymous. ANZ volume 40 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(2):b1–b2, October 1998. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-40-issue-2-cover-and-back-matter/F02FB116A949EF0873943524AA8086F6.

Jeyakumar:1999:P

[938] V. Jeyakumar and K. L. Teo. Preface. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):288, January 1999. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/preface/
93687443E0982706D16A6F354517A532.

Adly:1999:SGN

[939] S. Adly and W. Oettli. Solvability of generalized nonlinear symmetric variational inequalities. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):289–300, January 1999. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/solvability-of-generalized-
nonlinear-symmetric-variational-inequalities/AFDDF6B2C57EB02EB38A9EC117384A7E.

Borwein:1999:SRM

[940] Jonathan M. Borwein, Warren B. Moors, and Yongheng Shao. Subgradient representation of multifunctions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):301–313, January 1999. CODEN JAMMDU. ISSN 0334-2700. URL http:

//docserver.carma.newcastle.edu.au/184/;   http://journals.
cambridge.org/action/displayAbstract?fromPage=online&aid=3964380;
 https://www.cambridge.org/core/journals/anziam-journal/article/
subgradient-representation-of-multifunctions/A035439B7CEE8FBA04EAD6EB6888CDDB.

Volume in honour of B. Craven and B. Mond.

Teo:1999:CPE

[941] K. L. Teo, L. S. Jennings, H. W. J. Lee, and V. Rehbock. The control parameterization enhancing transform for constrained optimal control problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):314–335, January 1999. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/control-
parameterization-enhancing-transform-for-constrained-optimal-
control-problems/A0AC2BCCE5B25DE93AFE147330860B52.

Klatte:1999:SSN

[942] Diethard Klatte and Bernd Kummer. Strong stability in nonlinear programming revisited. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):336–352, January 1999. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/strong-stability-in-nonlinear-
programming-revisited/E3EA78848FA45F06E630991F74E4DA8C.

Zalinescu:1999:CCQ

[943] C. Z{ă \atop .}linescu. A comparison of constraint qualifications in infinitedimensional convex programming revisited. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (3):353–378, January 1999. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/comparison-of-constraint-qualifications-in-
infinitedimensional-convex-programming-revisited/061A5958E05E7D3BAA1B785437E376FC.

Xu:1999:SLS

[944] H. Xu, A. M. Rubinov, and B. M. Glover. Strict lower subdifferentiability and applications. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):379–391, January 1999. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/strict-lower-subdifferentiability-
and-applications/86409E7B3868EC255B1076B5429CC6D7.

Yang:1999:NOC

[945] X. Q. Yang and K. L. Teo. Necessary optimality conditions for bicriterion discrete optimal control problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (3):392–402, January 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/necessary-optimality-conditions-for-bicriterion-discrete-
optimal-control-problems/95E6B565E885DC5FEEC0B1448E9FA000.

Jeyakumar:1999:AHM

[946] V. Jeyakumar and X. Wang. Approximate Hessian matrices and secondorder optimality conditions for nonlinear programming problems with C^{1}-data. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):403–420, January 1999. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/approximate-hessian-matrices-and-secondorder-
optimality-conditions-for-nonlinear-programming-problems-with-
c1data/8AE049D7D3C173F1A42CF4A0CA028BD1.

Anonymous:1999:AVIa

[947] Anonymous. ANZ volume 40 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):f1–f2, January 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-40-issue-3-cover-and-front-matter/9A42DCD5EF0D263B3A6F83BC90551103.

Anonymous:1999:AVIb

[948] Anonymous. ANZ volume 40 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(3):b1–b2, January 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-40-issue-3-cover-and-back-matter/3C7482148F69B3A226AB9DFF9C19554C.

Tuck:1999:SFN

[949] E. O. Tuck and S. T. Simakov. Splash formation at the nose of a smoothly curved body in a stream. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (4):421–436, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/splash-formation-at-the-nose-of-a-smoothly-curved-body-
in-a-stream/311851C14985D07336988BB1C9BA76DC.

He:1999:WMS

[950] Z. He and M. W. Wong. Wavelet multipliers and signals. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):437–446, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/wavelet-multipliers-and-signals/21A6D131EE48D397A388738C61417C59.

Belward:1999:FNF

[951] Shaun R. Belward. Fully nonlinear flow over successive obstacles: hydraulic fall and supercritical flows. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (4):447–458, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fully-nonlinear-flow-over-successive-obstacles-hydraulic-
fall-and-supercritical-flows/45A5C430BB83A230AABA543F44CB5F17.

Lu:1999:SRM

[952] Linzhang Lu and C. E. M. Pearce. On the square-root method for continuous-time algebraic Riccati equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):459–468, April 1999. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-squareroot-method-for-continuoustime-algebraic-riccati-
equations/709BEA9A754E8C14F22D969F3749412E.

OMalley:1999:MIS

[953] Robert E. O’Malley. Mahony’s intriguing stiff equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):469–474, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/mahonys-intriguing-stiff-equations/CA17D226DEB25A5785FC7777E5A736FF.

Gumel:1999:NSD

[954] A. B. Gumel. On the numerical solution of the diffusion equation subject to the specification of mass. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):475–483, April 1999. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-numerical-solution-of-the-diffusion-equation-subject-to-
the-specification-of-mass/4E8B7F1F5DE167AD42041A1064ADD2CA.

Zhang:1999:IFM

[955] Jianzhong Zhang, Zhenhong Liu, and Zhongfan Ma. The inverse fractional matching problem. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):484–496, April 1999. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/inverse-
fractional-matching-problem/62D110210C2D6D9EEDD7E6A5C80B00F3.

Li:1999:OEP

[956] Mingyan Li and Cheng-Chew Lim. Observability-enhanced proportional navigation guidance with bearings-only measurements. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):497–512, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/observabilityenhanced-proportional-navigation-guidance-
with-bearingsonly-measurements/491048E9EF20684EFB5FB8F4703CD069.

Kim:1999:SOT

[957] Chang Ho Kim and U Jin Choi. Second-order time discretization with finite-element method for partial integro-differential equations with a weakly singular kernel. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):513–524, April 1999. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/secondorder-time-discretization-
with-finiteelement-method-for-partial-integrodifferential-equations-
with-a-weakly-singular-kernel/5909A97FCABF48497C7B235A59DBA86F.

Bondarev:1999:SLS

[958] E. A. Bondarev, V. A. Budugaeva, and E. L. Gusev. Synthesis of layered shells from a finite set of viscoelastic materials. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):525–534, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/synthesis-of-layered-shells-from-a-finite-set-of-viscoelastic-
materials/9C4CA32FD83245E87D8F7D381692EC4D.

Jardas:1999:SIE

[959] Cvetan Jardas, Josip Pecarić, Rajko Roki, and Nikola Sarapa. On some inequalities for entropies of discrete probability distributions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):535–541, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-some-inequalities-for-entropies-of-discrete-probability-
distributions/D14CCB243E0573CFF9E993E7C8A42C4D.

Zhang:1999:NTR

[960] Jianzhong Zhang and Detong Zhu. A nonmonotonic trust region method for constrained optimization problems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (4):542–567, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonmonotonic-trust-region-method-for-constrained-optimization-
problems/F7136C12A458B8B0727C84C801F71313.

Yang:1999:MAM

[961] X. Q. Yang and C. J. Goh. Models and algorithms for multiple criteria linear cost network programs. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40 (4):568–581, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/models-and-algorithms-for-multiple-criteria-linear-
cost-network-programs/D414E57F4F6CEEE6F44270747B927B5C.

Anonymous:1999:AVIc

[962] Anonymous. ANZ volume 40 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):f1–f2, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-40-issue-4-cover-and-front-matter/8E2D1525CF9105D24E6ACA8B4A38DB06.

Anonymous:1999:AVId

[963] Anonymous. ANZ volume 40 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 40(4):b1–b2, April 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-40-issue-4-cover-and-back-matter/971DF5B3137AB3D434B1C998C43FCE4A.

Qu:1999:RES

[964] Changzheng Qu. Reductions and exact solutions of some nonlinear partial differential equations under four types of generalized conditional symmetries. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):1–40, July 1999. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/reductions-and-exact-solutions-of-some-
nonlinear-partial-differential-equations-under-four-types-of-
generalized-conditional-symmetries/EE5B96C6A62ACD535B96EF74993F5B58.

Liu:1999:BMG

[965] Y. Liu and K. L. Teo. A bridging method for global optimization. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):41–57, July 1999. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/bridging-method-for-global-
optimization/ECB13282F057896C0B35A1661E982C15.

Shi:1999:CCT

[966] Peng Shi. Control of continuous-time systems with discrete jumps. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):58–82, July 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/control-of-continuoustime-systems-with-discrete-jumps/
086F0DB502E4B814DB1374717A541EA6.

Heidel:1999:WCS

[967] Jack Heidel and John Maloney. When can sigmoidal data be fit to a Hill curve? Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):83–92, July 1999. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/when-can-sigmoidal-data-be-fit-to-a-
hill-curve/6C399348CE328230A6622BD9B0484E06.

Subramaniam:1999:ESC

[968] R. Subramaniam and K. Balachandran. Existence of solutions of a class of stochastic Volterra integral equations with applications to chemotherapy. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):93–104, July 1999. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/existence-of-solutions-of-a-class-of-stochastic-
volterra-integral-equations-with-applications-to-chemotherapy/
4B6E59C7DC91241583F807D94726A095.

Ali:1999:GHT

[969] I. Ali and S. Kalla. A generalized Hankel transform and its use for solving certain partial differential equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41 (1):105–117, July 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalized-hankel-transform-and-its-use-for-solving-
certain-partial-differential-equations/7F2939201F75DC18C92F72902A19C089.

Makinde:1999:EUP

[970] O. D. Makinde. Extending the utility of perturbation series in problems of laminar flow in a porous pipe and a diverging channel. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):118–128, July 1999. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/
core/journals/anziam-journal/article/extending-the-utility-
of-perturbation-series-in-problems-of-laminar-flow-in-a-porous-
pipe-and-a-diverging-channel/3D746A6995889ACB945929F9C0AD2639.

Hird:1999:NSF

[971] L. D. Hird, P. F. Siew, and S. Wang. A numerical solution to the flow between eccentric rotating cylinders with a slotted sleeve. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):129–152, July 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-solution-to-the-flow-between-eccentric-rotating-
cylinders-with-a-slotted-sleeve/3575E6C1B35B477274E95F846A2E8CCD.

Anonymous:1999:AVIe

[972] Anonymous. ANZ volume 41 issue 1 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):f1–f2, July 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-41-issue-1-cover-and-front-matter/A982D55EDB0E4B006DDDA7154D466347.

Anonymous:1999:AVIf

[973] Anonymous. ANZ volume 41 issue 1 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(1):b1–b2, July 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-41-issue-1-cover-and-back-matter/2D471A34B4F52782E6FB545DB65FB75F.

Pearce:1999:P

[974] Charles Pearce. Preface. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):153, October 1999. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/preface/
DE6787E64D73987177B21FE74F09CC13.

Toh:1999:KG

[975] Tze-Chuen Toh and Malcolm R. Anderson. Knots and gravity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):154–160, October 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/knots-and-gravity/1FCBA2C3997C6A8D40E8724D44D00F02.

Chan:1999:FS

[976] Wai Kin Chan, Reynaldo Castillo, and King Fai Lai. Foliations in supergravity. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):161–166, October 1999. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/foliations-in-supergravity/
D48313B7D802C233B14087E578FB86B5.

Szekeres:1999:WSC

[977] Peter Szekeres and Anthony Lun. What is a shell-crossing singularity? Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):167–179, October 1999. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/what-is-a-shellcrossing-singularity/
D96D70A5FF67A5B7C62C65E23D2B06FC.

Anderson:1999:NFE

[978] Malcolm Anderson. Near-field expansion of the metric due to a cosmic string. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):180–197, October 1999. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/nearfield-expansion-of-the-
metric-due-to-a-cosmic-string/F915771C188C130CD172E48BD08FBA02.

Wiltshire:1999:DBH

[979] David L. Wiltshire. Dilaton black holes with a cosmological term. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):198–216, October 1999. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/dilaton-black-holes-with-a-
cosmological-term/2BC803CB2454FD601625B912F550FEEF.

Chow:1999:AHV

[980] E. W. M. Chow and A. W.-C. Lun. Apparent horizons in vacuum Robinson–Trautman spacetimes. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):217–230, October 1999. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/apparent-horizons-
in-vacuum-robinsontrautman-spacetimes/68C36DF148C06C5DFC2E55862701897A.

Klenk:1999:ESV

[981] Jürgen Klenk. Existence of stationary vacuum solutions of Einstein’s equations in an exterior domain. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41 (2):231–247, October 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-of-stationary-vacuum-solutions-of-einsteins-
equations-in-an-exterior-domain/84ADD05C616168A2721349A9B25DCE83.

Prince:1999:SHJ

[982] G. E. Prince, J. E. Aldridge, S. E. Godfrey, and G. B. Byrnes. The separation of the Hamilton–Jacobi equation for the Kerr metric. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):248–259, October 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/separation-of-the-hamiltonjacobi-equation-for-the-kerr-
metric/021A05353D8FA4AE9851F1599A39B6F0.

Fernandes:1999:ICB

[983] J. F. Q. Fernandes and A. W.-C. Lun. Integrability conditions for the Bianchi identities as transformations in Schwarzschild space–time. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):260–270, October 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/integrability-conditions-for-the-bianchi-identities-
as-transformations-in-schwarzschild-spacetime/B51AD89AF4F57668496035D1EBB3FAF4.

Prager:1999:NIA

[984] D. A. Prager and A. W.-C. Lun. Numerical integration of the axisymmetric Robinson–Trautman equation by a spectral method. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):271–280, October 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-integration-of-the-axisymmetric-robinsontrautman-
equation-by-a-spectral-method/03B1FC60DE8F9B9FEC8DECA52368FC8F.

Anonymous:1999:AVIg

[985] Anonymous. ANZ volume 41 issue 2 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):f1–f3, October 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-41-issue-2-cover-and-front-matter/517CB30F2B2816160084EEB5BD2780A5.

Anonymous:1999:AVIh

[986] Anonymous. ANZ volume 41 issue 2 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(2):b1–b2, October 1999. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-41-issue-2-cover-and-back-matter/077B159306CE1EBF3337D6CFE09CFF90.

Philip:2000:IPS

[987] J. R. Philip. Instantaneous point source solutions in nonlinear diffusion with nonlinear loss or gain. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41 (3):281–300, January 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/instantaneous-point-source-solutions-in-nonlinear-diffusion-
with-nonlinear-loss-or-gain/480DDB2A49FD220CC63C274910E78409.

Bu:2000:FCS

[988] Charles Bu. Forced cubic Schrödinger equation with Robin boundary data: continuous dependency result. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41 (3):301–311, January 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/forced-cubic-schrodinger-equation-with-robin-boundary-
data-continuous-dependency-result/575D7C415BAEB222D7E03AC3F55928C4.

Wang:2000:FII

[989] Weichung Wang. Final iterations in interior point methods — preconditioned conjugate gradients and modified search directions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(3):312–328, January 2000. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/final-iterations-in-interior-point-methods-
preconditioned-conjugate-gradients-and-modified-search-directions/
CA3DC13B833435449F7A596D552EB7EE.

Elezovic:2000:TLB

[990] N. Elezović, M. Matić, C. E. M. Pearce, and J. Pecarić. On two lemmas of Brown and Shepp having application to sum sets and fractals, III. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(3):329–337, January 2000. CODEN JAM-

MDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-two-lemmas-of-brown-and-
shepp-having-application-to-sum-sets-and-fractals-iii/580E4AD12EFBF2F7E66DCE5E6B54C58E.

Koerber:2000:ATT

[991] A. J. Koerber and L. K. Forbes. An analysis of two and three dimensional unsteady withdrawal flows, using shallow water theory. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(3):338–357, January 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/an-analysis-of-two-and-three-dimensional-unsteady-withdrawal-
flows-using-shallow-water-theory/35D67892E8CA8E859C86A4F0F2DE221A.

Barton:2000:PC

[992] J. C. Barton and C. J. Eliezer. On pursuit curves. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41 (3):358–371, January 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-pursuit-curves/8C87CB2301EE8DE3155A81C62BB5943E.

Wang:2000:AHJ

[993] Shihong Wang and Zuoyi Zhou. Averaging of the Hamilton–Jacobi equation in infinite dimensions and an application. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41 (3):372–385, January 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/averaging-of-the-hamiltonjacobi-equation-in-infinite-
dimensions-and-an-application/199D8CFBCD64C9ECF5410917F4B91C2E.

Kim:2000:GSS

[994] Hong Oh Kim and Jong Ha Park. The generalized Shannon system in wavelet space. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(3):386–400, January 2000. CODEN

JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/core/
journals/anziam-journal/article/generalized-shannon-system-
in-wavelet-space/57FF55207F819C72298E986C201F9C54.

Maloney:2000:RHT

[995] John Maloney, Jack Heidel, and Josip Pecarić. A reverse Hölder type inequality for the logarithmic mean and generalizations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(3):401–409, January 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reverse-holder-type-inequality-for-the-logarithmic-
mean-and-generalizations/6993582B91AA258F2C962AA469945076.

Heidel:2000:AFM

[996] Jack Heidel and John Maloney. An analysis of a fractal Michaelis–Menten curve. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(3):410–422, January 2000. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-analysis-of-a-fractal-michaelismenten-
curve/AF294EDF306405FC1ECBC494CAA5096A.

Anonymous:2000:AVIa

[997] Anonymous. ANZ volume 41 issue 3 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(3):f1–f2, January 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-41-issue-3-cover-and-front-matter/C8F26F02621F787196B3596AD5A3DA00.

Anonymous:2000:AVIb

[998] Anonymous. ANZ volume 41 issue 3 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(3):b1–b2, January 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-41-issue-3-cover-and-back-matter/A39307850852ECE2FB8BAFCDFBDAA6FF.

Coolen-Schrijner:2000:QCT

[999] Pauline Coolen-Schrijner, Andrew Hart, and Phil Pollett. Quasistationarity of continuous-time Markov chains with positive drift. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):423–441, April 2000. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/quasistationarity-of-continuoustime-markov-
chains-with-positive-drift/1F403FA9DF09759D448BE83DA529DCC9.

Nouri:2000:EPS

[1000] A. Nouri. An Euler–Poisson system in plasmas. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4): 442–450, April 2000. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
eulerpoisson-system-in-plasmas/5E2CFD2A4461F24CBFF67F072A0D755A.

Wendi:2000:DDP

[1001] Wang Wendi and Tang Chunlei. Dynamics of a delayed population model with feedback control. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):451–457, April 2000. CO-DEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.org/

core/journals/anziam-journal/article/dynamics-of-a-delayed-
population-model-with-feedback-control/157BD8A1097CAF47CF89590EAF5E1645.

Wiryanto:2000:OCF

[1002] L. H. Wiryanto and E. O. Tuck. An open-channel flow meeting a barrier and forming one or two jets. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):458–472, April 2000. CODEN JAMMDU. ISSN 0334-2700. URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
openchannel-flow-meeting-a-barrier-and-forming-one-or-two-jets/
F4F9AE5BDF211D8EE9B203B454DB8E33.

Cerone:2000:SSA

[1003] P. Cerone and A. Sofo. Summing series arising from integrodifferential-difference equations. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):473–486, April 2000. CODEN JAMMDU. ISSN 0334-2700. URL https://www.

cambridge.org/core/journals/anziam-journal/article/summing-
series-arising-from-integrodifferentialdifference-equations/
7573FA3CED618F9AAB2E951C570FFEF2.

Alam:2000:ASRa

[1004] Rafikul Alam, Rekha P. Kulkarni, and Balmohan V. Limaye. Accelerated spectral refinement. Part I: simple eigenvalue. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):487–507, April 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/accelerated-spectral-refinement-part-i-simple-eigenvalue/
3DD4F07C3E7CB531C81FA8909D77E236.

Young:2000:AOL

[1005] Brian Young. Analysis and optimisation of looped water distribution networks. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):508–526, April 2000. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/analysis-and-optimisation-of-looped-
water-distribution-networks/7E6F528F8735F36F3FE00A57C84BB7BB.

Liu:2000:EBS

[1006] Huan-Wen Liu. An expansion of bivariate spline functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):527–541, April 2000. CODEN JAMMDU. ISSN 0334-

2700. URL https://www.cambridge.org/core/journals/anziam-
journal/article/an-expansion-of-bivariate-spline-functions/
3EC510018FAE2AFF31444D3BDB2CD670.

Gao:2000:NCO

[1007] Hang Gao and Xunjing Li. Necessary conditions for optimal control of elliptic systems. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):542–567, April 2000. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/necessary-conditions-for-optimal-control-
of-elliptic-systems/1FC0309E930A2C9B956184CE1FCA7D24.

Graef:2000:GAD

[1008] J. R. Graef and C. Qian. Global attractivity in differential equations with variable delays. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):568–579, April 2000. CODEN JAMMDU.

ISSN 0334-2700. URL https://www.cambridge.org/core/journals/
anziam-journal/article/global-attractivity-in-differential-
equations-with-variable-delays/93D825DF0E85B6C099B70B8BFF0E45F0.

Anonymous:2000:CV

[1009] Anonymous. Content of volume 41. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):580–582, April 2000. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/content-of-volume-
41/7540737A0384C684E9CC2F6CF7ECACE6.

Anonymous:2000:AVIc

[1010] Anonymous. ANZ volume 41 issue 4 cover and front matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):f1–f2, April 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-41-issue-4-cover-and-front-matter/B42E886608824F4E5321A87DBA07CC89.

Anonymous:2000:AVId

[1011] Anonymous. ANZ volume 41 issue 4 cover and back matter. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 41(4):b1–b2, April 2000. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-journal/
article/anz-volume-41-issue-4-cover-and-back-matter/1449CADF0E7E5F7FB5F13368EFD34F57.

Pearce:2000:PV

[1012] Charles E. M. Pearce. Preface to this volume. The ANZIAM Journal, 42(1):1, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/preface-to-this-volume/3A3629ADAD4B815A040EBC0DD9823A8D.

Donaldson:2000:PI

[1013] John D. Donaldson. Preface to this issue. The ANZIAM Journal, 42(1):1–2, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/preface-to-this-issue/
89FA9C1CA9E5CB9CB3EADB665C9DC581.

Sloan:2000:MII

[1014] Ian H. Sloan. Multiple integration is intractable but not hopeless. The ANZIAM Journal, 42(1):3–8, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/multiple-integration-
is-intractable-but-not-hopeless/2E5E5748E4D654F9C724EFF68E918DF7.

Osborne:2000:SC

[1015] Michael R. Osborne. Scoring with constraints. The ANZIAM Journal, 42(1):9–25, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/scoring-with-constraints/7DB6FD389CFB92A0D6F14D1E16FFD7D7.

Anderssen:2000:MWD

[1016] R. S. Anderssen and M. Westcott. The molecular weight distribution problem and reptation mixing rules. The ANZIAM Journal, 42(1):26–40, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/molecular-weight-distribution-problem-and-reptation-
mixing-rules/CCD5EAD3A9462B42C809F9F9B233A91E.

Anh:2000:WHI

[1017] V. V. Anh, W. Grecksch, J. M. Angulo, and M. D. Ruiz-Medina. The Wiener–Hopf integral equation for fractional Riesz–Bessel motion. The ANZIAM Journal, 42(1):41–54, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/wienerhopf-
integral-equation-for-fractional-rieszbessel-motion/CA79B390A2CFC97FC517A35630AA7F0C.

Swift:2000:SNE

[1018] Adrian Swift and Easwaran Balakrishnan. Solution of nonlinear equations and computation of multiple solutions of a simple reaction– diffusion equation. The ANZIAM Journal, 42(1):55–64, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/solution-of-nonlinear-equations-and-computation-of-
multiple-solutions-of-a-simple-reactiondiffusion-equation/D84DCA48E64FE21D68E721C8CE13191B.

Kress:2000:IEM

[1019] Rainer Kress. Integral equation methods in inverse obstacle scattering. The ANZIAM Journal, 42(1):65–78, July 2000. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/integral-equation-methods-in-inverse-obstacle-scattering/
D4F49AB00C9D89921CBBE374C67A9AD1.

Phillips:2000:GBP

[1020] George M. Phillips. A generalization of the Bernstein polynomials based on the q-integers. The ANZIAM Journal, 42(1):79–86, July 2000.

CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalization-of-the-bernstein-polynomials-based-on-
the-qintegers/711CD84932645DD06DF9BA3F65B6F388.

Stenger:2000:SAC

[1021] Frank Stenger. Sinc approximation of Cauchy-type integrals over arcs. The ANZIAM Journal, 42(1):87–97, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/sinc-approximation-
of-cauchytype-integrals-over-arcs/F3E8E58B66C08F142BB6C63D61317703.

Papers in honour of David Elliott on the occasion of his sixty-fifth birthday.

Byrne:2000:LFG

[1022] Graeme J. Byrne, T. M. Mills, and Simon J. Smith. The Lebesgue function for generalized Hermite–Fejér interpolation on the Chebyshev nodes. The ANZIAM Journal, 42(1):98–109, July 2000. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/lebesgue-function-for-generalized-hermitefejer-interpolation-
on-the-chebyshev-nodes/59DA7F8D13A68000BB947490DBBA347F.

Hosking:2000:AEI

[1023] Roger J. Hosking. Approximate evaluation of integrals. The ANZIAM Journal, 42(1):110–118, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/approximate-evaluation-
of-integrals/5E2D65D551748A9AAE50125747FD9F82.

Champion:2000:VAS

[1024] R. Champion, C. T. Lenard, and T. M. Mills. A variational approach to splines. The ANZIAM Journal, 42(1):119–135, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/variational-approach-to-splines/21DBC999A17F5F68C6587000322EEFF5.

Andrew:2000:QEF

[1025] Alan L. Andrew. Quadrature errors in finite element eigenvalue computations. The ANZIAM Journal, 42(1):136–140, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/

article/quadrature-errors-in-finite-element-eigenvalue-computations/
CA0B0A15D07A9686A827C6990A1B9813.

Ainsworth:2000:DSS

[1026] Mark Ainsworth, Bill McLean, and Thanh Tran. Diagonal scaling of stiffness matrices in the Galerkin boundary element method. The ANZIAM Journal, 42(1):141–150, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/diagonal-scaling-of-stiffness-
matrices-in-the-galerkin-boundary-element-method/B0449020C0E00E242F7E1AA338E6935A.

Capobianco:2000:UCC

[1027] M. R. Capobianco, G. Criscuolo, P. Junghanns, and U. Luther. Uniform convergence of the collocation method for Prandtl’s integrodifferential equation. The ANZIAM Journal, 42(1):151–168, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/uniform-convergence-of-the-collocation-method-for-prandtls-
integrodifferential-equation/BCD7DAE215BBA8A1480F95820461BE30.

Pham:2000:AFG

[1028] Binh Pham. Aesthetic factors in geometric modelling. The ANZIAM Journal, 42(1):169–180, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/aesthetic-factors-
in-geometric-modelling/7C3F4FD0092BEAB07E951CED8F880A0C.

Lyness:2000:BSE

[1029] J. N. Lyness. A brief survey of extrapolation quadrature. The ANZIAM Journal, 42(1):181–185, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/brief-survey-of-extrapolation-
quadrature/29208D66FF5F1E2318463B8D49F733DD.

Anonymous:2000:AVIe

[1030] Anonymous. ANZ volume 42 issue 1 cover and front matter. The ANZIAM Journal, 42(1):f1–f5, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-42-issue-1-cover-and-front-matter/AA9B2BD1A2E533CE9A2EB1A58EF0085D.

Anonymous:2000:AVIf

[1031] Anonymous. ANZ volume 42 issue 1 back matter. The ANZIAM Journal, 42(1):b1, July 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-42-issue-1-
back-matter/02EBA03C15954AB75D40BB02BDA6B98C.

Bragg:2000:DTA

[1032] L. R. Bragg. Derivative-type ascent formulas for kernels of some halfspace Dirichlet problems. The ANZIAM Journal, 42(2):185–194, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/derivativetype-ascent-formulas-for-kernels-of-some-
halfspace-dirichlet-problems/5DDC20B96AE1D0835280B29AF22AF852.

vanVuuren:2000:CPW

[1033] Jan H. van Vuuren and John Norbury. Conditions for permanence in wellknown biological competition models. The ANZIAM Journal, 42(2):195– 223, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/conditions-for-permanence-in-wellknown-
biological-competition-models/42FD3DCDA10A017E8BC884BA425FF237.

Alam:2000:ASRb

[1034] Rafikul Alam, Rekha P. Kulkarni, and Balmohan V. Limaye. Accelerated spectral refinement. Part II: Cluster of eigenvalues. The ANZIAM Journal, 42(2):224–243, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/accelerated-spectral-refinement-
part-ii-cluster-of-eigenvalues/91EA6E763BD5A353D4F4F050553B55BE.

Boyd:2000:LVR

[1035] J. N. Boyd and P. N. Raychowdhury. Lattice vibrations with Rayleigh dissipation. The ANZIAM Journal, 42(2):244–253, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/lattice-vibrations-with-rayleigh-dissipation/CFD76E0D2A9D870C99DCB66067CC0B8A.

Jukic:2000:BLS

[1036] D. Jukić and R. Scitovski. The best least squares approximation problem for a 3-parametric exponential regression model. The ANZIAM Journal, 42(2):254–266, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/best-least-squares-approximation-
problem-for-a-3parametric-exponential-regression-model/7718BDC2113058BBEA73D0F97DEF76A6.

Varosanec:2000:SII

[1037] Sanja Varosanec and Josip Pecarić. Some integral inequalities with bounds for moments of distribution II. The ANZIAM Journal, 42(2): 267–276, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/some-integral-inequalities-
with-bounds-for-moments-of-distribution-ii/66D363D86E595321511714BC27757EA4.

Chakrabarti:2000:SPS

[1038] A. Chakrabarti. On the solution of the problem of scattering of surface water waves by a sharp discontinuity in the surface boundary conditions. The ANZIAM Journal, 42(2):277–286, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-solution-of-the-problem-of-scattering-of-surface-water-
waves-by-a-sharp-discontinuity-in-the-surface-boundary-conditions/
0785CE64F3C4F1A1FA882A09730A7196.

Jeon:2000:QMC

[1039] Youngmok Jeon. A quadrature method for constant-coefficient Cauchy singular integral equations on an interval. The ANZIAM Journal, 42(2):287–311, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/quadrature-method-
for-constantcoefficient-cauchy-singular-integral-equations-
on-an-interval/84A206809CFEFC8B05519070AC8F3C32.

Anonymous:2000:AVIg

[1040] Anonymous. ANZ volume 42 issue 2 cover and front matter. The ANZIAM Journal, 42(2):f1–f4, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-42-issue-2-cover-and-front-matter/13FBC9D0F1F1A5F4031C1A09F9AB0DE1.

Anonymous:2000:AVIh

[1041] Anonymous. ANZ volume 42 issue 2 back matter. The ANZIAM Journal, 42(2):b1–b2, October 2000. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-42-issue-2-
back-matter/E6511D2B54C0B63B86512122B99F6952.

Selvaratnam:2001:SCS

[1042] A. R. Selvaratnam, M. Vlieg-Hulstman, B. van Brunt, and W. D. Halford. On the solution of a class of second-order quasi-linear PDEs and the Gauss equation. The ANZIAM Journal, 42(3):312– 323, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-solution-of-a-class-
of-secondorder-quasilinear-pdes-and-the-gauss-equation/D0A7C34078E8A9196B5BE532207FC9D0.

Mays:2001:BPS

[1043] Laurence Mays and John Norbury. Bifurcation of positive solutions for a Neumann boundary value problem. The ANZIAM Journal, 42(3):324– 340, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/bifurcation-of-positive-solutions-for-
a-neumann-boundary-value-problem/C0B51289D6DB41AE55BF251832973CAE.

Bainov:2001:VLF

[1044] D. D. Bainov and I. M. Stamova. Vector Lyapunov functions and conditional stability for systems of impulsive differential-difference equations. The ANZIAM Journal, 42(3):341–353, January 2001. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/vector-lyapunov-functions-and-conditional-stability-
for-systems-of-impulsive-differentialdifference-equations/FDBCB157C434C0D06098573AAB0F9975.

McCue:2001:SAB

[1045] S. W. McCue and L. K. Forbes. Smoothly attaching bow flows with constant vorticity. The ANZIAM Journal, 42(3):354–371, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/smoothly-attaching-bow-flows-with-constant-
vorticity/779F55D6DA1F1FC2BD8D94ACE59AD403.

Gutierrez:2001:ANM

[1046] J. M. Gutiérrez and M. A. Hernández. An application of Newton’s method to differential and integral equations. The ANZIAM Journal, 42(3):372–386, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/an-application-of-
newtons-method-to-differential-and-integral-equations/06CA0C8686FD19A8CB5367EB1EA2752A.

Matic:2001:RSB

[1047] M. Matić, C. E. M. Pearce, and J. Pecarić. Refinements of some bounds in information theory. The ANZIAM Journal, 42(3):387– 398, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/refinements-of-some-bounds-
in-information-theory/6817AB039A6B90AB9E3B7DE0CF537452.

Christie:2001:CPL

[1048] J. R. Christie, K. Gopalsamy, and Jibin Li. Chaos in perturbed Lotka– Volterra systems. The ANZIAM Journal, 42(3):399–412, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/chaos-in-perturbed-lotkavolterra-systems/915D88955BDE6128D96107BC9B3DD215.

Celorrio:2001:ETC

[1049] Ricardo Celorrio and Francisco-Javier Sayas. Extrapolation techniques and the collocation method for a class of boundary integral equations. The ANZIAM Journal, 42(3):413–437, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/extrapolation-techniques-and-the-collocation-method-
for-a-class-of-boundary-integral-equations/8FA24AAD0C7C12E280A531FCFA27E7A8.

Guo:2001:ODE

[1050] Jong-Shenq Guo and Yung-Jen Lin Guo. An ordinary differential equation arising in the Ricci flow on the plane. The ANZIAM Journal, 42(3):438–444, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
ordinary-differential-equation-arising-in-the-ricci-flow-on-
the-plane/64F7809DFEB0A84AB67847FF896F14C4.

Brown:2001:FDL

[1051] A. Brown. Fast diffusion with loss at infinity-additional solutions. The ANZIAM Journal, 42(3):445–450, January 2001. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fast-diffusion-with-loss-at-infinityadditional-solutions/
CA0629A2916B0E486DEF04606EE70D0E.

Anonymous:2001:AVIa

[1052] Anonymous. ANZ volume 42 issue 3 cover and front matter. The ANZIAM Journal, 42(3):f1–f4, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-42-issue-3-cover-and-front-matter/B867890BBF777A304E0EAFD34650EA30.

Anonymous:2001:AVIb

[1053] Anonymous. ANZ volume 42 issue 3 back matter. The ANZIAM Journal, 42(3):b1–b2, January 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-42-issue-3-
back-matter/6CDDFF94A39E503CBCE5FAC637EB93A7.

Stewart:2001:TNE

[1054] David E. Stewart. Towards numerically estimating Hausdorff dimensions. The ANZIAM Journal, 42(4):451–461, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/towards-
numerically-estimating-hausdorff-dimensions/711A3BBBF930C6F8E5051DE5E2B695D3.

Rubinov:2001:EFB

[1055] A. M. Rubinov and B. M. Glover. Equilibrium with fixed budgets and superlinear connections. The ANZIAM Journal, 42(4):462–480, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/equilibrium-with-fixed-budgets-and-superlinear-connections/
7407A1608C3289245778284A744FF5C2.

Weir:2001:EQS

[1056] Graham J. Weir. Early quasi-steady electro-magnetic fields about conducting surfaces. The ANZIAM Journal, 42(4):481–493, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/early-quasisteady-electromagnetic-fields-about-conducting-
surfaces/4ABDBBF31DFE2374DF76EBCFB82B4278.

Avkhadiev:2001:BEC

[1057] F. G. Avkhadiev and A. M. Elizarov. Bilateral estimates of the critical Mach number for some classes of carrying wing profiles. The ANZIAM Journal, 42(4):494–503, April 2001. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/bilateral-estimates-of-the-critical-mach-number-for-
some-classes-of-carrying-wing-profiles/FE1DE6F363C5BAFCA2E826D90498A86C.

Bainov:2001:LSI

[1058] D. D. Bainov and I. M. Stamova. Lipschitz stability of impulsive functional-differential equations. The ANZIAM Journal, 42(4):504–514, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/lipschitz-stability-of-impulsive-functionaldifferential-
equations/686C18F3A2E4C61C2E84BF38E273154D.

Dragomir:2001:MSP

[1059] S. S. Dragomir and C. J. Goh. On monotonicity and superadditivity properties of the entropy function. The ANZIAM Journal, 42(4): 515–531, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-monotonicity-and-superadditivity-
properties-of-the-entropy-function/684D3DD6DE16650BB5FB4F2DF347950B.

Pan:2001:OCQ

[1060] Liping Pan and Jiongmin Yong. Optimal control for quasilinear retarded parabolic systems. The ANZIAM Journal, 42(4):532–551, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-control-for-quasilinear-retarded-parabolic-
systems/ED9DAEBFE15AC277A13B5F31688E7A6D.

Zhang:2001:CLO

[1061] B. G. Zhang and Jian-She Yu. Comparison and linearized oscillation theorems for a nonlinear partial difference equation. The ANZIAM Journal, 42(4):552–560, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/comparison-and-linearized-oscillation-
theorems-for-a-nonlinear-partial-difference-equation/320FA7BF238880C3826A872592874D29.

Hongliang:2001:ABN

[1062] Zhu Hongliang and Cui Jingan. Asymptotic behaviour of a nonautonomous cooperative system. The ANZIAM Journal, 42(4):561–568, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/asymptotic-behaviour-of-a-nonautonomous-cooperative-
system/3B54E6FA83FC0EFBBCAADEA87CD43994.

Li:2001:PMM

[1063] Yongkun Li. On a periodic mutualism model. The ANZIAM Journal, 42 (4):569–580, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-a-periodic-mutualism-model/
F73FA171C6A029D9A85811B5BBB7E456.

Anonymous:2001:CV

[1064] Anonymous. Contents of volume 42. The ANZIAM Journal, 42 (4):581–583, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/contents-of-volume-42/
EBA3A6116F011A9FD11F9539697542DF.

Anonymous:2001:AVIc

[1065] Anonymous. ANZ volume 42 issue 4 cover and front matter. The ANZIAM Journal, 42(4):f1–f4, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-42-issue-4-cover-and-front-matter/8610B31F73CE6B7FB761C6DE45AF89A8.

Anonymous:2001:AVId

[1066] Anonymous. ANZ volume 42 issue 4 back matter. The ANZIAM Journal, 42(4):b1–b2, April 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-42-issue-4-
back-matter/C8AD05632678422162205D7D380CD73F.

Weber:2001:CMH

[1067] Rodney Weber, Harvinder Sidhu, and Geoffrey Mercer. Combustion meeting in honour of Professor B. F. Gray. The ANZIAM Journal, 43(1):i, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/combustion-meeting-in-
honour-of-professor-b-f-gray/FC39BFF3E0FCC25E30D9AF03B2ACD395.

Pritchard:2001:HBG

[1068] Huw O. Pritchard. How Brian Gray became a combustion scientist. The ANZIAM Journal, 43(1):ii–iii, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/how-
brian-gray-became-a-combustion-scientist/6728C088F50F536BC17A36CBAD90A1EE.

Wake:2001:SPG

[1069] Graeme Wake. Speech by Professor G. C. Wake for the Professor B. F. Gray dinner. The ANZIAM Journal, 43(1):iv–vi, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/speech-by-professor-g-c-wake-for-the-professor-b-f-
gray-dinner/64249975C58B983C6C15E6A7000A210C.

Gray:2001:CCA

[1070] B. F. Gray. On the critical conditions for an assembly of interacting thermons. The ANZIAM Journal, 43(1):1–11, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-critical-conditions-for-an-assembly-of-interacting-thermons/
EA1D2E4A0ECDF14D9D2A734E90A510D5.

Macaskill:2001:RDM

[1071] C. Macaskill, M. J. Sexton, and B. F. Gray. A reaction–diffusion model of stored bagasse. The ANZIAM Journal, 43(1):13–34, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reactiondiffusion-model-of-stored-bagasse/57D5EE8A9807D249DAC7BD22E08B6053.

Forbes:2001:CWP

[1072] Lawrence K. Forbes and William Derrick. A combustion wave of permanent form in a compressible gas. The ANZIAM Journal, 43(1):35–58, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/combustion-wave-of-permanent-form-in-a-compressible-
gas/C0B3A7FBFACAFF4DF5E90D171DCFC643.

Mcintosh:2001:SNA

[1073] A. C. Mcintosh, B. F. Gray, G. C. Wake, and R. Ball. The stability of a near-adiabatic Endex batch CSTR reactor. The ANZIAM Journal, 43 (1):59–75, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/stability-of-a-nearadiabatic-endex-batch-
cstr-reactor/DBDE6FC3A62F349BF6FDF46E6906194C.

Weber:2001:CIU

[1074] R. O. Weber, G. C. Wake, H. S. Sidhu, G. N. Mercer, B. F. Gray, W. Derrick, and E. Balakrishnan. On the crossing of intermediate unstable steady state solutions for thermal ignition in a sphere. The ANZIAM Journal, 43(1):77–85, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-crossing-of-intermediate-unstable-steady-state-solutions-
for-thermal-ignition-in-a-sphere/8E8F3A36E095564B46DC1B9957D6D76A.

Nelson:2001:MSS

[1075] M. I. Nelson, G. C. Wake, X. D. Chen, and E. Balakrishnan. The multiplicity of steady-state solutions arising from microwave heating. I. Infinite Biot number and small penetration depth. The ANZIAM Journal, 43 (1):87–103, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/multiplicity-of-steadystate-solutions-
arising-from-microwave-heating-i-infinite-biot-number-and-small-
penetration-depth/D587BDF833D21BF5FFBF94F79913F231.

Nelson:2001:DSM

[1076] M. I. Nelson, H. S. Sidhu, R. O. Weber, and G. N. Mercer. A dynamical systems model of the limiting oxygen index test. The ANZIAM Journal, 43(1):105–117, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/dynamical-systems-model-of-
the-limiting-oxygen-index-test/CBD7470582B5F0B6E6CEFCB5FFDCC794.

Graham-Eagle:2001:CWR

[1077] J. Graham-Eagle and D. A. Schult. Combustion waves with reactant depletion. The ANZIAM Journal, 43(1):119–135, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/combustion-waves-with-reactant-depletion/91ACDC92A5C6350C95D3374B67F43BE9.

Marchant:2001:HTD

[1078] T. R. Marchant and B. Liu. On the heating of a two-dimensional slab in a microwave cavity: aperture effects. The ANZIAM Journal, 43(1):137–148, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/on-the-heating-of-
a-twodimensional-slab-in-a-microwave-cavity-aperture-effects/
3F9A20AB431E56CDA90CAC8614210823.

Brindley:2001:ICW

[1079] J. Brindley, J. F. Griffiths, A. C. McIntosh, and J. Zhang. Initiation of combustion waves in solids, and the effects of geometry. The ANZIAM Journal, 43(1):149–163, July 2001. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/initiation-of-combustion-waves-in-solids-and-the-effects-
of-geometry/5D31456780A1AD10AE093B563F302AB4.

Chen:2001:SHB

[1080] Xiao Dong Chen. Self-heating behaviour of low moisture content particles-modelling the basket-heating of solid particles and some aspects of the cross over behaviour using milk powder as an example. The ANZIAM Journal, 43(1):165–181, July 2001. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/selfheating-behaviour-of-low-moisture-content-particlesmodelling-
the-basketheating-of-solid-particles-and-some-aspects-of-the-
cross-over-behaviour-using-milk-powder-as-an-example/D25614FB476A13AEAFA09741661EA5AA.

Anonymous:2001:AVIe

[1081] Anonymous. ANZ volume 43 issue 1 cover and front matter. The ANZIAM Journal, 43(1):f1–f4, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-43-issue-1-cover-and-front-matter/B0C6CDF60F44F6BDA9808E7567D4F927.

Anonymous:2001:AVIf

[1082] Anonymous. ANZ volume 43 issue 1 cover and back matter. The ANZIAM Journal, 43(1):b1–b3, July 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-43-issue-1-cover-and-back-matter/67C8265F4D70237400EBBAA3589CFF8C.

Ivancevic:2001:TDH

[1083] V. Ivancevic and C. E. M. Pearce. Topological duality in humanoid robot dynamics. The ANZIAM Journal, 43(2):183–194, October 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/topological-duality-in-humanoid-robot-dynamics/CA0F6EE402B7C638F797EC62CDEE2AD1.

Smith:2001:WMC

[1084] William V. Smith. Wave motion in a conducting fluid with a layer adjacent to the boundary, II. Eigenfunction expansions. The ANZIAM Journal, 43(2):195–236, October 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/wave-motion-in-a-conducting-
fluid-with-a-layer-adjacent-to-the-boundary-ii-eigenfunction-
expansions/B9FD7A42DB1FF2B55F89694D36897676.

Belward:2001:SMO

[1085] Shaun R. Belward and Lawrence K. Forbes. A simple model for oilspill containment. The ANZIAM Journal, 43(2):237–246, October 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/simple-model-for-oilspill-containment/889E73230B0FB3C52BA0C5AACCD1EAFC.

Kuang:2001:GSP

[1086] Yang Kuang. Global stability and persistence in diffusive food chains. The ANZIAM Journal, 43(2):247–268, October 2001. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/global-stability-and-persistence-in-diffusive-food-
chains/1DDDD60C1972DB3E0F84DA1DAA8F4EAD.

Bainov:2001:SSI

[1087] D. D. Bainov and I. M. Stamova. Stability of the solutions of impulsive functional-differential equations by Lyapunov’s direct method. The ANZIAM Journal, 43(2):269–278, October 2001. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stability-of-the-solutions-of-impulsive-functionaldifferential-
equations-by-lyapunovs-direct-method/FD035D5A98FE259AF32AAFCA36FC3BCD.

Ma:2001:MMB

[1088] Liming Ma and Qianshun Chang. Multigrid methods for the biharmonic equation using some nonconforming plate elements. The ANZIAM Journal, 43(2):279–289, October 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/multigrid-methods-
for-the-biharmonic-equation-using-some-nonconforming-plate-
elements/97126B6C1E5D9F4EB0E0D743FF96BDB9.

Raina:2001:SSA

[1089] R. K. Raina, H. M. Srivastava, A. A. Kilbas, and M. Saigo. Solvability of some Abel-type integral equations involving the Gauss hypergeometric function as kernels in the spaces of summable functions. The ANZIAM Journal, 43(2):291–320, October 2001. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/solvability-of-some-abeltype-integral-equations-involving-
the-gauss-hypergeometric-function-as-kernels-in-the-spaces-
of-summable-functions/CB2E2A901D8BF04FBAE1C82064D31D9C.

Anonymous:2001:AVIg

[1090] Anonymous. ANZ volume 43 issue 2 cover and front matter. The ANZIAM Journal, 43(2):f1–f4, October 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-43-issue-2-cover-and-front-matter/04530FACC4CCD83A44622D2FAE593C50.

Anonymous:2001:AVIh

[1091] Anonymous. ANZ volume 43 issue 2 cover and back matter. The ANZIAM Journal, 43(2):b1–b3, October 2001. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-43-issue-2-cover-and-back-matter/34250534B2F0A5F17014DCC9EFFD38AB.

Kang:2002:SWV

[1092] Y. Kang and J.-M. Vanden-Broeck. Stern waves with vorticity. The ANZIAM Journal, 43(3):321–332, January 2002. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stern-waves-with-vorticity/DE99C9F5726C9CB64B394912939785D0.

Stott:2002:SBL

[1093] Jillian A. K. Stott and James P. Denier. The stability of boundary layers on curved heated plates. The ANZIAM Journal, 43 (3):333–358, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/stability-of-boundary-
layers-on-curved-heated-plates/FD7B66A37113705AA438F6C8A06DF5AD.

Liu:2002:ACP

[1094] Xinzhi Liu, S. Sivaloganathan, and Shenghai Zhang. Analysis of cell population PDE models with general maturation rates. The ANZIAM Journal, 43(3):359–374, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/analysis-of-cell-population-
pde-models-with-general-maturation-rates/1F1379E991C366632C594A05E0659C6A.

Snape-Jenkinson:2002:NSC

[1095] C. J. Snape-Jenkinson, S. Crozier, and L. K. Forbes. NMR shim coil design utilising a rapid spherical harmonic calculation method. The ANZIAM Journal, 43(3):375–386, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/nmr-
shim-coil-design-utilising-a-rapid-spherical-harmonic-calculation-
method/1155609A9DD82AE65A659E03B14CF4F5.

Hashim:2002:CBM

[1096] Ishak Hashim. On competition between modes at the onset of Bénard– Marangoni convection in a layer of fluid. The ANZIAM Journal, 43(3): 387–395, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-competition-between-modes-
at-the-onset-of-benardmarangoni-convection-in-a-layer-of-fluid/
3B03CA22DDA85F33FF696B43F50B70B4.

Kim:2002:GCR

[1097] Sangjeong Kim. Gevrey class regularity of the magnetohydrodynamics equations. The ANZIAM Journal, 43(3):397–408, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/gevrey-class-regularity-of-the-magnetohydrodynamics-
equations/C7839FDEF632B950D363C87B428F0780.

Saujani:2002:NES

[1098] S. Saujani, J. Drozd, and R. Mallier. Nonlinear evolution of singular disturbances to a tanh^{3} y mixing layer. The ANZIAM Journal, 43(3):409– 427, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/nonlinear-evolution-of-singular-disturbances-
to-a-tanh3y-mixing-layer/3C80D14410AC02ECEE6166224E6F2811.

Spalevic:2002:CCP

[1099] Miodrag M. Spalević. Calculation of Chakalov–Popoviciu quadratures of Radau and Lobatto type. The ANZIAM Journal, 43(3):429–447, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/calculation-of-chakalovpopoviciu-quadratures-of-radau-
and-lobatto-type/7E757B887962DF0F287A48B7B70FC9D7.

Guo:2002:FRO

[1100] Bao Zhu Guo. Further results for a one-dimensional linear thermoelastic equation with Dirichlet–Dirichlet boundary conditions. The ANZIAM Journal, 43(3):449–462, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/further-
results-for-a-onedimensional-linear-thermoelastic-equation-
with-dirichletdirichlet-boundary-conditions/7DD7C42F0E89138789F287D0346692BA.

Anonymous:2002:AVIa

[1101] Anonymous. ANZ volume 43 issue 3 cover and front matter. The ANZIAM Journal, 43(3):f1–f4, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-43-issue-3-cover-and-front-matter/9CB66F96BB90C4AA05CC54C057FFF1E9.

Anonymous:2002:AVIb

[1102] Anonymous. ANZ volume 43 issue 3 cover and back matter. The ANZIAM Journal, 43(3):b1–b2, January 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-43-issue-3-cover-and-back-matter/805FBEF6B5AFC99E5D4B13BD19F4F0FF.

Teo:2002:NSO

[1103] K. L. Teo, Y. Liu, W. R. Lee, L. S. Jennings, and S. Wang. Numerical solution of an optimal control problem with variable time points in the objective function. The ANZIAM Journal, 43 (4):463–478, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/numerical-solution-of-
an-optimal-control-problem-with-variable-time-points-in-the-
objective-function/862D2D195F5929EE746ECE98CE9208CD.

Acho:2002:PDS

[1104] Thomas M. Acho and Dominic P. Clemence. The parameter dependent Sturm–Liouville eigenproblem with an interior simple or double pole. The ANZIAM Journal, 43(4):479–491, April 2002. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/parameter-dependent-sturmliouville-eigenproblem-with-
an-interior-simple-or-double-pole/A60AD446677134863497324257CFF6A7.

Matic:2002:SRS

[1105] M. Matić, C. E. M. Pearce, and J. Pecarić. Some refinements of Shannon’s inequalities. The ANZIAM Journal, 43(4):493–511, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-refinements-of-shannons-inequalities/0716790A973267909605D9A7A1A14FF3.

Debsarma:2002:FON

[1106] Suma Debsarma and K. P. Das. Fourth order nonlinear evolution equations for gravity-capillary waves in the presence of a thin thermocline in deep water. The ANZIAM Journal, 43(4):513–524, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/fourth-order-nonlinear-evolution-equations-
for-gravitycapillary-waves-in-the-presence-of-a-thin-thermocline-
in-deep-water/E7785989CE242143E817F73D7790C74B.

Bainov:2002:PSS

[1107] D. D. Bainov, A. B. Dishliev, and I. M. Stamova. Practical stability of the solutions of impulsive systems of differentialdifference equations via the method of comparison and some applications to population dynamics. The ANZIAM Journal, 43(4): 525–539, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/practical-stability-of-the-
solutions-of-impulsive-systems-of-differentialdifference-equations-
via-the-method-of-comparison-and-some-applications-to-population-
dynamics/BF9334EC34CD2DEA81276D4D5282AB10.

Guo:2002:DSC

[1108] Xianping Guo and Weiping Zhu. Denumerable state continuous time Markov decision processes with unbounded cost and transition rates under average criterion. The ANZIAM Journal, 43 (4):541–557, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/denumerable-state-
continuous-time-markov-decision-processes-with-unbounded-cost-
and-transition-rates-under-average-criterion/A89CA26CF33825B0E3E500654EF5BAAB.

Stevic:2002:GTH

[1109] Stevo Stević. Growth theorems for homogeneous second-order difference equations. The ANZIAM Journal, 43(4):559–566, April 2002.

CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/growth-theorems-for-homogeneous-secondorder-difference-
equations/3CC011CABC75AF9D7179CF36FB40132D.

Srivastava:2002:CTB

[1110] H. M. Srivastava and Yeong-Nan Yeh. Certain theorems on bilinear and bilateral generating functions. The ANZIAM Journal, 43(4):567–574, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/certain-theorems-on-bilinear-and-bilateral-
generating-functions/46C12E284FA4B82B892A83A86FE0E716.

Abalo:2002:ETG

[1111] Kokou Y. Abalo and Michael M. Kostreva. Existence theory for games of pricing and technology. The ANZIAM Journal, 43(4): 575–585, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/existence-theory-for-games-
of-pricing-and-technology/137E8B6DE12B93E9F0570F9AE5CB1B76.

Anonymous:2002:CV

[1112] Anonymous. Contents of volume 43. The ANZIAM Journal, 43 (4):587–589, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/contents-of-volume-43/
E7AA7F953F01B849F8815030516CAE10.

Anonymous:2002:AVIc

[1113] Anonymous. ANZ volume 43 issue 4 cover and front matter. The ANZIAM Journal, 43(4):f1–f4, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-43-issue-4-cover-and-front-matter/6E04FDF2177EBA8990F5ACBCC7A822AD.

Anonymous:2002:AVId

[1114] Anonymous. ANZ volume 43 issue 4 cover and back matter. The ANZIAM Journal, 43(4):b1–b3, April 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-43-issue-4-cover-and-back-matter/709FDBC304ACC88A56E1DAAF7FF3A795.

Pearce:2002:PSI

[1115] Charles Pearce. Preface to this special issue. The ANZIAM Journal, 44 (1):i–ii, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/preface-to-this-special-issue/FDB4A563D3605CE658A8E0D8614BF30D.

Hietarinta:2002:TMS

[1116] Jarmo Hietarinta. Taming the movable singularities. The ANZIAM Journal, 44(1):1–9, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/taming-the-movable-singularities/
071BC8E1D4DF19021251B536C9013266.

Broadbridge:2002:HFE

[1117] P. Broadbridge, B. H. Bradshaw, G. R. Fulford, and G. K. Aldis. Huxley and Fisher equations for gene propagation: an exact solution. The ANZIAM Journal, 44(1):11–20, July 2002. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/huxley-and-fisher-equations-for-gene-propagation-an-
exact-solution/B2F157FDA758A2045F47AA146F86811C.

Ramani:2002:WDA

[1118] A. Ramani and B. Grammaticos. What is the discrete analogue of the Painlevé property? The ANZIAM Journal, 44(1):21–32, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/what-is-the-discrete-analogue-of-the-painleve-property/
5FE6108404908B97F3135D77FCABD97C.

Dewar:2002:ACT

[1119] R. L. Dewar. Asymptotology — a cautionary tale. The ANZIAM Journal, 44(1):33–40, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/asymptotologya-cautionary-tale/
75BD5B43E22D0D6917DE8069449D1BDF.

Witte:2002:IRM

[1120] N. S. Witte, P. J. Forrester, and Christopher M. Cosgrove. Integrability, random matrices and Painlevé transcendents. The ANZIAM Journal, 44 (1):41–50, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/

anziam-journal/article/integrability-random-matrices-and-painleve-
transcendents/5665EC1D517D44CEF3D2320E5CA0FA6F.

Booth:2002:NES

[1121] Hilary Booth. Nonlinear electron solutions and their characteristics at infinity. The ANZIAM Journal, 44(1):51–59, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinear-electron-solutions-and-their-characteristics-
at-infinity/06A9EB52156103569D02E30C3F12603F.

Nucci:2002:SAF

[1122] M. C. Nucci and P. G. L. Leach. Symmetry analysis of and first integrals for the continuum Heisenberg spin chain. The ANZIAM Journal, 44(1):61–72, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/symmetry-analysis-
of-and-first-integrals-for-the-continuum-heisenberg-spin-chain/
BA3D6DDD8145ABA0AEFD3BCD313AE5F4.

Pashaev:2002:BHS

[1123] Oktay K. Pashaev and Jyh-Hao Lee. Black holes and solitons of the quantized dispersionless NLS and DNLS equations. The ANZIAM Journal, 44 (1):73–81, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/black-holes-and-solitons-of-the-quantized-
dispersionless-nls-and-dnls-equations/E285D3B7563791ECC596FE82063DEFE7.

Vassiliou:2002:ISP

[1124] Peter J. Vassiliou. An integrable system of partial differential equations on the special linear group. The ANZIAM Journal, 44(1):83–93, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/an-integrable-system-of-partial-differential-equations-
on-the-special-linear-group/995B706A677A472ED21875285651FC62.

Marchant:2002:NSW

[1125] T. R. Marchant. Numerical solitary wave interaction: the order of the inelastic effect. The ANZIAM Journal, 44(1):95–102, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/numerical-solitary-wave-interaction-
the-order-of-the-inelastic-effect/F22A81E3478303288BD691F498FF416C.

Veerakumar:2002:PES

[1126] V. Veerakumar and M. Daniel. Propagation of an electromagnetic soliton in a ferromagnetic medium. The ANZIAM Journal, 44(1):103–110, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/propagation-of-an-
electromagnetic-soliton-in-a-ferromagnetic-medium/45951AD72F8D59332E3BF2D139195C69.

Takei:2002:EWA

[1127] Yoshitsugu Takei. On an exact WKB approach to Ablowitz– Segur’s connection problem for the second Painlevé equation. The ANZIAM Journal, 44(1):111–119, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
an-exact-wkb-approach-to-ablowitzsegurs-connection-problem-
for-the-second-painleve-equation/D724D3086E2DFDCB17E0B5713B893CE3.

Hu:2002:BTN

[1128] Xing-Biao Hu and Johan Springael. A Bäcklund transformation and nonlinear superposition formula for the Lotka–Volterra hierarchy. The ANZIAM Journal, 44(1):121–128, July 2002. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/backlund-transformation-and-nonlinear-superposition-
formula-for-the-lotkavolterra-hierarchy/F6987C08D08247248471E9B73E09618D.

Ma:2002:BCF

[1129] Wen-Xiu Ma and Yunbo Zeng. Binary constrained flows and separation of variables for soliton equations. The ANZIAM Journal, 44(1):129–139, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/binary-constrained-flows-and-separation-of-
variables-for-soliton-equations/7271853557BA04AF3BFC76BFC1B3C1C0.

Schief:2002:GPV

[1130] W. K. Schief. On the geometry of the Painlevé V equation and a Bäcklund transformation. The ANZIAM Journal, 44(1):141–148, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-the-geometry-of-the-painleve-v-equation-
and-a-backlund-transformation/AEFD2EFE04DBADABAAA8E9669FF4E642.

Kudryashov:2002:DEC

[1131] N. A. Kudryashov and M. B. Soukharev. Discrete equations corresponding to fourth-order differential equations of the P_{2} and K_{2} hierarchies. The ANZIAM Journal, 44(1):149–160, July 2002. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/discrete-equations-corresponding-to-fourthorder-differential-
equations-of-the-p2-and-k2-hierarchies/FEC8D76D79C1D29EE979F076E0E966D7.

Harmer:2002:ISM

[1132] M. S. Harmer. Inverse scattering for the matrix Schrödinger operator and Schrödinger operator on graphs with general selfadjoint boundary conditions. The ANZIAM Journal, 44(1):161– 168, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/inverse-scattering-for-the-
matrix-schrodinger-operator-and-schrodinger-operator-on-graphs-
with-general-selfadjoint-boundary-conditions/FCEFB32AFB935D0EAFAE5CA0F15D5BB6.

Guha:2002:DPS

[1133] Partha Guha. Diffeomorphisms on S^{1}, projective structures and integrable systems. The ANZIAM Journal, 44(1):169–180, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/diffeomorphisms-on-s1-projective-structures-and-integrable-
systems/3331891BD658CB6E194AB1540538AE8C.

Anonymous:2002:AVIe

[1134] Anonymous. ANZ volume 44 issue 1 cover and front matter. The ANZIAM Journal, 44(1):f1–f4, July 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-44-issue-1-cover-and-front-matter/8E1E6648A7F0D1863274B1973E186B14.

Hocking:2002:NWF

[1135] G. C. Hocking, J.-M. Vanden-Broeck, and L. K. Forbes. A note on withdrawal from a fluid of finite depth through a point sink. The ANZIAM Journal, 44(2):181–191, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
withdrawal-from-a-fluid-of-finite-depth-through-a-point-sink/
D3EFFC9CAF8C921617ABCCD8112C5DA1.

Franzblau:2002:OHC

[1136] D. S. Franzblau and A. Raychaudhuri. Optimal Hamiltonian completions and path covers for trees, and a reduction to maximum flow. The ANZIAM Journal, 44(2):193–204, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
hamiltonian-completions-and-path-covers-for-trees-and-a-reduction-
to-maximum-flow/5F80A57D0F4918546ECDFFC4E32E874B.

Shi:2002:EDR

[1137] Dong-Hua Shi and De-Xing Feng. Exponential decay rate of the energy of a Timoshenko beam with locally distributed feedback. The ANZIAM Journal, 44(2):205–220, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/exponential-decay-
rate-of-the-energy-of-a-timoshenko-beam-with-locally-distributed-
feedback/EA141A1D73B199E6CDF9BE22F0BF59F3.

Matic:2002:TPF

[1138] M. Matić, C. E. M. Pearce, and J. Pecarić. Two-point formulae of Euler type. The ANZIAM Journal, 44(2):221–245, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/twopoint-formulae-of-euler-type/F4D234B44927B8325C77E509C10AC62B.

Aizengendler:2002:NDM

[1139] Peter Aizengendler. Newton’s diagram method for nonlinear equations with several small parameters. The ANZIAM Journal, 44(2): 247–259, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/newtons-diagram-method-for-
nonlinear-equations-with-several-small-parameters/472E6FFB582C86CF92B769EB7A3EB00A.

Mohamad:2002:ESA

[1140] S. Mohamad and K. Gopalsamy. Extreme stability and almost periodicity in continuous and discrete neuronal models with finite delays. The ANZIAM Journal, 44(2):261–282, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/extreme-
stability-and-almost-periodicity-in-continuous-and-discrete-
neuronal-models-with-finite-delays/1B732727ADF6AB7682644574C63F7112.

Choi:2002:MGQ

[1141] Bong Dae Choi and Bara Kim. An M/G/l queueing system with fixed feedback policy. The ANZIAM Journal, 44(2):283–297, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/an-mgl-queueing-system-with-fixed-feedback-
policy/A3BD20EE49468604381E282C87058B6D.

Gupta:2002:SRG

[1142] K. C. Gupta, S. P. Goyal, and Rohit Mukherjee. Some results on generalised Voigt functions. The ANZIAM Journal, 44(2):299–303, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-results-on-generalised-voigt-functions/3717FE116F36EF412B4EB3B18067A810.

Liu:2002:DRB

[1143] Huan-Wen Liu and Song-Ping Zhu. The dual reciprocity boundary element method for magnetohydrodynamic channel flows. The ANZIAM Journal, 44(2):305–322, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/dual-reciprocity-boundary-
element-method-for-magnetohydrodynamic-channel-flows/5C625F627467B98D1574306ACEEE305E.

Anonymous:2002:AVIf

[1144] Anonymous. ANZ volume 44 issue 2 cover and front matter. The ANZIAM Journal, 44(2):f1–f4, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-44-issue-2-cover-and-front-matter/541F9D36D988B8242A9EC3254FC35697.

Anonymous:2002:AVIg

[1145] Anonymous. ANZ volume 44 issue 2 cover and back matter. The ANZIAM Journal, 44(2):b1–b3, October 2002. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-44-issue-2-cover-and-back-matter/CF36C464A5FA8E764926079E48BAEF8C.

Badora:2003:ASS

[1146] Roman Badora, Roman Ger, and Zsolt Páles. Additive selections and the stability of the Cauchy functional equation. The ANZIAM Journal, 44 (3):323–337, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/additive-selections-and-the-
stability-of-the-cauchy-functional-equation/A7874CB98314F15EBD9249CD521842F3.

Lai:2003:MFP

[1147] H. C. Lai and J. C. Liu. Minimax fractional programming involving generalised invex functions. The ANZIAM Journal, 44(3):339–354, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/minimax-fractional-programming-involving-generalised-
invex-functions/70A949E14415AFFAB835EFD508607D56.

Barnett:2003:QTI

[1148] N. S. Barnett, S. S. Dragomir, and C. E. M. Pearce. A quasi-trapezoid inequality for double integrals. The ANZIAM Journal, 44(3):355–364, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/quasitrapezoid-inequality-for-double-integrals/
FC0AEC7ED08DFA725A7EE1636C05F53C.

Weeks:2003:DFI

[1149] S. W. Weeks, G. C. Sander, and J.-Y. Parlange. n-dimensional first integral and similarity solutions for two-phase flow. The ANZIAM Journal, 44(3):365–380, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/ndimensional-first-integral-
and-similarity-solutions-for-twophase-flow/D7AEEF3B774DA95A058BDE8479AA3D28.

Gao:2003:GEG

[1150] Hongjun Gao and Keng-Huat Kwek. Global existence for the generalised 2D Ginzburg–Landau equation. The ANZIAM Journal, 44(3):381– 392, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/global-existence-for-the-generalised-
2d-ginzburglandau-equation/8411F682F41C606024039032E05B08CE.

Franco:2003:NCE

[1151] Daniel Franco and Rodrigo L. Pouso. Nonresonance conditions and extremal solutions for first-order impulsive problems under weak assumptions. The ANZIAM Journal, 44(3):393–407, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonresonance-conditions-and-extremal-solutions-for-
firstorder-impulsive-problems-under-weak-assumptions/7BDD309CB6DBECDB6AD6D22E30D82295.

Le:2003:SNV

[1152] Vy Khoi Le. On some noncoercive variational inequalities containing degenerate elliptic operators. The ANZIAM Journal, 44(3):409–430, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-some-noncoercive-variational-inequalities-
containing-degenerate-elliptic-operators/CFA4A2EBBEF8833E1E297CEC5CDF49AA.

Argyropoulos:2003:CNI

[1153] E. Argyropoulos, D. Gintides, and K. Kiriaki. On the condition number of integral equations in linear elasticity using the modified Green’s function. The ANZIAM Journal, 44(3):431– 446, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-condition-number-of-
integral-equations-in-linear-elasticity-using-the-modified-
greens-function/841F03C308CFA6783E53EE9114A98846.

Kumar:2003:RRS

[1154] Rajneesh Kumar, Sushil K. Tomar, and Asha Chopra. Reflection/refraction of SH-waves at a corrugated interface between two different anisotropic and vertically heterogeneous elastic solid halfspaces. The ANZIAM Journal, 44(3):447–460, January 2003. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reflectionrefraction-of-shwaves-at-a-corrugated-interface-
between-two-different-anisotropic-and-vertically-heterogeneous-
elastic-solid-halfspaces/32AB27DA993C1830BBDF164E6D9AF6D8.

Chaudhuri:2003:EAR

[1155] P. K. Chaudhuri and Subhankar Ray. Effects of an axisymmetric rigid punch on a nonhomogeneous transversely isotropic half-space. The ANZIAM Journal, 44(3):461–474, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/effects-
of-an-axisymmetric-rigid-punch-on-a-nonhomogeneous-transversely-
isotropic-halfspace/4C642004C9BB15E2779ED0D85872AF5A.

Srivastava:2003:CCG

[1156] H. M. Srivastava, M. A. Pathan, and M. G. Bin-Saad. A certain class of generating functions involving bilateral series. The ANZIAM Journal, 44(3):475–483, January 2003. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/certain-
class-of-generating-functions-involving-bilateral-series/160E558DD60FBAE9C9072D3AA7AD2EEC.

Anonymous:2003:AVIa

[1157] Anonymous. ANZ volume 44 issue 3 cover and front matter. The ANZIAM Journal, 44(3):f1–f4, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-44-issue-3-cover-and-front-matter/5577629969FD238CC485E964D7717B0D.

Anonymous:2003:AVIb

[1158] Anonymous. ANZ volume 44 issue 3 cover and back matter. The ANZIAM Journal, 44(3):b1–b3, January 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-44-issue-3-cover-and-back-matter/FF5C41E54BA17F59BE2AC18511FB9B4C.

Howlett:2003:OLF

[1159] P. G. Howlett, C. E. M. Pearce, and A. P. Torokhti. An optimal linear filter for random signals with realisations in a separable Hilbert space. The ANZIAM Journal, 44(4):485–500, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
optimal-linear-filter-for-random-signals-with-realisations-
in-a-separable-hilbert-space/2B1FAFA7F0F27DDAFE1F759C366EC306.

Cullen:2003:SCC

[1160] R. M. Cullen, A. Korobeinikov, and W. J. Walker. Seasonality and critical community size for infectious diseases. The ANZIAM Journal, 44(4): 501–512, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/seasonality-and-critical-community-size-
for-infectious-diseases/F913A555E84A24F38AE7FBC939D4EE74.

Agarwal:2003:SFO

[1161] R. P. Agarwal and S. Heikkilä. On the solvability of first-order discontinuous scalar initial and boundary value problems. The ANZIAM Journal, 44(4):513–538, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/on-the-solvability-of-
firstorder-discontinuous-scalar-initial-and-boundary-value-
problems/0056BDAA339343B95E55665467EDDE54.

Chen:2003:MCE

[1162] Qihong Chen. Minimax control of an elliptic variational bilateral problem. The ANZIAM Journal, 44(4):539–559, April 2003. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/minimax-control-of-an-elliptic-variational-bilateral-
problem/B981E3FBEF20AF43196922E0BC58EB1F.

Stokes:2003:DRD

[1163] Y. M. Stokes. Determining rotational deformity in broken forearms. The ANZIAM Journal, 44(4):561–568, April 2003. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/determining-rotational-deformity-in-broken-forearms/
9788529356D40CF01B8E1DED0F5EF4DC.

Wen:2003:GAP

[1164] Xianzhang Wen and Zhicheng Wang. Global attractivity of the periodic Kolmogorov system. The ANZIAM Journal, 44(4):569–581, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/global-attractivity-of-the-periodic-
kolmogorov-system/23474DE75F5CA45364814ABF2B45B15A.

Wu:2003:VCF

[1165] Hanzhong Wu. A variation-of-constants formula for a linear abstract evolution equation in Hilbert space. The ANZIAM Journal, 44(4):583–590, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/variationofconstants-formula-for-a-linear-abstract-
evolution-equation-in-hilbert-space/605AD93B194FA1F700F2DB8966284DDE.

Habets:2003:ENS

[1166] Patrick Habets and Rodrigo L. Pouso. Examples of the nonexistence of a solution in the presence of upper and lower solutions. The ANZIAM Journal, 44(4):591–594, April 2003. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/examples-of-the-nonexistence-of-a-solution-in-the-presence-
of-upper-and-lower-solutions/5FAEF37F530F02E4BF32963B566660A6.

Yang:2003:NTR

[1167] Y. F. Yang. A new trust region method for nonsmooth equations. The ANZIAM Journal, 44(4):595–607, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/new-
trust-region-method-for-nonsmooth-equations/89C35B7220BC753F81686E8861CB49E5.

Alzer:2003:IBF

[1168] Horst Alzer. Inequalities for the beta function of n variables. The ANZIAM Journal, 44(4):609–623, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/inequalities-
for-the-beta-function-of-n-variables/8D5958EDD198A51B905ABD3FA6F16FF9.

Guan:2003:OBR

[1169] Xinping Guan, Yichang Liu, Cailian Chen, and Peng Shi. Observer-based robust H_{∞}, control for uncertain time-delay systems. The ANZIAM Journal, 44(4):625–634, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/observerbased-robust-h-
control-for-uncertain-timedelay-systems/59170FE96A8980C66E9106A59EC07D04.

Anonymous:2003:CV

[1170] Anonymous. Contents of volume 44. The ANZIAM Journal, 44 (4):635–638, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/contents-of-volume-44/
1E0E9BCB4E63A7A1F1B31F556A6A0775.

Anonymous:2003:AVIc

[1171] Anonymous. ANZ volume 44 issue 4 cover and front matter. The ANZIAM Journal, 44(4):f1–f4, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-44-issue-4-cover-and-front-matter/ABFB82DE4E0CB9E10BC47C57E768281C.

Anonymous:2003:AVId

[1172] Anonymous. ANZ volume 44 issue 4 cover and back matter. The ANZIAM Journal, 44(4):b1–b3, April 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-44-issue-4-cover-and-back-matter/56A3E3E8C3FD53B8E21165140EEFA70E.

Zhu:2003:FST

[1173] Songping Zhu and Yinglong Zhang. A flat ship theory on bow and stern flows. The ANZIAM Journal, 45(1):1–15, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/flat-ship-theory-on-bow-and-stern-flows/B41F5F9B32A819E416C6E51E2C0F8681.

Wong:2003:TLO

[1174] M. W. Wong and Zhaohui Zhang. Traces of localisation operators with two admissible wavelets. The ANZIAM Journal, 45(1):17–25, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/traces-of-localisation-operators-with-two-admissible-
wavelets/0F24A1DA150D40FB45E4175C48F32FE6.

Pleasants:2003:DPD

[1175] A. B. Pleasants, G. C. Wake, and C. C. Daly. Derivation of the probability density function for ultimate muscle pH in slaughtered animals. The ANZIAM Journal, 45(1):27–34, July 2003. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/derivation-of-the-probability-density-function-for-
ultimate-muscle-ph-in-slaughtered-animals/9A5B43327F28D66ABDCCD507F0FCBD4A.

Al-Refai:2003:SEE

[1176] M. Al-Refai and K. K. Tam. Sequential eigenfunction expansion for a problem in combustion theory. The ANZIAM Journal, 45(1):35–48, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/sequential-eigenfunction-expansion-for-a-problem-in-
combustion-theory/99B41769E333F7FD3266908BB1EB351E.

Yang:2003:CSP

[1177] H. Yang, G. Yin, K. Yin, and Q. Zhang. Control of singularly perturbed Markov chains: a numerical study. The ANZIAM Journal, 45(1):49–74, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/control-of-singularly-
perturbed-markov-chains-a-numerical-study/F14FE9011A5E534ABF9C342FE07CADFE.

Jimenez:2003:NLP

[1178] B. Jiménez and V. Novo. A notion of local proper efficiency in the Borwein sense in vector optimisation. The ANZIAM Journal, 45(1):75–89, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/notion-of-local-proper-efficiency-in-the-borwein-
sense-in-vector-optimisation/54FFD574CEC237456161E9A9A39D8D3B.

Li:2003:NAC

[1179] Z. F. Li, M. R. Osborne, and T. Prvan. Numerical algorithms for constrained maximum likelihood estimation. The ANZIAM Journal, 45(1): 91–114, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/numerical-algorithms-for-constrained-
maximum-likelihood-estimation/1958473DF1B00779E060E7E7F462E480.

Lou:2003:EOC

[1180] Hongwei Lou. Existence of optimal controls for semilinear elliptic equations without Cesari-type conditions. The ANZIAM Journal, 45(1): 115–131, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/existence-of-optimal-controls-for-semilinear-
elliptic-equations-without-cesaritype-conditions/95114EB8B6CF91743657F5D64039820B.

Mohr:2003:NCE

[1181] G. A. Mohr and A. S. Power. Natural cubic element formulation and infinite domain modelling for potential flow problems. The ANZIAM Journal, 45(1):133–143, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/natural-cubic-element-
formulation-and-infinite-domain-modelling-for-potential-flow-
problems/D56CBF660C88DB98E12485C301AB5ED1.

El-Sayed:2003:TSI

[1182] Salah M. El-Sayed. A two-sided iterative method for computing positive definite solutions of a nonlinear matrix equation. The ANZIAM Journal, 45(1):145–152, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/twosided-iterative-method-for-
computing-positive-definite-solutions-of-a-nonlinear-matrix-
equation/8BC0221C9B935A461E67263CF58CC37C.

Anonymous:2003:AVIe ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-45-issue-1-cover-and-front-matter/F52C73E56A59B4BEE01B0F0FFFE36FF0.

Anonymous:2003:AVIf

[1184] Anonymous. ANZ volume 45 issue 1 cover and back matter. The ANZIAM Journal, 45(1):b1–b3, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-45-issue-1-cover-and-back-matter/EC347F0B72866C9E8450FFCBDBBA7A30.

Norhayati:2003:SSN

[1185] Norhayati and G. C. Wake. The solution and the stability of a nonlinear age-structured population model. The ANZIAM Journal, 45(2): 153–165, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/solution-and-the-stability-
of-a-nonlinear-agestructured-population-model/4E334AF6855D8A180B05623194E9B66A.

Agarwal:2003:SPM

[1186] Ravi P. Agarwal and Donal O’Regan. Singular problems modelling phenomena in the theory of pseudoplastic fluids. The ANZIAM Journal, 45(2):167–179, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/singular-problems-
modelling-phenomena-in-the-theory-of-pseudoplastic-fluids/1701C8C6903B86BDE9B6FBA2B2DA2B0F.

Moghadas:2003:EMT

[1187] S. M. Moghadas and A. B. Gumel. An epidemic model for the transmission dynamics of HIV and another pathogen. The ANZIAM Journal, 45(2):181–193, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/an-epidemic-model-for-
the-transmission-dynamics-of-hiv-and-another-pathogen/CF432E181E280E2DD9C9599654710537.

Ahn:2003:DRB

[1188] Young Joon Ahn. Degree reduction of Bézier curves using constrained Chebyshev polynomials of the second kind. The ANZIAM Journal, 45 (2):195–205, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/degree-reduction-of-beezier-
curves-using-constrained-chebyshev-polynomials-of-the-second-
kind/979DDDB5A3913E15B0ADC66048F1218F.

Caraballo:2003:DAN

[1189] T. Caraballo, J. A. Langa, and J. Valero. The dimension of attractors of nonautonomous partial differential equations. The ANZIAM Journal, 45 (2):207–222, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/dimension-of-attractors-of-
nonautonomous-partial-differential-equations/E0535BA0FA6735DCD576BD0F27A194B1.

Jankowski:2003:MMS

[1190] Tadeusz Jankowski. Minimal and maximal solutions to systems of differential equations with a singular matrix. The ANZIAM Journal, 45(2): 223–231, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/minimal-and-maximal-solutions-
to-systems-of-differential-equations-with-a-singular-matrix/
0EB0710609696F7CE381A6A88E6F2A30.

Zhang:2003:PST

[1191] Zhengqiu Zhang and Zhicheng Wang. Periodic solutions of a two-species ratio-dependent predator–prey system with time delay in a two-patch environment. The ANZIAM Journal, 45(2): 233–244, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/periodic-solutions-of-
a-twospecies-ratiodependent-predatorprey-system-with-time-delay-
in-a-twopatch-environment/F155978DB68D7C5C444DD089ED2F7BF4.

Muthu:2003:IWP

[1192] P. Muthu, B. V. Rathish Kumar, and Peeyush Chandra. On the influence of wall properties in the peristaltic motion of micropolar fluid. The ANZIAM Journal, 45(2):245–260, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-influence-of-wall-properties-in-the-peristaltic-motion-
of-micropolar-fluid/FF860502D444EFF1C419D74048B5A442.

Maloney:2003:AFK

[1193] John Maloney and Jack Heidel. An analysis of a fractal kinetics curve of Savageau. The ANZIAM Journal, 45(2):261–269, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/an-analysis-of-a-fractal-kinetics-curve-of-
savageau/4271546103E4DD3FB32993CECF8FDAAB.

Liu:2003:GAC

[1194] Yuji Liu and Binggen Zhang. Global attractivity of a class of delay differential equations with impulses. The ANZIAM Journal, 45(2):271– 284, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/global-attractivity-of-a-class-of-delay-
differential-equations-with-impulses/6C180D299DBC9D5F9F41F8C098EC5386.

Pearce:2003:SSI

[1195] C. E. M. Pearce and J. Pecarić. On some switching inequalities of Brenner and Alzer. The ANZIAM Journal, 45(2):285–288, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
some-switching-inequalities-of-brenner-and-alzer/4B0EACEDA110B8C6BBBFC006CEF82926.

Kamarujjama:2003:UPG

[1196] M. Kamarujjama. A unified presentation of generalised Voigt functions. The ANZIAM Journal, 45(2):289–293, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/unified-
presentation-of-generalised-voigt-functions/8C348A7E72A9DCCAB885BBDF0E708407.

Deeba:2003:DMS

[1197] Elias Deeba, Ghassan Dibeh, Suheil Khuri, and Shishen Xie. Decomposition method for solving a nonlinear business cycle model. The ANZIAM Journal, 45(2):295–302, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/decomposition-method-
for-solving-a-nonlinear-business-cycle-model/A7CAE6CA03A073E756AE6EEFBAF6C4FE.

Anonymous:2003:AVIg

[1198] Anonymous. ANZ volume 45 issue 2 cover and front matter. The ANZIAM Journal, 45(2):f1–f4, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-45-issue-2-cover-and-front-matter/06A3B39022DE19E5186A3D573EA81E88.

Anonymous:2003:AVIh

[1199] Anonymous. ANZ volume 45 issue 2 cover and back matter. The ANZIAM Journal, 45(2):b1–b2, October 2003. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-45-issue-2-cover-and-back-matter/48C94967905BB5B679A1D08D78F0B59A.

Nelson:2004:BPO

[1200] M. I. Nelson and H. S. Sidhu. Bifurcation phenomena for an oxidation reaction in a continuously stirred tank reactor. II. Diabatic operation. The ANZIAM Journal, 45(3):303–326, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/bifurcation-phenomena-for-an-oxidation-reaction-in-
a-continuously-stirred-tank-reactor-ii-diabatic-operation/61A342AE65F791B6750C3047AFDDF3D3.

Torres:2004:SRN

[1201] Pedro J. Torres. Some remarks on a Neumann boundary value problem arising in fluid dynamics. The ANZIAM Journal, 45 (3):327–332, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/some-remarks-on-a-
neumann-boundary-value-problem-arising-in-fluid-dynamics/0FE43B9618D22558D1DB9974B56AE13A.

VanBrunt:2004:CLS

[1202] B. Van Brunt, D. Pidgeon, M. Vlieg-Hulstman, and W. D. Halford. Conservation laws for second-order parabolic partial differential equations. The ANZIAM Journal, 45(3):333–348, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/conservation-laws-for-secondorder-parabolic-partial-
differential-equations/639A09DEC4FF704158319C88507E85BC.

Li:2004:MAH

[1203] Hongxu Li and Falun Huang. Minimal approximate Hessians for continuously Gâteaux differentiable functions. The ANZIAM Journal, 45(3): 349–359, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/minimal-approximate-hessians-
for-continuously-gateaux-differentiable-functions/5D3FFA07E44F535AA9624BF1479A17DC.

Harris:2004:IEH

[1204] P. J. Harris, H. Al-Awadi, and W. K. Soh. An investigation into the effects of heat transfer on the motion of a spherical bubble. The ANZIAM Journal, 45(3):361–371, January 2004. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/an-investigation-into-the-effects-of-heat-transfer-
on-the-motion-of-a-spherical-bubble/612ADAFD8DF2972FE433360A851DD06C.

Cox:2004:LIT

[1205] G. M. Cox and J. M. Hill. The limiting ideal theory for shearindex cohesionless granular materials. The ANZIAM Journal, 45 (3):373–392, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/limiting-ideal-theory-
for-shearindex-cohesionless-granular-materials/E26E3718A9188E1989617D6364FDCA0C.

Belen:2004:RGI

[1206] Selma Belen and C. E. M. Pearce. Rumours with general initial conditions. The ANZIAM Journal, 45(3):393–400, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/rumours-with-general-initial-conditions/851C5D225935BAF641DD28BCCC7C1D80.

Lucas:2004:MOO

[1207] S. K. Lucas. Maximising output from oil reservoirs without water breakthrough. The ANZIAM Journal, 45(3):401–422, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/maximising-output-from-oil-reservoirs-without-water-
breakthrough/2013A24B9E9D15C78B76AE40BE7CDF70.

Penesis:2004:PFG

[1208] I. Penesis, J. J. Shepherd, and H. J. Connell. The pressure field in the gas-lubricated step slider bearing. The ANZIAM Journal, 45(3):423– 442, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/pressure-field-in-the-gaslubricated-
step-slider-bearing/32D8B0050D00D3224CCAEE7131D2E4E6.

Kar:2004:HTP

[1209] T. K. Kar and K. S. Chaudhuri. Harvesting in a two-prey one-predator fishery: a bioeconomic model. The ANZIAM Journal, 45(3):443–456, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/harvesting-in-a-twoprey-onepredator-
fishery-a-bioeconomic-model/769626433C249917D91489997F9F74AF.

Anonymous:2004:AVIa

[1210] Anonymous. ANZ volume 45 issue 3 cover and front matter. The ANZIAM Journal, 45(3):f1–f4, January 2004. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-45-issue-3-cover-and-front-matter/BA7D3A5102ED73A6B328729F5F0BAABF.

Anonymous:2004:AVIb

[1211] Anonymous. ANZ volume 45 issue 3 cover and back matter. The ANZIAM Journal, 45(3):b1–b3, January 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-45-issue-3-cover-and-back-matter/1356D8FCEF5B8CA8DAECF1E97130464A.

Hunt:2004:PAF

[1212] Emma Hunt. A probabilistic algorithm for finding the rate matrix of a block-GI/M/1 Markov chain. The ANZIAM Journal, 45(4):457– 475, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/probabilistic-algorithm-for-finding-
the-rate-matrix-of-a-blockgim1-markov-chain/B62B1D12CA0A1D1E388932354F717C8D.

Li:2004:SMC

[1213] S. J. Li, X. Q. Yang, K. L. Teo, and S. Y. Wu. A solution method for combined semi-infinite and semi-definite programming. The ANZIAM Journal, 45(4):477–494, April 2004. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/solution-method-for-combined-semiinfinite-and-semidefinite-
programming/76BBA9FCE5A87AC7D5CBE10A0BE0C0F5.

Bulte:2004:PRE

[1214] D. P. Bulte, L. K. Forbes, and S. Crozier. Phase-retardation effects at radio frequencies in flat-plate conductors. The ANZIAM Journal, 45(4): 495–510, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/phaseretardation-effects-at-radio-frequencies-
in-flatplate-conductors/332A69869413318E994333395AE1B6A8.

Byatt:2004:PVB

[1215] D. Byatt, I. D. Coope, and C. J. Price. Performance of various BFGS implementations with limited precision second-order information. The ANZIAM Journal, 45(4):511–522, April 2004. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/performance-of-various-bfgs-implementations-with-limited-
precision-secondorder-information/E485B6EC98F980D603F61DC63F320E89.

Sach:2004:EGC

[1216] Pham Huu Sach, Gue Myung Lee, and Do Sang Kim. Efficiency and generalised convexity in vector optimisation problems. The ANZIAM Journal, 45(4):523–546, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/efficiency-and-generalised-
convexity-in-vector-optimisation-problems/175E5A19DFB67BF9BFFBC831290608D8.

Gaffney:2004:MMS

[1217] J. M. Gaffney and C. E. M. Pearce. Memory, market stability and the nonlinear cobweb theorem. The ANZIAM Journal, 45(4):547–555, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/memory-market-stability-and-the-nonlinear-cobweb-theorem/
DCFBAD01915416D2B3F227CF9A761904.

Sun:2004:PSS

[1218] Yan Sun, Lishan Liu, and Yeol Je Cho. Positive solutions of singular nonlinear Sturm–Liouville boundary value problems. The ANZIAM Journal, 45(4):557–571, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/positive-solutions-of-
singular-nonlinear-sturmliouville-boundary-value-problems/B7D42B5BAAAB56E805E146ED272A970C.

Mohr:2004:FRD

[1219] G. A. Mohr. Flow ratio design of primal and dual network models of distribution. The ANZIAM Journal, 45(4):573–583, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/flow-ratio-design-of-primal-and-dual-network-models-
of-distribution/46BAF6B680B6967A0F2D5437F1D4463A.

Guo:2004:SSE

[1220] Jong-Shenq Guo, Chu-Pin Lo, and Je-Chiang Tsai. The structure of solutions for equations related to the motions of plane curves. The ANZIAM Journal, 45(4):585–592, April 2004. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/structure-of-solutions-for-equations-related-to-the-
motions-of-plane-curves/49F813A914144AECBBB8BE66CD90A2CE.

Zhao:2004:OFO

[1221] Aimin Zhao, Xianhua Tang, and Jurang Yan. Oscillation of first-order delay differential equations. The ANZIAM Journal, 45(4):593–599, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/oscillation-of-firstorder-delay-differential-equations/
17C02CD492BF0ED80C51317ECAEC686B.

Anonymous:2004:CV

[1222] Anonymous. Contents of volume 45. The ANZIAM Journal, 45 (4):601–604, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/contents-of-volume-45/
371E120FCCC275169DD6D665C5D46E38.

Anonymous:2004:AVIc

[1223] Anonymous. ANZ volume 45 issue 4 cover and front matter. The ANZIAM Journal, 45(4):f1–f4, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-45-issue-4-cover-and-front-matter/61518D1A76B089BD952C609BBB0B3CBB.

Anonymous:2004:AVId

[1224] Anonymous. ANZ volume 45 issue 4 cover and back matter. The ANZIAM Journal, 45(4):b1–b4, April 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-45-issue-4-cover-and-back-matter/513A6C7EF1277BCCC425089A3ECF8D63.

McIntosh:2004:NAC

[1225] A. C. McIntosh, R. O. Weber, and G. N. Mercer. Non-adiabatic combustion waves for general Lewis numbers: wave speed and extinction conditions. The ANZIAM Journal, 46(1):1–16, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonadiabatic-combustion-waves-for-general-lewis-numbers-
wave-speed-and-extinction-conditions/082D8FE098D402729946C24BDFEEE969.

Bohner:2004:OSD

[1226] Martin Bohner and Ondrej Doslý. Oscillation of symplectic dynamic systems. The ANZIAM Journal, 46(1):17–32, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/oscillation-of-symplectic-dynamic-systems/9FD5859E6FA9A4A3B04602ECB543B99E.

Cabada:2004:EST

[1227] Alberto Cabada and Seppo Heikkilä. Existence of solutions of third-order functional problems with nonlinear boundary conditions. The ANZIAM Journal, 46(1):33–44, July 2004. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-of-solutions-of-thirdorder-functional-problems-
with-nonlinear-boundary-conditions/C74A90FC74E39E6BD9F5085D7BD3AB97.

Zhu:2004:ASI

[1228] Detong Zhu. An affine scaling interior point backtracking algorithm for nonlinear constrained optimisation. The ANZIAM Journal, 46 (1):45–66, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/an-affine-scaling-interior-
point-backtracking-algorithm-for-nonlinear-constrained-optimisation/
1CF7101577476D5C7F1525163C6B9D16.

Krishnarayalu:2004:SPM

[1229] M. S. Krishnarayalu. Singular perturbation methods for a class of initial and boundary value problems in multi-parameter classical digital control systems. The ANZIAM Journal, 46(1):67–77, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/singular-perturbation-methods-for-a-
class-of-initial-and-boundary-value-problems-in-multiparameter-
classical-digital-control-systems/CE6B9989249FEBCED478D7EF3338C70E.

Gaffney:2004:NCD

[1230] J. M. Gaffney and C. E. M. Pearce. Nonlinear-cobweb dynamics in the approach to equilibrium. The ANZIAM Journal, 46(1):79–83, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinearcobweb-dynamics-in-the-approach-to-equilibrium/
BABA0E0C56E12B89B3BD3F6C787F4173.

Reich:2004:RLI

[1231] Edgar Reich. E. R. Love’s integral equation for the circular plate condenser. The ANZIAM Journal, 46(1):85–93, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:

//www.cambridge.org/core/journals/anziam-journal/article/e-
r-loves-integral-equation-for-the-circular-plate-condenser/
D29362022CE31A269777D1B476C2ED11.

Zhao:2004:EUR

[1232] Chunshan Zhao and Kaitai Li. On existence, uniqueness and L^{r}- exponential stability for stationary solutions to the MHD equations in three-dimensional domains. The ANZIAM Journal, 46(1):95–109, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-existence-uniqueness-and-lrexponential-stability-
for-stationary-solutions-to-the-mhd-equations-in-threedimensional-
domains/F99071B1F514758A7F4EB4F957CEAF3E.

Jiang:2004:NEF

[1233] Zhuhan Jiang and Xiling Guo. A note on the extension of a family of biorthogonal Coifman wavelet systems. The ANZIAM Journal, 46 (1):111–120, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/note-on-the-extension-of-a-
family-of-biorthogonal-coifman-wavelet-systems/A85DC0FC2D8A0B312504065C8D9240C2.

Xu:2004:PGS

[1234] Rui Xu, Lansun Chen, and M. A. J. Chaplain. Persistence and global stability in a delayed predator–prey system with Hollingtype functional response. The ANZIAM Journal, 46(1):121–141, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/persistence-and-global-stability-in-
a-delayed-predatorprey-system-with-hollingtype-functional-response/
6F5A7A8FE7EBAF4164F2FEA6C881721E.

Rao:2004:EVS

[1235] R. Raghavendra Rao and K. R. Prasad. Effects of velocity-slip and viscosity variation on journal bearings. The ANZIAM Journal, 46(1):143–155, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/effects-of-velocityslip-and-viscosity-variation-
on-journal-bearings/CAE4ED6151FF15ADEFC23BDB5CD36EFE.

Stevic:2004:ABS

[1236] Stevo Stević. Asymptotic behaviour of second-order difference equations. The ANZIAM Journal, 46(1):157–170, July 2004. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-behaviour-of-secondorder-difference-equations/
0D562DF94C873153666F83519A07E64D.

Anonymous:2004:AVIe

[1237] Anonymous. ANZ volume 46 issue 1 cover and front matter. The ANZIAM Journal, 46(1):f1–f4, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-46-issue-1-cover-and-front-matter/407435E6484CEF95FDB07CFB62B67F80.

Anonymous:2004:AVIf

[1238] Anonymous. ANZ volume 46 issue 1 cover and back matter. The ANZIAM Journal, 46(1):b1–b2, July 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-46-issue-1-cover-and-back-matter/2A71B150D491D6073558A55AD4F89A97.

Lee:2004:PMP

[1239] Mi Jin Lee and Jong Yeoul Park. Pontryagin’s maximum principle for optimal control of a non-well-posed parabolic differential equation involving a state constraint. The ANZIAM Journal, 46(2):171–184, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/pontryagins-maximum-principle-for-optimal-control-of-
a-nonwellposed-parabolic-differential-equation-involving-a-
state-constraint/019CB10E2B3D3B7041CBDD9ED75DEAEC.

Shouzhi:2004:FAC

[1240] Yang Shouzhi. A fast algorithm for constructing orthogonal multiwavelets. The ANZIAM Journal, 46(2):185–200, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fast-algorithm-for-constructing-orthogonal-multiwavelets/
D9C680CF9D58C8FAB3E79ECCAD47183E.

Kulkarni:2004:SRU

[1241] Rekha P. Kulkarni and N. Gnaneshwar. Spectral refinement using a new projection method. The ANZIAM Journal, 46(2):203–224, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/spectral-refinement-using-a-new-projection-
method/8BB1A4A85BF39E6DBA3589DABE518C6F.

Amini:2004:SRO

[1242] Keyvan Amini. Solving rank one revised linear systems by the scaled ABS method. The ANZIAM Journal, 46(2):225–236, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/solving-rank-one-revised-linear-systems-by-
the-scaled-abs-method/4F0C35FDC2AA04C3FE0401CDDAC153CA.

Attia:2004:NNC

[1243] Hazem A. Attia and Mohamed E. S. Ahmed. Non-Newtonian conducting fluid flow and heat transfer due to a rotating disk. The ANZIAM Journal, 46(2):237–248, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/nonnewtonian-conducting-
fluid-flow-and-heat-transfer-due-to-a-rotating-disk/A03E4383B2F451CF55884F80E8463AA1.

Brearley:2004:RRB

[1244] Maurice N. Brearley and Neville J. de Mestre. Rolling of a rigid ball on a horizontal deformable surface. The ANZIAM Journal, 46(2): 249–264, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/rolling-of-a-rigid-ball-on-
a-horizontal-deformable-surface/3BD3E45457768883C4F2D31AE7E1EB41.

Leipnik:2004:INI

[1245] Roy B. Leipnik and Charles E. M. Pearce. Independent non-identical fiveparameter gamma-Weibull variates and their sums. The ANZIAM Journal, 46(2):265–271, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/independent-nonidentical-
fiveparameter-gammaweibull-variates-and-their-sums/5D5668A29B6623590E3AEC1EF79849D9.

Suarez:2004:NSH

[1246] Antonio Suárez. Nonnegative solutions for a heterogeneous degenerate competition model. The ANZIAM Journal, 46(2):273–297, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonnegative-solutions-for-a-heterogeneous-degenerate-
competition-model/4A5B90A225DCA3BA9B284310F25E3933.

Purohit:2004:BMN

[1247] D. Purohit and K. S. Chaudhuri. A bioeconomic model of nonselective harvesting of two competing fish species. The ANZIAM Journal, 46(2):299–308, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/bioeconomic-model-
of-nonselective-harvesting-of-two-competing-fish-species/039F71DDDA4D622FB069FA1E19DDFCD4.

Anonymous:2004:AVIg

[1248] Anonymous. ANZ volume 46 issue 2 cover and front matter. The ANZIAM Journal, 46(2):f1–f4, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-46-issue-2-cover-and-front-matter/4D4CDAFD7473D2728F7AFD048DCD1A1C.

Anonymous:2004:AVIh

[1249] Anonymous. ANZ volume 46 issue 2 cover and back matter. The ANZIAM Journal, 46(2):b1–b3, October 2004. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-46-issue-2-cover-and-back-matter/6373AE76F2923137B2BA5A666DC95FAD.

Zhou:2005:DVN

[1250] Yong Zhou. Direction of vorticity and a new regularity criterion for the Navier–Stokes equations. The ANZIAM Journal, 46(3):309– 316, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/direction-of-vorticity-and-
a-new-regularity-criterion-for-the-navierstokes-equations/EE901023FA35DB5ACC8763578B355948.

Huang:2005:TFD

[1251] F. Huang and F. Liu. The time fractional diffusion equation and the advection–dispersion equation. The ANZIAM Journal, 46(3):317– 330, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/time-fractional-diffusion-equation-and-
the-advectiondispersion-equation/E79BE2CCFA92200CDDEBB2490BF15547.

Hou:2005:SSL

[1252] Zhenting Hou, Jiaowan Luo, and Peng Shi. Stochastic stability of linear systems with semi-Markovian jump parameters. The ANZIAM Journal, 46(3):331–340, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/stochastic-stability-of-
linear-systems-with-semimarkovian-jump-parameters/675F51831762E1C9D0EF3D941ACEC28C.

Sastry:2005:CBP

[1253] Challa S. Sastry and P. C. Das. A convolution back projection algorithm for local tomography. The ANZIAM Journal, 46(3):341–360, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/convolution-back-projection-algorithm-for-local-tomography/
FC59C793383E162FCD565A856792DF4B.

Matache:2005:QSM

[1254] Mihaela T. Matache and Valentin Matache. Queueing systems for multiple FBM-based traffic models. The ANZIAM Journal, 46(3):361– 377, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/queueing-systems-for-multiple-fbmbased-
traffic-models/D205EDD025A8487FFB89FA22913A287D.

Belen:2005:ICR

[1255] Selma Belen, C. Yalçin Kaya, and C. E. M. Pearce. Impulsive control of rumours with two broadcasts. The ANZIAM Journal, 46(3):379–391, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/impulsive-control-
of-rumours-with-two-broadcasts/F0A6699A623DA57634C0BF7AA6D0AF96.

Scott:2005:DLM

[1256] Carlton H. Scott and Thomas R. Jefferson. Duality for linear multiplicative programs. The ANZIAM Journal, 46(3):393–397, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/duality-for-linear-multiplicative-programs/
E2CF7BE3B45356ABBDAC1CB42D6FD7C1.

Nelson:2005:BPO

[1257] M. I. Nelson. Bifurcation phenomena for an oxidation reaction in a continuously stirred tank reactor. III. The inhibiting effect of an inert species. The ANZIAM Journal, 46(3):399–416, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/bifurcation-phenomena-for-an-oxidation-reaction-in-
a-continuously-stirred-tank-reactor-iii-the-inhibiting-effect-
of-an-inert-species/441F6A9BE3270271A12352B3C2EBF78F.

Dedic:2005:EMF

[1258] Lj. Dedić, M. Matić, and J. Pecarić. On Euler midpoint formulae. The ANZIAM Journal, 46(3):417–438, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
euler-midpoint-formulae/4587A4D5A559ED8889A53181534438EF.

Stevic:2005:GES

[1259] Stevo Stević. Growth estimates for solutions of nonlinear second-order difference equations. The ANZIAM Journal, 46(3):439–448, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/growth-estimates-for-solutions-of-nonlinear-secondorder-
difference-equations/9A32B8FB6017D7B7EF75198937AA2F86.

Anonymous:2005:AVIa

[1260] Anonymous. ANZ volume 46 issue 3 cover and front matter. The ANZIAM Journal, 46(3):f1–f4, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-46-issue-3-cover-and-front-matter/39511A43E9744D90FE01B839CDC17098.

Anonymous:2005:AVIb

[1261] Anonymous. ANZ volume 46 issue 3 cover and back matter. The ANZIAM Journal, 46(3):b1–b3, January 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-46-issue-3-cover-and-back-matter/652D57DA9771B05A3F891625161F4F5E.

Slodicka:2005:RLS

[1262] Marian Slodicka. A robust linearisation scheme for a nonlinear elliptic boundary value problem: Error estimates. The ANZIAM Journal, 46(4): 449–470, April 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/robust-linearisation-scheme-for-a-nonlinear-
elliptic-boundary-value-problem-error-estimates/9AD6713181A0E37324609DA7117CFEE6.

Xu:2005:RHS

[1263] Honglei Xu, Xinzhi Liu, and Kok Lay Teo. Robust H_{∞} stabilisation with definite attenuance of an uncertain impulsive switched system. The ANZIAM Journal, 46(4):471–484, April 2005. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/robust-h-stabilisation-with-definite-attenuance-of-
an-uncertain-impulsive-switche-system/40BE2E6EE5A7D04629684E2DB30AFD56.

Hunt:2005:RDP

[1264] Emma Hunt. Ramaswami’s duality and probabilistic algorithms for determining the rate matrix for a structured GI/M/1 Markov chain. The ANZIAM Journal, 46(4):485–493, April 2005. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/ramaswamis-duality-and-probabilistic-algorithms-for-
determining-the-rate-matrix-for-a-structured-gim1-markov-chain/
7197007744AAB93E80CC6991A5F0C137.

Wilson:2005:DTC

[1265] D. P. Wilson and D. L. S. McElwain. Diffusion theory can be applied to antibodies attaching to ligand sites. The ANZIAM Journal, 46(4): 495–505, April 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/diffusion-theory-can-be-applied-to-antibodies-
attaching-to-ligand-sites/96ECB4F5EFF3FE94D3A8D114F3C433FE.

Watson:2005:SCH

[1266] C. E. Watson and S. R. Otto. The stability of a curved, heated boundary layer: linear and nonlinear problems. The ANZIAM Journal, 46(4):507–543, April 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/stability-of-a-curved-
heated-boundary-layer-linear-and-nonlinear-problems/844E3F302695A14BEA563AAB743F209B.

Yan:2005:OID

[1267] Jurang Yan, Aimin Zhao, and Linping Peng. Oscillation of impulsive delay differential equations and applications to population dynamics. The ANZIAM Journal, 46(4):545–554, April 2005. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/oscillation-of-impulsive-delay-differential-equations-
and-applications-to-population-dynamics/900D0B5E981725176320E656825533E6.

Pecaric:2005:SII

[1268] J. Pecarić, I. Perić, and A. Vukelić. Sharp integral inequalities based on general Euler two-point formulae. The ANZIAM Journal, 46(4):555–574, April 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/sharp-integral-inequalities-based-on-general-
euler-twopoint-formulae/D944DCCB6595C9E2963D406A2FFB7C74.

Attia:2005:PJC

[1269] Hazem Ali Attia. Point-joint coordinate formulation for the dynamic analysis of generalised planar linkages. The ANZIAM Journal, 46(4):575–589, April 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/pointjoint-coordinate-
formulation-for-the-dynamic-analysis-of-generalised-planar-
linkages/4FE70B2D02550D1BC6A26EF2C163F4EF.

Anonymous:2005:CV

[1270] Anonymous. Contents of volume 46. The ANZIAM Journal, 46 (4):591–593, April 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/contents-of-volume-46/
E3BB937A34E242CA00FF4297F4B58538.

Anonymous:2005:AVIc

[1271] Anonymous. ANZ volume 46 issue 4 cover and front matter. The ANZIAM Journal, 46(4):f1–f4, April 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-46-issue-4-cover-and-front-matter/D1160FEB17609C7341468DF36CB3A9D1.

Anonymous:2005:AVId

[1272] Anonymous. ANZ volume 46 issue 4 cover and back matter. The ANZIAM Journal, 46(4):b1–b2, April 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-46-issue-4-cover-and-back-matter/5BF6422BD60126A56919F15D734AB9EB.

Hunt:2005:MHB

[1273] Emma Hunt and Charles Pearce. In memoriam: Hilary Booth. The ANZIAM Journal, 47(1):i, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/in-
memoriam-hilary-booth/7CAF74B3420B8AD3AF5677CDF35F7E27.

Penot:2005:SPA

[1274] Jean-Paul Penot and Constantin Z{ă \atop .}linescu. Some problems about the representation of monotone operators by convex functions. The ANZIAM Journal, 47(1):1–20, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/some-problems-about-the-
representation-of-monotone-operators-by-convex-functions/F6BF4DABB104B3ABDA09278DBA23905A.

McBain:2005:PPTa

[1275] G. D. McBain. Plane poloidal-toroidal decomposition of doubly periodic vector fields. Part 1. Fields with divergence. The ANZIAM Journal, 47 (1):21–38, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/plane-poloidaltoroidal-decomposition-
of-doubly-periodic-vector-fields-part-1-fields-with-divergence/
48C94CA8C79A7A29E06F6F8AC12E1297.

McBain:2005:PPTb

[1276] G. D. McBain. Plane poloidal-toroidal decomposition of doubly periodic vector fields. Part 2. The Stokes equations. The ANZIAM Journal, 47 (1):39–50, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/plane-poloidaltoroidal-decomposition-
of-doubly-periodic-vector-fields-part-2-the-stokes-equations/
3F97C530A5F5CA5BEC74E01130CB96BC.

Parks:2005:ESO

[1277] Jong Yeoul Parks and Sun Hye Park. Existence of solutions and optimal control problems for hyperbolic hemivariational inequalities. The ANZIAM Journal, 47(1):51–63, July 2005. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-of-solutions-and-optimal-control-problems-
for-hyperbolic-hemivariational-inequalities/EA6DD8868BA0F49D2C1E615439841247.

Fakhar:2005:SAR

[1278] K. Fakhar, Zu-Chi Chen, and Xiaoda Ji. Symmetry analysis of rotating fluid. The ANZIAM Journal, 47(1):65–74, July 2005. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/symmetry-analysis-of-rotating-fluid/DCFF040070016819EE699B00C3D1899F.

Emamizadeh:2005:ESS

[1279] B. Emamizadeh and F. Bahrami. Existence of seamount steady vortex flows. The ANZIAM Journal, 47(1):75–88, July 2005. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-of-seamount-steady-vortex-flows/913A5A836421CBB04B5FD339C2443317.

Zhou:2005:EPS

[1280] Zhan Zhou, Jianshe Yu, and Zhiming Guo. The existence of periodic and subharmonic solutions to subquadratic discrete Hamiltonian systems. The ANZIAM Journal, 47(1):89–102, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-of-periodic-and-subharmonic-solutions-to-
subquadratic-discrete-hamiltonian-systems/41535E12C291B4FA17A978736DA8EBFF.

Haese:2005:CAP

[1281] P. M. Haese. Cavities at atmospheric pressure behind two-dimensional bodies at an angle of attack. The ANZIAM Journal, 47(1):103–119, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/cavities-at-atmospheric-pressure-behind-twodimensional-
bodies-at-an-angle-of-attack/E60A761212B75D1F9F91E39F65FFC771.

Mazumdar:2005:EPA

[1282] J. Mazumdar, A. Ghosh, J. S. Hewitt, and P. K. Bhattacharya. Elastic plastic analysis of shallow shells — a new approach. The ANZIAM Journal, 47(1):121–130, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/elastic-
plastic-analysis-of-shallow-shellsa-new-approach/747A12071B996AFC787C633142B61A2A.

Irk:2005:NIR

[1283] Dursun Irk, Idris Dag, and Abdülkadir Dogan. Numerical integration of the RLW equation using cubic splines. The ANZIAM Journal, 47(1):131–142, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/numerical-integration-
of-the-rlw-equation-using-cubic-splines/9BAB1A67B270B1AC5699985CA34585D8.

Anonymous:2005:AVIe ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-47-issue-1-cover-and-front-matter/BF1C1FF78F23739BBC0EB205BB8B4EE7.

Anonymous:2005:AVIf

[1285] Anonymous. ANZ volume 47 issue 1 cover and back matter. The ANZIAM Journal, 47(1):b1–b3, July 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-47-issue-1-cover-and-back-matter/172699AAC82730D0A049705BDC0374FA.

Craven:2005:OCI

[1286] B. D. Craven. Optimal control on an infinite domain. The ANZIAM Journal, 47(2):143–153, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
control-on-an-infinite-domain/B480340967DA17634E088584D46853BA.

Antczak:2005:SPC

[1287] Tadeusz Antczak. Saddle point criteria and duality in multiobjective programming via an η-approximation method. The ANZIAM Journal, 47(2):155–172, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/saddle-point-criteria-
and-duality-in-multiobjective-programming-via-an-approximation-
method/CED590EF4870F5A9848DC7251FEC110A.

Nelson:2005:AMM

[1288] M. I. Nelson, X. D. Chen, and M. J. Sexton. Analysis of the Michaelis–Menten mechanism in an immobilised enzyme reactor. The ANZIAM Journal, 47(2):173–184, October 2005. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/analysis-of-the-michaelismenten-mechanism-in-an-immobilised-
enzyme-reactor/7C7935A2B2EE7B291B0F6026E64BAAA6.

Stokes:2005:UFI

[1289] T. E. Stokes, G. C. Hocking, and L. K. Forbes. Unsteady flow induced by a withdrawal point beneath a free surface. The ANZIAM Journal, 47(2): 185–202, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/unsteady-flow-induced-by-a-
withdrawal-point-beneath-a-free-surface/CA556D798865DA8E9B13BAE37841F83B.

Liu:2005:CPS

[1290] Xinzhi Liu, Xuemin Shen, and Yi Zhang. A comparison principle and stability for large-scale impulsive delay differential systems. The ANZIAM Journal, 47(2):203–235, October 2005. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/comparison-principle-and-stability-for-largescale-impulsive-
delay-differential-systems/F009021EF7110D3F56FDCFC8CF58DFFB.

Dal:2005:BVP

[1291] F. Dal and G. Sh. Guseinov. A boundary value problem for second-order nonlinear difference equations on the integers. The ANZIAM Journal, 47 (2):237–248, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/boundary-value-problem-for-
secondorder-nonlinear-difference-equations-on-the-integers/
5149D1DC51CB3174686455A8ACE913B5.

Zhang:2005:PSN

[1292] Zhengqiu Zhang, Yusen Zhu, and Biwen Li. Periodic solutions of a nonlinear oscillatory system with two degrees of freedom. The ANZIAM Journal, 47(2):249–263, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/periodic-solutions-of-
a-nonlinear-oscillatory-system-with-two-degrees-of-freedom/
098C09615DE0075DEB02CC4BE56D7D3E.

Narain:2005:FNI

[1293] Laxmi Narain and P. C. Bagga. Flowshop/no-idle scheduling to minimise the mean flowtime. The ANZIAM Journal, 47(2):265–275, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/flowshopnoidle-scheduling-to-minimise-the-mean-flowtime/
9EDFB434D63C300FB437A108ACC40021.

Leipnik:2005:DSM

[1294] Roy B. Leipnik and C. E. M. Pearce. Diversity sensitivity and multimodal Bayesian statistical analysis by relative entropy. The ANZIAM Journal, 47(2):277–287, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/diversity-sensitivity-
and-multimodal-bayesian-statistical-analysis-by-relative-entropy/
86E33D49AD8A34110B46C47633401D45.

Anonymous:2005:AVIg

[1295] Anonymous. ANZ volume 47 issue 2 cover and front matter. The ANZIAM Journal, 47(2):f1–f4, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-47-issue-2-cover-and-front-matter/486F41405445DF41B508B8FC76254333.

Anonymous:2005:AVIh

[1296] Anonymous. ANZ volume 47 issue 2 cover and back matter. The ANZIAM Journal, 47(2):b1–b3, October 2005. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-47-issue-2-cover-and-back-matter/39D068B0C22577182C9B78C8FCE9A89B.

Baxter:2006:HPG

[1297] R. J. Baxter. Hyperelliptic parametrisation of the generalised order parameter of the N = 3 chiral Potts model. The ANZIAM Journal, 47(3): 289–307, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/hyperelliptic-parametrisation-
of-the-generalised-order-parameter-of-the-n-3-chiral-potts-
model/D76551C482B587977C9F1F8BF0D5239F.

Agarwal:2006:CSP

[1298] Ravi P. Agarwal, Donal O’Regan, and Patricia J. Y. Wong. On constantsign periodic solutions in modelling the spread of interdependent epidemics. The ANZIAM Journal, 47(3):309–332, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
constantsign-periodic-solutions-in-modelling-the-spread-of-
interdependent-epidemics/F007E7538D9EB6C3E4CB3D838801BBFC.

Munro:2006:IIS

[1299] Angus I. S. Munro and Larry K. Forbes. Including ionisation in a simple model of single-bubble sonoluminescence. The ANZIAM Journal, 47(3): 333–358, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/including-ionisation-in-a-simple-
model-of-singlebubble-sonoluminescence/CA73B06A27FE5CA408EEEDA776B78B2E.

Berenhaut:2006:EBT

[1300] Kenneth S. Berenhaut, Eva G. Goedhart, and Stevo Stević. Explicit bounds for third-order difference equations. The ANZIAM Journal, 47 (3):359–366, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/explicit-bounds-for-thirdorder-
difference-equations/15F5AF30CF75026207146F696500FB92.

Pecaric:2006:GCS

[1301] J. Pecarić and I. Franjić. Generalisation of a corrected Simpson’s formula. The ANZIAM Journal, 47(3):367–385, January 2006. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalisation-of-a-corrected-simpsons-formula/BAF3F7A604F8821513AF6A09C6891041.

Duff:2006:GER

[1302] G. F. D. Duff, R. B. Leipnik, and C. E. M. Pearce. Guide expansions for the recursive parametric solution of polynomial dynamical systems. The ANZIAM Journal, 47(3):387–396, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/guide-expansions-for-the-recursive-parametric-solution-
of-polynomial-dynamical-systems/62A98D47601711466A154F23C81C7D27.

Anestopoulos:2006:LFA

[1303] C. N. Anestopoulos and E. E. Argyropoulos. On the low frequency asymptotics for the 2-D electromagnetic transmission problem. The ANZIAM Journal, 47(3):397–411, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-low-frequency-asymptotics-for-the-2d-electromagnetic-transmission-
problem/6ADBE03230E52C0A63DB851D3FC5A2A3.

Rao:2006:SAR

[1304] V. Sree Hari Rao and P. Raja Sekhara Rao. Stability analysis of resource-consumer dynamic models. The ANZIAM Journal, 47 (3):413–438, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/stability-analysis-of-
resourceconsumer-dynamic-models/7E8CB9E860BA2279926C485D8D9A592D.

Anonymous:2006:AVIa

[1305] Anonymous. ANZ volume 47 issue 3 cover and front matter. The ANZIAM Journal, 47(3):f1–f4, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-47-issue-3-cover-and-front-matter/5EF389E75219B3783D04624F339F6543.

Anonymous:2006:AVIb

[1306] Anonymous. ANZ volume 47 issue 3 cover and back matter. The ANZIAM Journal, 47(3):b1–b3, January 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-47-issue-3-cover-and-back-matter/0350CD5FDBF35C841F4FFDE300383CB3.

Huy:2006:SGO

[1307] N. Q. Huy, V. Jeyakumar, and G. M. Lee. Sufficient global optimality conditions for multi-extremal smooth minimisation problems with bounds and linear matrix inequality constraints. The ANZIAM Journal, 47(4):439–450, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/sufficient-global-
optimality-conditions-for-multiextremal-smooth-minimisation-
problems-with-bounds-and-linear-matrix-inequality-constraints/
B153AEA2F1B6EF8A7C3BF570E6ED80EF.

inYun:2006:SCS

[1308] Beong in Yun. Sigmoidal cosine series on the interval. The ANZIAM Journal, 47(4):451–475, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/sigmoidal-cosine-series-
on-the-interval/E09B36D981066AB8AD422EC69DE60F54.

Zhu:2006:CFA

[1309] Song-Ping Zhu. A closed-form analytical solution for the valuation of convertible bonds with constant dividend yield. The ANZIAM Journal, 47(4):477–494, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/closedform-analytical-
solution-for-the-valuation-of-convertible-bonds-with-constant-
dividend-yield/DE68297EB9649769185D2CB841E8BE28.

Key:2006:PPH

[1310] E. S. Key, M. M. Klosek, and D. Abbott. On Parrondo’s paradox: how to construct unfair games by composing fair games. The ANZIAM Journal, 47(4):495–511, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-

parrondos-paradox-how-to-construct-unfair-games-by-composing-
fair-games/B1BF6673651973BF0FDBFA875868350D.

Shouzhi:2006:ACB

[1311] Yang Shouzhi. An algorithm for constructing biorthogonal multiwavelets with higher approximation orders. The ANZIAM Journal, 47(4):513–526, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/an-algorithm-for-constructing-biorthogonal-
multiwavelets-with-higher-approximation-orders/219D70670C80EF7C154953CEFD311A81.

Trenham:2006:CTT

[1312] Claire E. Trenham and Larry K. Forbes. A comparison of two- and three-variable models for combustion in sealed containers. The ANZIAM Journal, 47(4):527–540, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/comparison-of-two-and-
threevariable-models-for-combustion-in-sealed-containers/FDA885B2D9A7610DD5E3EBF89FCBD8C6.

Munzir:2006:CMS

[1313] S. Munzir, L. S. Jennings, and M. T. Koh. Computational models satisfying relative angle constraints for 2-dimensional segmented bodies. The ANZIAM Journal, 47(4):541–554, April 2006. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/computational-models-satisfying-relative-angle-constraints-
for-2dimensional-segmented-bodies/392B3DF105BF26374A867E58DC0F6768.

Guo:2006:RRS

[1314] Faming Guo, Bin Tang, and Falun Huang. Robustness with respect to small delays for exponential stability of abstract differential equations in Banach spaces. The ANZIAM Journal, 47(4): 555–568, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/robustness-with-respect-to-
small-delays-for-exponential-stability-of-abstract-differential-
equations-in-banach-spaces/8E1D7CDF0A46BBAE014F3A2269D9C079.

Fu:2006:RGC

[1315] Yan-Ming Fu, Di Wu, and Guang-Ren Duan. Robust guaranteed cost control for descriptor systems with Markov jumping parameters and state delays. The ANZIAM Journal, 47 (4):569–580, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/robust-guaranteed-
cost-control-for-descriptor-systems-with-markov-jumping-parameters-
and-state-delays/623CF3B1A6136A20CB6558B74EE20A18.

Leach:2006:NST

[1316] J. A. Leach. A note on the small-time development of the solution to a coupled, nonlinear, singular reaction–diffusion system. The ANZIAM Journal, 47(4):581–591, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/note-on-the-smalltime-
development-of-the-solution-to-a-coupled-nonlinear-singular-
reactiondiffusion-system/1D69CA79E81C1FE70455A120BC6E3290.

Anonymous:2006:CV

[1317] Anonymous. Contents of volume 47. The ANZIAM Journal, 47 (4):593–595, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/contents-of-volume-47/
D6A5055ABE0F9D47BEA843C4693373AC.

Anonymous:2006:AVIc

[1318] Anonymous. ANZ volume 47 issue 4 cover and front matter. The ANZIAM Journal, 47(4):f1–f4, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-47-issue-4-cover-and-front-matter/F15B994A682E0E1CAACEFE391AAED677.

Anonymous:2006:AVId

[1319] Anonymous. ANZ volume 47 issue 4 cover and back matter. The ANZIAM Journal, 47(4):b1–b2, April 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-47-issue-4-cover-and-back-matter/7A537FA8E094564931BE96AF18E91A15.

Baake:2006:FFG

[1320] Michael Baake, Uwe Grimm, and Harald Jockusch. Freely forming groups: Trying to be rare. The ANZIAM Journal, 48(1):1–10, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/freely-forming-groups-trying-to-be-rare/FCC32F4808B73479A23FA21D40B192A2.

Garcke:2006:PSG

[1321] Jochen Garcke, Markus Hegland, and Ole Nielsen. Parallelisation of sparse grids for large scale data analysis. The ANZIAM Journal, 48(1): 11–22, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/parallelisation-of-sparse-grids-for-
large-scale-data-analysis/39A59F768A5F4983F8773B0733063750.

Orea-Flores:2006:WDC

[1322] I. Orea-Flores, M. A. Acevedo, and J. López-Bonilla. Wavelet and discrete cosine transforms for inserting information into BMP images. The ANZIAM Journal, 48(1):23–35, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/wavelet-
and-discrete-cosine-transforms-for-inserting-information-into-
bmp-images/1CF04D3972EC784D67DA8D147A9D1A5F.

Xiang:2006:BPS

[1323] Zhaoyin Xiang, Qiong Chen, and Chunlai Mu. Blowup properties for several diffusion systems with localised sources. The ANZIAM Journal, 48(1):37–56, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/blowup-properties-for-
several-diffusion-systems-with-localised-sources/2FDF475F9E28E959B7C49E76103FD703.

Lin:2006:TWP

[1324] Guojian Lin and Rong Yuan. Travelling waves for the population genetics model with delay. The ANZIAM Journal, 48(1):57–71, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/travelling-waves-for-the-population-genetics-model-
with-delay/0710AD421C75F1504C8ADD7EFA227325.

Yueh:2006:EEI

[1325] Wen-Chyuan Yueh and Sui Sun Cheng. Explicit eigenvalues and inverses of several Toeplitz matrices. The ANZIAM Journal, 48(1):73–97, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/explicit-eigenvalues-and-inverses-of-several-toeplitz-
matrices/A5BF912462D961E7A8017FA5729BC350.

Zafer:2006:ECM

[1326] A. Zafer. The exponential of a constant matrix on time scales. The ANZIAM Journal, 48(1):99–106, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/exponential-
of-a-constant-matrix-on-time-scales/0D54F0AAFE9085F5652224E0A1B4DF11.

Mandal:2006:WWS

[1327] B. N. Mandal and Soumen De. Water wave scattering by two submerged nearly vertical barriers. The ANZIAM Journal, 48 (1):107–117, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/water-wave-scattering-
by-two-submerged-nearly-vertical-barriers/FB9BD32247ABBC7783AF6C16AC1CC102.

Yan:2006:SEM

[1328] Ping Yan and Shengqiang Liu. SEIR epidemic model with delay. The ANZIAM Journal, 48(1):119–134, July 2006. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/seir-epidemic-model-with-delay/906A072AA2ACB250ADB0D1D9DCC38D8F.

Uslu:2006:DAS

[1329] Kemal Uslu. On the discrete asymptotic stability conditions of perturbed linear discrete systems with periodic coefficients. The ANZIAM Journal, 48(1):135–142, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-discrete-asymptotic-
stability-conditions-of-perturbed-linear-discrete-systems-with-
periodic-coefficients/014F9FFCA12FCAB250501EBC3E32DCF2.

Qureshi:2006:SMG

[1330] M. I. Qureshi, M. Sadiq Khan, M. A. Pathan, and N. U. Khan. Some multivariable Gaussian hypergeometric extensions of the Preece theorem. The ANZIAM Journal, 48(1):143–150, July 2006. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/some-multivariable-gaussian-hypergeometric-extensions-
of-the-preece-theorem/6F0DE368CB96667763E623E727A83FF5.

Anonymous:2006:AVIe ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-48-issue-1-cover-and-front-matter/2BCB3BD5329FE812365C318AF632DABA.

Anonymous:2006:AVIf

[1332] Anonymous. ANZ volume 48 issue 1 cover and back matter. The ANZIAM Journal, 48(1):b1–b3, July 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-48-issue-1-cover-and-back-matter/D86A7614094ECF4061F99DB00BEC71C0.

Szekeres:2006:CSP

[1333] George Szekeres and Lindsay Peters. Computer solution to the 17-point Erdős–Szekeres problem. The ANZIAM Journal, 48(2): 151–164, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/computer-solution-
to-the-17point-erdosszekeres-problem/0EC7876789232266D60439A4C00D86D9.

Hocking:2006:SPB

[1334] G. C. Hocking. Steady Prandtl–Batchelor flows past a circular cylinder. The ANZIAM Journal, 48(2):165–177, October 2006. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/steady-prandtlbatchelor-flows-past-a-circular-cylinder/
BA6E0CE9A2777519B0F88F3EB1BFEB5D.

Le:2006:SSV

[1335] Vy Khoi Le. Sub-supersolutions in a variational inequality related to a sandpile problem. The ANZIAM Journal, 48(2):179–197, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/subsupersolutions-in-a-variational-inequality-related-
to-a-sandpile-problem/6472D2BC122B074C023C9E4C4AFC324E.

Sterling:2006:DAT

[1336] G. Sterling, G. D. McBain, J. A. Harris, and M. Boland. Drainage after total knee replacement. The ANZIAM Journal, 48(2):199–210, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/drainage-after-total-knee-replacement/A5B91297E5B6D7F875ADAE3DE293FA07.

Berger:2006:CNM

[1337] A. Berger and T. P. Hill. A characterisation of Newton maps. The ANZIAM Journal, 48(2):211–223, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/characterisation-
of-newton-maps/0CAA37E889EEDEA90B3FCDDB9C751173.

Ozugurlu:2006:NSW

[1338] E. Özugurlu and J.-M. Vanden-Broeck. A note on solitary waves with variable surface tension in water of infinite depth. The ANZIAM Journal, 48(2):225–235, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
solitary-waves-with-variable-surface-tension-in-water-of-infinite-
depth/BB9A78B664B49814AF3EC93EF0E781B8.

Cabada:2006:SDF

[1339] A. Cabada, J. Ángel Cid, and S. Heikkilä. Solvability of discontinuous functional differential systems in I_{∞}(M). The ANZIAM Journal, 48(2): 237–243, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/solvability-of-discontinuous-
functional-differential-systems-in-lm/B1082BC27FC4FEBBBBF24D087C703B05.

Galewska:2006:SSP

[1340] Elzbieta Galewska and Marek Galewski. On the stability of solutions for the p(x)-Laplacian equation and some applications to optimisation problems with state constraints. The ANZIAM Journal, 48(2):245–257, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-the-stability-of-solutions-for-the-pxlaplacian-
equation-and-some-applications-to-optimisation-problems-with-
state-constraints/3E26C22913991A3D6B3992C0D13F3353.

Liu:2006:RSI

[1341] Xinzhi Liu and Hongtao Zhang. Robust stability of impulsive switched systems with disturbance. The ANZIAM Journal, 48(2):259–270, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/robust-stability-of-impulsive-switched-systems-
with-disturbance/685FCEB166AF2DDAC3D2B81CE46C3778.

Xu:2006:SAK

[1342] Houbao Xu. Stability analysis of a k-out-of-N: G reparable system. The ANZIAM Journal, 48(2):271–284, October 2006. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stability-analysis-of-a-koutofng-reparable-system/46BDCF43E0A713EEAB94A7EF72396708.

Iyai:2006:ENC

[1343] Davies Iyai. Euclidean null controllability of infinite neutral differential systems. The ANZIAM Journal, 48(2):285–293, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/euclidean-null-controllability-of-infinite-neutral-
differential-systems/F986BCD7A9DBB629419FC595AC9005B6.

Anonymous:2006:C

[1344] Anonymous. Corrigendum. The ANZIAM Journal, 48(2):295, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/corrigendum/4E1EBAF8CB781AE955BA17AE108C164C.

Anonymous:2006:AVIg

[1345] Anonymous. ANZ volume 48 issue 2 cover and front matter. The ANZIAM Journal, 48(2):f1–f4, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-48-issue-2-cover-and-front-matter/34DD29D99538F83A3AA54900B9E90BE1.

Anonymous:2006:AVIh

[1346] Anonymous. ANZ volume 48 issue 2 cover and back matter. The ANZIAM Journal, 48(2):b1–b3, October 2006. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-48-issue-2-cover-and-back-matter/D2F08068BED13AAD66C5F2B2C6BB5489.

Abramov:2007:MQS

[1347] Vyacheslav M. Abramov. Multiserver queueing systems with retrials and losses. The ANZIAM Journal, 48(3):297–314, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/multiserver-queueing-systems-with-retrials-
and-losses/861FB57041DB4648990FF204F18AE5CD.

Ugon:2007:FLC

[1348] J. Ugon, S. Kouhbor, M. Mammadov, A. Rubinov, and A. Kruger. Facility location via continuous optimization with discontinuous objective functions. The ANZIAM Journal, 48(3):315–325, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/facility-location-via-continuous-optimization-with-
discontinuous-objective-functions/CD7085703ECF6DB2BCB9A00F0F51C3FB.

Leipnik:2007:MFB

[1349] Roy B. Leipnik and Charles E. M. Pearce. The multivariate Faà di Bruno formula and multivariate Taylor expansions with explicit integral remainder term. The ANZIAM Journal, 48(3):327–341, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/multivariate-faa-di-bruno-formula-and-multivariate-
taylor-expansions-with-explicit-integral-remainder-term/BA960C9CAB40DE52DB84799BEB0C730A.

Amat:2007:DFT

[1350] Sergio Amat, Sonia Busquier, and Sergio Plaza. On the dynamics of a family of third-order iterative functions. The ANZIAM Journal, 48(3): 343–359, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/on-the-dynamics-of-a-family-
of-thirdorder-iterative-functions/0FF7E7D409C335E7C5B7EC4A77951055.

Doha:2007:ESG

[1351] E. H. Doha and A. H. Bhrawy. Efficient spectral-Galerkin algorithms for direct solution of the integrated forms of second-order equations using ultraspherical polynomials. The ANZIAM Journal, 48(3):361– 386, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/efficient-spectralgalerkin-
algorithms-for-direct-solution-of-the-integrated-forms-of-secondorder-
equations-using-ultraspherical-polynomials/B66E4E2A8A81BD83DB07A796033D77BB.

Bakula:2007:GHI

[1352] M. Klaricić Bakula and J. Pecarić. Generalized Hadamard’s inequalities based on general Euler 4-point formulae. The ANZIAM Journal, 48(3):387–404, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/generalized-hadamards-
inequalities-based-on-general-euler-4point-formulae/AFA01CA1B683AD5F6E4BCB47A94670C8.

Nadarajah:2007:DLP

[1353] Saralees Nadarajah and Samuel Kotz. On a distribution of Leipnik and Pearce. The ANZIAM Journal, 48(3):405–407, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
a-distribution-of-leipnik-and-pearce/0BEAE4EBC857526BB64017965B57378C.

Gani:2007:NTS

[1354] J. Gani and L. Stals. A note on three stochastic processes with immigration. The ANZIAM Journal, 48(3):409–418, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-three-stochastic-processes-with-immigration/
FD0ED0D97BE40604A1D32C2ECC4B238F.

Krnic:2007:HIE

[1355] Mario Krnić and Josip Pecarić. A Hilbert inequality and an Euler– Maclaurin summation formula. The ANZIAM Journal, 48(3):419–431, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/hilbert-inequality-and-an-eulermaclaurin-
summation-formula/95A1A033E4CEF38EDC7070B53DC5E1D9.

Singh:2007:RCE

[1356] Baljeet Singh. Reflection coefficients and energy ratios in a micropolar thermoelastic medium without energy dissipation. The ANZIAM Journal, 48(3):433–447, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/reflection-coefficients-
and-energy-ratios-in-a-micropolar-thermoelastic-medium-without-
energy-dissipation/961D1F6DA95356CE91922D2CCB79367A.

Anonymous:2007:AVIa

[1357] Anonymous. ANZ volume 48 issue 3 cover and front matter. The ANZIAM Journal, 48(3):f1–f4, January 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-48-issue-3-cover-and-front-matter/3DA47FA2999240836E86A6F704999E2E.

Anonymous:2007:AVIb

[1358] Anonymous. ANZ volume 48 issue 3 cover and back matter. The ANZIAM Journal, 48(3):b1–b3, January 2007. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-48-issue-3-cover-and-back-matter/AA93142BC0009A669298B3A29F22F975.

Weir:2007:IDA

[1359] Graham Weir. Initial deformation about convex surfaces formed from identical, rough elastic-plastic bodies which approach along their normal at first contact. The ANZIAM Journal, 48(4):449–459, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/initial-deformation-about-convex-surfaces-formed-from-
identical-rough-elasticplastic-bodies-which-approach-along-
their-normal-at-first-contact/206EA0F14F504B23495ECB46A3AE0FCF.

Lawrie:2007:UMP

[1360] J. Lawrie and J. W. Hearne. Using Modified Proper Orthogonal Decomposition (MPOD) for reducing ecosystem models. The ANZIAM Journal, 48(4):461–473, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/using-modified-proper-
orthogonal-decomposition-mpod-for-reducing-ecosystem-models/
DC360FF894CCEB645177B62E583F89E6.

Dunstall:2007:RCS

[1361] Simon Dunstall and Graham Mills. Robustness of cyclic schedules for the charging of batteries. The ANZIAM Journal, 48(4):475–492, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/robustness-of-cyclic-schedules-for-the-charging-of-
batteries/8CC4AE520DE3B12453085590AFC4B6A9.

Leipnik:2007:TMM

[1362] R. B. Leipnik and C. E. M. Pearce. Thermodynamics, mnemonic matrices and generalized inverses. The ANZIAM Journal, 48(4):493–501, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/thermodynamics-mnemonic-matrices-and-generalized-inverses/
368B059567DC0C115E6599F01274DA37.

Haynes:2007:EST

[1363] P. D. Haynes and S. K. Lucas. Extension of a short-time solution of the diffusion equation with application to micropore diffusion in a finite system. The ANZIAM Journal, 48(4):503–521, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/extension-of-a-shorttime-solution-of-the-diffusion-
equation-with-application-to-micropore-diffusion-in-a-finite-
system/7E423168650520E5551647344955024F.

Qi:2007:CMF

[1364] Feng Qi. A completely monotonic function involving the divided difference of the psi function and an equivalent inequality involving sums. The ANZIAM Journal, 48(4):523–532, April 2007. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/completely-monotonic-function-involving-the-divided-
difference-of-the-psi-function-and-an-equivalent-inequality-
involving-sums/96EF740B9CE5392AA8E1C44667CDB3AF.

Grundy:2007:HIV

[1365] R. E. Grundy. Hermite interpolation visits ordinary two-point boundary value problems. The ANZIAM Journal, 48(4):533–552, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/hermite-interpolation-visits-ordinary-twopoint-boundary-
value-problems/D2078E37532B31B6A06994FD53D4AED5.

Tseng:2007:GWT

[1366] Kuei-Lin Tseng, Gou-Sheng Yang, and Sever S. Dragomir. Generalizations of a weighted trapezoidal inequality for monotonic functions and applications. The ANZIAM Journal, 48(4):553–566, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/generalizations-of-a-weighted-trapezoidal-inequality-
for-monotonic-functions-and-applications/A31E2C48F098B7B5B34667CFB7980D39.

Milovanovic:2007:MET

[1367] Gradimir V. Milovanović and Miodrag M. Spalević. Monotonicity of the error term in Gauss–Turán quadratures for analytic function. The ANZIAM Journal, 48(4):567–581, April 2007. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/monotonicity-of-the-error-term-in-gaussturan-quadratures-
for-analytic-function/49F7D1B1524B4D7E4CAED9170D699F04.

Tchuenche:2007:PDE

[1368] Jean M. Tchuenche. Patient-dependent effects in disease control: a mathematical model. The ANZIAM Journal, 48(4):583– 596, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/patientdependent-effects-in-disease-
control-a-mathematical-model/A0D44663C88476BF352EB98313642F6F.

Anonymous:2007:CV

[1369] Anonymous. Contents of volume 48. The ANZIAM Journal, 48 (4):597–599, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/contents-of-volume-48/
A9262F62401BCDDAD3D55C340CE2372A.

Anonymous:2007:AVIc

[1370] Anonymous. ANZ volume 48 issue 4 cover and front matter. The ANZIAM Journal, 48(4):f1–f4, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-48-issue-4-cover-and-front-matter/0756FC047A073327285F6A2A1EFA7EEA.

Anonymous:2007:AVId

[1371] Anonymous. ANZ volume 48 issue 4 cover and back matter. The ANZIAM Journal, 48(4):b1–b2, April 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-48-issue-4-cover-and-back-matter/3DDE6CA63358A384FE8538618FAC625B.

Hintermuller:2007:MIF

[1372] M. Hintermüller. Mesh independence and fast local convergence of a primal-dual active-set method for mixed control-state constrained elliptic control problems. The ANZIAM Journal, 49(1): 1–38, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/mesh-independence-and-fast-
local-convergence-of-a-primaldual-activeset-method-for-mixed-
controlstate-constrained-elliptic-control-problems/CBC3D9DE8FD643E3A81A660886FCF460.

VanBrunt:2007:EEH

[1373] B. Van Brunt and M. Vlieg Hulstman. Evolution equations having conservation laws with flux characteristics. The ANZIAM Journal, 49(1):39–52, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/evolution-equations-having-
conservation-laws-with-flux-characteristics/691B4F636C8EB43D06AF2B22BB714D4E.

Lasky:2007:IVF

[1374] P. D. Lasky, A. W. C. Lun, and R. B. Burston. Initial value formalism for Lemaître–Tolman–Bondi collapse. The ANZIAM Journal, 49(1):53–73, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/initial-value-formalism-for-lemaitretolmanbondi-
collapse/1A034B703917C9C424C1AB24E297BA12.

Galewski:2007:NSA

[1375] Marek Galewski. A note on the stability and the approximation of solutions for a Dirichlet problem with p(x)-Laplacian. The ANZIAM Journal, 49(1):75–83, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/note-on-the-stability-
and-the-approximation-of-solutions-for-a-dirichlet-problem-
with-pxlaplacian/0FD7E80BA81B58951CF00D48FAA0F85E.

Shouzhi:2007:BIM

[1376] Yang Shouzhi. Biorthogonal interpolatory multiscaling functions and corresponding multiwavelets. The ANZIAM Journal, 49(1): 85–97, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/biorthogonal-interpolatory-
multiscaling-functions-and-corresponding-multiwavelets/0F9E6F7FB62BFA068BF335B79B748C38.

Martin:2007:CNF

[1377] P. A. Martin. Calculating the near field of a line of sources using Mellin transforms. The ANZIAM Journal, 49(1):99–109, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/calculating-the-near-field-of-a-line-of-sources-using-
mellin-transforms/705BBFB3819CF38A04A65C955DD48F32.

Ma:2007:DDS

[1378] Shuping Ma, Xinzhi Liu, and Chenghui Zhang. Delay-dependent stability and stabilization of uncertain discrete-time Markovian jump singular systems with time delay. The ANZIAM Journal, 49(1):111–129, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-

journal/article/delaydependent-stability-and-stabilization-
of-uncertain-discretetime-markovian-jump-singular-systems-with-
time-delay/219649B1788C7814AB668B5E341C67EB.

Zhu:2007:CSD

[1379] Song-Ping Zhu and Yinglong Zhang. A comparative study of the direct boundary element method and the dual reciprocity boundary element method in solving the Helmholtz equation. The ANZIAM Journal, 49(1):131–150, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/comparative-study-
of-the-direct-boundary-element-method-and-the-dual-reciprocity-
boundary-element-method-in-solving-the-helmholtz-equation/B23FCE31CFF5B36E0C9786FC94603814.

Anonymous:2007:AVIe

[1380] Anonymous. ANZ volume 49 issue 1 cover and front matter. The ANZIAM Journal, 49(1):f1–f4, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-49-issue-1-cover-and-front-matter/17AD78C049104856F1625CB106064B0A.

Anonymous:2007:AVIf

[1381] Anonymous. ANZ volume 49 issue 1 cover and back matter. The ANZIAM Journal, 49(1):b1–b3, July 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-49-issue-1-cover-and-back-matter/6605505C5D32F17EEA57E6EBDC460A46.

Chapman:2007:NMA

[1382] S. J. Chapman, M. J. Plank, A. James, and B. Basse. A nonlinear model of age and size-structured populations with applications to cell cycles. The ANZIAM Journal, 49(2):151–169, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinear-model-of-age-and-sizestructured-populations-
with-applications-to-cell-cycles/C5033A1436501A784CE6572C0BFAEE50.

Cherid:2007:DDI

[1383] A. Cherid, M. A. El-Gebeily, Donal O’Regan, and Ravi Agarwal. Deblurring and denoising of images with minimization of variation and negative norms. The ANZIAM Journal, 49(2):171–185, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/deblurring-and-denoising-of-images-with-minimization-
of-variation-and-negative-norms/C262E7D3EDEB4E57B939898B0E3ABCB2.

Li:2007:TCL

[1384] R. Li, Z. G. Feng, K. L. Teo, and G. R. Duan. Tracking control of linear switched systems. The ANZIAM Journal, 49(2):187–203, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/tracking-control-of-linear-switched-
systems/41F2DAEC28BC0499018A2E9A0852DEC1.

Verma:2007:GPS

[1385] Ram U. Verma. General projection systems and relaxed cocoercive nonlinear variational inequalities. The ANZIAM Journal, 49(2):205– 212, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/general-projection-systems-
and-relaxed-cocoercive-nonlinear-variational-inequalities/30723147887C3E1E6B5EDEAD6A2E8BFE.

Fernandezbonder:2007:BST

[1386] J. Fernandezbonder, R. Orive, and J. D. Rossi. The best Sobolev trace constant in domains with holes for critical or subcritical exponents. The ANZIAM Journal, 49(2):213–230, October 2007. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/best-sobolev-trace-constant-in-domains-with-holes-for-
critical-or-subcritical-exponents/FF3AF25067DE3B6ED051BB59D0D61F0A.

Hou:2007:ASD

[1387] Zhenting Hou, Hailing Dong, and Peng Shi. Asymptotic stability in the distribution of nonlinear stochastic systems with semi-Markovian switching. The ANZIAM Journal, 49(2):231–241, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-stability-in-the-distribution-of-nonlinear-
stochastic-systems-with-semimarkovian-switching/08D45F29F4049A21973B6B8141932D6B.

Chen:2007:SNR

[1388] Kung Yu Chen, Shouh Jung Liu, and H. M. Srivastava. Some new results for the Lagrange polynomials in several variables. The ANZIAM Journal, 49(2):243–258, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/some-new-results-for-the-
lagrange-polynomials-in-several-variables/5C691BDC54D0AB12076E95B5239DAEC5.

Aminis:2007:NPF

[1389] Keyvan Aminis and Arash Haseli. A new proximity function generating the best known iteration bounds for both large-update and small-update interior-point methods. The ANZIAM Journal, 49(2):259–270, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/new-proximity-function-generating-the-best-
known-iteration-bounds-for-both-largeupdate-and-smallupdate-
interiorpoint-methods/D3187878192C6227AA23905A39912F00.

Yang:2007:ICP

[1390] Runsheng Yang and Yunhua Ou. Inverse coefficient problems for nonlinear elliptic equations. The ANZIAM Journal, 49(2):271–279, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/inverse-coefficient-problems-for-nonlinear-
elliptic-equations/17A9C95C69646BD674C8B98661791DF8.

Wu:2007:MAT

[1391] Shiliang Wu and Tingzhu Huang. A modified AOR-type iterative method for L-matrix linear systems. The ANZIAM Journal, 49(2): 281–292, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/modified-aortype-iterative-
method-for-lmatrix-linear-systems/2699130FA4B1A48493413A1D58DAC796.

Moriya:2007:UCT

[1392] K. Moriya and T. Nodera. Usage of the convergence test of the residual norm in the Tsuno–Nodera version of the GMRES algorithm. The ANZIAM Journal, 49(2):293–308, October 2007. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/usage-of-the-convergence-test-of-the-residual-norm-
in-the-tsunonodera-version-of-the-gmres-algorithm/19D1431EBE5864039D2E41EBAACD105B.

Anonymous:2007:AVIg

[1393] Anonymous. ANZ volume 49 issue 2 cover and front matter. The ANZIAM Journal, 49(2):f1–f4, October 2007. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-49-issue-2-cover-and-front-matter/10E54F7A8C415D4753BF3FE3A81DF360.

Anonymous:2007:AVIh

[1394] Anonymous. ANZ volume 49 issue 2 cover and back matter. The ANZIAM Journal, 49(2):b1–b3, October 2007. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-49-issue-2-cover-and-back-matter/F23EF017AC4BCB9E21C500EAC7539359.

Antczak:2008:AMN

[1395] Tadeusz Antczak. An η-approximation method for nonsmooth multiobjective programming problems. The ANZIAM Journal, 49(3):309– 323, January 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-approximation-method-for-nonsmooth-
multiobjective-programming-problems/1072E11A3FE330964A6DF323F8E1A41B.

Lim:2008:SDM

[1396] Leng Leng Lim, Winston L. Sweatman, and Robert McKibbin. A simple deterministic model for volcanic ashfall deposition. The ANZIAM Journal, 49(3):325–336, January 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/simple-deterministic-model-
for-volcanic-ashfall-deposition/ABDF76D0379601A963C10072F70C627D.

Li:2008:TNS

[1397] Jianli Li and Jianhua Shen. Three nonnegative solutions for secondorder impulsive differential equations with a three-point boundary value problem. The ANZIAM Journal, 49(3):337–346, January 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/three-nonnegative-solutions-for-secondorder-impulsive-
differential-equations-with-a-threepoint-boundary-value-problem/
7A3337F8E76BBD3C2875C6BADD5F1E0B.

Mcdonald:2008:VEF

[1398] N. Robb Mcdonald. Vortex equilibria in flow past a plate. The ANZIAM Journal, 49(3):347–359, January 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/vortex-equilibria-
in-flow-past-a-plate/8E2EE27666F756FBBD3C554B72BA4865.

Yueh:2008:EEI

[1399] Wen-Chyuan Yueh and Sui Sun Cheng. Explicit eigenvalues and inverses of tridiagonal Toeplitz matrices with four perturbed corners. The ANZIAM Journal, 49(3):361–387, January 2008. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/explicit-eigenvalues-and-inverses-of-tridiagonal-toeplitz-
matrices-with-four-perturbed-corners/006A7DE595A047355C6C3178AA8E0CBA.

Saka:2008:QBS

[1400] Bülent Saka, Idris Dag, and Dursun Irk. Quintic B-spline collocation method for numerical solution of the RLW equation. The ANZIAM Journal, 49(3):389–410, January 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/quintic-
bspline-collocation-method-for-numerical-solution-of-the-rlw-
equation/BD9AD43B8738521BC23A991DCAFACF3F.

Rafiee:2008:MGN

[1401] Ashkan Rafiee. Modelling of generalized Newtonian lid-driven cavity flow using an SPH method. The ANZIAM Journal, 49(3):411–422, January 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/modelling-of-generalized-newtonian-liddriven-cavity-
flow-using-an-sph-method/94013F3143911A7A1EEC04CE1B872EFC.

Liu:2008:SIO

[1402] Zheng Liu. A sharp L_{2} inequality of Ostrowski type. The ANZIAM Journal, 49(3):423–429, January 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/sharp-l2-inequality-of-
ostrowski-type/75F43518FD1890606D704F6BFAFD696A.

inYun:2008:STS

[1403] Beong in Yun. Sigmoidal-type series expansion. The ANZIAM Journal, 49(3):431–450, January 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/sigmoidaltype-series-expansion/
D25FCD4C4C1DA503D3AAD828B4E9E4DE.

Wu:2008:PTF

[1404] Chi-Ye Wu and Ting-Zhu Huang. Perturbation theory for the LU and QR factorizations. The ANZIAM Journal, 49(4):451–461, April 2008.

CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/perturbation-theory-for-the-lu-and-qr-factorizations/
C321CB2CCD02EBB387F8B50560F7413B.

Liu:2008:CSE

[1405] Yuanyuan Liu, Hanjun Zhang, and Yiqiang Zhao. Computable strongly ergodic rates of convergence for continuous-time Markov chains. The ANZIAM Journal, 49(4):463–478, April 2008. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/computable-strongly-ergodic-rates-of-convergence-for-
continuoustime-markov-chains/C54B26C5BD7AFDC3D7CFA3B0E75F3D6F.

Barkatou:2008:SRS

[1406] Mohammed Barkatou and Samira Khatmi. Symmetry result for some overdetermined value problems. The ANZIAM Journal, 49(4):479–494, April 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/symmetry-result-for-some-overdetermined-value-
problems/66324D82A117D8278CE6F0EDE64A6718.

Lacayo:2008:EBP

[1407] Ramon Lacayo. Expansion in bell polynomials of the distribution of the total claim amount with weibull-distributed claim sizes. The ANZIAM Journal, 49(4):495–501, April 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/expansion-in-bell-
polynomials-of-the-distribution-of-the-total-claim-amount-with-
weibulldistributed-claim-sizes/C0585988F5414DEEDC9037EC25522EB4.

Barton:2008:EHE

[1408] N. G. Barton. An evaporation heat engine and condensation heat pump. The ANZIAM Journal, 49(4):503–524, April 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
evaporation-heat-engine-and-condensation-heat-pump/737FC28588CC1B89BABE725D52241B20.

Frajzadeh:2008:GMQ

[1409] Ali P. Frajzadeh and Muhammad Aslam Noor. Generalized mixed quasicomplementarity problems in topological vector spaces. The ANZIAM Journal, 49(4):525–531, April 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/generalized-mixed-
quasicomplementarity-problems-in-topological-vector-spaces/
37DEBDEEE959BDAA05B61403E05883B9.

Ko:2008:FCL

[1410] Mi-Hwa Ko, Hyun-Chull Kim, and Tae-Sung Kim. On functional central limit theorems for linear random fields with dependent innovations. The ANZIAM Journal, 49(4):533–541, April 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
functional-central-limit-theorems-for-linear-random-fields-
with-dependent-innovations/55F7749145EB0A2D7530B4A2A0D7D1A9.

Tari:2008:CMS

[1411] A. Tari and S. Shahmorad. A computational method for solving two-dimensional linear Fredholm integral equations of the second kind. The ANZIAM Journal, 49(4):543–549, April 2008. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/computational-method-for-solving-twodimensional-linear-
fredholm-integral-equations-of-the-second-kind/DE36FF437743C20BDA74E80FA86F4447.

Sun:2008:SIP

[1412] Bo Sun, Xiangkui Zhao, and Weigao Ge. Successive iteration and positive solutions for a p-Laplacian multipoint boundary value problem. The ANZIAM Journal, 49(4):551–560, April 2008. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/successive-iteration-and-positive-solutions-for-a-plaplacian-
multipoint-boundary-value-problem/808D266E350B93FAD4A47183644589CE.

Leu:2008:NHH

[1413] Ming-Guang Leu. A note on the Hu–Hwang–Wang conjecture for group testing. The ANZIAM Journal, 49(4):561–571, April 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-the-huhwangwang-conjecture-for-group-testing/
1A57EDF93A3B819AA210B0AD7D1C8CB5.

Shi:2008:MCD

[1414] Y. J. Shi and G. F. Ma. Missile control design based on the linear multiple sliding mode recursive method. The ANZIAM Journal, 49(4):573–587, April 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/missile-control-design-based-on-the-linear-
multiple-sliding-mode-recursive-method/2759F340D5E578B44C1347BCF4B90AD0.

Shu:2008:SLH

[1415] Shichang Shu. Space-like hypersurfaces in locally symmetric Lorentz space. The ANZIAM Journal, 50(1):1–11, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/spacelike-
hypersurfaces-in-locally-symmetric-lorentz-space/56C4BA1034BB4FF3587935928D26F8A0.

Zhou:2008:GEB

[1416] Jun Zhou and Chunlai Mu. Global existence and blow-up for a non-Newton polytropic filtration system with nonlocal source. The ANZIAM Journal, 50(1):13–29, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/global-existence-and-blowup-
for-a-nonnewton-polytropic-filtration-system-with-nonlocal-
source/02E4BBA956E4226D95618ADD0A8CBC1F.

Liu:2008:SDN

[1417] Zaiming Liu and Jun Peng. Stability in distribution of nonlinear systems with time-varying delays and semi-Markovian switching. The ANZIAM Journal, 50(1):31–44, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/stability-in-distribution-
of-nonlinear-systems-with-timevarying-delays-and-semimarkovian-
switching/2E154771F7367C6F9BB0874AA40C2571.

Chen:2008:FNS

[1418] J. Chen, F. Liu, I. Turner, and V. Anh. The fundamental and numerical solutions of the Riesz space-fractional reaction–dispersion equation. The ANZIAM Journal, 50(1):45–57, July 2008. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fundamental-and-numerical-solutions-of-the-riesz-spacefractional-
reactiondispersion-equation/B0EDAF431DF87BC977B5358B303E7157.

Dosiyev:2008:OME

[1419] A. A. Dosiyev and S. Cival Buranay. On the order of maximum error of the finite difference solutions of Laplace’s equation on rectangles. The ANZIAM Journal, 50(1):59–73, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-order-of-maximum-error-of-the-finite-difference-solutions-
of-laplaces-equation-on-rectangles/9B7645475CA3431270F94D0792907B71.

See correction [1469].

Eliasi:2008:HWI

[1420] Mehdi Eliasi and Bijn Taeri. Hyper-Wiener index of zigzag polyhex nanotubes. The ANZIAM Journal, 50(1):75–86, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/hyperwiener-index-of-zigzag-polyhex-nanotubes/509B3975BF1FB470B744D1A40125696E.

Chu:2008:PRR

[1421] E. K.-W. Chu, W.-W. Lin, and C.-S. Wang. Perturbation results related to palindromic eigenvalue problems. The ANZIAM Journal, 50(1): 87–100, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/perturbation-results-related-to-palindromic-
eigenvalue-problems/24FF8278A28FFF33B292751A7B2AB8AF.

Hocking:2008:NWT

[1422] G. C. Hocking and H. Zhang. A note on withdrawal from a two-layer fluid through a line sink in a porous medium. The ANZIAM Journal, 50(1):101–110, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
withdrawal-from-a-twolayer-fluid-through-a-line-sink-in-a-porous-
medium/409D4F073916FEC9D033D2C83B8CFD3C.

Cho:2008:NGO

[1423] Yeol Je Cho, Young-Ho Kim, and Josip Pecarić. New Gronwall– Ou-Iang type integral inequalities and their applications. The ANZIAM Journal, 50(1):111–127, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/new-
gronwallouiang-type-integral-inequalities-and-their-applications/
D5FF60662859A9403AF4D0BDAC184CCE.

Liu:2008:ASI

[1424] Zheng Liu. Another sharp L_{2} inequality of Ostrowski type. The ANZIAM Journal, 50(1):129–136, July 2008. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/another-
sharp-l2-inequality-of-ostrowski-type/962ABB8CD47B1F915F663DB49050F959.

Mortici:2008:AMV

[1425] Cristinel Mortici. Arithmetic mean of values and value at mean of arguments for convex functions. The ANZIAM Journal, 50(1):137– 141, July 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/arithmetic-mean-of-values-and-value-
at-mean-of-arguments-for-convex-functions/37346E6EC2BA222DD856D551D36B8674.

Szekeres:2008:STS

[1426] George Szekeres and Lindsay Peters. Space–time structure and spinor geometry. The ANZIAM Journal, 50(2):143–176, October 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/spacetime-structure-and-spinor-geometry/1AE7F3EED7DEAC3D9B74757047455D89.

Kenny:2008:ADF

[1427] Brian G. Kenny and Tony W. Dixon. Ambiguity in the determination of the free energy associated with the critical circle map. The ANZIAM Journal, 50(2):177–184, October 2008. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/ambiguity-in-the-determination-of-the-free-energy-associated-
with-the-critical-circle-map/386973B3CCF7B4AAF29742F970331989.

Sabo:2008:BLA

[1428] Kristian Sabo and Rudolf Scitovski. The best least absolute deviations line — properties and two efficient methods for its derivation. The ANZIAM Journal, 50(2):185–198, October 2008. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/best-least-absolute-deviations-line-properties-and-
two-efficient-methods-for-its-derivation/5B233AD13F34B93FEFB0A454766514C6.

Wilson:2008:CCD

[1429] Miles Wilson, John R. Blake, and Peter M. Haese. Cloud cavitation dynamics. The ANZIAM Journal, 50(2):199–208, October 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/cloud-cavitation-dynamics/0B09926D05E06A69EE4BEE0863437388.

Zhu:2008:DDR

[1430] Shuqian Zhu, Chenghui Zhang, Xinzhi Liu, and Zhenbo Li. Delaydependent robust H_{∞} control for singular systems with multiple delays. The ANZIAM Journal, 50(2):209–230, October 2008. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/delaydependent-robust-h-control-for-singular-systems-
with-multiple-delays/DE5EA88DAE23035A1FFA2A8569E580B2.

Singh:2008:TMR

[1431] Jitender Singh and Renu Bajaj. Temperature modulation in Rayleigh– Bénard convection. The ANZIAM Journal, 50(2):231–245, October 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/temperature-modulation-in-rayleighbenard-convection/
3055C9C4DB3B7EFB503D9444D2C23E5F.

Gavrea:2008:GUB

[1432] B. Gavrea, J. Jaksetić, and J. Pecarić. On a global upper bound for Jessen’s inequality. The ANZIAM Journal, 50(2):246–257, October 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-a-global-upper-bound-for-jessens-
inequality/D0E84FA11718222320D1AD3487979942.

Farajzadeh:2008:PSV

[1433] A. P. Farajzadeh. On pseudomonotone set-valued mappings in topological vector spaces. The ANZIAM Journal, 50(2):258–265, October 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-pseudomonotone-setvalued-mappings-in-topological-
vector-spaces/506BB3F66FB8E1B7CE9995F3690CBA7E.

Mackenzie:2008:SHS

[1434] Scott Mackenzie and Graham Mills. Scheduling a heterogeneous set of trains over a single line track using Lagrangian relaxation. The ANZIAM Journal, 50(2):266–281, October 2008. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/scheduling-a-heterogeneous-
set-of-trains-over-a-single-line-track-using-lagrangian-relaxation/
CAA9B0F0ADB9060809549C286AED4A35.

Burnell:2009:PSI

[1435] John Burnell and Graham Weir. Preface to this special issue. The ANZIAM Journal, pages i–iv, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/preface-
to-this-special-issue/1F67AA0B2578B6BE8FC66E30542BC356.

Luo:2009:NDC

[1436] W. Luo, G. C. Wake, and C. W. Hawk. Numerical determination of critical conditions for thermal ignition. The ANZIAM Journal, pages 283– 305, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/numerical-determination-of-critical-
conditions-for-thermal-ignition/2FFCA31CF16A1F7179454F5C37B39D5B.

Mcguinness:2009:MSI

[1437] Mark J. Mcguinness. Modelling sea ice growth. The ANZIAM Journal, pages 306–319, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/modelling-sea-ice-
growth/D2BCBA472CB037A3E0CB0B6A924EDD66.

ONeale:2009:RTI

[1438] Dion R. J. O’Neale and Robert I. Mclachlan. Reconsidering trigonometric integrators. The ANZIAM Journal, pages 320–332, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reconsidering-trigonometric-integrators/C59AAB92584336C2E272285A31E9EFDC.

Butcher:2009:PRK

[1439] J. C. Butcher. Practical Runge–Kutta methods for scientific computation. The ANZIAM Journal, pages 333–342, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/practical-rungekutta-methods-for-scientific-computation/
F356AA1AE59BED1D14B298ED18B851D7.

Weir:2009:MML

[1440] Graham Weir. A mathematical model for the large-scale transport of heat and water in the Taupo volcanic zone of New Zealand. The ANZIAM Journal, pages 343–354, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/mathematical-model-
for-the-largescale-transport-of-heat-and-water-in-the-taupo-
volcanic-zone-of-new-zealand/08A43B6D5FDB7ACABDCF20098305A03F.

Burnell:2009:MIT

[1441] John Burnell. A model of ion transport in conjugated polymers. The ANZIAM Journal, pages 355–364, January 2009. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/model-of-ion-transport-in-conjugated-polymers/61CF1114B94C7313FF277C9EFF543CBD.

McKibbin:2009:CPM

[1442] Robert McKibbin, Thomasin A. Smith, and Luke Fullard. Components and phases: modelling progressive hydrothermal eruptions. The ANZIAM Journal, pages 365–380, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/components-
and-phases-modelling-progressive-hydrothermal-eruptions/DC8DF49EC807BDF2D14AE1315E1C39DE.

Lund:2009:ESL

[1443] Nat J. Lund and Shaun C. Hendy. Effective slip length of nanoscale mixed-slip surfaces. The ANZIAM Journal, pages 381–394, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/effective-slip-length-of-nanoscale-mixedslip-surfaces/
8971BA20B0B0C622AB17DF67F9CDE9CD.

Mackay:2009:MGZ

[1444] Jade R. Mackay, Stephen P. White, and Shaun C. Hendy. Modelling the growth of zinc oxide nanostructures. The ANZIAM Journal, pages 395– 406, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/modelling-the-growth-of-zinc-oxide-nanostructures/
5663674A454BBFF91E132A94935C621C.

Young:2009:ASO

[1445] Roger Young. An analytic solution for one-dimensional dissipational strain-gradient plasticity. The ANZIAM Journal, pages 407–420, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/an-analytic-solution-for-onedimensional-dissipational-
straingradient-plasticity/09BC2C1448FA650011CE870E078534F0.

McNabb:2009:CMO

[1446] Alex McNabb. A conceptual model for the origins of geothermal and volcanic activity in the North Island of New Zealand. The ANZIAM Journal, pages 421–425, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/conceptual-model-for-the-
origins-of-geothermal-and-volcanic-activity-in-the-north-island-
of-new-zealand/BE30028A5F325B6FB143A0C371C710E2.

Grant:2009:MMW

[1447] Malcolm A. Grant. Mathematical modelling of Wairakei geothermal field. The ANZIAM Journal, pages 426–434, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/mathematical-modelling-of-wairakei-geothermal-field/
1CAE6640D2A66482796D3477E35B21E1.

Withers:2009:OFZ

[1448] C. S. Withers. Orthogonal functions and Zernike polynomials — a random variable interpretation. The ANZIAM Journal, pages 435–444, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/orthogonal-functions-and-zernike-polynomialsa-
random-variable-interpretation/7FAA2E658743471945E88E5C40E8B32E.

Anonymous:2009:AVIa

[1449] Anonymous. ANZ volume 50 issue 3 cover and front matter. The ANZIAM Journal, pages f1–f2, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-50-issue-3-cover-and-front-matter/7BD7761C76157482445AB18A183F5A9E.

Anonymous:2009:AVIb

[1450] Anonymous. ANZ volume 50 issue 3 cover and back matter. The ANZIAM Journal, pages b1–b2, January 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-50-issue-3-cover-and-back-matter/CE8358FC41A3326255F864A832F2BA91.

De-la-Pena:2009:IED

[1451] Victor De la Peña, Gerardo Hernández-Del-Valle, and Carlos G. Pacheco-González. An integral equation for the distribution of the first exit time of a reflected Brownian motion. The ANZIAM Journal, 50(4):445–454, April 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/an-integral-equation-for-the-distribution-of-
the-first-exit-time-of-a-reflected-brownian-motion/AB56C5BBA4B1241D77E6E0EFE8F4D531.

Wan:2009:CMD

[1452] Zhong Wan, K. L. Teo, LingShuang Kong, and Chunhua Yang. A class of mix design problems: formulation, solution methods and applications. The ANZIAM Journal, 50(4):455–474, April 2009. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/class-of-mix-design-problems-formulation-solution-methods-
and-applications/2E99FCCE9612F1A5562914D8641774FF.

Skipper:2009:VFM

[1453] Max Skipper and Peter Buchen. A valuation formula for multi-asset, multi-period binaries in a Black–Scholes economy. The ANZIAM Journal, 50(4):475–485, April 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/valuation-formula-for-
multiasset-multiperiod-binaries-in-a-blackscholes-economy/7B83C60DA8B1CC85725F56AE603669DE.

Ren:2009:RBS

[1454] Yong Ren and Xiliang Fan. Reflected backward stochastic differential equations driven by a Lévy process. The ANZIAM Journal, 50(4):486– 500, April 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/reflected-backward-stochastic-differential-
equations-driven-by-a-levy-process/02845AF2E151D9E91690F4C95FD7B5E6.

Zheng:2009:ABS

[1455] Wenjun Zheng and Zhiqin Zhao. Analysis of block-sor iteration for the three-dimensional Laplacian. The ANZIAM Journal, 50(4):501–512, April 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/analysis-of-blocksor-iteration-for-the-threedimensional-
laplacian/6C60B4DC015A5E824142540ADBE9575B.

Liu:2009:FRP

[1456] Yuji Liu. Further results on positive periodic solutions of impulsive functional differential equations and applications. The ANZIAM Journal, 50(4):513–533, April 2009. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/further-
results-on-positive-periodic-solutions-of-impulsive-functional-
differential-equations-and-applications/7BEA2D2490E29760FA71DF572591AA9F.

Brearley:2009:MIO

[1457] Maurice N. Brearley. A method of improving oar efficiency. The ANZIAM Journal, 50(4):534–540, April 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/method-of-improving-oar-
efficiency/EE6500BD9ED57F9376FC7828318BB1BB.

VanGorder:2009:UBL

[1458] Robert A. Van Gorder and K. Vajravelu. Unsteady boundary layers: convective heat transfer over a vertical flat plate. The ANZIAM Journal, 50(4):541–549, April 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/unsteady-boundary-
layers-convective-heat-transfer-over-a-vertical-flat-plate/
16389F9D3A88DFCBD6BB8D5386757E61.

Kerman:2009:EED

[1459] R. Kerman, M. L. Huang, and M. Brannan. Error estimates for Dominici’s Hermite function asymptotic formula and some applications. The ANZIAM Journal, 50(4):550–561, April 2009. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/error-estimates-for-dominicis-hermite-function-asymptotic-
formula-and-some-applications/A77F7793E454440C1BD7EC2ADCCADFF2.

Anonymous:2009:AVIc

[1460] Anonymous. ANZ volume 50 issue 4 cover and front matter. The ANZIAM Journal, 50(4):f1–f2, April 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-50-issue-4-cover-and-front-matter/E2A6CF9DAD4539A26169939E0FEB944F.

Anonymous:2009:AVId

[1461] Anonymous. ANZ volume 50 issue 4 cover and back matter. The ANZIAM Journal, 50(4):b1–b9, April 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-50-issue-4-cover-and-back-matter/1052E0F4826AAE73C3E314BF3E0C96EF.

Moriya:2009:AMI

[1462] Kentaro Moriya, Linjie Zhang, and Takashi Nodera. An approximate matrix inversion procedure by parallelization of the Sherman– Morrison formula. The ANZIAM Journal, 51(1):1–9, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
approximate-matrix-inversion-procedure-by-parallelization-of-
the-shermanmorrison-formula/44A9FCF98BBA392C0435499D620121C5.

Liao:2009:SSL

[1463] Shijun Liao. Series solution of large deformation of a beam with arbitrary variable cross section under an axial load. The ANZIAM Journal, 51(1): 10–33, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/series-solution-of-large-deformation-
of-a-beam-with-arbitrary-variable-cross-section-under-an-axial-
load/5FCC386F289A848821214193FAED1F68.

Qian:2009:RPU

[1464] Yiping Qian and Xiang Lin. Ruin probabilities under an optimal investment and proportional reinsurance policy in a jump diffusion risk process. The ANZIAM Journal, 51(1):34–48, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/ruin-probabilities-under-an-optimal-investment-and-
proportional-reinsurance-policy-in-a-jump-diffusion-risk-process/
09A728026258BF80491C975ECCD81F6B.

Yu:2009:TWS

[1465] Zhi-Xian Yu and Rong Yuan. Travelling wave solutions in nonlocal reaction–diffusion systems with delays and applications. The ANZIAM Journal, 51(1):49–66, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/travelling-wave-solutions-
in-nonlocal-reactiondiffusion-systems-with-delays-and-applications/
9710EC287535831D721B9CB16E3D8D40.

Aljinovic:2009:SII

[1466] A. Aglić Aljinović, J. Pecarić, and M. Ribicić Penava. Sharp integral inequalities based on general two-point formulae via an extension of Montgomery’s identity. The ANZIAM Journal, 51(1):67–101, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/sharp-integral-inequalities-based-on-general-twopoint-
formulae-via-an-extension-of-montgomerys-identity/548A59E9A0772F5842B80D1ED0BD62B7.

Whiten:2009:SAD

[1467] Bill Whiten. A simple algorithm for deduction. The ANZIAM Journal, 51(1):102–122, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/simple-algorithm-for-deduction/
84120CF18361995C5029E2FB2134BF06.

Li:2009:MDD

[1468] Minghao Li, Wuneng Zhou, Ziwei Ni, and Mingjun Wang. Mixed delaydependent stability of high-order neural networks based on a weak coupling LMI set. The ANZIAM Journal, 51(1):123–140, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/mixed-delaydependent-stability-of-highorder-neural-
networks-based-on-a-weak-coupling-lmi-set/6932AA4823759B720D8980E2ACD6DD5F.

Dosiyev:2009:COM

[1469] A. A. Dosiyev and S. Cival Buranay. Correction to ‘On the order of maximum error of the finite difference solutions to Laplace’s equation on rectangles’. The ANZIAM Journal, 51(1):141, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/correction-to-on-the-order-of-maximum-error-of-the-
finite-difference-solutions-to-laplaces-equation-on-rectangles/
EE0EAD2CDC9F70545ED3589F149D7229. See [1419].

Anonymous:2009:AVIe

[1470] Anonymous. ANZ volume 51 issue 1 cover and front matter. The ANZIAM Journal, 51(1):f1–f2, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-51-issue-1-cover-and-front-matter/52C10D41236C7F94FB0622161301683B.

Anonymous:2009:AVIf

[1471] Anonymous. ANZ volume 51 issue 1 cover and back matter. The ANZIAM Journal, 51(1):b1–b5, July 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-51-issue-1-cover-and-back-matter/707C235866CEFCAE6884ECED0149916C.

Chen:2009:OEP

[1472] Wen-Ting Chen and Song-Ping Zhu. Optimal exercise price of American options near expiry. The ANZIAM Journal, 51(2):145–161, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/optimal-exercise-price-of-american-options-
near-expiry/E5CD0BA42E6F7A30FDDDB6E46E432732.

Li:2009:ECA

[1473] B. Li, K. L. Teo, G. H. Zhao, and G. R. Duan. An efficient computational approach to a class of minmax optimal control problems with applications. The ANZIAM Journal, 51(2):162–177, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-efficient-computational-approach-
to-a-class-of-minmax-optimal-control-problems-with-applications/
A5465DDBABA4E59E6A6CB38882F0E2CB.

Leach:2009:LTS

[1474] J. A. Leach and Andrew P. Bassom. The large-time solution of a nonlinear fourth-order equation initial-value problem I. Initial data with a discontinuous expansive step. The ANZIAM Journal, 51 (2):178–190, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/largetime-solution-of-
a-nonlinear-fourthorder-equation-initialvalue-problem-i-initial-
data-with-a-discontinuous-expansive-step/A6F05C2771A1864B6870F7796EF21767.

Mahale:2009:ILR

[1475] P. Mahale and M. T. Nair. Iterated Lavrentiev regularization for nonlinear ill-posed problems. The ANZIAM Journal, 51(2):191–217, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/iterated-lavrentiev-regularization-for-nonlinear-illposed-
problems/434D17E9650CB36E1C55BFEC9B70D487.

Han:2009:OFD

[1476] Chunyan Han and Huanshui Zhang. Optimal filtering in discretetime systems with time delays and Markovian jump parameters. The ANZIAM Journal, 51(2):218–233, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-

filtering-in-discretetime-systems-with-time-delays-and-markovian-
jump-parameters/2D0168E1CFA401891572A9D9CB865748.

Gani:2009:DSM

[1477] J. Gani and R. J. Swift. Deterministic and stochastic models for the spread of cholera. The ANZIAM Journal, 51(2):234–240, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/deterministic-and-stochastic-models-for-the-spread-
of-cholera/A27C0B0ACA7B6F82BFB3DB6F5BD5577C.

Manam:2009:SMC

[1478] S. R. Manam. Scattering of membrane coupled gravity waves by partial vertical barriers. The ANZIAM Journal, 51(2):241–260, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/scattering-of-membrane-coupled-gravity-waves-by-partial-
vertical-barriers/CEE350067D9BFA53E1BE4889BAF00EA1.

Li:2009:TFO

[1479] R. Li and Y. J. Shi. A time-fuel optimal control problem of a cruise missile based on an improved sliding mode variable structure model. The ANZIAM Journal, 51(2):261–276, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/timefuel-optimal-control-problem-of-a-cruise-missile-
based-on-an-improved-sliding-mode-variable-structure-model/
6B05C53A9286FDACEA99F0981C9475E4.

Srinivasacharya:2009:MFM

[1480] D. Srinivasacharya and Mekonnen Shiferaw. Magnetohydrodynamic flow of a micropolar fluid in a circular pipe with hall effects. The ANZIAM Journal, 51(2):277–285, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/magnetohydrodynamic-
flow-of-a-micropolar-fluid-in-a-circular-pipe-with-hall-effects/
94F5A5D6B63555823176D2A5A24A19AE.

Salahi:2009:SRN

[1481] M. Salahi. A self-regular Newton based algorithm for linear optimization. The ANZIAM Journal, 51(2):286–301, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/selfregular-newton-based-algorithm-for-linear-optimization/
358B71D19872A10FC271A57CE82A2B44.

Ahmadabadi:2009:ADM

[1482] M. Nili Ahmadabadi and F. M. Maalek Ghaini. An Adomian decomposition method for solving Liénard equations in general form. The ANZIAM Journal, 51(2):302–308, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/an-{
Adomian}-decomposition-method-for-solving-lienard-equations-
in-general-form/5479B2913F2D35D1AF24B900AA1493F0.

Anonymous:2009:AVIg

[1483] Anonymous. ANZ volume 51 issue 2 cover and front matter. The ANZIAM Journal, 51(2):f1–f2, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-51-issue-2-cover-and-front-matter/5E92E432648B204EC877C510E7946940.

Anonymous:2009:AVIh

[1484] Anonymous. ANZ volume 51 issue 2 cover and back matter. The ANZIAM Journal, 51(2):b1–b6, October 2009. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-51-issue-2-cover-and-back-matter/50588BF8D69DB9632A2974054A6C2700.

Bassom:2010:Ea

[1485] Andrew Bassom and Graeme Hocking. Editorial. The ANZIAM Journal, 51(3):i–ii, January 2010. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/editorial/48FD28AA9369C00B3CF3B4245913A234.

Baxter:2010:PDF

[1486] R. J. Baxter. Proof of the determinantal form of the spontaneous magnetization of the superintegrable chiral Potts model. The ANZIAM Journal, 51(3):309–316, January 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/proof-of-the-determinantal-
form-of-the-spontaneous-magnetization-of-the-superintegrable-
chiral-potts-model/08E23227AB3BAC2CC4AAA05BCA53F680.

Ozugurlu:2010:NNA

[1487] E. Özugurlu. A note on the numerical approach for the reaction– diffusion problem with a free boundary condition. The ANZIAM Journal, 51(3):317–330, January 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/note-on-the-numerical-
approach-for-the-reactiondiffusion-problem-with-a-free-boundary-
condition/5FDEF9B56E6057F7BC61BF64CF2B848E.

Gao:2010:OCP

[1488] Rui Gao and Xinzhi Liu. Optimal control problems for general global hybrid dynamical systems with matrix cost functional. The ANZIAM Journal, 51(3):331–349, January 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
control-problems-for-general-global-hybrid-dynamical-systems-
with-matrix-cost-functional/F491EF2A32C3112656E3A634D2DD1A11.

Zhang:2010:NPS

[1489] Lie-Hui Zhang and Yong Wang. A note on periodic solutions of a forced Liénard-type equation. The ANZIAM Journal, 51(3):350–368, January 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-periodic-solutions-of-a-forced-lienardtype-
equation/6081C0351EE20A37A12E06CE0E768C50.

Li:2010:AIU

[1490] Jian-Lei Li, Ting-Zhu Huang, and Liang Li. Analysis of the inexact Uzawa algorithms for nonlinear saddle-point problems. The ANZIAM Journal, 51(3):369–382, January 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/analysis-of-the-inexact-
uzawa-algorithms-for-nonlinear-saddlepoint-problems/9B31CE11E63BA41D95BF90FEF405408B.

Anonymous:2010:AVIa

[1491] Anonymous. ANZ volume 51 issue 3 cover and front matter. The ANZIAM Journal, 51(3):f1–f2, January 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-51-issue-3-cover-and-front-matter/D17961B035B433FFD492D4E09E945C6B.

Anonymous:2010:AVIb

[1492] Anonymous. ANZ volume 51 issue 3 cover and back matter. The ANZIAM Journal, 51(3):b1–b6, January 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-51-issue-3-cover-and-back-matter/EEA079BE872C4EECD6C1C61A37DBBCE8.

VanBrunt:2010:EPI

[1493] Bruce Van Brunt and M. Vlieg-Hulstman. An eigenvalue problem involving a functional differential equation arising in a cell growth model. The ANZIAM Journal, 51(4):383–393, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
eigenvalue-problem-involving-a-functional-differential-equation-
arising-in-a-cell-growth-model/D5D0F7803169374D82DDC6604C88AAB4.

Agusto:2010:AIO

[1494] F. B. Agusto and O. R. Ogunye. Avian influenza optimal seasonal vaccination strategy. The ANZIAM Journal, 51(4):394–405, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/avian-influenza-optimal-seasonal-vaccination-strategy/
4670B1D5B6F9BBB67EF40E2706710B35.

Sellier:2010:NAB

[1495] M. Sellier and R. D. Lenz. A note on approximate benchmark solutions for viscous two-layer flows. The ANZIAM Journal, 51(4):406–415, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-approximate-benchmark-solutions-for-viscous-
twolayer-flows/9AA26D400EC22E875A3CF15990A71181.

Akhtar:2010:ESP

[1496] W. Akhtar, Corina Fetecau, and A. U. Awan. Exact solutions for the Poiseuille flow of a generalized Maxwell fluid induced by timedependent shear stress. The ANZIAM Journal, 51(4):416–429, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/exact-solutions-for-the-poiseuille-flow-of-
a-generalized-maxwell-fluid-induced-by-timedependent-shear-
stress/029416956D3242B1BE3999CAC17C1A56.

Lauko:2010:CNA

[1497] M. Lauko and D. Sevcovic. Comparison of numerical and analytical approximations of the early exercise boundary of American put options. The ANZIAM Journal, 51(4):430–448, April 2010. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/comparison-of-numerical-and-analytical-approximations-
of-the-early-exercise-boundary-of-american-put-options/E1393501C1DAAE2CDC345E06D8F77958.

Liang:2010:OPR

[1498] Zhibin Liang and Junyi Guo. Optimal proportional reinsurance under two criteria: maximizing the expected utility and minimizing the value at risk. The ANZIAM Journal, 51(4):449–463, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/optimal-proportional-reinsurance-under-two-
criteria-maximizing-the-expected-utility-and-minimizing-the-
value-at-risk/FA3B3AAF1651EEE6016FFD21787420C0.

Sweilam:2010:CPS

[1499] N. H. Sweilam and M. M. Khader. A Chebyshev pseudo-spectral method for solving fractional-order integro-differential equations. The ANZIAM Journal, 51(4):464–475, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/chebyshev-pseudospectral-
method-for-solving-fractionalorder-integrodifferential-equations/
D54540C52DE837C74F79EDB24A32FE71.

Cho:2010:PDI

[1500] G. M. Cho, Y. Y. Cho, and Y. H. Lee. A primal-dual interior-point algorithm based on a new kernel function. The ANZIAM Journal, 51(4): 476–491, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/primaldual-interiorpoint-algorithm-based-
on-a-new-kernel-function/AF6E4BB96C5826190925CEEE52799078.

Anonymous:2010:AVIc

[1501] Anonymous. ANZ volume 51 issue 4 cover and front matter. The ANZIAM Journal, 51(4):f1–f2, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-51-issue-4-cover-and-front-matter/697C39A6B412072F5CF27FCE919A59E7.

Anonymous:2010:AVId

[1502] Anonymous. ANZ volume 51 issue 4 cover and back matter. The ANZIAM Journal, 51(4):b1–b13, April 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-51-issue-4-cover-and-back-matter/60627069FCA37FAC5072B74402EFB92D.

Bassom:2010:Eb

[1503] Andrew Bassom and Graeme Hocking. Editorial. The ANZIAM Journal, 52(1):i, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/editorial/AF991FED780D42C28DAA3C91D57D6ACF.

Grimshaw:2010:TFP

[1504] R. Grimshaw. Transcritical flow past an obstacle. The ANZIAM Journal, 52(1):2–26, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/transcritical-flow-past-an-obstacle/
C2732A6E7A83BE8CBC9D0EFFCFA0796E.

Thomas:2010:MST

[1505] Emma G. Thomas, Hannah E. Barrington, Kamalini M. Lokuge, and Geoffry N. Mercer. Modelling the spread of tuberculosis, including drug resistance and HIV: a case study in Papua New Guinea’s Western Province. The ANZIAM Journal, 52(1):26–45, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/modelling-the-spread-of-tuberculosis-including-drug-
resistance-and-hiv-a-case-study-in-papua-new-guineas-western-
province/C2BBAFB69AC59AEB25B63E95C1C471F3.

VanBrunt:2010:EAF

[1506] Bruce Van Brunt and M. Vlieg-Hulstman. Eigenfunctions arising from a first-order functional differential equation in a cell growth model. The ANZIAM Journal, 52(1):46–58, July 2010. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/eigenfunctions-arising-from-a-firstorder-functional-
differential-equation-in-a-cell-growth-model/83B0C8514D721A55CA92FAA13D513152.

Jiracheewanun:2010:CNC

[1507] S. Jiracheewanun, S. W. Armfield, and M. Behnia. Combined natural convection cooling of a drink can. The ANZIAM Journal, 52(1):

59–68, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/combined-natural-convection-cooling-
of-a-drink-can/5667F6098ADF03997A8E6E56F0577297.

Clements:2010:ACP

[1508] David L. Clements. On an antiplane crack problem for functionally graded elastic materials. The ANZIAM Journal, 52(1):69–86, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-an-antiplane-crack-problem-for-functionally-graded-
elastic-materials/009D7A42BAB6D11B805B35E477B5B45E.

Le:2010:PMS

[1509] Jiang Le, Huang Jin, Xiao-Guang Lv, and Qing-Song Cheng. A preconditioned method for the solution of the Robbins problem for the Helmholtz equation. The ANZIAM Journal, 52(1):87–100, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/preconditioned-method-for-the-solution-of-the-robbins-
problem-for-the-helmholtz-equation/0246DD99D3EAEC3CF9932FE0520D7398.

Mansour:2010:TWS

[1510] M. B. A. Mansour. Travelling wave solutions for doubly degenerate reaction–diffusion equations. The ANZIAM Journal, 52(1):101–109, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/travelling-wave-solutions-for-doubly-degenerate-reactiondiffusion-
equations/3157335003DCA174DF09C6C36E09C5CF.

Khani:2010:NES

[1511] F. Khani, M. T. Darvishi, A. Farmany, and L. Kavitha. New exact solutions of coupled (2 + 1)-dimensional nonlinear systems of Schrödinger equations. The ANZIAM Journal, 52(1):110–121, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/new-exact-solutions-of-coupled-21dimensional-nonlinear-
systems-of-schrodinger-equations/0A6184EC45C92A3A0355FCF1EE588878.

Anonymous:2010:AVIe ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-52-issue-1-cover-and-front-matter/665708DD5E51BF9295A9200738F95775.

Anonymous:2010:AVIf

[1513] Anonymous. ANZ volume 52 issue 1 cover and back matter. The ANZIAM Journal, 52(1):b1–b7, July 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-52-issue-1-cover-and-back-matter/7900E43BFCFD284735C5C686BE645DEC.

Mclean:2010:RST

[1514] William Mclean. Regularity of solutions to a time-fractional diffusion equation. The ANZIAM Journal, 52(2):123–138, October 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/regularity-of-solutions-to-a-timefractional-diffusion-
equation/4F1835286CD1F88A08377C136D5ACDDE.

Merdan:2010:SAL

[1515] Hüseyin Merdan. Stability analysis of a Lotka–Volterra type predator– prey system involving Allee effects. The ANZIAM Journal, 52 (2):139–145, October 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/stability-analysis-of-
a-lotkavolterra-type-predatorprey-system-involving-allee-effects/
B16D971CAB5D279B7955653CC785A067.

Sellier:2010:STR

[1516] M. Sellier and S. Panda. Surface temperature reconstruction based on the thermocapillary effect. The ANZIAM Journal, 52(2):146–159, October 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/surface-temperature-reconstruction-based-on-the-thermocapillary-
effect/48F542C97A7E824A72CD2CECCF8F08BF.

Matei:2010:CPN

[1517] A. Matei and R. Ciurcea. Contact problems for nonlinearly elastic materials: weak solvability involving dual Lagrange multipliers. The ANZIAM Journal, 52(2):160–178, October 2010. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/contact-problems-for-nonlinearly-elastic-materials-
weak-solvability-involving-dual-lagrange-multipliers/A86755564DCA8574556271242FC867D1.

Liu:2010:ESS

[1518] Zeng Liu, Martin Oberlack, Vladimir N. Grebenev, and Shi-Jun Liao. Explicit series solution of a closure model for the von Kármán– Howarth equation. The ANZIAM Journal, 52(2):179–202, October 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/explicit-series-solution-of-a-closure-model-for-the-
von-karmanhowarth-equation/3069F3028EEEAB2748874E06F795930F.

Guo:2010:RAS

[1519] L. N. Guo, H. B. Xu, C. Gao, and G. T. Zhu. Reliability analysis of a simple repairable system. The ANZIAM Journal, 52(2):203–217, October 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reliability-analysis-of-a-simple-repairable-system/
28D2E25FA1BA2AE88A1C34F4A137BE25.

Hussain:2010:BHD

[1520] Sabir Hussain and Josip Pecarić. Bounds for hardy differences. The ANZIAM Journal, 52(2):218–224, October 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/bounds-for-
hardy-differences/0388BB9E8797EE8E9535EA7391117776.

Anonymous:2010:AVIg

[1521] Anonymous. ANZ volume 52 issue 2 cover and front matter. The ANZIAM Journal, 52(2):f1–f2, October 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-52-issue-2-cover-and-front-matter/8E4B9E52F080C79ABD8297225F99451F.

Anonymous:2010:AVIh

[1522] Anonymous. ANZ volume 52 issue 2 cover and back matter. The ANZIAM Journal, 52(2):b1–b4, October 2010. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-52-issue-2-cover-and-back-matter/40C66C149463AAF9D2012D19E74069AD.

Champneys:2011:MHC

[1523] A. R. Champneys, D. Avitabile, M. Homer, and R. Szalai. The mechanics of hearing: a comparative case study in bio-mathematical modelling. The ANZIAM Journal, 52(3):225–249, January 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/mechanics-of-hearing-a-comparative-case-study-in-biomathematical-
modelling/4CD69C9D3DA5E5D42F7703B15B9A1B46.

Lin:2011:OIR

[1524] Xiang Lin and Peng Yang. Optimal investment and reinsurance in a jump diffusion risk model. The ANZIAM Journal, 52(3):250–262, January 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-investment-and-reinsurance-in-a-jump-diffusion-
risk-model/BAE800D1241E84CDD9E3A0465CF36CF0.

Gates:2011:SST

[1525] David J. Gates. Symmetric solutions for two-body dynamics in a collision prevention model. The ANZIAM Journal, 52(3):263–288, January 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/symmetric-solutions-for-twobody-dynamics-in-a-collision-
prevention-model/AE57CEBD84B798A519C4C95BC8CF7C58.

Srinivasacharya:2011:CFP

[1526] D. Srinivasacharya and M. Krishna Prasad. Creeping flow past a porous approximately spherical shell: stress jump boundary condition. The ANZIAM Journal, 52(3):289–300, January 2011. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/creeping-flow-past-a-porous-approximately-spherical-
shell-stress-jump-boundary-condition/5D6434D63902CBD5D87C2D45DC9B6950.

Ashraf:2011:NSM

[1527] Muhammad Ashraf and M. Anwar Kamal. Numerical simulation of MHD stagnation point flow towards a heated axisymmetric surface. The ANZIAM Journal, 52(3):301–308, January 2011. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-simulation-of-mhd-stagnation-point-flow-towards-
a-heated-axisymmetric-surface/A665D457A26F0E8110A3FC3AE8070BF2.

Franjic:2011:GTP

[1528] Iva Franjić, Josip Pecarić, and Ivan Perić. General three-point quadrature formulas of Euler type. The ANZIAM Journal, 52(3):309–317, January 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-

journal/article/general-threepoint-quadrature-formulas-of-euler-
type/E5461DF68BC82D31A8358D10F832C490.

Anonymous:2011:AVIa

[1529] Anonymous. ANZ volume 52 issue 3 cover and front matter. The ANZIAM Journal, 52(3):f1–f2, January 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-52-issue-3-cover-and-front-matter/D5C5A9BF24103C0605E5F1A5820B3C36.

Anonymous:2011:AVIb

[1530] Anonymous. ANZ volume 52 issue 3 cover and back matter. The ANZIAM Journal, 52(3):b1–b5, January 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-52-issue-3-cover-and-back-matter/AD83DE31D98BE34A2013263361588BBF.

Moloney:2011:DKC

[1531] D. Moloney, N. Sukhorukova, P. Vamplew, J. Ugon, G. Li, G. Beliakov, C. Philippe, H. Amiel, and A. Ugon. Detecting k-complexes for sleep stage identification using nonsmooth optimization. The ANZIAM Journal, 52(4):319–332, April 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/detecting-kcomplexes-for-
sleep-stage-identification-using-nonsmooth-optimization/436EBE82D932F1A2093DFBCE7D4BDAFE.

Hickson:2011:CCT

[1532] R. I. Hickson, S. I. Barry, H. S. Sidhu, and G. N. Mercer. A comparison of critical time definitions in multilayer diffusion. The ANZIAM Journal, 52(4):333–358, April 2011. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/comparison-of-critical-time-
definitions-in-multilayer-diffusion/40E3BD109EEBDEEB98375AF5D4C219E0.

Singh:2011:EIS

[1533] S. S. Singh. Effect of initial stresses on incident qSV-waves in prestressed elastic half-spaces. The ANZIAM Journal, 52(4):359–371, April 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/effect-of-initial-stresses-on-incident-qsvwaves-in-
prestressed-elastic-halfspaces/929968EF5160E0BE37BAE83F48E2A057.

Nguyen:2011:SSJ

[1534] Dung Tien Nguyen, Xuerong Mao, G. Yin, and Chenggui Yuan. Stability of singular jump-linear systems with a large state space: a twotime-scale approach. The ANZIAM Journal, 52(4):372–390, April 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stability-of-singular-jumplinear-systems-with-a-large-
state-space-a-twotimescale-approach/B923D791028A9A60E1372F1BD5158F82.

Anonymous:2011:AVIc

[1535] Anonymous. ANZ volume 52 issue 4 cover and front matter. The ANZIAM Journal, 52(4):f1–f2, April 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-52-issue-4-cover-and-front-matter/0A51C1DC6F1FD500F028E427B078F5BB.

Anonymous:2011:AVId

[1536] Anonymous. ANZ volume 52 issue 4 cover and back matter. The ANZIAM Journal, 52(4):b1–b10, April 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-52-issue-4-cover-and-back-matter/3D6E24227F2D2D0DE0D9CB4B192F5DCB.

Kuo:2011:QMC

[1537] F. Y. Kuo, Ch. Schwab, and I. H. Sloan. Quasi-Monte Carlo methods for high-dimensional integration: the standard (weighted Hilbert space) setting and beyond. The ANZIAM Journal, 53(1):1–37, July 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/quasimonte-carlo-methods-for-highdimensional-integration-
the-standard-weighted-hilbert-space-setting-and-beyond/0F2D532AF2A533AA02975520D9341A27.

See corrections [1553, 1581].

Plank:2011:EPD

[1538] M. J. Plank. Effects of predator diet breadth on stability of size spectra. The ANZIAM Journal, 53(1):38–47, July 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/effects-
of-predator-diet-breadth-on-stability-of-size-spectra/8A504501C11ABA838098DBA44B622CAB.

Kheirfam:2011:FNS

[1539] B. Kheirfam. A full NT-step infeasible interior-point algorithm for semidefinite optimization based on a self-regular proximity. The ANZIAM Journal, 53(1):48–67, July 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/full-ntstep-
infeasible-interiorpoint-algorithm-for-semidefinite-optimization-
based-on-a-selfregular-proximity/76A968ADB30384000BB3ACD85F175106.

Gui:2011:OCM

[1540] W. H. Gui, X. Y. Shen, N. Chen, C. H. Yang, and L. Y. Wang. Optimal control of multiple-time delayed systems based on the control parameterization method. The ANZIAM Journal, 53(1):68–86, July 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-control-of-multipletime-delayed-systems-based-
on-the-control-parameterization-method/9A6399ADFE3374FF7A99369702B1E67A.

Anonymous:2011:AVIe

[1541] Anonymous. ANZ volume 53 issue 1 cover and front matter. The ANZIAM Journal, 53(1):f1–f2, July 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-53-issue-1-cover-and-front-matter/BA2F25AC45C1B252FC2040368DDA3220.

Anonymous:2011:AVIf

[1542] Anonymous. ANZ volume 53 issue 1 cover and back matter. The ANZIAM Journal, 53(1):b1–b4, July 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-53-issue-1-cover-and-back-matter/053ACF52855B5F7D3FAFA183ABE9CC73.

Forbes:2011:RTI

[1543] Lawrence K. Forbes. Rayleigh–Taylor instabilities in axi-symmetric outflow from a point source. The ANZIAM Journal, 53(2):87–121, October 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/rayleightaylor-instabilities-in-axisymmetric-outflow-
from-a-point-source/160F4C7D816C4A8EF6D3C7D0AA6C6C8E.

Binder:2011:MPU

[1544] Benjamin J. Binder, Emily J. Hackett-Jones, Jonathan Tuke, and Kerry A. Landman. A modified Pólya urn process and an index for spatial distributions with volume exclusion. The ANZIAM Journal, 53(2): 122–133, October 2011. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/modified-polya-urn-process-
and-an-index-for-spatial-distributions-with-volume-exclusion/
D9CE6B51E2844F0BF41A4BD1F2CBCD17.

Mclean:2011:ISS

[1545] William Mclean and Vidar Thomée. Iterative solution of shifted positive-definite linear systems arising in a numerical method for the heat equation based on Laplace transformation and quadrature. The ANZIAM Journal, 53(2):134–155, October 2011. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/iterative-solution-of-shifted-positivedefinite-linear-
systems-arising-in-a-numerical-method-for-the-heat-equation-
based-on-laplace-transformation-and-quadrature/C4C660766E68DB1D8F5509E15C1D4430.

Fijavz:2011:EAM

[1546] Marjeta Kramar Fijavz, Mitja Lakner, and Marjeta Skapin Rugelj. An equal-area method for scalar conservation laws. The ANZIAM Journal, 53(2):156–170, October 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/an-equalarea-method-for-
scalar-conservation-laws/7FD83390FF458FBD0A8A066743F3BF3B.

Anonymous:2011:AVIg

[1547] Anonymous. ANZ volume 53 issue 2 cover and front matter. The ANZIAM Journal, 53(2):f1–f2, October 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-53-issue-2-cover-and-front-matter/F376394C6549B26AD82AB73A8A20AF01.

Anonymous:2011:AVIh

[1548] Anonymous. ANZ volume 53 issue 2 cover and back matter. The ANZIAM Journal, 53(2):b1–b6, October 2011. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-53-issue-2-cover-and-back-matter/06FAA51267EB2F860E3EEC35038025CD.

Crowdy:2012:CSM

[1549] Darren Crowdy. Conformal slit maps in applied mathematics. The ANZIAM Journal, 53(3):171–189, January 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/conformal-
slit-maps-in-applied-mathematics/91FE5DCF0CB1360F517135D80A1998E6.

Tuck:2012:TSB

[1550] E. O. Tuck and Y. M. Stokes. On thin or slender bodies. The ANZIAM Journal, 53(3):190–212, January 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/on-thin-
or-slender-bodies/14CF97A05CE1DABD0FD2ED6065FF2695.

Ndii:2012:MIW

[1551] Meksianis Z. Ndii, Roslyn I. Hickson, and Geoffry N. Mercer. Modelling the introduction of Wolbachia into Aedes aegypti mosquitoes to reduce dengue transmission. The ANZIAM Journal, 53(3):213–227, January 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/modelling-the-introduction-of-wolbachia-into-aedes-
aegypti-mosquitoes-to-reduce-dengue-transmission/462B67A054CFC051E98E947F9BE421DA.

Letchford:2012:IVM

[1552] Nicholas A. Letchford, Lawrence K. Forbes, and Graeme C. Hocking. Inviscid and viscous models of axisymmetric fluid jets or plumes. The ANZIAM Journal, 53(3):228–250, January 2012. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/inviscid-and-viscous-models-of-axisymmetric-fluid-jets-
or-plumes/2F66A32BF3F601E7F824EA5712B8F109.

Kuo:2012:CQM

[1553] F. Y. Kuo, Ch. Schwab, and I. H. Sloan. Correction to ‘Quasi-Monte Carlo methods for high-dimensional integration: the standard (weighted Hilbert space) setting and beyond’. The ANZIAM Journal, 53(3): 251, January 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/correction-to-quasimonte-carlo-methods-
for-highdimensional-integration-the-standard-weighted-hilbert-
space-setting-and-beyond/5631D3CA3C06ABDD277816D3E200455D. See

[1537].

Anonymous:2012:AVIa

[1554] Anonymous. ANZ volume 53 issue 3 cover and front matter. The ANZIAM Journal, 53(3):f1–f2, January 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-53-issue-3-cover-and-front-matter/4A7C437C16A56A59D9F96AA8BEF56351.

Anonymous:2012:AVIb

[1555] Anonymous. ANZ volume 53 issue 3 cover and back matter. The ANZIAM Journal, 53(3):b1–b6, January 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-53-issue-3-cover-and-back-matter/B89B684FA96E0805CB0B2821EDCC7205.

Cosgrove:2012:SWT

[1556] Jason M. Cosgrove and Lawrence K. Forbes. Selective withdrawal of a two-layer viscous fluid. The ANZIAM Journal, 53(4):253–277, April 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/selective-withdrawal-of-a-twolayer-viscous-
fluid/AF4E764D16D56F3A750333B7339A179A.

Chen:2012:NSA

[1557] C. X. Chen, Y. Ding, and J. A. Gear. Numerical simulation of atherosclerotic plaque growth using two-way fluid-structural interaction. The ANZIAM Journal, 53(4):278–291, April 2012. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/numerical-simulation-of-atherosclerotic-plaque-growth-
using-twoway-fluidstructural-interaction/1044F4169A2F6B6CF38C8F5561BF1EAC.

Wong:2012:OCS

[1558] K. H. Wong and W. M. Tang. Optimal control of switched impulsive systems with time delay. The ANZIAM Journal, 53(4):292–307, April 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-control-of-switched-impulsive-systems-with-
time-delay/AE5F4321D755BC52EE9FA2647EB07D7F.

Kasimova:2012:TSH

[1559] R. G. Kasimova and Yu. V. Obnosov. Topology of steady heat conduction in a solid slab subject to a nonuniform boundary condition: the Carslaw– Jaeger solution revisited. The ANZIAM Journal, 53(4):308–320, April 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/topology-of-steady-heat-conduction-in-a-solid-slab-
subject-to-a-nonuniform-boundary-condition-the-carslawjaeger-
solution-revisited/2BBE9FDE59B498B724DF2A40B0647610.

Matthews:2012:ANS

[1560] Miccal T. Matthews and Karen M. Hastie. An analytical and numerical study of unsteady channel flow with slip. The ANZIAM Journal, 53(4):321–336, April 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/an-analytical-and-
numerical-study-of-unsteady-channel-flow-with-slip/DD3D26DE616DF43D211CA74A1FB8460D.

Anonymous:2012:AVIc

[1561] Anonymous. ANZ volume 53 issue 4 cover and back matter. The ANZIAM Journal, 53(4):b1–b4, April 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-53-issue-4-cover-and-back-matter/16C8190406D2ADA3EB939986B54A51DF.

Anonymous:2012:AVId

[1562] Anonymous. ANZ volume 53 issue 4 cover and front matter. The ANZIAM Journal, 53(4):f1–f2, April 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-53-issue-4-cover-and-front-matter/A85CD16E20B383565BBC8BB82576662A.

Simpson:2012:EID

[1563] Matthew J. Simpson and Geoffry N. Mercer. Editorial: infectious disease modelling. The ANZIAM Journal, 54(1–2):1–2, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/editorial-infectious-disease-modelling/424F1814389B4F267FA9F8DFEF46A6D9.

Bartlett:2012:EDR

[1564] J. Bartlett and M. J. Plank. Epidemic dynamics on random and scalefree networks. The ANZIAM Journal, 54(1–2):3–22, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/epidemic-dynamics-on-random-and-scalefree-networks/
C9E4B59C3BEAC49124D8C0B934A03F11.

Waters:2012:SHS

[1565] E. K. Waters, H. S. Sidhu, and G. N. Mercer. Spatial heterogeneity in simple deterministic SIR models assessed ecologically. The ANZIAM Journal, 54(1–2):23–36, October 2012. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/spatial-
heterogeneity-in-simple-deterministic-sir-models-assessed-ecologically/
C24D3E55E075B29059AF654BB6D84576.

Binder:2012:HMS

[1566] Benjamin J. Binder, Joshua V. Ross, and Matthew J. Simpson. A hybrid model for studying spatial aspects of infectious diseases. The ANZIAM Journal, 54(1–2):37–49, October 2012. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/hybrid-model-for-studying-spatial-aspects-of-infectious-
diseases/028CC7B302F28AAC52660F1608CB53D9.

Dafilis:2012:IIL

[1567] M. P. Dafilis, F. Frascoli, J. G. Wood, and J. M. Mccaw. The influence of increasing life expectancy on the dynamics of SIRS systems with immune boosting. The ANZIAM Journal, 54(1–2):50–63, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/influence-of-increasing-life-expectancy-on-
the-dynamics-of-sirs-systems-with-immune-boosting/4FFD0E2AC37E5494F8DC00E2372598C9.

Yakob:2012:LHO

[1568] L. Yakob and P. J. Mumby. Life histories offer a clue to the future of infectious disease on coral reefs. The ANZIAM Journal, 54(1–2):64–73, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/life-histories-offer-a-clue-to-the-future-of-
infectious-disease-on-coral-reefs/5BD8D97B659D6FE43EE43A6E29927AF2.

Simpson:2012:MEV

[1569] J. L. Simpson and M. G. Roberts. Modelling the effect of vaccination on the meningococcal B epidemic in New Zealand. The ANZIAM Journal, 54(1–2):74–88, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/modelling-the-effect-of-
vaccination-on-the-meningococcal-b-epidemic-in-new-zealand/
03D274F84EABEE71CB7BB86B75644E1C.

Beeton:2012:DSA

[1570] N. J. Beeton and L. K. Forbes. Dynamical systems analysis of a model describing Tasmanian Devil facial tumour disease. The ANZIAM Journal, 54(1–2):89–107, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/dynamical-systems-
analysis-of-a-model-describing-tasmanian-devil-facial-tumour-
disease/37A1411D5249EF813105A21010D57A4C.

Roberts:2012:TSE

[1571] M. G. Roberts. A two-strain epidemic model with uncertainty in the interaction. The ANZIAM Journal, 54(1–2):108–115, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/twostrain-epidemic-model-with-uncertainty-in-the-interaction/
83D32823B3FBABCE9DF851FD93B2EC58.

Anonymous:2012:AVIe

[1572] Anonymous. ANZ volume 54 issue 1–2 cover and front matter. The ANZIAM Journal, 54(1–2):f1–f2, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-54-issue-12-cover-and-front-matter/6E1AD72C522F21D48AC3ABC4D1FF6259.

Anonymous:2012:AVIf

[1573] Anonymous. ANZ volume 54 issue 1–2 cover and back matter. The ANZIAM Journal, 54(1–2):b1–b7, October 2012. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-54-issue-12-cover-and-back-matter/1CA32B07C254358A90CF616A5FD196B9.

Wake:2013:CPM

[1574] G. C. Wake and H. M. Byrne. Calculus from the past: multiple delay systems arising in cancer cell modelling. The ANZIAM Journal, 54(3):117–126, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/calculus-from-the-
past-multiple-delay-systems-arising-in-cancer-cell-modelling/
84F3DA96115F8F637AE22FAAAF52CDE8.

Simpson:2013:CTT

[1575] Matthew J. Simpson, Adam J. Ellery, Scott W. McCue, and Ruth E. Baker. Critical timescales and time intervals for coupled linear processes. The ANZIAM Journal, 54(3):127–142, January 2013. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/critical-timescales-and-time-intervals-for-coupled-
linear-processes/E7AF5F8DCB0A06E5A57F96205475A53A.

Huang:2013:MPC

[1576] Shuai Huang, Zhong Wan, and Songhai Deng. A modified projected conjugate gradient algorithm for unconstrained optimization problems. The ANZIAM Journal, 54(3):143–152, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/modified-projected-conjugate-gradient-algorithm-for-
unconstrained-optimization-problems/F0E474D62AB843A18F726670E5F399B1.

Xu:2013:NWE

[1577] Runzhang Xu, Yanbing Yang, Shaohua Chen, Jia Su, Jihong Shen, and Shaobin Huang. Nonlinear wave equations and reaction–diffusion equations with several nonlinear source terms of different signs at high energy level. The ANZIAM Journal, 54(3):153–170, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonlinear-wave-equations-and-reactiondiffusion-equations-
with-several-nonlinear-source-terms-of-different-signs-at-high-
energy-level/39B49C280DF92E9CF87CD7FA00FF72E6.

Solekhudin:2013:DRB

[1578] I. Solekhudin and K. C. Ang. A dual-reciprocity boundary element method for steady infiltration problems. The ANZIAM Journal, 54(3): 171–180, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/dualreciprocity-boundary-element-
method-for-steady-infiltration-problems/C87AB70B070847FC8F7CBEAE1C4012E4.

Bratsos:2013:ISO

[1579] A. G. Bratsos. An improved second-order numerical method for the generalized Burgers–Fisher equation. The ANZIAM Journal, 54(3):181– 199, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-improved-secondorder-numerical-method-
for-the-generalized-burgersfisher-equation/3C63C8E777102CF15541C45AE43E0D39.

Li:2013:IPC

[1580] R. Li, Y. J. Shi, and H. L. Xu. Integrated PID controller design for an unmanned aerial vehicle with static stability. The ANZIAM Journal, 54 (3):200–215, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/integrated-pid-controller-design-
for-an-unmanned-aerial-vehicle-with-static-stability/3CFBF3E6A859422D6018A62C84003E0C.

Kuo:2013:CQM

[1581] F. Y. Kuo, Ch. Schwab, and I. H. Sloan. Correction to ’Quasi-Monte Carlo methods for high-dimensional integration: the standard (weighted Hilbert space) setting and beyond’. The ANZIAM Journal, 54(3):216– 219, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/correction-to-quasimonte-carlo-methods-
for-highdimensional-integration-the-standard-weighted-hilbert-
space-setting-and-beyond/DFA027EBF7C3A34229A9E22CFF1D515C. See

[1537].

Anonymous:2013:AVIa

[1582] Anonymous. ANZ volume 54 issue 3 cover and back matter. The ANZIAM Journal, 54(3):b1–b9, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-54-issue-3-cover-and-back-matter/901509A3DAAD81DA14D3A4FD5515FA71.

Anonymous:2013:AVIb

[1583] Anonymous. ANZ volume 54 issue 3 cover and front matter. The ANZIAM Journal, 54(3):f1–f2, January 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-54-issue-3-cover-and-front-matter/105D934D87EECFB5165898EDF0E5706F.

Baowan:2013:DJR

[1584] D. Baowan, B. J. Cox, and J. M. Hill. Determination of join regions between carbon nanostructures using variational calculus. The ANZIAM Journal, 54(4):221–247, April 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/determination-of-join-
regions-between-carbon-nanostructures-using-variational-calculus/
31B4894B55ABB2F5C00A1197090E1692.

Leduc:2013:EOG

[1585] Guillaume Leduc. A European option general first-order error formula. The ANZIAM Journal, 54(4):248–272, April 2013. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/european-option-general-firstorder-error-formula/CE263CF8AC0BC71D676B7C5EA33176CE.

Nelson:2013:BPS

[1586] M. I. Nelson, T. Nicholls, and N. Hamzah. A biological process subject to noncompetitive substrate inhibition in a generalized flow reactor. The ANZIAM Journal, 54(4):273–290, April 2013. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/biological-process-subject-to-noncompetitive-substrate-
inhibition-in-a-generalized-flow-reactor/4873A5F5FB0B92677BDD2E818A7AF7A1.

Park:2013:BOC

[1587] Jong Yeoul Park, Sun Hye Park, and Yong Han Kang. Bilinear optimal control of the velocity term in a von Kármán plate equation. The ANZIAM Journal, 54(4):291–305, April 2013. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/bilinear-optimal-control-of-the-velocity-term-in-a-
von-karman-plate-equation/1ACC9E08455139E96936CD1AFF98BE61.

Zhong:2013:AKF

[1588] Min Zhong, R. J. Loy, and R. S. Anderssen. Approximating the Kohlrausch function by sums of exponentials. The ANZIAM Journal, 54(4):306–323, April 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/approximating-the-
kohlrausch-function-by-sums-of-exponentials/1F2BD299466198D202D9D2355E34116F.

Anonymous:2013:AVIc

[1589] Anonymous. ANZ volume 54 issue 4 cover and front matter. The ANZIAM Journal, 54(4):f1–f2, April 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-54-issue-4-cover-and-front-matter/008B26503833FF3FB238409B73294227.

Anonymous:2013:AVId

[1590] Anonymous. ANZ volume 54 issue 4 cover and back matter. The ANZIAM Journal, 54(4):b1–b7, April 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-54-issue-4-cover-and-back-matter/49E25AD1911ED311B0BF3F89C64DC1EE.

Forbes:2013:NTW

[1591] Lawrence K. Forbes. A note on travelling waves in competitive reaction systems. The ANZIAM Journal, 55(1):1–13, July 2013.

CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-travelling-waves-in-competitive-reaction-systems/
E91C99FED6D484DBA171EB872585A754.

Kilianova:2013:TMS

[1592] S. Kilianová and D. Sevcovic. A transformation method for solving the Hamilton–Jacobi–Bellman equation for a constrained dynamic stochastic optimal allocation problem. The ANZIAM Journal, 55(1):14–38, July 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/transformation-method-
for-solving-the-hamiltonjacobibellman-equation-for-a-constrained-
dynamic-stochastic-optimal-allocation-problem/3493295BE8CD62E0CB7034FE37CC6FE2.

Molano:2013:LST

[1593] Luis Alejandro Molano Molano. On Laguerre–Sobolev type orthogonal polynomials: zeros and electrostatic interpretation. The ANZIAM Journal, 55(1):39–54, July 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
laguerresobolev-type-orthogonal-polynomials-zeros-and-electrostatic-
interpretation/7C80F10B35FC316150A9D4450008F06A.

Slade:2013:SIT

[1594] Paul F. Slade. Some inequalities for theoretical spatial ecology. The ANZIAM Journal, 55(1):55–68, July 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/some-inequalities-
for-theoretical-spatial-ecology/ACF7D12D1E7F27077BA582CE8FDECC67.

Khader:2013:NFA

[1595] M. M. Khader. A new formula for Adomian polynomials and the analysis of its truncated series solution for fractional non-differentiable initial value problems. The ANZIAM Journal, 55(1):69–92, July 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/new-formula-for-{Adomian}-polynomials-and-the-analysis-
of-its-truncated-series-solution-for-fractional-nondifferentiable-
initial-value-problems/B2B61A77F4960DAAB93946DC552E4BCA.

Anonymous:2013:AVIe

[1596] Anonymous. ANZ volume 55 issue 1 cover and back matter. The ANZIAM Journal, 55(1):b1–b6, July 2013. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-55-issue-1-cover-and-back-matter/0B002E678AFF2BA3E0929399F372351B.

Anonymous:2013:AVIf

[1597] Anonymous. ANZ volume 55 issue 1 cover and front matter. The ANZIAM Journal, 55(1):f1–f2, July 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-55-issue-1-cover-and-front-matter/0AA91A7C1586AFBC1D30D5ED106DDAF1.

Hywood:2013:BRW

[1598] Jack D. Hywood and Kerry A. Landman. Biased random walks, partial differential equations and update schemes. The ANZIAM Journal, 55(2): 93–108, October 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/biased-random-walks-partial-differential-
equations-and-update-schemes/5C6B22B2EC76014F2F312A1B4AA56894.

Robertson:2013:CMB

[1599] B. L. Robertson, C. J. Price, and M. Reale. A CARTOPT method for bound-constrained global optimization. The ANZIAM Journal, 55(2): 109–128, October 2013. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/cartopt-method-for-boundconstrained-
global-optimization/542108026B5472E28D3A55F077E2BFB3.

Jin:2013:ODP

[1600] Zhuo Jin and George Yin. An optimal dividend policy with delayed capital injections. The ANZIAM Journal, 55(2):129–150, October 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/an-optimal-dividend-policy-with-delayed-capital-injections/
C62587C0152A4F03C486A44507740F41.

Frauendiener:2013:BND

[1601] Jörg Frauendiener and Ralf Peter. Blow-up of the nonequivariant (2+1)- dimensional wave map. The ANZIAM Journal, 55(2):151–161, October 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/blowup-of-the-nonequivariant-21-dimensional-wave-map/
75AE69E7664E9F7A79E3153BFF248670.

Asif:2013:ECB

[1602] Mohammad Asif, Emad Ali, and Abdelhamid Ajbar. On the existence of chaotic behaviour in pure and simple microbial competition: the role of Contois kinetics. The ANZIAM Journal, 55(2):162–174, October 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-the-existence-of-chaotic-behaviour-in-pure-and-simple-
microbial-competition-the-role-of-contois-kinetics/ECC67ACBAFA395B16E2F1185B6736C53.

Dhillon:2013:TDW

[1603] Harpreet Dhillon and B. N. Mandal. Three-dimensional wave-free potentials in the theory of water waves. The ANZIAM Journal, 55(2):175– 195, October 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/threedimensional-wavefree-potentials-
in-the-theory-of-water-waves/F544F02586F2AD43E3E4A30502908472.

Anonymous:2013:AVIg

[1604] Anonymous. ANZ volume 55 issue 2 cover and front matter. The ANZIAM Journal, 55(2):f1–f2, October 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-55-issue-2-cover-and-front-matter/C06FB9ECC6668D4454D962FECDCFFFEE.

Anonymous:2013:AVIh

[1605] Anonymous. ANZ volume 55 issue 2 cover and back matter. The ANZIAM Journal, 55(2):b1–b11, October 2013. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-55-issue-2-cover-and-back-matter/D0442450F460E019CB7CE8E611649338.

Willmott:2014:MRP

[1606] G. R. Willmott and B. G. Smith. Modelling of resistive pulse sensing: flexible methods for submicron particles. The ANZIAM Journal, 55(3): 197–213, January 2014. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/modelling-of-resistive-pulse-
sensing-flexible-methods-for-submicron-particles/288F04002ABBAA4A91663B29F9CB9F33.

Karaoglu:2014:HBA

[1607] E. Karaoglu and H. Merdan. Hopf bifurcation analysis for a ratiodependent predator–prey system involving two delays. The ANZIAM Journal, 55(3):214–231, January 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/hopf-bifurcation-analysis-
for-a-ratiodependent-predatorprey-system-involving-two-delays/
2248FF1BFDBBB0F8B14417AFCB4A1EEE.

Forbes:2014:ESI

[1608] Lawrence K. Forbes and Michael A. Brideson. Exact solutions for interfacial outflows with straining. The ANZIAM Journal, 55(3):232–244, January 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/exact-solutions-for-interfacial-outflows-with-
straining/F19DF7DC20364EB61FA07D688BED7597.

Goswami:2014:OES

[1609] Deepjyoti Goswami, Amiya K. Pani, and Sangita Yadav. Optimal L^{2} estimates for the semidiscrete Galerkin method applied to parabolic integro-differential equations with nonsmooth data. The ANZIAM Journal, 55(3):245–266, January 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/optimal-def-xmlpi-
1def-mathsfbi-1boldsymbol-mathsf-1let-le-leqslant-let-leq-leqslant-
let-ge-geqslant-let-geq-geqslant-def-pr-mathit-prdef-fr-mathit-
frdef-rey-mathit-rel2-estimates-for-the-semidiscrete-galerkin-
method-applied-to-parabolic-integrodifferential-equations-with-
nonsmooth-data/9BD823EB0433DDDE473C6A1666C43C4C.

Manam:2014:MCG

[1610] S. R. Manam and R. B. Kaligatla. Membrane-coupled gravity wave scattering by a vertical barrier with a gap. The ANZIAM Journal, 55(3):267– 288, January 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/membranecoupled-gravity-wave-scattering-
by-a-vertical-barrier-with-a-gap/0BD3FF861F1B8F60D69EA5354D2B3036.

Hsiao:2014:NDN

[1611] Hsi-Yue Hsiao, Chih-Yao Hsieh, Xi Chen, Yongyi Gong, Xiaonan Luo, and Guojun Liao. New development of nonrigid registration. The ANZIAM Journal, 55(3):289–297, January 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/new-
development-of-nonrigid-registration/71CAD8370FAA16583C4D2444D3DC6424.

Anonymous:2014:AVIa

[1612] Anonymous. ANZ volume 55 issue 3 cover and front matter. The ANZIAM Journal, 55(3):f1–f2, January 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-55-issue-3-cover-and-front-matter/C4D181EADACA975CE596AEDE410FBF2C.

Anonymous:2014:AVIb

[1613] Anonymous. ANZ volume 55 issue 3 cover and back matter. The ANZIAM Journal, 55(3):b1–b5, January 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-55-issue-3-cover-and-back-matter/1A23D74E94229D9438B0316307C43E69.

AragonArtacho:2014:DRF

[1614] Francisco J. Aragón Artacho, Jonathan M. Borwein, and Matthew K. Tam. Douglas–Rachford feasibility methods for matrix completion problems. The ANZIAM Journal, 55(4):299–326, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL http:

//arxiv.org/abs/1308.4243;   http://journals.cambridge.org/
action/displayAbstract?fromPage=online&aid=9347336;   https:
//www.cambridge.org/core/journals/anziam-journal/article/douglasrachford-
feasibility-methods-for-matrix-completion-problems/0DCB430BF7CF1187A7A1DB5B9C3C2BCC.

Hocking:2014:NAS

[1615] G. C. Hocking and H. Zhang. A note on axisymmetric supercritical coning in a porous medium. The ANZIAM Journal, 55(4):327–335, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-axisymmetric-supercritical-coning-in-a-porous-
medium/1B9218CD96D5367DCAF3E146FDE2FD37.

Kasbawati:2014:ETD

[1616] Kasbawati, A. Y. Gunawan, R. Hertadi, and K. A. Sidarto. Effects of time delay on the dynamics of a kinetic model of a microbial fermentation process. The ANZIAM Journal, 55(4):336–356, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/effects-of-time-delay-on-the-dynamics-of-a-kinetic-
model-of-a-microbial-fermentation-process/E4998C8424A0524A597B31ED9245EB4F.

Eggar:2014:ORR

[1617] M. H. Eggar. On the organization of round robin tournaments with constraints. The ANZIAM Journal, 55(4):357–361, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-organization-of-round-robin-tournaments-with-constraints/
67C7FE33CC4835109353531A15161D74.

Zhang:2014:CFP

[1618] Li-Wei Zhang. A closed-form pricing formula for variance swaps with mean-reverting Gaussian volatility. The ANZIAM Journal, 55(4): 362–382, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/closedform-pricing-formula-
for-variance-swaps-with-meanreverting-gaussian-volatility/3E7A1B1026964FF06F27DFE5F01C28E8.

Saad:2014:MSS

[1619] E. I. Saad. Motion of a slip sphere in a nonconcentric fictitious spherical envelope of micropolar fluid. The ANZIAM Journal, 55(4):383–401, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/motion-of-a-slip-sphere-in-a-nonconcentric-
fictitious-spherical-envelope-of-micropolar-fluid/3E77283A2A798F97CE10D878C77511E5.

Anonymous:2014:AVIc

[1620] Anonymous. ANZ volume 55 issue 4 cover and front matter. The ANZIAM Journal, 55(4):f1–f2, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-55-issue-4-cover-and-front-matter/DC957F9886DEE54611A0DC8B6BD17865.

Anonymous:2014:AVId

[1621] Anonymous. ANZ volume 55 issue 4 cover and back matter. The ANZIAM Journal, 55(4):b1–b7, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-55-issue-4-cover-and-back-matter/1F1FEFF40ECF942D0A72448C7F0B5BC2.

Rujivan:2014:SCF

[1622] Sanae Rujivan and Song-Ping Zhu. A simple closed-form formula for pricing discretely-sampled variance swaps under the Heston model. The ANZIAM Journal, 56(1):1–27, July 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/simple-closedform-
formula-for-pricing-discretelysampled-variance-swaps-under-
the-heston-model/2CE2D740ECDE49182735925664E33388.

Forbes:2014:TMT

[1623] Lawrence K. Forbes. On turbulence modelling and the transition from laminar to turbulent flow. The ANZIAM Journal, 56(1):28–47, July 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-turbulence-modelling-and-the-transition-
from-laminar-to-turbulent-flow/365CACDD09AFEF1860E584BB84659535.

Mehrpouya:2014:CAC

[1624] M. A. Mehrpouya, M. Shamsi, and M. Razzaghi. A combined adaptive control parametrization and homotopy continuation technique for the numerical solution of bang-bang optimal control problems. The ANZIAM Journal, 56(1):48–65, July 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/combined-adaptive-
control-parametrization-and-homotopy-continuation-technique-
for-the-numerical-solution-of-bangbang-optimal-control-problems/
03858661C30055551C58A3FBE2BB04C6.

Liang:2014:OTC

[1625] Xiaoqing Liang, Lihua Bai, and Junyi Guo. Optimal time-consistent portfolio and contribution selection for defined benefit pension schemes under mean-variance criterion. The ANZIAM Journal, 56(1):66–90, July 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-timeconsistent-portfolio-and-contribution-selection-
for-defined-benefit-pension-schemes-under-meanvariance-criterion/
05BE803BAC2A2D068E05164B506A7C96.

Tantawy:2014:SSC

[1626] S. F. Tantawy. Solving a special class of multiple objective linear fractional programming problems. The ANZIAM Journal, 56(1):91– 103, July 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/solving-a-special-class-of-multiple-
objective-linear-fractional-programming-problems/09A4A89C63E862DA3F21E4DEDE1120A6.

Anonymous:2014:AVIe

[1627] Anonymous. ANZ volume 56 issue 1 cover and front matter. The ANZIAM Journal, 56(1):f1–f2, July 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-56-issue-1-cover-and-front-matter/A5A1C955BF1E00CE02174B6006D3449D.

Anonymous:2014:AVIf

[1628] Anonymous. ANZ volume 56 issue 1 cover and back matter. The ANZIAM Journal, 56(1):b1–b7, July 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-56-issue-1-cover-and-back-matter/D43084206B6F2CFAC4AAA343609CB72C.

Jarvis:2014:AIT

[1629] P. D. Jarvis and J. G. Sumner. Adventures in invariant theory. The ANZIAM Journal, 56(2):105–115, October 2014. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/adventures-in-invariant-theory/F0317B06FABD49D27B9F1AA9A15A9DFF.

Jiang:2014:CFO

[1630] Le Jiang, Jin Huang, Jun Liu, and Xiao-Guang Lv. A combined first-order and second-order variation approach for multiplicative noise removal. The ANZIAM Journal, 56(2):116–137, October 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/combined-firstorder-and-secondorder-
variation-approach-for-multiplicative-noise-removal/63F126080041039E51B5D6D75361DFF5.

Liu:2014:FPC

[1631] Yanqing Liu and Fei Liu. Feedback predictive control of nonhomogeneous Markov jump systems with nonsymmetric constraints. The ANZIAM Journal, 56(2):138–149, October 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/feedback-predictive-
control-of-nonhomogeneous-markov-jump-systems-with-nonsymmetric-
constraints/3DA86C4E178152F36C797EF6801FCC1F.

Hocking:2014:NSF

[1632] G. C. Hocking, L. K. Forbes, and T. E. Stokes. A note on steady flow into a submerged point sink. The ANZIAM Journal, 56(2):150–159, October 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/note-on-steady-flow-into-a-submerged-point-
sink/1A09CE8B6787D11DBBEACF3632459FAD.

Li:2014:DPG

[1633] Jueyou Li, Changzhi Wu, Zhiyou Wu, Qiang Long, and Xiangyu Wang. Distributed proximal-gradient method for convex optimization with inequality constraints. The ANZIAM Journal, 56(2):160– 178, October 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/distributed-proximalgradient-method-
for-convex-optimization-with-inequality-constraints/7FE45FCC8FB02098BCC5A10ED14992C3.

Nazemi:2014:SIH

[1634] Alireza Nazemi and Neda Mahmoudy. Solving infinite-horizon optimal control problems using the Haar wavelet collocation method. The ANZIAM Journal, 56(2):179–191, October 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/solving-
infinitehorizon-optimal-control-problems-using-the-haar-wavelet-
collocation-method/A5989C840040E50EDCAAFA00BF08873A.

Lin:2014:LIS

[1635] Yinwei Lin, Tzon-Tzer Lu, and Cha’o-Kuang Chen. Large interval solution of the Emden–Fowler equation using a modified Adomian decomposition method with an integrating factor. The ANZIAM Journal, 56(2): 192–208, October 2014. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/large-interval-solution-of-
the-emdenfowler-equation-using-a-modified-{Adomian}-decomposition-
method-with-an-integrating-factor/726EB43D481BB3D9BA4BE08BD5665C11.

Anonymous:2014:AVIg

[1636] Anonymous. ANZ volume 56 issue 2 cover and front matter. The ANZIAM Journal, 56(2):f1–f2, October 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-56-issue-2-cover-and-front-matter/0D64E86244F521A4D2C94EE1DABB8A5E.

Anonymous:2014:AVIh

[1637] Anonymous. ANZ volume 56 issue 2 cover and back matter. The ANZIAM Journal, 56(2):b1–b6, October 2014. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-56-issue-2-cover-and-back-matter/31EE3A71E24842E45C58AD338A5E409F.

Dattu:2015:NSS

[1638] H. Dattu and M. Subbiah. A note on the stability of swirling flows with radius-dependent density with respect to infinitesimal azimuthal disturbances. The ANZIAM Journal, 56(3):209–232, January 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/note-on-the-stability-of-swirling-flows-with-radiusdependent-
density-with-respect-to-infinitesimal-azimuthal-disturbances/
0BDB509699CF92389BF71C4F755E3AFE.

Paul:2015:TWO

[1639] Rhys A. Paul and Lawrence K. Forbes. Travelling waves and oscillations in Sal’nikov’s combustion reaction in a compressible gas. The ANZIAM Journal, 56(3):233–247, January 2015. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/travelling-waves-and-oscillations-in-salnikovs-combustion-
reaction-in-a-compressible-gas/D4D1D0D3A3C1A33D3467C0F1B75C2E77.

Panda:2015:TLF

[1640] Srikumar Panda, S. C. Martha, and A. Chakrabarti. Three-layer fluid flow over a small obstruction on the bottom of a channel. The ANZIAM Journal, 56(3):248–274, January 2015. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/threelayer-fluid-flow-over-a-small-obstruction-on-the-
bottom-of-a-channel/4B224B6DE050D938E50DB6796A004748.

Korostil:2015:FON

[1641] I. A. Korostil and S. R. Clarke. Fourth-order numerical methods for the coupled Korteweg–de Vries equations. The ANZIAM Journal, 56(3):275–285, January 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/fourthorder-numerical-
methods-for-the-coupled-kortewegde-vries-equations/87887DE8711D94C3F44456FC5EFDCA8B.

Chakraborty:2015:OWS

[1642] Rumpa Chakraborty and B. N. Mandal. Oblique wave scattering by a rectangular submarine trench. The ANZIAM Journal, 56(3):

286–298, January 2015. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/oblique-wave-scattering-by-
a-rectangular-submarine-trench/B52F4DDD9F241487D15E779F0AECC191.

Chen:2015:LSM

[1643] Yu Chen and Zhong Wan. A locally smoothing method for mathematical programs with complementarity constraints. The ANZIAM Journal, 56(3):299–315, January 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/locally-smoothing-
method-for-mathematical-programs-with-complementarity-constraints/
D88A352157829EE0ACF9D279C3D6291A.

Anonymous:2015:AVIa

[1644] Anonymous. ANZ volume 56 issue 3 cover and front matter. The ANZIAM Journal, 56(3):f1–f2, January 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-56-issue-3-cover-and-front-matter/99E52CEC502C2E3A6188D17ACE5F36A6.

Anonymous:2015:AVIb

[1645] Anonymous. ANZ volume 56 issue 3 cover and back matter. The ANZIAM Journal, 56(3):b1–b7, January 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-56-issue-3-cover-and-back-matter/3D1E17E0DC4B0F98C26CB7FD9853E1CC.

Forbes:2015:KHC

[1646] Lawrence K. Forbes, Rhys A. Paul, Michael J. Chen, and David E. Horsley. Kelvin–Helmholtz creeping flow at the interface between two viscous fluids. The ANZIAM Journal, 56(4):317–358, April 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/kelvinhelmholtz-creeping-flow-at-the-interface-between-
two-viscous-fluids/7AA9A36BD9789C392F9FDB7A9C1A9F07.

Shevchenko:2015:HEE

[1647] Pavel V. Shevchenko. Holder-extendible European option: corrections and extensions. The ANZIAM Journal, 56(4):359–372, April 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/holderextendible-european-option-corrections-and-extensions/
4A32DD1F16068C9BF479012902E7DA8B.

Woodcock:2015:PPM

[1648] S. Woodcock, B. Manojlovic, M. E. Baird, and P. J. Ralph. A Poisson–Pareto model of chlorophyll — a fluorescence signals in marine environments. The ANZIAM Journal, 56(4):373–380, April 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735

(electronic).  URL https://www.cambridge.org/core/journals/
anziam-journal/article/poissonpareto-model-of-chlorophylla-
fluorescence-signals-in-marine-environments/239B12B77C99C663671C0AF631541584.

Singh:2015:TEW

[1649] S. S. Singh. Transmission of elastic waves in anisotropic nematic elastomers. The ANZIAM Journal, 56(4):381–396, April 2015. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/transmission-of-elastic-waves-in-anisotropic-nematic-
elastomers/EA93DBF4356D2E95E26E05F096E473F8.

Alqahtani:2015:BTI

[1650] Rubayyi T. Alqahtani, Mark I. Nelson, and Annette L. Worthy. A biological treatment of industrial wastewaters: Contois kinetics. The ANZIAM Journal, 56(4):397–415, April 2015. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/biological-treatment-of-industrial-wastewaters-contois-
kinetics/88A4D05DAF8F72632FCFD1EEA8F4E7F8.

Anonymous:2015:AVIc

[1651] Anonymous. ANZ volume 56 issue 4 cover and front matter. The ANZIAM Journal, 56(4):f1–f2, April 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-56-issue-4-cover-and-front-matter/5CC89735757C48DC6584CB9B6B8C980B.

Anonymous:2015:AVId

[1652] Anonymous. ANZ volume 56 issue 4 cover and back matter. The ANZIAM Journal, 56(4):b1–b7, April 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-56-issue-4-cover-and-back-matter/8D81A049B89CB3D1D7D43971FBAE647E.

Cox:2015:EMA

[1653] Barry J. Cox and Shaun C. Hendy. Editorial: modelling approach to nanoscale science and technology. The ANZIAM Journal, 57(1):1–2, July 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/editorial-modelling-approach-to-nanoscale-science-and-
technology/E160B6E0EDCA346DDD83BBF010959BC0.

Winkler:2015:DSL

[1654] R. Winkler and U. Zülicke. Discrete symmetries of low-dimensional Dirac models: a selective review with a focus on condensed-matter realizations. The ANZIAM Journal, 57(1):3–17, July 2015. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/discrete-symmetries-of-lowdimensional-dirac-models-
a-selective-review-with-a-focus-on-condensedmatter-realizations/
A217602F3172EACD57108CBAB059A1A8.

Baowan:2015:PSG

[1655] D. Baowan. Penetration of spherical gold nanoparticle into a lipid bilayer. The ANZIAM Journal, 57(1):18–28, July 2015. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/penetration-of-spherical-gold-nanoparticle-into-a-lipid-
bilayer/EC6E5AEF2C273520555FF5DC219505B0.

Lee:2015:DNM

[1656] Richard K. F. Lee and James M. Hill. Design of a nanotorimetallofullerene logic gate. The ANZIAM Journal, 57(1):29–42, July 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/design-of-a-nanotorimetallofullerene-logic-
gate/5129281390B5DA8DB80754973941643C.

Lim:2015:ARR

[1657] Wei-Xian Lim and Aaron W. Thornton. Analytical representations of regular-shaped nanostructures for gas storage applications. The ANZIAM Journal, 57(1):43–61, July 2015. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/analytical-representations-of-regularshaped-nanostructures-
for-gas-storage-applications/CC8516751C72AE6EAC52ABC3A45661BF.

Thamwattana:2015:MIW

[1658] N. Thamwattana. Modelling ion, water and ion-water cluster entering peptide nanotubes. The ANZIAM Journal, 57(1):62–78, July 2015.

CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/modelling-ion-water-and-ionwater-cluster-entering-peptide-
nanotubes/74EA7E2B0209F0181068656618EA628C.

Zhang:2015:ESL

[1659] Xingyou (Philip) Zhang, Nat J. Lund, and Shaun C. Hendy. Effective slip length: some analytical and numerical results. The ANZIAM Journal, 57(1):79–88, July 2015. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/effective-slip-length-some-
analytical-and-numerical-results/0E675095A0DDC93C4ACBADF7DC7B02C3.

Anonymous:2015:AVIe

[1660] Anonymous. ANZ volume 57 issue 1 cover and front matter. The ANZIAM Journal, 57(1):f1–f2, July 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-57-issue-1-cover-and-front-matter/CB72CAFCBFEEFAD1F39AC87904ACD460.

Anonymous:2015:AVIf

[1661] Anonymous. ANZ volume 57 issue 1 cover and back matter. The ANZIAM Journal, 57(1):b1–b10, July 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-57-issue-1-cover-and-back-matter/AC7BAF5881B6DC0903F82BF59547B052.

Forbes:2015:TTP

[1662] L. K. Forbes. Transition to turbulence from plane Couette flow. The ANZIAM Journal, 57(2):89–113, October 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/transition-
to-turbulence-from-plane-couette-flow/B1BCC409EE68B9A7B1162A112CDA3B10.

Li:2015:NSA

[1663] Miao Li, Yong Zhang, Hong Zhang, and Hong Guan. Numerical stability and accuracy of the scaled boundary finite element method in engineering applications. The ANZIAM Journal, 57(2):114–137, October 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/numerical-stability-and-accuracy-of-
the-scaled-boundary-finite-element-method-in-engineering-applications/
5A38465135932FC8FE4C62C9D34BB193.

VanBrunt:2015:CGM

[1664] B. Van Brunt, S. Gul, and G. C. Wake. A cell growth model adapted for the minimum cell size division. The ANZIAM Journal, 57(2):138–149, October 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/cell-growth-model-
adapted-for-the-minimum-cell-size-division/C8DC4B0FDECA3FB174683E1DB478613E.

Amore:2015:HSD

[1665] Paolo Amore. Heterogeneous systems in d dimensions: lower spectrum. The ANZIAM Journal, 57(2):150–165, October 2015. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/heterogeneous-systems-in-d-dimensions-lower-spectrum/
D198D666B0091E09C17FE688C3A8D471.

Charkhgard:2015:EAT

[1666] H. Charkhgard and M. Savelsbergh. Efficient algorithms for travelling salesman problems arising in warehouse order picking. The ANZIAM Journal, 57(2):166–174, October 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/efficient-algorithms-
for-travelling-salesman-problems-arising-in-warehouse-order-
picking/680A237A0B82F4BA38C39D5A77D02D28.

Fowkes:2015:BPG

[1667] Neville D. Fowkes and Andrew P. Bassom. Batch processing in a glass furnace. The ANZIAM Journal, 57(2):175–188, October 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/batch-processing-in-a-glass-furnace/
A2CE8B39006CDDFFE1EB89BF7D083B23.

Saha:2015:EST

[1668] S. Saha and S. N. Bora. Effects of surface tension on trapped modes in a two-layer fluid. The ANZIAM Journal, 57(2):189– 203, October 2015. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/effects-of-surface-tension-
on-trapped-modes-in-a-twolayer-fluid/8C844E4002A3EF97A0AF899B59FE54F3.

Anonymous:2015:AVIg

[1669] Anonymous. ANZ volume 57 issue 2 cover and front matter. The ANZIAM Journal, 57(2):f1–f2, October 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-57-issue-2-cover-and-front-matter/B38CB29F9A5477B50DBA34BAEE774E00.

Anonymous:2015:AVIh

[1670] Anonymous. ANZ volume 57 issue 2 cover and back matter. The ANZIAM Journal, 57(2):b1–b7, October 2015. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-57-issue-2-cover-and-back-matter/B7D871A7854E22D2D0734AEA82B71AFE.

Zhu:2016:ESC

[1671] Song-Ping Zhu and Alexander Novikov. Editorial: stochastic and computational methods in finance. The ANZIAM Journal, 57(3):205–206, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/editorial-stochastic-and-computational-methods-
in-finance/EF1B4F8736698A0A4C1658D887AC3F71.

Yang:2016:CFT

[1672] X. Yang, J. Liang, and Y. Wu. CPDO with finite termination: maximal return under cash-in and cash-out conditions. The ANZIAM Journal, 57 (3):207–221, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/cpdo-with-finite-termination-
maximal-return-under-cashin-and-cashout-conditions/CFA888F0DE6859B2D32F87E82073C0CF.

Goard:2016:ASB

[1673] Joanna Goard. Approximate solutions for the British put option and its optimal exercise boundary. The ANZIAM Journal, 57 (3):222–243, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/approximate-solutions-
for-the-british-put-option-and-its-optimal-exercise-boundary/
2F3D1A8DB1731218D4A51BB0D2542E81.

Rujivan:2016:NAA

[1674] Sanae Rujivan. A novel analytical approach for pricing discretely sampled gamma swaps in the Heston model. The ANZIAM Journal, 57(3):244– 268, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/novel-analytical-approach-for-pricing-
discretely-sampled-gamma-swaps-in-the-heston-model/2D307871D6F970414A37E43B3BDFB217.

Le:2016:ASP

[1675] Nhat-Tan Le, Xiaoping Lu, and Song-Ping Zhu. An analytical solution for Parisian up-and-in calls. The ANZIAM Journal, 57(3):269– 279, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-analytical-solution-for-parisian-
upandin-calls/04324F2AE7E0E77801EE333D9FF373FD.

Lai:2016:SMA

[1676] Yongzeng Lai and Haixiang Yao. Simulation of multi-asset option Greeks under a special Lévy model by Malliavin calculus. The ANZIAM Journal, 57(3):280–298, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/simulation-of-multiasset-
option-greeks-under-a-special-levy-model-by-malliavin-calculus/
B4B3E01F392117F6B24BF5B59189CB0E.

Alexander:2016:BPA

[1677] Scott Alexander, Alexander Novikov, and Nino Kordzakhia. Bounds on prices for Asian options via Fourier methods. The ANZIAM Journal, 57 (3):299–318, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/bounds-on-prices-for-asian-
options-via-fourier-methods/9B99B3B4699C11CBE54557F34792FC02.

Ling:2016:HBL

[1678] T. G. Ling and P. V. Shevchenko. Historical backtesting of local volatility model using AUD/usd vanilla options. The ANZIAM Journal, 57(3):319– 338, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/historical-backtesting-of-local-volatility-
model-using-audusd-vanilla-options/BEB0C4C0E5C789FD8FF496B0A45D1595.

Hinz:2016:SDS

[1679] J. Hinz and N. Yap. Solutions and diagnostics of switching problems with linear state dynamics. The ANZIAM Journal, 57(3):339–351, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/solutions-and-diagnostics-of-switching-problems-with-
linear-state-dynamics/5DE34317AB540F030C2F1B734DFA906C.

Zhu:2016:OPR

[1680] Huiming Zhu, Ya Huang, Jieming Zhou, Xiangqun Yang, and Chao Deng. Optimal proportional reinsurance and investment problem with constraints on risk control in a general jump-diffusion financial market. The ANZIAM Journal, 57(3):352–368, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
proportional-reinsurance-and-investment-problem-with-constraints-
on-risk-control-in-a-general-jumpdiffusion-financial-market/
32577BC9D20FC01729C4C13815F38020.

Maisano:2016:LMD

[1681] J. Maisano, A. Radchik, and T. Ling. A lognormal model for demand forecasting in the national electricity market. The ANZIAM Journal, 57 (3):369–383, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/lognormal-model-for-demand-
forecasting-in-the-national-electricity-market/2AE8584C7B69DA742D7711D1130771C9.

Anonymous:2016:AVIa

[1682] Anonymous. ANZ volume 57 issue 3 cover and front matter. The ANZIAM Journal, 57(3):f1–f2, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-57-issue-3-cover-and-front-matter/B4B09EEFEF07BBCFB1A1C179849D1C41.

Anonymous:2016:AVIb

[1683] Anonymous. ANZ volume 57 issue 3 cover and back matter. The ANZIAM Journal, 57(3):b1–b7, January 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-57-issue-3-cover-and-back-matter/B56FBCFEFA34CEA6A37F3055F5F0C753.

Zoppou:2016:SCL

[1684] C. Zoppou, S. G. Roberts, and J. Pitt. A solution of the conservation law form of the Serre equations. The ANZIAM Journal, 57(4):385–394, April 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/solution-of-the-conservation-law-form-of-the-serre-
equations/1CEE9D62A1C8D208FC5067E813D36124.

Cosgrove:2016:FLA

[1685] Jason M. Cosgrove and Lawrence K. Forbes. The formation of largeamplitude fingers in atmospheric vortices. The ANZIAM Journal, 57(4): 395–416, April 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/formation-of-largeamplitude-fingers-
in-atmospheric-vortices/C9F073AD0184202B1A078302724C779C.

Hocking:2016:EST

[1686] G. C. Hocking, H. H. N. Nguyen, L. K. Forbes, and T. E. Stokes. The effect of surface tension on free-surface flow induced by a point sink. The ANZIAM Journal, 57(4):417–428, April 2016. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/effect-of-surface-tension-on-freesurface-flow-induced-
by-a-point-sink/52562EAA58DC37452C55C7B8CD596A04.

Mcculloch:2016:EAE

[1687] K. McCulloch, M. G. Roberts, and C. R. Laing. Exact analytical expressions for the final epidemic size of an SIR model on small networks. The ANZIAM Journal, 57(4):429–444, April 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/exact-analytical-expressions-for-the-final-
epidemic-size-of-an-sir-model-on-small-networks/C5352260B2D077BEFC8C690AC072F1F9.

Celik:2016:BAL

[1688] Canan Çelik and Gökçen Çekiç. Bifurcation analysis of a logistic predator–prey system with delay. The ANZIAM Journal, 57(4):445–460, April 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/bifurcation-analysis-of-a-logistic-predatorprey-
system-with-delay/B6513B68B61A5B436758DC2F08D01BB2.

Mortezaee:2016:WCS

[1689] Marziyeh Mortezaee and Alireza Nazemi. A wavelet collocation scheme for solving some optimal path planning problems. The ANZIAM Journal, 57(4):461–481, April 2016. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/wavelet-collocation-scheme-
for-solving-some-optimal-path-planning-problems/D0195C6A313DE96EC0FF0E86C2B8853B.

Zhang:2016:HOU

[1690] Qian Zhang, Jinliang Yan, and Zhiyue Zhang. High-order upwind finite volume element method for first-order hyperbolic optimal control problems. The ANZIAM Journal, 57(4):482–498, April 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/highorder-upwind-finite-volume-element-method-for-firstorder-
hyperbolic-optimal-control-problems/B567BD3147CF3E1DEE3019FBBFA89903.

Anonymous:2016:AVIc

[1691] Anonymous. ANZ volume 57 issue 4 cover and front matter. The ANZIAM Journal, 57(4):f1–f2, April 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-57-issue-4-cover-and-front-matter/B61A82B4A9CB3344942525E59FCAAC18.

Anonymous:2016:AVId

[1692] Anonymous. ANZ volume 57 issue 4 cover and back matter. The ANZIAM Journal, 57(4):b1–b8, April 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-57-issue-4-cover-and-back-matter/F80A3DEDE9084F7CF941E951FF8F3483.

Anderssen:2016:SJI

[1693] R. S. Anderssen, A. R. Davies, F. R. de Hoog, and R. J. Loy. Simple joint inversion localized formulae for relaxation spectrum recovery. The ANZIAM Journal, 58(1):1–9, July 2016. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/simple-joint-inversion-localized-formulae-for-relaxation-
spectrum-recovery/CFE95820BD4DD1E6415C80AC940FE05E.

Albrecht:2016:CTC

[1694] Amie Albrecht, Phil Howlett, and Peter Pudney. The cost–time curve for an optimal train journey on level track. The ANZIAM Journal, 58 (1):10–32, July 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/costtime-curve-for-an-optimal-train-
journey-on-level-track/15DF103AF47A83AA3BB67343589B9EF4.

VanGoethem:2016:DEI

[1695] Nicolas Van Goethem. Direct expression of incompatibility in curvilinear systems. The ANZIAM Journal, 58(1):33–50, July 2016. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/direct-expression-of-incompatibility-in-curvilinear-
systems/EDB3CD0C3862BB29398E0E1855B5B0F7.

Berres:2016:PRC

[1696] S. Berres, A. Coronel, R. Lagos, and M. Sepúlveda. Performance of a real coded genetic algorithm for the calibration of scalar conservation laws. The ANZIAM Journal, 58(1):51–77, July 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735

(electronic).  URL https://www.cambridge.org/core/journals/
anziam-journal/article/performance-of-a-real-coded-genetic-
algorithm-for-the-calibration-of-scalar-conservation-laws/14B312460F86955603EBBFFCA097FE7A.

Casas:2016:APH

[1697] J. M. Casas, M. Ladra, B. A. Omirov, and R. Turdibaev. On the algebraic properties of the human ABO-blood group inheritance pattern. The ANZIAM Journal, 58(1):78–95, July 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-algebraic-properties-of-the-human-aboblood-group-inheritance-
pattern/9E5F237F16DC6EF2B5B69F5DCE52F28E.

Zaman:2016:UTL

[1698] Akbar Zaman, Nasir Ali, O. Anwar Beg, and M. Sajid. Unsteady twolayered blood flow through a w-shaped stenosed artery using the generalized Oldroyd-B fluid model. The ANZIAM Journal, 58(1):96–118, July 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/unsteady-twolayered-blood-flow-through-a-w-
shaped-stenosed-artery-using-the-generalized-oldroydb-fluid-
model/5749E3C3D767A403BD25A636880E7C72.

Anonymous:2016:AVIe

[1699] Anonymous. ANZ volume 58 issue 1 cover and front matter. The ANZIAM Journal, 58(1):f1–f2, July 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-58-issue-1-cover-and-front-matter/8C0454F3CB6ADD000D9BFF54633C3EE8.

Anonymous:2016:AVIf

[1700] Anonymous. ANZ volume 58 issue 1 cover and back matter. The ANZIAM Journal, 58(1):b1–b4, July 2016. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-58-issue-1-cover-and-back-matter/F89E231F10780F789154BB9C29D6A3DF.

Boland:2016:MSV

[1701] J. Boland, P. Howlett, J. Piantadosi, and R. Zakaria. Modelling and simulation of volumetric rainfall for a catchment in the Murray– Darling basin. The ANZIAM Journal, 58(2):119–142, October 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/modelling-and-simulation-of-volumetric-rainfall-for-
a-catchment-in-the-murraydarling-basin/7CF2686E8BAFDC484E0801A80E9CEEF7.

Debsarma:2016:CME

[1702] Suma Debsarma and K. P. Das. Current-modified evolution equation for a broader bandwidth capillary-gravity wave packet. The ANZIAM Journal, 58(2):143–161, October 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/currentmodified-evolution-
equation-for-a-broader-bandwidth-capillarygravity-wave-packet/
13D75EA3E90E8B531C45EEDD23879B59.

Ming:2016:OMV

[1703] Zhiqin Ming, Zhibin Liang, and Caibin Zhang. Optimal mean-variance reinsurance with common shock dependence. The ANZIAM Journal, 58 (2):162–181, October 2016. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/optimal-meanvariance-reinsurance-
with-common-shock-dependence/76C6E31EAF2A86EEC31CC63F294AB4DF.

Alghalith:2016:NNA

[1704] Moawia Alghalith. A note on a new approach to both price and volatility jumps: an application to the portfolio model. The ANZIAM Journal, 58(2):182–186, October 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
a-new-approach-to-both-price-and-volatility-jumps-an-application-
to-the-portfolio-model/0726360176C2D3F8AD1404198B0FAE67.

Chen:2016:ODM

[1705] Xingran Chen, Yongmei Liu, and Zhong Wan. Optimal decision making for online and offline retailers under BOPS mode. The ANZIAM Journal, 58(2):187–208, October 2016. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/optimal-
decision-making-for-online-and-offline-retailers-under-bops-
mode/41EFF70E037B2569EA59FDFF79A8448C.

Anonymous:2016:AVIg

[1706] Anonymous. ANZ volume 58 issue 2 cover and front matter. The ANZIAM Journal, 58(2):f1–f2, October 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-58-issue-2-cover-and-front-matter/E799A443809D6E24DC035F0BCB3B67F0.

Anonymous:2016:AVIh

[1707] Anonymous. ANZ volume 58 issue 2 cover and back matter. The ANZIAM Journal, 58(2):b1–b8, October 2016. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-58-issue-2-cover-and-back-matter/46A4DAAFF530CDE89A52FEEF3E4531F9.

Broadbridge:2017:EMM

[1708] Phil Broadbridge, Mark Nelson, and Justin Wang. Editorial: mathematical methods for applications. The ANZIAM Journal, 58(3–4):209–210, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/editorial-mathematical-methods-for-applications/
A77CDCE50BD95BDF139A648547D0ACDB.

Wang:2017:AMC

[1709] Sen Wang, Qingxiang Fang, and Jun-E Feng. An alternative method of concept learning. The ANZIAM Journal, 58(3–4):211–219, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/an-alternative-method-of-concept-learning/EA517FC0E7ECA68BD9E30BB8DCC44F92.

Dong:2017:LGN

[1710] Xue-Mei Dong. Learning gradients from nonidentical data. The ANZIAM Journal, 58(3–4):220–230, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/learning-gradients-
from-nonidentical-data/E636A1BA09561BD83D74155096709632.

Wang:2017:SED

[1711] Cheng Wang and Feilong Cao. A study on the error of distributed algorithms for big data classification with SVM. The ANZIAM Journal, 58(3–4):231–237, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/study-on-the-error-of-
distributed-algorithms-for-big-data-classification-with-svm/
11C64B2A284CA5609C8FE5F464BCCBFC.

Chen:2017:ASN

[1712] Zhixiang Chen and Feilong Cao. Approximation by spherical neural networks with zonal functions. The ANZIAM Journal, 58(3–4):238–246, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/approximation-by-spherical-neural-networks-
with-zonal-functions/AA46E7935230B8015F3FCB7D12F8313D.

Shen:2017:IIP

[1713] Yi Shen, Bin Han, and Elena Braverman. Image inpainting from partial noisy data by directional complex tight framelets. The ANZIAM Journal, 58(3–4):247–255, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/image-inpainting-from-
partial-noisy-data-by-directional-complex-tight-framelets/8F79ED004D4AE6C29616341B6932B00A.

Yang:2017:MCP

[1714] Jianwei Yang, Liang Zhang, and Zhengda Lu. The Mellin central projection transform. The ANZIAM Journal, 58(3–4):256–264, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/mellin-central-projection-transform/
0B5F6571DBFABE43E3DB855BA67E274E.

Fan:2017:CBF

[1715] Tai-He Fan and Meng-Ke Bian. Characterizations of the Borel σ-fields of the fuzzy number space. The ANZIAM Journal, 58 (3–4):265–275, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/characterizations-
of-the-borel-unicodestixx1d70e-fields-of-the-fuzzy-number-space/
43783641007DC4622F38D0DBB5D1D4A0.

Pei:2017:NGI

[1716] D. Pei and Y. Zhu. New generalized h-implications. The ANZIAM Journal, 58(3–4):276–286, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/new-generalized-h-
implications/1F62C4CC8ADBF9D498070930F0C1E8E8.

Chen:2017:CPC

[1717] Xiaolong Chen, Xiangbo Meng, Xiaoshi Song, and Chun Shan. Coverage performance of cognitive radio networks powered by renewable energy. The ANZIAM Journal, 58(3–4):287–305, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/coverage-performance-of-cognitive-radio-networks-powered-
by-renewable-energy/8CF660A98D1EEDFBDCE23A65D949A0C3.

Dong:2017:PMS

[1718] J. M. Dong, X. S. Wang, L. L. Wang, and J. L. Hu. Parallel machine scheduling with job delivery coordination. The ANZIAM Journal, 58(3– 4):306–313, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/parallel-machine-scheduling-
with-job-delivery-coordination/839A8684CFFF679E21CEADA5DB0567A1.

Jiang:2017:STP

[1719] Yiwei Jiang, Ping Zhou, Huijuan Wang, and Jueliang Hu. Scheduling on two parallel machines with two dedicated servers. The ANZIAM Journal, 58(3–4):314–323, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/scheduling-on-two-parallel-
machines-with-two-dedicated-servers/9EF4369719180DBB7F6819D04BF2B7D9.

Fang:2017:PSF

[1720] Q. Fang and J. Peng. Pinning synchronization of fractional-order complex networks by a single controller. The ANZIAM Journal, 58(3– 4):324–332, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/pinning-synchronization-of-
fractionalorder-complex-networks-by-a-single-controller/BFCADCEDC6BA2F80425ABB8106D0555B.

Wu:2017:CML

[1721] Longshu Wu, Qin Wang, and Xiaobing Yang. Computational methods for logistics problems related to optimal trees. The ANZIAM Journal, 58(3–4):333–341, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/computational-methods-
for-logistics-problems-related-to-optimal-trees/787FCB641A8DFEF5B042830F1B674263.

Wang:2017:NDM

[1722] Qin Wang and Longshu Wu. Network design for minimum spanning trees under Hamming distance. The ANZIAM Journal, 58(3–4):342– 349, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/network-design-for-minimum-spanning-
trees-under-hamming-distance/47A75FEB8221EBF3E145AD1C919198D5.

Li:2017:ISN

[1723] Fengwei Li, Qingfang Ye, and Xiaoyan Zhang. Isolated scattering number of split graphs and graph products. The ANZIAM Journal, 58(3–4): 350–358, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/isolated-scattering-number-of-split-
graphs-and-graph-products/685CC2E67A10191116EFD496DF3DF271.

Xia:2017:DPA

[1724] Zhinan Xia. Dynamics of pseudo almost periodic solution for impulsive neoclassical growth model. The ANZIAM Journal, 58(3–4):359–367, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/dynamics-of-pseudo-almost-periodic-solution-
for-impulsive-neoclassical-growth-model/361AEEE601B772F6BE9A8E3095A6005F.

Han:2017:CAI

[1725] Shuguang Han, Jueliang Hu, and Diwei Zhou. Competitive analysis of interrelated price online inventory problems with demands. The ANZIAM Journal, 58(3–4):368–378, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/competitive-analysis-
of-interrelated-price-online-inventory-problems-with-demands/
97E69FAD729E3CED384F8AFEF64426E3.

Ishimura:2017:CDP

[1726] N. Ishimura and N. Yoshida. On the convergence of discrete processes with multiple independent variables. The ANZIAM Journal, 58(3–4):379–385, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/on-the-convergence-of-
discrete-processes-with-multiple-independent-variables/77A8CFFA5ED7FA8B6709ABC18B927AA8.

Tao:2017:MAS

[1727] Xiangxing Tao and Yafeng Shi. On multi-asset spread option pricing in a Wick–Itô–Skorohod integral framework. The ANZIAM Journal, 58(3–4): 386–396, April 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-multiasset-spread-option-pricing-
in-a-wickitoskorohod-integral-framework/6E9DD756F5630B499B8401E5A5E30104.

Forbes:2017:MTT

[1728] Lawrence K. Forbes and Michael A. Brideson. On modelling the transition to turbulence in pipe flow. The ANZIAM Journal, 59(1):1–34, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/on-modelling-the-transition-to-turbulence-in-pipe-flow/
D9A14F2D76B0AC0734407CCF0AFDC793.

Fowkes:2017:TSB

[1729] Neville Fowkes, Steve Durkin, and Andrew P. Bassom. Truck safety barriers for mining sites. The ANZIAM Journal, 59(1):35–50, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/truck-safety-barriers-for-mining-sites/671389CD09068A15B5698CEF4922BA13.

Hogan:2017:CDN

[1730] A. B. Hogan, K. Glass, and R. S. Anderssen. Complex demodulation: a novel time series method for analysing seasonal infectious diseases. The ANZIAM Journal, 59(1):51–60, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/complex-
demodulation-a-novel-time-series-method-for-analysing-seasonal-
infectious-diseases/5126126085F54DDC4084C3874790DB57.

Le:2017:FEA

[1731] Kim Ngan Le, William Mclean, and Bishnu Lamichhane. Finite element approximation of a time-fractional diffusion problem for a domain with a re-entrant corner. The ANZIAM Journal, 59(1):61–82, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/finite-element-approximation-of-a-timefractional-diffusion-
problem-for-a-domain-with-a-reentrant-corner/94309FB03F7E0EB5AA55A2B9CB42B29C.

Cao:2017:AAV

[1732] Jian-Peng Cao and Yan-Bing Fang. An analytical approach for variance swaps with an Ornstein–Uhlenbeck process. The ANZIAM Journal, 59 (1):83–102, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/an-analytical-approach-for-variance-
swaps-with-an-ornsteinuhlenbeck-process/EE61B8C62B4240CD4A91B1E1DF29F2EF.

Chowdhury:2017:FOE

[1733] Dipankar Chowdhury and Suma Debsarma. Fifth-order evolution equation of gravity-capillary waves. The ANZIAM Journal, 59(1):103–114, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/fifthorder-evolution-equation-of-gravitycapillary-
waves/04CD919A3F2849312DAB40422FFE7897.

Mohamed:2017:IMO

[1734] K. Mohamed, A. Mehdi, and M. Abdelkader. An iterative model order reduction method for large-scale dynamical systems. The ANZIAM Journal, 59(1):115–133, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
iterative-model-order-reduction-method-for-largescale-dynamical-
systems/68C287CFF398A63A6E361BBDD56EABBE.

Anonymous:2017:AVIa

[1735] Anonymous. ANZ volume 59 issue 1 cover and front matter. The ANZIAM Journal, 59(1):f1–f2, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-59-issue-1-cover-and-front-matter/27CF37DBC292EB92017B85068B6188E3.

Anonymous:2017:AVIb

[1736] Anonymous. ANZ volume 59 issue 1 cover and back matter. The ANZIAM Journal, 59(1):b1–b5, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-59-issue-1-cover-and-back-matter/5CFBFC0B86A8E963852C33791BDAEDB6.

Nelson:2017:DD

[1737] Mark Ian Nelson, Peter Hagedoorn, and Annette L. Worthy. The demon drink. The ANZIAM Journal, 59(2):135–154, October 2017.

CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/demon-drink/AE458878FBA06C940487E152C3F99737.

Promrak:2017:PPM

[1738] J. Promrak, G. C. Wake, and C. Rattanakul. Predator–prey model with age structure. The ANZIAM Journal, 59(2):155–166, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/predatorprey-model-with-age-structure/78D80DEA683D4B2746404F6D7DF34318.

Alfifi:2017:SAS

[1739] H. Y. Alfifi. Semi-analytical solutions for the Brusselator reaction– diffusion model. The ANZIAM Journal, 59(2):167–182, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/semianalytical-solutions-for-the-brusselator-reactiondiffusion-
model/CFEFAC24AC7FA09A2AE9DC05112839F8.

Zeng:2017:PEO

[1740] X. C. Zeng, I. Guo, and S. P. Zhu. Pricing European options on regime-switching assets: a comparative study of Monte Carlo and finitedifference approaches. The ANZIAM Journal, 59(2):183–199, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/pricing-european-options-on-regimeswitching-assets-
a-comparative-study-of-monte-carlo-and-finitedifference-approaches/
5A15C82EE2006D03A339EB693BDC90E3.

Amore:2017:BSW

[1741] Paolo Amore, John P. Boyd, Francisco M. Fernández, Martin Jacobo, and Petr Zhevandrov. Bound states in weakly deformed waveguides: numerical versus analytical results. The ANZIAM Journal, 59(2):200– 214, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/bound-states-in-weakly-deformed-waveguides-
numerical-versus-analytical-results/32830311BA0089B13944399F5433589C.

Price:2017:DSQ

[1742] C. J. Price. A direct search quasi-Newton method for nonsmooth unconstrained optimization. The ANZIAM Journal, 59(2):215–231, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/direct-search-quasinewton-method-for-nonsmooth-unconstrained-
optimization/C8A76A8779EBAC861A533C6BAD8C21AA.

Lamichhane:2017:NMP

[1743] B. P. Lamichhane. A new minimization principle for the Poisson equation leading to a flexible finite element approach. The ANZIAM Journal, 59(2):232–239, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/new-
minimization-principle-for-the-poisson-equation-leading-to-
a-flexible-finite-element-approach/BD2F06FFD2624E6E33BF3BBDAEF63B5F.

Sumner:2017:MCM

[1744] Jeremy G. Sumner. Multiplicatively closed Markov models must form Lie algebras. The ANZIAM Journal, 59(2):240–246, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/multiplicatively-closed-markov-models-must-
form-lie-algebras/00D742D434F61A87C5243E268CB15E70.

Li:2017:BLP

[1745] Yinxue Li, Zhong Wan, and Jingjing Liu. Bi-level programming approach to optimal strategy for vendor-managed inventory problems under random demand. The ANZIAM Journal, 59(2):247–270, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/bilevel-programming-approach-to-optimal-strategy-for-
vendormanaged-inventory-problems-under-random-demand/3CE95BD0E415454C288E97197E012D5A.

Maioli:2017:NCI

[1746] D. S. Maioli, C. Lavor, and D. S. Gonçalves. A note on computing the intersection of spheres in R^{n}. The ANZIAM Journal, 59 (2):271–279, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/note-on-computing-the-
intersection-of-spheres-in-mathbbrn/D15FB22917024962409980AC7D3C086D.

Anonymous:2017:AVIc

[1747] Anonymous. ANZ volume 59 issue 2 cover and front matter. The ANZIAM Journal, 59(2):f1–f2, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-59-issue-2-cover-and-front-matter/ACF4857B6016DE2AE9E9E0A0A16A4FD8.

Anonymous:2017:AVId

[1748] Anonymous. ANZ volume 59 issue 2 cover and back matter. The ANZIAM Journal, 59(2):b1–b9, October 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-59-issue-2-cover-and-back-matter/E28B240B6FA4A2161C36DAC65913AA9F.

Farrow:2018:MHC

[1749] D. E. Farrow, G. C. Hocking, S. J. Cringle, and D.-Y. Yu. Modelling hydrogen clearance from the retina. The ANZIAM Journal, 59(3):281– 292, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/modelling-hydrogen-clearance-from-the-
retina/407830A245A6230820A2A69097F2FAEB.

VanBrunt:2018:CDE

[1750] B. Van Brunt, A. Almalki, T. Lynch, and A. Zaidi. On a cell division equation with a linear growth rate. The ANZIAM Journal, 59(3):293– 312, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/on-a-cell-division-equation-with-a-linear-
growth-rate/61D6649517D4B7BB36D4C149DA9BFD2D.

Alotaibi:2018:CMP

[1751] Hammad Alotaibi, Barry Cox, and A. J. Roberts. Couple microscale periodic patches to simulate macroscale emergent dynamics. The ANZIAM Journal, 59(3):313–334, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/couple-microscale-
periodic-patches-to-simulate-macroscale-emergent-dynamics/1456CC4A82566CCF3CE75284FC5DA32D.

Hill:2018:NNS

[1752] J. M. Hill and Y. M. Stokes. A note on Navier–Stokes equations with nonorthogonal coordinates. The ANZIAM Journal, 59(3):335– 348, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/note-on-navierstokes-equations-with-
nonorthogonal-coordinates/2172D844DB90D0C4511739EA8202C2A1.

Ke:2018:AAP

[1753] Ziwie Ke, Joanna Goard, and Song-Ping Zhu. An appropriate approach to pricing European-style options with the Adomian decomposition method. The ANZIAM Journal, 59(3):349–369, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
appropriate-approach-to-pricing-europeanstyle-options-with-
the-{Adomian}-decomposition-method/7AE29FFA1B6A1A84BCD6FD6F8DCC13F8.

Agrawal:2018:DRB

[1754] Rashmi Agrawal, Debaldev Jana, Ranjit Kumar Upadhyay, and V. Sree Hari Rao. Dynamic relationship between the mutual interference and gestation delays of a hybrid tritrophic food chain model. The ANZIAM Journal, 59(3):370–401, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/dynamic-
relationship-between-the-mutual-interference-and-gestation-
delays-of-a-hybrid-tritrophic-food-chain-model/ED78AB1220CF62AEF76E26D932B58447.

Luo:2018:CCP

[1755] Quanbing Luo, Dong Liang, Ting Ren, and Jian Zhang. Calculation of critical parameters for spontaneous combustion for some complex geometries using an indirect numerical method. The ANZIAM Journal, 59(3): 402–412, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/calculation-of-critical-parameters-
for-spontaneous-combustion-for-some-complex-geometries-using-
an-indirect-numerical-method/76752C3118003321394A8E73D043502F.

Anonymous:2018:AVIa

[1756] Anonymous. ANZ volume 59 issue 3 cover and front matter. The ANZIAM Journal, 59(3):f1–f2, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-59-issue-3-cover-and-front-matter/680DF0983A46EB085E2B028E875B58A8.

Anonymous:2018:AVIb

[1757] Anonymous. ANZ volume 59 issue 3 cover and back matter. The ANZIAM Journal, 59(3):b1–b5, January 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-59-issue-3-cover-and-back-matter/F8704037554F5747B527BDCD983355EC.

Dyson:2018:EMB

[1758] Rosemary J. Dyson and J. Edward F. Green. Editorial: mechanics in biology. The ANZIAM Journal, 59(4):413–415, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/editorial-mechanics-in-biology/C1FB11B04C3E3C54ACF64EB02E3D0FE0.

Smith:2018:BFM

[1759] David J. Smith. Biological fluid mechanics under the microscope: a tribute to John Blake. The ANZIAM Journal, 59(4): 416–442, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/biological-fluid-mechanics-
under-the-microscope-a-tribute-to-john-blake/58735BC04B1D9C456AE65C780A5A3372.

Gagnon:2018:TSS

[1760] D. A. Gagnon and T. D. Montenegro-Johnson. Thrifty swimming with shear-thinning: a note on out-of-plane effects for undulatory locomotion through shear-thinning fluids. The ANZIAM Journal, 59(4):443–454, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/thrifty-swimming-with-shearthinning-a-note-
on-outofplane-effects-for-undulatory-locomotion-through-shearthinning-
fluids/1866D73B7DC19CDB4DCBFDDB1DE21902.

Clarke:2018:PMS

[1761] R. J. Clarke. Photofocusing of microorganisms swimming in a flow with shear. The ANZIAM Journal, 59(4):455–471, April 2018. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/photofocusing-of-microorganisms-swimming-in-a-flow-
with-shear/F30EC9E99BBC5BB884029DFD6461C64E.

Copos:2018:PVM

[1762] Calina A. Copos and Robert D. Guy. A porous viscoelastic model for the cell cytoskeleton. The ANZIAM Journal, 59(4):472–498, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/porous-viscoelastic-model-for-the-cell-cytoskeleton/
6881609876803536915F877BF591918B.

Holden:2018:MMM

[1763] E. C. Holden, J. Collis, B. S. Brook, and R. D. O’Dea. A multiphase multiscale model for nutrient limited tissue growth. The ANZIAM Journal, 59(4):499–532, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/multiphase-multiscale-
model-for-nutrient-limited-tissue-growth/30C4ACA604627DED7FE469DEEBF01E85.

Ovenden:2018:NBF

[1764] N. C. Ovenden and F. T. Smith. Nonsymmetric branching of fluid flows in 3D vessels. The ANZIAM Journal, 59(4):533–561, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/nonsymmetric-branching-of-fluid-flows-in-3d-vessels/
01B00D8945B3A61B98A5FF028E8F6A57.

Clark:2018:TSH

[1765] A. R. Clark and M. H. Tawhai. Temporal and spatial heterogeneity in pulmonary perfusion: a mathematical model to predict interactions between macro- and micro-vessels in health and disease. The ANZIAM Journal, 59(4):562–580, April 2018. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/temporal-and-spatial-heterogeneity-in-pulmonary-perfusion-
a-mathematical-model-to-predict-interactions-between-macro-
and-microvessels-in-health-and-disease/B5DCA34FC17A9ABCCEEA76766F6AC73A.

Erlich:2018:MFS

[1766] A. Erlich, R. Howell, A. Goriely, R. Chirat, and D. E. Moulton. Mechanical feedback in seashell growth and form. The ANZIAM Journal, 59(4): 581–606, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/mechanical-feedback-in-seashell-growth-
and-form/0E57126EDB7CD3D6D3A4D0BCFEFF7D42.

Anonymous:2018:AVIc

[1767] Anonymous. ANZ volume 59 issue 4 cover and front matter. The ANZIAM Journal, 59(4):f1–f2, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-59-issue-4-cover-and-front-matter/A99FA79338C1FE5ABA00BFA88104E56F.

Anonymous:2018:AVId

[1768] Anonymous. ANZ volume 59 issue 4 cover and back matter. The ANZIAM Journal, 59(4):b1–b8, April 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-59-issue-4-cover-and-back-matter/C553B2CED8BFF195F9754D1564C42CA5.

Trefethen:2018:SSL

[1769] Lloyd N. Trefethen. Series solution of Laplace problems. The ANZIAM Journal, 60(1):1–26, July 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/series-solution-of-laplace-
problems/1C03350BF99653D580BD4AD1E0048BEF.

Dipierro:2018:PIP

[1770] Serena Dipierro, Luca Lombardini, Pietro Miraglio, and Enrico Valdinoci. The Phillip Island penguin parade (a mathematical treatment). The ANZIAM Journal, 60(1):27–54, July 2018. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/phillip-island-penguin-parade-a-mathematical-treatment/
80990E6616EAC880239DFD0C174CABB2.

Falletta:2018:SST

[1771] J. Falletta and S. Woodcock. A simulation study of Texas hold ’em poker: what Taylor Swift understands and James Bond doesn’t. The ANZIAM Journal, 60(1):55–64, July 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/simulation-study-of-texas-
hold-em-poker-what-taylor-swift-understands-and-james-bond-
doesnt/CB91B0B6D5964E647D383749F1901DD9.

Mallier:2018:IEF

[1772] R. Mallier and J. Goard. Integral equation formulation for shout options. The ANZIAM Journal, 60(1):65–85, July 2018. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/integral-equation-formulation-for-shout-options/3A2C1D5DDF9C925779E144EFE11CDE3E.

Athanasiadis:2018:RRC

[1773] C. E. Athanasiadis, E. S. Athanasiadou, and S. Dimitroula. Reciprocity relations for a conductive scatterer with a chiral core in quasi-static form. The ANZIAM Journal, 60(1):86–94, July 2018.

CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/reciprocity-relations-for-a-conductive-scatterer-with-
a-chiral-core-in-quasistatic-form/74201A71D0C25BA6DE6FA80B51909774.

Li:2018:CRA

[1774] L. Li, G. Q. Wang, and J. L. Zhang. On the o(1/k) convergence rate of the alternating direction method of multipliers in a complex domain. The ANZIAM Journal, 60(1):95–117, July 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-o1k-convergence-rate-of-the-alternating-direction-method-
of-multipliers-in-a-complex-domain/1C4C332C69432C49E88AB6532A213ADD.

Kundu:2018:EUW

[1775] Sumana Kundu, Suma Debsarma, and K. P. Das. Effect of uniform wind flow on modulational instability of two crossing waves over finite depth water. The ANZIAM Journal, 60(1):118–136, July 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/effect-of-uniform-wind-flow-on-modulational-instability-
of-two-crossing-waves-over-finite-depth-water/0E6C555729D996D482DDD205450F3DC1.

Anonymous:2018:AVIe

[1776] Anonymous. ANZ volume 60 issue 1 cover and front matter. The ANZIAM Journal, 60(1):f1–f2, July 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-60-issue-1-cover-and-front-matter/3E96ED20AC390566A307A5935E5F0026.

Anonymous:2018:AVIf

[1777] Anonymous. ANZ volume 60 issue 1 cover and back matter. The ANZIAM Journal, 60(1):b1–b10, July 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-60-issue-1-cover-and-back-matter/AFAA9E883525394A9894DD97598E1CCE.

Howlett:2018:TTS

[1778] Phil Howlett. The two-train separation problem on level track with discrete control. The ANZIAM Journal, 60(2):137–174, October 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/twotrain-separation-problem-on-level-track-
with-discrete-control/9CD8DFFF22309E08C9CE30EB2BF7B89F.

Chen:2018:OPU

[1779] Wenting Chen and Sha Lin. Option pricing under the KOBOL model. The ANZIAM Journal, 60(2):175–190, October 2018. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/option-pricing-under-the-kobol-model/63FD2E1571C12D7BDBAAFCB8F2E8541B.

Nelson:2018:MMR

[1780] Mark I. Nelson, Rubayyi T. Alqahtani, and Faisal I. Hai. Mathematical modelling of the removal of organic micropollutants in the activated sludge process: a linear biodegradation model. The ANZIAM Journal, 60 (2):191–229, October 2018. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/mathematical-modelling-of-the-
removal-of-organic-micropollutants-in-the-activated-sludge-
process-a-linear-biodegradation-model/C274E72F852E8ABB0CD7E65321439AEB.

Zhan:2018:CDS

[1781] T. Zhan and S. P. Ma. The controller design for singular fractionalorder systems with fractional order 0 < α < 1. The ANZIAM Journal, 60(2):230–248, October 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/controller-design-for-
singular-fractionalorder-systems-with-fractional-order-0-1/
3537D8FFFFAA47033C0284DD8E5A9278.

Panda:2018:FSD

[1782] S. Panda, K. K. Patra, and M. Sellier. Free-surface dynamics of thin second-grade fluid over an unsteady stretching sheet. The ANZIAM Journal, 60(2):249–268, October 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/freesurface-dynamics-of-
thin-secondgrade-fluid-over-an-unsteady-stretching-sheet/DFF0783A13FF36AFEA7E7351C05FD84A.

Anonymous:2018:AVIg

[1783] Anonymous. ANZ volume 60 issue 2 cover and front matter. The ANZIAM Journal, 60(2):f1–f2, October 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-60-issue-2-cover-and-front-matter/2A732FA93235579BAAE1238D650DEA2E.

Anonymous:2018:AVIh

[1784] Anonymous. ANZ volume 60 issue 2 cover and back matter. The ANZIAM Journal, 60(2):b1–b6, October 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-60-issue-2-cover-and-back-matter/D0CAE233CE901D503B94565B76DE8B22.

Harding:2019:RRM

[1785] Brendan Harding. A Rayleigh–Ritz method for Navier–Stokes flow through curved ducts. The ANZIAM Journal, 61(1):1–22, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/rayleighritz-method-for-navierstokes-flow-through-curved-
ducts/CD4D056C19BE322436F4C8CFB6AFDBD2.

Lamichhane:2019:APA

[1786] Bishnu P. Lamichhane, Scott B. Lindstrom, and Brailey Sims. Application of projection algorithms to differential equations: boundary value problems. The ANZIAM Journal, 61(1):23–46, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/application-of-projection-algorithms-to-differential-
equations-boundary-value-problems/71166EBBA4240D8C170F467644C8348B.

Sivanesan:2019:WWS

[1787] M. Sivanesan and S. R. Manam. Water wave scattering by a vertical porous barrier with two gaps. The ANZIAM Journal, 61(1):47–63, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/water-wave-scattering-by-a-vertical-porous-
barrier-with-two-gaps/C5F69BD5DCB5547FD5DA114538DAA333.

Charkhgard:2019:NAS

[1788] Hadi Charkhgard and Ali Eshragh. A new approach to select the best subset of predictors in linear regression modelling: bi-objective mixed integer linear programming. The ANZIAM Journal, 61(1): 64–75, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/new-approach-to-select-the-
best-subset-of-predictors-in-linear-regression-modelling-biobjective-
mixed-integer-linear-programming/A48905B2999A024AD834BEE1D7374EFE.

Li:2019:NAB

[1789] Ting Li and Zhong Wan. New adaptive Barzilai–Borwein step size and its application in solving large-scale optimization problems. The ANZIAM Journal, 61(1):76–98, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/new-
adaptive-barzilaiborwein-step-size-and-its-application-in-solving-
largescale-optimization-problems/ACD9681C2B8C9F20A84A62658C3DEAB8.

Li:2019:OIC

[1790] L. Li and H. Mi. Optimal investment and consumption with stochastic factor and delay. The ANZIAM Journal, 61(1):99–117, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-investment-and-consumption-with-stochastic-
factor-and-delay/EA9A06171661717EDB5BE5189C6EB8B2.

Anonymous:2019:AVIa

[1791] Anonymous. ANZ volume 61 issue 1 cover and front matter. The ANZIAM Journal, 61(1):f1–f2, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-61-issue-1-cover-and-front-matter/2821A0AF77D9C100364EEFC45D9C504D.

Anonymous:2019:AVIb

[1792] Anonymous. ANZ volume 61 issue 1 cover and back matter. The ANZIAM Journal, 61(1):b1–b5, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-61-issue-1-cover-and-back-matter/3385C7C78832F954DF81A39E3C878E5B.

Allwright:2019:APV

[1793] Emma J. Allwright, L. K. Forbes, and S. J. Walters. Axisymmetric plumes in viscous fluids. The ANZIAM Journal, 61(2):119–147, April 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/axisymmetric-plumes-in-viscous-fluids/BC1CC19B224CCC99A5434B5575DBA8A5.

Sandor:2019:TDV

[1794] Balázs Sándor, Péter Torma, K. Gábor Szabó, and Hong Zhang. On the topography-driven vorticity production in shallow lakes. The ANZIAM Journal, 61(2):148–160, April 2019. CODEN AJNOA2.

ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-topographydriven-vorticity-production-in-shallow-lakes/
1F076DE4E75772ED1DDFFE9E3A6B65B8.

Amaral:2019:PIL

[1795] L. R. Amaral and A. Papanicolaou. Price impact of large orders using Hawkes processes. The ANZIAM Journal, 61(2):161–194, April 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/price-impact-of-large-orders-using-hawkes-processes/
9E5FEA23A880A5FEB9452E7A74A0DAF0.

Aminifard:2019:MAD

[1796] Z. Aminifard and S. Babaie-Kafaki. Matrix analyses on the Dai–Liao conjugate gradient method. The ANZIAM Journal, 61(2):195–203, April 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/matrix-analyses-on-the-dailiao-conjugate-gradient-method/
5F56CBAEDC828D7DCAA8B107CAE4FC61.

Cui:2019:NLC

[1797] Jin Cui, Wenjun Cai, Chaolong Jiang, and Yushun Wang. A new linear and conservative finite difference scheme for the Gross–Pitaevskii equation with angular momentum rotation. The ANZIAM Journal, 61(2): 204–232, April 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/new-linear-and-conservative-finite-difference-
scheme-for-the-grosspitaevskii-equation-with-angular-momentum-
rotation/AE80FCDB63D973FE07D64C37CF341F3B.

Debsarma:2019:EPR

[1798] S. Debsarma and D. Chowdhury. Evolution of a pair of random inhomogeneous wave systems over infinite-depth water. The ANZIAM Journal, 61(2):233–247, April 2019. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/evolution-of-a-pair-of-random-
inhomogeneous-wave-systems-over-infinitedepth-water/A43C3BF07E2F951E8F8B42C2F3DC1EC4.

Anonymous:2019:AVIc

[1799] Anonymous. ANZ volume 61 issue 2 cover and front matter. The ANZIAM Journal, 61(2):f1–f2, April 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-61-issue-2-cover-and-front-matter/B24F8537E2CA77E907967C90150A5BDA.

Anonymous:2019:AVId

[1800] Anonymous. ANZ volume 61 issue 2 cover and back matter. The ANZIAM Journal, 61(2):b1–b6, April 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-61-issue-2-cover-and-back-matter/CC9E79DFA828B94356417E200F69E582.

Al-Ali:2019:CSC

[1801] S. Al-Ali, G. C. Hocking, and D. E. Farrow. Critical surface coning due to a line sink in a vertical drain containing a porous medium. The ANZIAM Journal, 61(3):249–269, July 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/critical-surface-coning-due-to-a-line-sink-
in-a-vertical-drain-containing-a-porous-medium/D19F0F9BC4457FFC1D9DEC8AF2567C8F.

Edson:2019:LEK

[1802] Russell A. Edson, J. E. Bunder, Trent W. Mattner, and A. J. Roberts. Lyapunov exponents of the Kuramoto–Sivashinsky PDE. The ANZIAM Journal, 61(3):270–285, July 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/lyapunov-exponents-of-
the-kuramotosivashinsky-pde/FE2E4CBD4AAFC9BE25C7453AD6FB2575.

Walters:2019:FRT

[1803] S. J. Walters and L. K. Forbes. Fully 3D Rayleigh–Taylor instability in a Boussinesq fluid. The ANZIAM Journal, 61(3):286–304, July 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/fully-{$3$d}-rayleightaylor-instability-in-
a-boussinesq-fluid/A2161B6272EF7E66C101E50158F717FE.

Danet:2019:EUW

[1804] Cristian-Paul Danet. Existence and uniqueness of weak and classical solutions for a fourth-order semilinear boundary value problem. The ANZIAM Journal, 61(3):305–319, July 2019. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/existence-and-uniqueness-of-weak-and-classical-solutions-
for-a-fourthorder-semilinear-boundary-value-problem/54433AF8FFA94835AE6F3B9CD544DF3D.

Stewart:2019:SEO

[1805] Peter S. Stewart and Alexander J. E. Foss. Self-excited oscillations in a collapsible channel with applications to retinal venous pulsation. The ANZIAM Journal, 61(3):320–348, July 2019. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/selfexcited-oscillations-in-a-collapsible-channel-with-
applications-to-retinal-venous-pulsation/037DD41817DB06D6863BB4609ABEA148.

Anonymous:2019:AVIe

[1806] Anonymous. ANZ volume 61 issue 3 cover and front matter. The ANZIAM Journal, 61(3):f1–f2, July 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-61-issue-3-cover-and-front-matter/573DE2DDD9E0AFBB6013534E8300A2F8.

Anonymous:2019:AVIf

[1807] Anonymous. ANZ volume 61 issue 3 cover and back matter. The ANZIAM Journal, 61(3):b1–b6, July 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-61-issue-3-cover-and-back-matter/6400CCD53B9E821740CBCF2F03D3A302.

Fackrell:2019:VCC

[1808] M. Fackrell, C. Li, P. G. Taylor, and J. Wang. The value of communication and cooperation when servers are strategic. The ANZIAM Journal, 61(4):349–367, October 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/value-of-communication-
and-cooperation-when-servers-are-strategic/06A7BF65F871B219915B13D0C8C42D5F.

Holden:2019:MMM

[1809] E. C. Holden, S. J. Chapman, B. S. Brook, and R. D. O’Dea. A multiphase multiscale model for nutrient-limited tissue growth, Part II: a simplified description. The ANZIAM Journal, 61(4): 368–381, October 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/multiphase-multiscale-
model-for-nutrientlimited-tissue-growth-part-ii-a-simplified-
description/CA0DA4642716CBAD894AA4D715527243.

Ibrahim:2019:PHE

[1810] S. N. I. Ibrahim, A. DÍAz-HernÁNdez, J. G. O’Hara, and N. Constantinou. Pricing holder-extendable call options with mean-reverting stochastic volatility. The ANZIAM Journal, 61(4):382–397, October 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/pricing-holderextendable-call-options-with-
meanreverting-stochastic-volatility/44F8FB787C2FC522430403AB9A740737.

Mungkasi:2019:NEP

[1811] Sudi Mungkasi and Stephen Gwyn Roberts. Numerical entropy production as smoothness indicator for shallow water equations. The ANZIAM Journal, 61(4):398–415, October 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/numerical-entropy-
production-as-smoothness-indicator-for-shallow-water-equations/
0DAEC9C0C58886C9E3E1FDB8B80196A5.

Loy:2019:ACM

[1812] R. J. Loy and R. S. Anderssen. Approximation of and by completely monotone functions. The ANZIAM Journal, 61(4):416–430, October 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/approximation-of-and-by-completely-monotone-functions/
980C8FD22BFA3CC721824347C5D7538E.

He:2019:SAP

[1813] Xin-Jiang He and Sha Lin. A semi-analytical pricing formula for European options under the rough Heston–CIR model. The ANZIAM Journal, 61(4):431–445, October 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/semianalytical-pricing-
formula-for-european-options-under-the-rough-hestoncir-model/
8A34440AA057C9820F2FE61FC32AB47F.

Anonymous:2019:AVIg

[1814] Anonymous. ANZ volume 61 issue 4 cover and front matter. The ANZIAM Journal, 61(4):f1–f2, October 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-61-issue-4-cover-and-front-matter/E4D8B9205BA25A3FD53D6B9DC4F0362E.

Anonymous:2019:AVIh

[1815] Anonymous. ANZ volume 61 issue 4 cover and back matter. The ANZIAM Journal, 61(4):b1–b9, October 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-61-issue-4-cover-and-back-matter/EF07859765D2260461E784486A48545E.

Bassom:2020:E

[1816] Andrew Bassom and Graeme Hocking. Editorial. The ANZIAM Journal, 62(1):1–2, January 2020. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/editorial/E374D3469B5CD65958F7FCE7A13E3202.

Simpson:2020:CLS

[1817] Matthew J. Simpson. Critical length for the spreading-vanishing dichotomy in higher dimensions. The ANZIAM Journal, 62(1):3–17, January 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/critical-length-for-the-spreadingvanishing-
dichotomy-in-higher-dimensions/7227448793C266052066E39AF7204E60.

Nolan:2020:SSM

[1818] Tui H. Nolan and Matt P. Wand. Streamlined solutions to multilevel sparse matrix problems. The ANZIAM Journal, 62(1):18–41, January 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/streamlined-solutions-to-multilevel-sparse-
matrix-problems/606582E7E1A4E0A60DAD9971B7EA3874.

Wang:2020:MIF

[1819] Q. Wang and Z. Zhang. A modified immersed finite volume element method for elliptic interface problems. The ANZIAM Journal, 62(1):42–61, January 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/modified-immersed-
finite-volume-element-method-for-elliptic-interface-problems/
8CC3038E5EC93F65EC398224F55B8C88.

Ouyang:2020:GMO

[1820] Ying Ouyang, Zhaoman Wan, and Zhong Wan. Game model for online and offline retailers under buy-online and pick-up-in-store mode with delivery cost and random demand. The ANZIAM Journal, 62(1):62–88, January 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/game-model-for-online-
and-offline-retailers-under-buyonline-and-pickupinstore-mode-
with-delivery-cost-and-random-demand/8771B83D8BE31517B9132C77813765AC.

Panda:2020:SCM

[1821] Gopinath Panda and Veena Goswami. Strategic customers in Markovian queues with vacations and synchronized abandonment. The ANZIAM Journal, 62(1):89–120, January 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/strategic-customers-
in-markovian-queues-with-vacations-and-synchronized-abandonment/
9E46BC821BC8623451EEC5FB0ED3E55F.

Anonymous:2020:AVIa

[1822] Anonymous. ANZ volume 62 issue 1 cover and front matter. The ANZIAM Journal, 62(1):f1–f2, January 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-62-issue-1-cover-and-front-matter/F4E1ADC9FAE41DB0BA92B6EE44D3CC11.

Anonymous:2020:AVIb

[1823] Anonymous. ANZ volume 62 issue 1 cover and back matter. The ANZIAM Journal, 62(1):b1–b6, January 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-62-issue-1-cover-and-back-matter/0332E82D021168043864160016AFB045.

Mclean:2020:IHO

[1824] William Mclean. Implementation of high-order, discontinuous Galerkin time stepping for fractional diffusion problems. The ANZIAM Journal, 62(2):121–147, April 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/implementation-of-
highorder-discontinuous-galerkin-time-stepping-for-fractional-
diffusion-problems/5D406DB07FE4C86968EBCE34E4E9ACB4.

Albrecht:2020:EDD

[1825] Amie Albrecht, Konstantin Avrachenkov, Phil Howlett, and Geetika Verma. Evolutionary dynamics in discrete time for the perturbed positive definite replicator equation. The ANZIAM Journal, 62(2):148–184, April 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/evolutionary-dynamics-in-discrete-time-for-
the-perturbed-positive-definite-replicator-equation/574E7DD657BCDD6F3C2C6990AC66F234.

Li:2020:ABS

[1826] Haijiao Li and Kuan Yang. Asymptotic behaviour of the stochastic Maki–Thompson model with a forgetting mechanism on open populations. The ANZIAM Journal, 62(2):185–208, April 2020. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-behaviour-of-the-stochastic-makithompson-
model-with-a-forgetting-mechanism-on-open-populations/6C7FF1B0AA6D11C9B7BFCF6C1AE42913.

Chiu:2020:MVE

[1827] Mei Choi Chiu. Mean-variance equilibrium asset-liability management strategy with cointegrated assets. The ANZIAM Journal, 62(2):209–234, April 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/meanvariance-equilibrium-assetliability-management-
strategy-with-cointegrated-assets/0F682918F2C821E1E564215F6EE7298B.

Anonymous:2020:AVIc

[1828] Anonymous. ANZ volume 62 issue 2 cover and front matter. The ANZIAM Journal, 62(2):f1–f2, April 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-62-issue-2-cover-and-front-matter/DAB5459D92E16EAB547C8A985EB860B4.

Anonymous:2020:AVId

[1829] Anonymous. ANZ volume 62 issue 2 cover and back matter. The ANZIAM Journal, 62(2):b1–b8, April 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-62-issue-2-cover-and-back-matter/714F6196E8183090AC48AA9C82584FAC.

Marsland:2020:CIR

[1830] S. Marsland, R. I. Mclachlan, and M. Y. Tufail. Conformal image registration based on constrained optimization. The ANZIAM Journal, 62(3): 235–255, July 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/conformal-image-registration-based-on-
constrained-optimization/AD2E8AFB8BD0ADE3BBFADCF93B17ACD8.

Yan:2020:LIE

[1831] J. L. Yan, L. H. Zheng, L. Zhu, and F. Q. Lu. Linearly implicit energy-preserving Fourier pseudospectral schemes for the complex modified Korteweg–de Vries equation. The ANZIAM Journal, 62(3):256–273, July 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/linearly-implicit-energypreserving-fourier-
pseudospectral-schemes-for-the-complex-modified-kortewegde-
vries-equation/F7B85D9D262EE5F79B4FD1782E4B0A47.

Howlett:2020:OLF

[1832] Phil Howlett and Anatoli Torokhti. An optimal linear filter for estimation of random functions in Hilbert space. The ANZIAM Journal, 62(3):274–301, July 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/an-optimal-linear-
filter-for-estimation-of-random-functions-in-hilbert-space/
0D30C24122356AA505DA67EFDE899F96.

Ivanitskiy:2020:PRM

[1833] A. Y. Ivanitskiy, V. V. Ejov, and F. P. Vasilyev. Pointwise residual method for solving primal and dual ill-posed linear programming problems with approximate data. The ANZIAM Journal, 62(3):302–317, July 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/pointwise-residual-method-for-solving-primal-and-dual-
illposed-linear-programming-problems-with-approximate-data/
A54583A935C36F188568EC3907235932.

Zulkarnaen:2020:MHC

[1834] Diny Zulkarnaen and Marianito R. Rodrigo. Modelling human carrying capacity as a function of food availability. The ANZIAM Journal, 62(3): 318–333, July 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/modelling-human-carrying-capacity-as-
a-function-of-food-availability/378183C3E8DBBDE947C8AEDDBD949C50.

Sirinanda:2020:OLU

[1835] K. G. Sirinanda, M. Brazil, P. A. Grossman, J. H. Rubinstein, and D. A. Thomas. Optimal location of an underground connector using discounted Steiner tree theory. The ANZIAM Journal, 62(3):334–351, July 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-location-of-an-underground-connector-using-
discounted-steiner-tree-theory/C986F6AC6D670DB0DEA105D3346900E1.

Anonymous:2020:AVIe

[1836] Anonymous. ANZ volume 62 issue 3 cover and front matter. The ANZIAM Journal, 62(3):f1–f2, July 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-62-issue-3-cover-and-front-matter/A2908BC01F9882E1EF8D3EB1DCD8D52D.

Anonymous:2020:AVIf

[1837] Anonymous. ANZ volume 62 issue 3 cover and back matter. The ANZIAM Journal, 62(3):b1–b5, July 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-62-issue-3-cover-and-back-matter/BD41C1754D8239816EBD6406ACF85A21.

Nelson:2020:ESI

[1838] Mark I. Nelson and Rodney O. Weber. Editorial: special issue celebrating the achievements of Professor G. C. Wake. The ANZIAM Journal, 62(4):353–354, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/editorial-special-
issue-celebrating-the-achievements-of-professor-g-c-wake/86EE57970295AD29746E3946E50BDCC5.

Howell:2020:DWS

[1839] P. D. Howell, H. Ockendon, and J. R. Ockendon. Draping woven sheets. The ANZIAM Journal, 62(4):355–385, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/draping-
woven-sheets/199B59F4E830E1CBF976C993B4CF2FF3.

Weir:2020:MFA

[1840] Graham Weir, George Chisholm, and Jerome Leveneur. The magnetic field about a three-dimensional block neodymium magnet. The ANZIAM Journal, 62(4):386–405, October 2020. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/magnetic-field-about-a-threedimensional-block-neodymium-
magnet/2394AA69DE851BFFA4C06234C26A6404.

Fradkin:2020:AKA

[1841] L. J. Fradkin, A. K. Djakou, C. Prior, M. Darmon, S. Chatillon, and P.-F. Calmon. The alternative Kirchhoff approximation in elastodynamics with applications in ultrasonic nondestructive testing. The ANZIAM Journal, 62(4):406–422, October 2020. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/alternative-kirchhoff-approximation-in-elastodynamics-
with-applications-in-ultrasonic-nondestructive-testing/7615EFE657DE2942E0320E5A4796C4E1.

Watt:2020:ODC

[1842] S. D. Watt, Z. Huang, H. S. Sidhu, A. C. Mcintosh, and J. Brindley. One-dimensional chaotic laminar flow with competitive exothermic and endothermic reactions. The ANZIAM Journal, 62(4):423– 445, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/onedimensional-chaotic-laminar-flow-
with-competitive-exothermic-and-endothermic-reactions/818B8B08FB15070E787127EC8F600462.

Harper:2020:AGH

[1843] J. F. Harper. Asymptotics of a Gauss hypergeometric function with two large parameters: a new case. The ANZIAM Journal, 62(4):446–452, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/asymptotics-of-a-gauss-hypergeometric-function-
with-two-large-parameters-a-new-case/32B40986E7DB85F500FC9024F846E527.

Korobeinikov:2020:PES

[1844] Andrei Korobeinikov, Elena Shchepakina, and Vladimir Sobolev. The paradox of enrichment, spatial heterogeneity, community effects and the phenomenon of apparent disappearance in the marine bacteriophage dynamics. The ANZIAM Journal, 62(4):453–468, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/paradox-of-enrichment-spatial-heterogeneity-community-
effects-and-the-phenomenon-of-apparent-disappearance-in-the-
marine-bacteriophage-dynamics/C60CED0872FFEDF84C3D1EBBDFFB6491.

Taylor:2020:EAS

[1845] Stephen Taylor and Xueshan Yang. Estimates for approximate solutions to a functional differential equation model of cell division. The ANZIAM Journal, 62(4):469–488, October 2020. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/estimates-for-approximate-solutions-to-a-functional-
differential-equation-model-of-cell-division/47C9AAE64986F5FD482BBD21A432A985.

Mohsin:2020:EUS

[1846] M. Mohsin and A. A. Zaidi. On existence and uniqueness of solutions to a pantograph type equation. The ANZIAM Journal, 62(4):489–512, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-existence-and-uniqueness-of-solutions-to-
a-pantograph-type-equation/85EE1600721A8628F03A40FC57FF7D1C.

Anonymous:2020:AVIg

[1847] Anonymous. ANZ volume 62 issue 4 cover and front matter. The ANZIAM Journal, 62(4):f1–f2, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-62-issue-4-cover-and-front-matter/91DE37685D7E86E9915DA51F73AC7134.

Anonymous:2020:AVIh

[1848] Anonymous. ANZ volume 62 issue 4 cover and back matter. The ANZIAM Journal, 62(4):b1–b5, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-62-issue-4-cover-and-back-matter/0EE2B22A499D23F0D47E444020A72857.

Iyaniwura:2021:AAM

[1849] S. Iyaniwura and M. J. Ward. Asymptotic analysis for the mean first passage time in finite or spatially periodic 2d domains with a cluster of small traps. The ANZIAM Journal, 63(1):1–22, January 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymptotic-analysis-for-the-mean-first-passage-time-
in-finite-or-spatially-periodic-2d-domains-with-a-cluster-of-
small-traps/72BBBC71722360D3FC702DBC18BE36A6.

Galloway:2021:SBI

[1850] David Galloway and David Ivers. Slow-burning instabilities of Dufort– Frankel finite differencing. The ANZIAM Journal, 63(1):23–38, January 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/slowburning-instabilities-of-dufortfrankel-
finite-differencing/A7505AC9995FBC851797C42D32467B09.

Li:2021:SMP

[1851] Zuoxun Li and Kai Zhang. Stochastic model predictive control for spacecraft rendezvous and docking via a distributionally robust optimization approach. The ANZIAM Journal, 63(1):39–57, January 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/stochastic-model-predictive-control-for-spacecraft-
rendezvous-and-docking-via-a-distributionally-robust-optimization-
approach/E9C193C1D3A86997F209F77B1ADB5198.

Akers:2021:ECC

[1852] B. F. Akers and D. M. Ambrose. Efficient computation of coordinatefree models of flame fronts. The ANZIAM Journal, 63(1):58–69, January 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/efficient-computation-of-coordinatefree-models-of-flame-
fronts/D591F8EDB0A4407797D62592C7E0744F.

Zaidi:2021:ACD

[1853] A. A. Zaidi and B. Van Brunt. Asymmetrical cell division with exponential growth. The ANZIAM Journal, 63(1):70–83, January 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/asymmetrical-cell-division-with-exponential-growth/
1BDA256E08C5AB609DAD964F11B862C8.

Maulana:2021:ACT

[1854] A. S. Maulana and S. R. Pudjaprasetya. Analysis of cell transmission model for traffic flow simulation with application to network traffic. The ANZIAM Journal, 63(1):84–99, January 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/analysis-of-cell-transmission-model-for-traffic-flow-
simulation-with-application-to-network-traffic/5BBE7B04F3F8FC95A38C94D815F03316.

Anonymous:2021:AVIa

[1855] Anonymous. Anz volume 63 issue 1 cover and front matter. The ANZIAM Journal, 63(1):f1–f2, January 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-63-issue-1-
cover-and-front-matter/126A1F4D9227FD193E1204A7B318F006.

Anonymous:2021:AVIb

[1856] Anonymous. Anz volume 63 issue 1 cover and back matter. The ANZIAM Journal, 63(1):b1–b6, January 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-63-issue-1-
cover-and-back-matter/E1833CD91477045782B07536E2C091E8.

Zhu:2021:ESI

[1857] Song-Ping Zhu, Xiaoping Lu, and Xin-Jiang He. Editorial: special issue on financial mathematics and quantitative finance. The ANZIAM Journal, 63(2):101–103, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.

org/core/journals/anziam-journal/article/editorial-special-
issue-on-financial-mathematics-and-quantitative-finance/3EEE4C70061C46CA7FCC41B8572BEDD9.

Fukasawa:2021:OTP

[1858] Masaaki Fukasawa, Hitomi Maeda, and Jun Sekine. On optimal thresholds for pairs trading in a one-dimensional diffusion model. The ANZIAM Journal, 63(2):104–122, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/on-optimal-thresholds-
for-pairs-trading-in-a-onedimensional-diffusion-model/9D74C2412859519FE8009615C64C7A4F.

Han:2021:OPU

[1859] Y. Han, Z. Li, and C. Liu. Option pricing under the fractional stochastic volatility model. The ANZIAM Journal, 63(2):123–142, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/option-pricing-under-the-fractional-
stochastic-volatility-model/B13092728199907AD5BDEADC3CD6D51E.

He:2021:AAF

[1860] Xin-Jiang He and Sha Lin. An analytical approximation formula for the pricing of credit default swaps with regime switching. The ANZIAM Journal, 63(2):143–162, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
analytical-approximation-formula-for-the-pricing-of-credit-
default-swaps-with-regime-switching/0B310F5F74EC89DCD99ADF5A92EAFFEE.

Lu:2021:FMA

[1861] Xiaoping Lu and Endah R. M. Putri. Finite maturity American-style stock loans with regime-switching volatility. The ANZIAM Journal, 63 (2):163–177, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/finite-maturity-americanstyle-
stock-loans-with-regimeswitching-volatility/EF4A5D9CA9D67371DF580C209F6B8D84.

Nonsoong:2021:AOP

[1862] P. Nonsoong, K. Mekchay, and S. Rujivan. An analytical option pricing formula for mean-reverting asset with time-dependent parameter. The ANZIAM Journal, 63(2):178–202, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
analytical-option-pricing-formula-for-meanreverting-asset-with-
timedependent-parameter/A7B34779368554FED1F62D635786661A.

Thakoor:2021:LRB

[1863] N. Thakoor. Localized radial basis functions for no-arbitrage pricing of options under stochastic alpha-beta-rho dynamics. The ANZIAM Journal, 63(2):203–227, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/localized-radial-basis-
functions-for-noarbitrage-pricing-of-options-under-stochastic-
alphabetarho-dynamics/5223363C7BAF659E91BAA3B5D9F755CF.

Tour:2021:SAO

[1864] Geraldine Tour, Nawdha Thakoor, and DÉSirÉ Yannick Tangman. Spectrally accurate option pricing under the time-fractional Black– Scholes model. The ANZIAM Journal, 63(2):228–248, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/spectrally-accurate-option-pricing-under-the-timefractional-
blackscholes-model/7967251C82E48FA7241BDCA26B7B1FC7.

Wang:2021:PTO

[1865] Xuhui Wang, Sheng-Jhih Wu, and Xingye Yue. Pricing timer options: second-order multiscale stochastic volatility asymptotics. The ANZIAM Journal, 63(2):249–267, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/pricing-
timer-options-secondorder-multiscale-stochastic-volatility-
asymptotics/45739BD15064BF6C3DD794783E3BA553.

Anonymous:2021:AVIc

[1866] Anonymous. Anz volume 63 issue 2 cover and front matter. The ANZIAM Journal, 63(2):f1–f2, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-63-issue-2-
cover-and-front-matter/2DF96193A4C277C20F397844D1D9C4AB.

Anonymous:2021:AVId

[1867] Anonymous. Anz volume 63 issue 2 cover and back matter. The ANZIAM Journal, 63(2):b1–b6, April 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-63-issue-2-
cover-and-back-matter/8EB3472F7E10F7EE54C95A3610218809.

Morrow:2021:ROP

[1868] Liam C. Morrow, Timothy J. Moroney, Michael C. Dallaston, and Scott W. Mccue. A review of one-phase Hele–Shaw flows and a level-set method for nonstandard configurations. The ANZIAM Journal, 63(3): 269–307, July 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/review-of-onephase-heleshaw-flows-and-
a-levelset-method-for-nonstandard-configurations/5635F5A2D0A79B7537B75237A26F460A.

Zhang:2021:OPC

[1869] Caibin Zhang, Zhibin Liang, and Kam Chuen Yuen. Optimal portfolio and consumption for a Markovian regime-switching jump-diffusion process. The ANZIAM Journal, 63(3):308–332, July 2021. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/optimal-portfolio-and-consumption-for-a-markovian-regimeswitching-
jumpdiffusion-process/7B6A5233C88126866FE0CFD120F4C099.

Patkowski:2021:NAD

[1870] Alexander E. Patkowski. A note on the axisymmetric diffusion equation. The ANZIAM Journal, 63(3):333–341, July 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.

cambridge.org/core/journals/anziam-journal/article/note-on-
the-axisymmetric-diffusion-equation/838E1BD8FD008F431BBCCAE130AFD001.

Jena:2021:ISS

[1871] J. Jena and S. Mittal. Interaction of a singular surface with a strong shock in the interstellar gas clouds. The ANZIAM Journal, 63(3):342–358, July 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/interaction-of-a-singular-surface-with-a-strong-shock-
in-the-interstellar-gas-clouds/EE86C2AB93BC18E2382293A2DC3CC185.

Mijiddorj:2021:ACI

[1872] R. Mijiddorj and T. Zhanlav. Algorithm to construct integro splines. The ANZIAM Journal, 63(3):359–375, July 2021. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/algorithm-to-construct-integro-splines/62FDB3DFDC4CA29554805B6BBE35476B.

Anonymous:2021:AVIe

[1873] Anonymous. Anz volume 63 issue 3 cover and front matter. The ANZIAM Journal, 63(3):f1–f2, July 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-63-issue-3-
cover-and-front-matter/44D7262DD175351AE69E42DBAE51D649.

Anonymous:2021:AVIf

[1874] Anonymous. Anz volume 63 issue 3 cover and back matter. The ANZIAM Journal, 63(3):b1–b6, July 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-63-issue-3-
cover-and-back-matter/38881927161E20E3A8EA6AD8CCAC436B.

Forbes:2021:IPF

[1875] Lawrence K. Forbes, Stephen J. Walters, and Graeme C. Hocking. Ideal planar fluid flow over a submerged obstacle: review and extension. The ANZIAM Journal, 63(4):377–419, October 2021. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/ideal-planar-fluid-flow-over-a-submerged-obstacle-review-
and-extension/B051AFF35F8924D37BAB46BD4530F642.

Maldon:2021:NSF

[1876] Benjamin Maldon, Bishnu Prasad Lamichhane, and Ngamta Thamwattana. Numerical solutions to a fractional diffusion equation used in modelling dye-sensitized solar cells. The ANZIAM Journal, 63(4):420– 433, October 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/numerical-solutions-to-a-fractional-
diffusion-equation-used-in-modelling-dyesensitized-solar-cells/
6EAA195EA8444D7734F4DFC14887F9AA.

Ndisabiye:2021:DPE

[1877] D. Ndisabiye, E. K. Waters, R. Gore, and H. Sidhu. Do poor environmental conditions drive trachoma transmission in Burundi? A mathematical modelling study. The ANZIAM Journal, 63(4): 434–447, October 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/do-poor-environmental-
conditions-drive-trachoma-transmission-in-burundi-a-mathematical-
modelling-study/CF47C724A5F08E71F88A6B0B8CF6A132.

Rodrigo:2021:BCT

[1878] Marianito R. Rodrigo. Bounds on the critical times for the general Fisher–KPP equation. The ANZIAM Journal, 63(4):448–468, October 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/bounds-on-the-critical-times-for-the-general-fisherkpp-
equation/37BC9FCEFDB71B73C13CDCFF581DAC67.

Assari:2021:MLG

[1879] P. Assari, F. Asadi-Mehregan, and M. Dehghan. A meshless local Galerkin integral equation method for solving a type of Darboux problems based on radial basis functions. The ANZIAM Journal, 63(4):469– 492, October 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/meshless-local-galerkin-integral-equation-
method-for-solving-a-type-of-darboux-problems-based-on-radial-
basis-functions/672085E24DB76E601E82D989B1684830.

Anonymous:2021:I

[1880] Anonymous. Index. The ANZIAM Journal, 63(4):493–494, October 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/index/142F8A7F595FCF2A26CADE02E1C8A05E.

Anonymous:2021:AVIg

[1881] Anonymous. Anz volume 63 issue 4 cover and front matter. The ANZIAM Journal, 63(4):f1–f2, October 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-63-issue-4-
cover-and-front-matter/48EF73948E4068E44D3448E5D16BE7E0.

Anonymous:2021:AVIh

[1882] Anonymous. Anz volume 63 issue 4 cover and back matter. The ANZIAM Journal, 63(4):b1–b6, October 2021. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/anz-volume-63-issue-4-
cover-and-back-matter/8E2F13D2A05CAA561D59E54C7DF3CE22.

Mansoor:2022:DHR

[1883] W. F. Mansoor, G. C. Hocking, and D. E. Farrow. Dispersal of hydrogen in the retina – a three-layer model. The ANZIAM Journal, 64(1):1–22, January 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/dispersal-of-hydrogen-
in-the-retinaa-threelayer-model/2F9360935F6C872508897500DC86485B.

Celik:2022:HBA

[1884] C. Celik and K. Degerli. Hopf bifurcation analysis of a fractional-order Holling–Tanner predator-prey model with time delay. The ANZIAM Journal, 64(1):23–39, January 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/hopf-bifurcation-analysis-
of-a-fractionalorder-hollingtanner-predatorprey-model-with-
time-delay/06FE7A8278ABDA4B68E77AD8E2A20F21.

Farhane:2022:TDA

[1885] M. Farhane, O. Alehyane, and O. Souhar. Three-dimensional analytical solution of the advection–diffusion equation for air pollution dispersion. The ANZIAM Journal, 64(1):40–53, January 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/threedimensional-analytical-solution-of-the-advectiondiffusion-
equation-for-air-pollution-dispersion/533FE5C693AAFFE0B6494450C2E72B2C.

Asanjarani:2022:SMA

[1886] Azam Asanjarani and Yoni Nazarathy. Stationary Markovian arrival processes: results and open problems. The ANZIAM Journal, 64(1): 54–68, January 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/stationary-markovian-arrival-processes-
results-and-open-problems/B037C27C281FB3B22AA591EEB0B5CF99.

Ramazani:2022:ENS

[1887] P. Ramazani, A. Abdi, G. Hojjati, and A. Moradi. Explicit Nordsieck second derivative general linear methods for ODEs. The ANZIAM Journal, 64(1):69–88, January 2022. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/explicit-nordsieck-second-derivative-
general-linear-methods-for-odes/AC3DE8ABE3934F147C2ED4C587A597C8.

Anonymous:2022:AVIa

[1888] Anonymous. ANZ volume 64 issue 1 cover and front matter. The ANZIAM Journal, 64(1):f1–f2, January 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-64-issue-1-cover-and-front-matter/D49812AEC623AEFE5F05D96978CB2069.

Anonymous:2022:AVIb

[1889] Anonymous. ANZ volume 64 issue 1 cover and back matter. The ANZIAM Journal, 64(1):b1–b6, January 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-64-issue-1-cover-and-back-matter/A17455627034CD7ADF62A4D7C934350A.

Bassom:2022:E

[1890] Andrew P. Bassom. Editorial. The ANZIAM Journal, 64(2):89, April 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/editorial/892C9EC49B3B6567D7A31BC20A420087.

Broadbridge:2022:MSR

[1891] Philip Broadbridge, Ravindi Nanayakkara, and Andriy Olenko. On multifractionality of spherical random fields with cosmological applications. The ANZIAM Journal, 64(2):90–118, April 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
multifractionality-of-spherical-random-fields-with-cosmological-
applications/D4B93E44BE2177796411597F734163AA.

Charkhgard:2022:MNS

[1892] Hadi Charkhgard, Kimia Keshanian, Rasul Esmaeilbeigi, and Parisa Charkhgard. The magic of Nash social welfare in optimization: do not sum, just multiply! The ANZIAM Journal, 64(2):119–134, April 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/magic-of-nash-social-welfare-in-optimization-do-not-
sum-just-multiply/A694F280B86DE52F9BF9A77A941B682F.

Goard:2022:AAC

[1893] Joanna Goard. An analytical approximation for convertible bonds. The ANZIAM Journal, 64(2):135–148, April 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
analytical-approximation-for-convertible-bonds/84D20455B8D1D17E8854A3114B46B400.

Forbes:2022:FFO

[1894] Lawrence K. Forbes and Stephen J. Walters. Fully 3d fluid outflow from a spherical source. The ANZIAM Journal, 64(2):149–182, April 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/fully-3d-fluid-outflow-from-a-spherical-source/3014567EA260400E05B2FAB99E51999C.

Jain:2022:BTB

[1895] Ekta Jain, Kalpana Dahiya, and Vanita Verma. Branching technique for a bi-objective two-stage assignment problem. The ANZIAM Journal, 64 (2):183–204, April 2022. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/branching-technique-for-a-biobjective-
twostage-assignment-problem/361CC417AF26A95A3A436B1A473F051A.

Anonymous:2022:AVIc

[1896] Anonymous. ANZ volume 64 issue 2 cover and front matter. The ANZIAM Journal, 64(2):f1–f2, April 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-64-issue-2-cover-and-front-matter/2CEA361ECCD1ADA513C9A3066F772BAD.

Anonymous:2022:AVId

[1897] Anonymous. ANZ volume 64 issue 2 cover and back matter. The ANZIAM Journal, 64(2):b1–b6, April 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-64-issue-2-cover-and-back-matter/D54A56A4BE0D52CF7443BA1C409A7D96.

Lamichhane:2022:LPS

[1898] Bishnu P. Lamichhane and Jordan A. Shaw-Carmody. A local projection stabilization for convection–diffusion–reaction equations using biorthogonal systems. The ANZIAM Journal, 64(3):205–226, July 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/local-projection-stabilization-for-convectiondiffusionreaction-
equations-using-biorthogonal-systems/0114DBB3537D7238420DFF0CE1CE1F82.

Walters:2022:CER

[1899] Stephen J. Walters, Ross J. Turner, and Lawrence K. Forbes. A comparison of explicit Runge–Kutta methods. The ANZIAM Journal, 64(3): 227–249, July 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/comparison-of-explicit-rungekutta-methods/
323422936943F471AC1F1E77401F87C9.

He:2022:VSV

[1900] Xin-Jiang He and Sha Lin. Volatility swaps valuation under a modified risk-neutralized Heston model with a stochastic long-run variance level. The ANZIAM Journal, 64(3):250–263, July 2022. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/volatility-swaps-valuation-under-a-modified-riskneutralized-
heston-model-with-a-stochastic-longrun-variance-level/E57F0F4DF1FB72E2917F423797D9C4D9.

Moradi:2022:HOE

[1901] A. Moradi, A. Abdi, and G. Hojjati. High order explicit second derivative methods with strong stability properties based on Taylor series conditions. The ANZIAM Journal, 64(3):264–291, July 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/high-order-explicit-second-derivative-methods-with-
strong-stability-properties-based-on-taylor-series-conditions/
EEB2EE3CA24E8E3E4E89DEFA05BD3252.

Halder:2022:MSE

[1902] Sourav Halder and Asoke Kumar Dhar. A modification to the Schrödinger equation for broader bandwidth gravity–capillary waves on deep water with depth-uniform current. The ANZIAM Journal, 64(3): 292–313, July 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/modification-to-the-schrodinger-equation-
for-broader-bandwidth-gravitycapillary-waves-on-deep-water-
with-depthuniform-current/872A42D9D9CA62B5D1B2ECB39BE19726.

Anonymous:2022:AVIe

[1903] Anonymous. ANZ volume 64 issue 3 cover and front matter. The ANZIAM Journal, 64(3):f1–f2, July 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-64-issue-3-cover-and-front-matter/64D21455499E64A14C318FBB31D54156.

Anonymous:2022:AVIf

[1904] Anonymous. ANZ volume 64 issue 3 cover and back matter. The ANZIAM Journal, 64(3):b1–b5, July 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-64-issue-3-cover-and-back-matter/DD86C47A008D81709BC052406F23A8BC.

Mccaw:2022:RMS

[1905] James M. Mccaw and Michael J. Plank. The role of the mathematical sciences in supporting the COVID-19 response in Australia and New Zealand. The ANZIAM Journal, 64(4):315–337, October 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/role-of-the-mathematical-sciences-in-supporting-the-
covid19-response-in-australia-and-new-zealand/BF3ACAFB57E833C3D0054B8A28103724.

Broadbridge:2022:ESH

[1906] P. Broadbridge and J. Goard. Exact solutions of hyperbolic reaction– diffusion equations in two dimensions. The ANZIAM Journal, 64(4):338– 354, October 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-

8735 (electronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/exact-solutions-of-hyperbolic-reactiondiffusion-
equations-in-two-dimensions/B319971BF6D799812CC2E79A0744057F.

Tran-Duc:2022:NSL

[1907] Thien Tran-Duc, Michael H. Meylan, and Ngamta Thamwattana. Numerical simulations for largely deformed beams and rings adopting a nontensile smoothed particle hydrodynamics algorithm. The ANZIAM Journal, 64(4):355–379, October 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/numerical-simulations-
for-largely-deformed-beams-and-rings-adopting-a-nontensile-
smoothed-particle-hydrodynamics-algorithm/817BC97B3D31F9200FEA99646D886B02.

Chen:2022:IBA

[1908] Wenting Chen and Xiaoying Jiang. An IMEX-based approach for the pricing of equity warrants under fractional Brownian motion models. The ANZIAM Journal, 64(4):380–393, October 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/an-
imexbased-approach-for-the-pricing-of-equity-warrants-under-
fractional-brownian-motion-models/A2DB9784A2C32F874A8AA309C4FCC980.

Shaikhet:2022:NSC

[1909] Leonid Shaikhet and Syed Abbas. Novel stability conditions for some generalization of Nicholson’s blowflies model with stochastic perturbations. The ANZIAM Journal, 64(4):394–405, October 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/novel-stability-conditions-for-some-generalization-
of-nicholsons-blowflies-model-with-stochastic-perturbations/
40735DE74090E2B6B328ACF9B573F5C2.

Anonymous:2022:I

[1910] Anonymous. Index. The ANZIAM Journal, 64(4):406–407, October 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/index/8627C490C43A0302EE86907D5FDF9890.

Anonymous:2022:AVIg

[1911] Anonymous. ANZ volume 64 issue 4 cover and front matter. The ANZIAM Journal, 64(4):f1–f2, October 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-64-issue-4-cover-and-front-matter/27B52D8FC0FB4B03104323A701C4E164.

Anonymous:2022:AVIh

[1912] Anonymous. ANZ volume 64 issue 4 cover and back matter. The ANZIAM Journal, 64(4):b1–b5, October 2022. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-64-issue-4-cover-and-back-matter/9D91CE3D25705A5B214C9EFF7795B76B.

Bassom:2023:ESI

[1913] Andrew P. Bassom and Michael H. Meylan. Editorial: special issue in honour of Professor Graeme Hocking. The ANZIAM Journal, 65 (1–2):1–2, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/core/
journals/anziam-journal/article/editorial-special-issue-in-
honour-of-professor-graeme-hocking/5F10F070A90FD929537BEC13E468990A.

Forbes:2023:CFH

[1914] Lawrence K. Forbes and Stephen J. Walters. Couette flow over a heat island. The ANZIAM Journal, 65(1–2):3–28, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/couette-flow-over-a-heat-island/61DA21D1AFC6B1447B9BAE5B214E5F3D.

Fowkes:2023:DLC

[1915] N. D. Fowkes and D. P. Mason. Double layered compressible masks. The ANZIAM Journal, 65(1–2):29–54, January 2023. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/double-layered-compressible-masks/F085F52BB4D39A3822247EBFAE49F5EC.

Mason:2023:WSM

[1916] D. P. Mason, N. D. Fowkes, R. M. Yemata, C. A Onyeagoziri, and H. Yilmaz. Wall stabilization in mines by spray-on liners. The ANZIAM Journal, 65(1–2):55–78, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/wall-stabilization-in-
mines-by-sprayon-liners/E1BB6062FF14E13B7066B6E744AA8A6B.

Mitchell:2023:SSB

[1917] S. L. Mitchell and T. G. Myers. On the safe storage of bagasse. The ANZIAM Journal, 65(1–2):79–92, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https:/

/www.cambridge.org/core/journals/anziam-journal/article/on-
the-safe-storage-of-bagasse/596DFF141BBA8F6CB5E7FE6224408539.

Fowler:2023:BSR

[1918] A. C. Fowler and M. J. Mcguinness. Bursting solutions of the Rössler equations. The ANZIAM Journal, 65(1–2):93–110, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/bursting-solutions-of-the-rossler-equations/0A15E45019AF4A2BEB32FF67378B8409.

Paea:2023:MMB

[1919] Lata I. Paea, Sione Paea, and Mark J. Mcguinness. Modelling microwaves in bauxite. The ANZIAM Journal, 65(1–2):111–134, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/modelling-microwaves-in-bauxite/531E5CE8DA0636FDDC3A61AC1170B67E.

Maldon:2023:FLI

[1920] B. Maldon and Michael H. Meylan. Flight limitations imposed on single rotor and coaxial helicopters by the lift equation. The ANZIAM Journal, 65(1–2):135–154, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/flight-limitations-imposed-
on-single-rotor-and-coaxial-helicopters-by-the-lift-equation/
2D6BE7B234C138960C01B240078F4489.

Stokes:2023:ELW

[1921] Y. M. Stokes, W. L. Sweatman, and G. C. Hocking. An examination of the ‘Lanier wing’ design. The ANZIAM Journal, 65 (1–2):155–177, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/an-examination-of-the-
lanier-wing-design/43322F9070C7025883498D6DE06E475B.

Zhang:2023:NAA

[1922] H. Zhang, D. A. Barry, B. Seymour, and G. Hocking. Numerical analysis of apparatus-induced dispersion for density-dependent solute transport in porous media. The ANZIAM Journal, 65(1– 2):178–194, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/numerical-analysis-of-
apparatusinduced-dispersion-for-densitydependent-solute-transport-
in-porous-media/CE8FA2DC92420FBB84BB49E884CE7520.

Anonymous:2023:AVIa

[1923] Anonymous. ANZ volume 65 issue 1-2 cover and front matter. The ANZIAM Journal, 65(1–2):f1–f2, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-65-issue-12-cover-and-front-matter/FA6AC399E4CCCA0F6DF9DF9D5B0F8931.

Anonymous:2023:AVIb

[1924] Anonymous. ANZ volume 65 issue 1-2 cover and back matter. The ANZIAM Journal, 65(1–2):b1–b8, January 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-65-issue-12-cover-and-back-matter/485E3F466B074D6B9F43BC73BB439F24.

Germano:2023:ART

[1925] Domenic P. J. Germano, Stephanie Khuu, Adrianne L. Jenner, James M. Osborne, Mary R. Myerscough, and Mark B. Flegg. Active remodelling of tissues to describe biphasic rheological responses. The ANZIAM Journal, 65(3):195–214, July 2023. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-journal/
article/active-remodelling-of-tissues-to-describe-biphasic-
rheological-responses/08AA9942DA80600911A62AA1A6CC8F86.

Wang:2023:MOP

[1926] Wen-Hua Wang, Zheng-Li Chen, and Wei Li. The metric operators for pseudo-Hermitian Hamiltonian. The ANZIAM Journal, 65(3):215–228, July 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/metric-operators-for-pseudohermitian-hamiltonian/
58DC3753495166AFAE88CDDE75F9B680.

Han:2023:APD

[1927] Y. Han and X. Zheng. Approximate pricing of derivatives under fractional stochastic volatility model. The ANZIAM Journal, 65(3):229–247, July 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/approximate-pricing-of-derivatives-under-fractional-
stochastic-volatility-model/9AF149DAC48777104FFFDD16F2497C2C.

Pal:2023:NSM

[1928] Tanmoy Pal and Asoke Kumar Dhar. Nonlinear self-modulation of gravity-capillary waves on shear currents in finite depth. The ANZIAM Journal, 65(3):248–272, July 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://www.cambridge.org/

core/journals/anziam-journal/article/nonlinear-selfmodulation-
of-gravitycapillary-waves-on-shear-currents-in-finite-depth/
77AC8082A9B7E45F4170B4E422D928DF.

Mansour:2023:MWQ

[1929] Mahmoud B. A. Mansour. On modelling water quality with stochastic differential equations. The ANZIAM Journal, 65(3):273–284, July 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/anziam-
journal/article/on-modelling-water-quality-with-stochastic-
differential-equations/3DCDA8A8747172FA45C0CFC0288CFD3C.

Anonymous:2023:AVIc

[1930] Anonymous. ANZ volume 65 issue 3 cover and front matter. The ANZIAM Journal, 65(3):f1–f2, July 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-65-issue-3-cover-and-front-matter/8C30933CBAB9D1155D42DE6AEEED832F.

Anonymous:2023:AVId

[1931] Anonymous. ANZ volume 65 issue 3 cover and back matter. The ANZIAM Journal, 65(3):b1–b6, July 2023. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/anz-
volume-65-issue-3-cover-and-back-matter/63A19947A51ABF3748A38A83B8159EA6.